GNU Emacs Manual

GNU Emacs Manual

Thirteenth Edition, Updated for Emacs Version 20.3 Richard Stallman Copyright \begin{matrix}c \\ ◯ \end{matrix} 1985, 1986, 1987, 1993, 1994, 1995, 1996, 1997, 1998 Free Software Foundation, Inc.

Thirteenth Edition Updated for Emacs Version 20.3, August 1998 ISBN 1-882114-06-X

Published by the Free Software Foundation 59 Temple Place, Suite 330 Boston, MA 02111-1307 USA

Permission is granted to make and distribute verbatim copies of this manual provided the copyright notice and this permission notice are preserved on all copies. Permission is granted to copy and distribute modified versions of this manual under the conditions for verbatim copying, provided also that the sections entitled “The GNU Manifesto”, “Distribution” and “GNU General Public License” are included exactly as in the original, and provided that the entire resulting derived work is distributed under the terms of a permission notice identical to this one. Permission is granted to copy and distribute translations of this manual into another language, under the above conditions for modified versions, except that the sections entitled “The GNU Manifesto”, “Distribution” and “GNU General Public License” may be included in a translation approved by the Free Software Foundation instead of in the original English.

Cover art by Etienne Suvasa.

Preface

This manual documents the use and simple customization of the Emacs editor. The reader is not expected to be a programmer; simple customizations do not require programming skill. But the user who is not interested in customizing can ignore the scattered customization hints.

This is primarily a reference manual, but can also be used as a primer. For complete beginners, it is a good idea to start with the on-line, learn-by-doing tutorial, before reading the manual. To run the tutorial, start Emacs and type C-h t. This way you can learn Emacs by using Emacs on a specially designed file which describes commands, tells you when to try them, and then explains the results you see.

On first reading, just skim chapters 1 and 2, which describe the notational conventions of the manual and the general appearance of the Emacs display screen. Note which questions are answered in these chapters, so you can refer back later. After reading chapter 4, you should practice the commands there. The next few chapters describe fundamental techniques and concepts that are used constantly. You need to understand them thoroughly, experimenting with them if necessary.

Chapters 14 through 19 describe intermediate-level features that are useful for all kinds of editing. Chapter 20 and following chapters describe features that you may or may not want to use; read those chapters when you need them.

Read the Trouble chapter if Emacs does not seem to be working properly. It explains how to cope with some common problems (see Section 32.2 [Lossage], page 418), as well as when and how to report Emacs bugs (see Section 32.3 [Bugs], page 422).

To find the documentation on a particular command, look in the index. Keys (character commands) and command names have separate indexes. There is also a glossary, with a cross reference for each term.

This manual is available as a printed book and also as an Info file. The Info file is for on-line perusal with the Info program, which will be the principal way of viewing documentation on-line in the GNU system. Both the Info file and the Info program itself are distributed along with GNU Emacs. The Info file and the printed book contain substantially the same text and are generated from the same source files, which are also distributed along with GNU Emacs.

GNU Emacs is a member of the Emacs editor family. There are many Emacs editors, all sharing common principles of organization. For information on the underlying philosophy of Emacs and the lessons learned from its development, write for a copy of AI memo 519a, “Emacs, the Extensible, Customizable Self-Documenting Display Editor,” to Publications Department, Artificial Intelligence Lab, 545 Tech Square, Cambridge, MA 02139, USA. At last report they charge $2.25 per copy. Another useful publication is LCS TM-165, “A Cookbook for an Emacs,” by Craig Finseth, available from Publications Department, Laboratory for Computer Science, 545 Tech Square, Cambridge, MA 02139, USA. The price today is $3.

This edition of the manual is intended for use with GNU Emacs installed on GNU and Unix systems. GNU Emacs can also be used on VMS, MS-DOS (also called MS-DOG), Windows NT, and Windows 95 systems. Those systems use different file name syntax; in addition, VMS and MS-DOS do not support all GNU Emacs features. We don’t try to describe VMS usage in this manual. See Appendix C [MS-DOS], page 455, for information about using Emacs on MS-DOS.

Distribution

GNU Emacs is free software; this means that everyone is free to use it and free to redistribute it on certain conditions. GNU Emacs is not in the public domain; it is copyrighted and there are restrictions on its distribution, but these restrictions are designed to permit everything that a good cooperating citizen would want to do. What is not allowed is to try to prevent others from further sharing any version of GNU Emacs that they might get from you. The precise conditions are found in the GNU General Public License that comes with Emacs and also appears following this section.

One way to get a copy of GNU Emacs is from someone else who has it. You need not ask for our permission to do so, or tell any one else; just copy it. If you have access to the Internet, you can get the latest distribution version of GNU Emacs by anonymous FTP; see the file ‘etc/FTP’ in the Emacs distribution for more information.

You may also receive GNU Emacs when you buy a computer. Computer manufacturers are free to distribute copies on the same terms that apply to everyone else. These terms require them to give you the full sources, including whatever changes they may have made, and to permit you to redistribute the GNU Emacs received from them under the usual terms of the General Public License. In other words, the program must be free for you when you get it, not just free for the manufacturer.

You can also order copies of GNU Emacs from the Free Software Foundation on CD-ROM. This is a convenient and reliable way to get a copy; it is also a good way to help fund our work. (The Foundation has always received most of its funds in this way.) An order form is included in the file ‘etc/ORDERS’ in the Emacs distribution, and on our web site in

http://www.gnu.org/order/order.html. For further information, write

to Free Software Foundation 59 Temple Place, Suite 330 Boston, MA 02111-1307 USA USA The income from distribution fees goes to support the foundation’s pur-

pose: the development of new free software, and improvements to our existing programs including GNU Emacs.

If you find GNU Emacs useful, please send a donation to the Free Software Foundation to support our work. Donations to the Free Software Foundation are tax deductible in the US. If you use GNU Emacs at your workplace, please suggest that the company make a donation. If company policy is unsympathetic to the idea of donating to charity, you might instead suggest ordering a CD-ROM from the Foundation occasionally, or subscribing to periodic updates.

Contributors to GNU Emacs include Per Abrahamsen, Jay K. Adams, Joe Arceneaux, Boaz Ben-Zvi, Jim Blandy, Terrence Brannon, Frank Bresz, Peter Breton, Kevin Broadey, Vincent Broman, David M. Brown, Hans Chalupsky, Bob Chassell, James Clark, Mike Clarkson, Andrew Csillag, Doug Cutting, Michael DeCorte, Gary Delp, Matthieu Devin, Eri Ding, Carsten Dominik, Scott Draves, Viktor Dukhovni, John Eaton, Rolf Ebert, Stephen Eglen, Torbjörn Einarsson, Tsugumoto Enami, Hans Henrik Eriksen, Michael Ernst, Ata Etemadi, Frederick Farnback, Fred Fish, Karl Fogel, Gary Foster, Noah Friedman, Keith Gabryelski, Kevin Gallagher, Kevin Gallo, Howard Gayle, Stephen Gildea, David Gillespie, Boris Goldowsky, Michelangelo Grigni, Michael Gschwind, Henry Guillaume, Doug Gwyn, Ken’ichi Handa , Chris Hanson, K. Shane Hartman, John Heidemann, Markus Heritsch, Karl Heuer, Manabu Higashida, Anders Holst, Kurt Hornik, Tom Houlder, Lars Ingebrigtsen, Andrew Innes, Michael K. Johnson, Kyle Jones, Tomoji Kagatani, Brewster Kahle, David Kaufman, Henry Kautz, Howard Kaye, Michael Kifer, Richard King, Larry K. Kolodney, Robert Krawitz, Sebastian Kremer, Geoff Kuenning, David Kågedal, Daniel LaLiberte, Aaron Larson, James R. Larus, Frederic Lepied, Lars Lindberg, Neil M. Mager, Ken Manheimer, Bill Mann, Brian Marick, Simon Marshall, Bengt Martensson, Charlie Martin, Thomas May, Roland McGrath, David Megginson, Wayne Mesard, Richard Mlynarik, Keith Moore, Erik Naggum, Thomas Neumann, Mike Newton, Jurgen Nickelsen, Jeff Norden, Andrew Norman, Jeff Peck, Damon Anton Permezel, Tom Perrine, Daniel Pfeiffer, Fred Pierresteguy, Christian Plaunt, Francesco A. Potorti, Michael D. Prange, Ashwin Ram, Eric S. Raymond, Paul Reilly, Edward M. Reingold, Rob Riepel, Roland B. Roberts, John Robinson, Danny Roozendaal, William Rosenblatt, Guillermo J. Rozas, Ivar Rummelhoff, Wolfgang Rupprecht, James B. Salem, Masahiko Sato, William Schelter, Ralph Schleicher, Gregor Schmid, Michael Schmidt, Ronald S. Schnell, Philippe Schnoebelen, Stephen Schoef, Randal Schwartz, Stanislav Shalunov, Mark Shapiro, Olin Shivers, Sam Shteingold, Espen Skoglund, Rick Sladkey, Lynn Slater, Chris Smith, David Smith, William Sommerfeld, Michael Staats, Ake Stenhoff, Peter Stephenson, Jonathan Stigelman, Steve Strassman, Jens T. Berger Thielemann, Spencer Thomas, Jim Thompson, Masanobu Umeda, Neil W. Van Dyke, Ulrik Vieth, Geoffrey Voelker, Johan Vromans, Barry Warsaw, Morten Welinder, Joseph Brian Wells, Ed Wilkinson, Mike Williams, Steven A. Wood, Dale R. Worley, Felix S. T. Wu, Tom Wurgler, Eli Zaretskii, Jamie Zawinski, and Neal Ziring.

GNU GENERAL PUBLIC LICENSE

Version 2, June 1991

Copyright \begin{matrix}c \\ ◯ \end{matrix} 1989, 1991 Free Software Foundation, Inc. 59 Temple Place, Suite 330, Boston, MA 02111-1307 USA

Everyone is permitted to copy and distribute verbatim copies of this license document, but changing it is not allowed.

Preamble

The licenses for most software are designed to take away your freedom to share and change it. By contrast, the GNU General Public License is intended to guarantee your freedom to share and change free software—to make sure the software is free for all its users. This General Public License applies to most of the Free Software Foundation’s software and to any other program whose authors commit to using it. (Some other Free Software Foundation software is covered by the GNU Library General Public License instead.) You can apply it to your programs, too.

When we speak of free software, we are referring to freedom, not price. Our General Public Licenses are designed to make sure that you have the freedom to distribute copies of free software (and charge for this service if you wish), that you receive source code or can get it if you want it, that you can change the software or use pieces of it in new free programs; and that you know you can do these things.

To protect your rights, we need to make restrictions that forbid anyone to deny you these rights or to ask you to surrender the rights. These restrictions translate to certain responsibilities for you if you distribute copies of the software, or if you modify it.

For example, if you distribute copies of such a program, whether gratis or for a fee, you must give the recipients all the rights that you have. You must make sure that they, too, receive or can get the source code. And you must show them these terms so they know their rights.

We protect your rights with two steps: (1) copyright the software, and (2) offer you this license which gives you legal permission to copy, distribute and/or modify the software.

Also, for each author’s protection and ours, we want to make certain that everyone understands that there is no warranty for this free software. If the software is modified by someone else and passed on, we want its recipients to know that what they have is not the original, so that any problems introduced by others will not reflect on the original authors’ reputations.

Finally, any free program is threatened constantly by software patents. We wish to avoid the danger that redistributors of a free program will individually obtain patent licenses, in effect making the program proprietary.

To prevent this, we have made it clear that any patent must be licensed for everyone’s free use or not licensed at all.

The precise terms and conditions for copying, distribution and modification follow.

TERMS AND CONDITIONS FOR COPYING, DISTRIBUTION AND MODIFICATION

0. This License applies to any program or other work which contains a notice placed by the copyright holder saying it may be distributed under the terms of this General Public License. The “Program”, below, refers to any such program or work, and a “work based on the Program” means either the Program or any derivative work under copyright law: that is to say, a work containing the Program or a portion of it, either verbatim or with modifications and/or translated into another language. (Hereinafter, translation is included without limitation in the term “modification”.) Each licensee is addressed as “you”. Activities other than copying, distribution and modification are not covered by this License; they are outside its scope. The act of running the Program is not restricted, and the output from the Program is covered only if its contents constitute a work based on the Program (independent of having been made by running the Program). Whether that is true depends on what the Program does.

1. You may copy and distribute verbatim copies of the Program’s source code as you receive it, in any medium, provided that you conspicuously and appropriately publish on each copy an appropriate copyright notice and disclaimer of warranty; keep intact all the notices that refer to this License and to the absence of any warranty; and give any other recipients of the Program a copy of this License along with the Program. You may charge a fee for the physical act of transferring a copy, and you may at your option offer warranty protection in exchange for a fee.

2. You may modify your copy or copies of the Program or any portion of it, thus forming a work based on the Program, and copy and distribute such modifications or work under the terms of Section 1 above, provided that you also meet all of these conditions: a. You must cause the modified files to carry prominent notices stating

that you changed the files and the date of any change. b. You must cause any work that you distribute or publish, that in

whole or in part contains or is derived from the Program or any part thereof, to be licensed as a whole at no charge to all third parties under the terms of this License.

c. If the modified program normally reads commands interactively when run, you must cause it, when started running for such interactive use in the most ordinary way, to print or display an announcement including an appropriate copyright notice and a notice that there is no warranty (or else, saying that you provide a warranty) and that users may redistribute the program under these conditions, and telling the user how to view a copy of this License. (Exception: if the Program itself is interactive but does not normally print such an announcement, your work based on the Program is not required to print an announcement.)

These requirements apply to the modified work as a whole. If identifiable sections of that work are not derived from the Program, and can be reasonably considered independent and separate works in themselves, then this License, and its terms, do not apply to those sections when you distribute them as separate works. But when you distribute the same sections as part of a whole which is a work based on the Program, the distribution of the whole must be on the terms of this License, whose permissions for other licensees extend to the entire whole, and thus to each and every part regardless of who wrote it. Thus, it is not the intent of this section to claim rights or contest your rights to work written entirely by you; rather, the intent is to exercise the right to control the distribution of derivative or collective works based on the Program. In addition, mere aggregation of another work not based on the Program with the Program (or with a work based on the Program) on a volume of a storage or distribution medium does not bring the other work under the scope of this License.

3. You may copy and distribute the Program (or a work based on it, under Section 2) in object code or executable form under the terms of Sections 1 and 2 above provided that you also do one of the following: a. Accompany it with the complete corresponding machine-readable

source code, which must be distributed under the terms of Sections 1 and 2 above on a medium customarily used for software interchange; or,

b. Accompany it with a written offer, valid for at least three years, to give any third party, for a charge no more than your cost of physically performing source distribution, a complete machine-readable copy of the corresponding source code, to be distributed under the terms of Sections 1 and 2 above on a medium customarily used for software interchange; or,

c. Accompany it with the information you received as to the offer to distribute corresponding source code. (This alternative is allowed only for noncommercial distribution and only if you received the program in object code or executable form with such an offer, in accord with Subsection b above.)

The source code for a work means the preferred form of the work for making modifications to it. For an executable work, complete source code means all the source code for all modules it contains, plus any associated interface definition files, plus the scripts used to control compilation and installation of the executable. However, as a special exception, the source code distributed need not include anything that is normally distributed (in either source or binary form) with the major components (compiler, kernel, and so on) of the operating system on which the executable runs, unless that component itself accompanies the executable. If distribution of executable or object code is made by offering access to copy from a designated place, then offering equivalent access to copy the source code from the same place counts as distribution of the source code, even though third parties are not compelled to copy the source along with the object code.

4. You may not copy, modify, sublicense, or distribute the Program except as expressly provided under this License. Any attempt otherwise to copy, modify, sublicense or distribute the Program is void, and will automatically terminate your rights under this License. However, parties who have received copies, or rights, from you under this License will not have their licenses terminated so long as such parties remain in full compliance.

5. You are not required to accept this License, since you have not signed it. However, nothing else grants you permission to modify or distribute the Program or its derivative works. These actions are prohibited by law if you do not accept this License. Therefore, by modifying or distributing the Program (or any work based on the Program), you indicate your acceptance of this License to do so, and all its terms and conditions for copying, distributing or modifying the Program or works based on it.

6. Each time you redistribute the Program (or any work based on the Program), the recipient automatically receives a license from the original licensor to copy, distribute or modify the Program subject to these terms and conditions. You may not impose any further restrictions on the recipients’ exercise of the rights granted herein. You are not responsible for enforcing compliance by third parties to this License.

7. If, as a consequence of a court judgment or allegation of patent infringement or for any other reason (not limited to patent issues), conditions are imposed on you (whether by court order, agreement or otherwise) that contradict the conditions of this License, they do not excuse you from the conditions of this License. If you cannot distribute so as to satisfy simultaneously your obligations under this License and any other pertinent obligations, then as a consequence you may not distribute the Program at all. For example, if a patent license would not permit royalty-free redistribution of the Program by all those who receive copies directly or indirectly through you, then the only way you could satisfy both it and this License would be to refrain entirely from distribution of the Program.

If any portion of this section is held invalid or unenforceable under any particular circumstance, the balance of the section is intended to apply and the section as a whole is intended to apply in other circumstances. It is not the purpose of this section to induce you to infringe any patents or other property right claims or to contest validity of any such claims; this section has the sole purpose of protecting the integrity of the free software distribution system, which is implemented by public license practices. Many people have made generous contributions to the wide range of software distributed through that system in reliance on consistent application of that system; it is up to the author/donor to decide if he or she is willing to distribute software through any other system and a licensee cannot impose that choice. This section is intended to make thoroughly clear what is believed to be a consequence of the rest of this License.

8. If the distribution and/or use of the Program is restricted in certain countries either by patents or by copyrighted interfaces, the original copyright holder who places the Program under this License may add an explicit geographical distribution limitation excluding those countries, so that distribution is permitted only in or among countries not thus excluded. In such case, this License incorporates the limitation as if written in the body of this License.

9. The Free Software Foundation may publish revised and/or new versions of the General Public License from time to time. Such new versions will be similar in spirit to the present version, but may differ in detail to address new problems or concerns. Each version is given a distinguishing version number. If the Program specifies a version number of this License which applies to it and “any later version”, you have the option of following the terms and conditions either of that version or of any later version published by the Free Software Foundation. If the Program does not specify a version number of this License, you may choose any version ever published by the Free Software Foundation.

10. If you wish to incorporate parts of the Program into other free programs whose distribution conditions are different, write to the author to ask for permission. For software which is copyrighted by the Free Software Foundation, write to the Free Software Foundation; we sometimes make exceptions for this. Our decision will be guided by the two goals of preserving the free status of all derivatives of our free software and of promoting the sharing and reuse of software generally.

NO WARRANTY

11. BECAUSE THE PROGRAM IS LICENSED FREE OF CHARGE, THERE IS NO WARRANTY FOR THE PROGRAM, TO THE EX-TENT PERMITTED BY APPLICABLE LAW. EXCEPT WHEN OTHERWISE STATED IN WRITING THE COPYRIGHT HOLDERS AND/OR OTHER PARTIES PROVIDE THE PROGRAM “AS IS” WITHOUT WARRANTY OF ANY KIND, EITHER EXPRESSED OR IMPLIED, INCLUDING, BUT NOT LIMITED TO, THE IMPLIED WARRANTIES OF MERCHANTABILITY AND FITNESS FOR A PARTICULAR PURPOSE. THE ENTIRE RISK AS TO THE QUAL-ITY AND PERFORMANCE OF THE PROGRAM IS WITH YOU. SHOULD THE PROGRAM PROVE DEFECTIVE, YOU ASSUME THE COST OF ALL NECESSARY SERVICING, REPAIR OR COR-RECTION.

12. IN NO EVENT UNLESS REQUIRED BY APPLICABLE LAW OR AGREED TO IN WRITING WILL ANY COPYRIGHT HOLDER, OR ANY OTHER PARTY WHO MAY MODIFY AND/OR REDIS-TRIBUTE THE PROGRAM AS PERMITTED ABOVE, BE LIABLE TO YOU FOR DAMAGES, INCLUDING ANY GENERAL, SPECIAL, INCIDENTAL OR CONSEQUENTIAL DAMAGES ARISING OUT OF THE USE OR INABILITY TO USE THE PROGRAM (INCLUD-ING BUT NOT LIMITED TO LOSS OF DATA OR DATA BEING RENDERED INACCURATE OR LOSSES SUSTAINED BY YOU OR THIRD PARTIES OR A FAILURE OF THE PROGRAM TO OPER-ATE WITH ANY OTHER PROGRAMS), EVEN IF SUCH HOLDER OR OTHER PARTY HAS BEEN ADVISED OF THE POSSIBILITY OF SUCH DAMAGES.

END OF TERMS AND CONDITIONS How to Apply These Terms to Your New Programs

If you develop a new program, and you want it to be of the greatest possible use to the public, the best way to achieve this is to make it free software which everyone can redistribute and change under these terms.

To do so, attach the following notices to the program. It is safest to attach them to the start of each source file to most effectively convey the exclusion of warranty; and each file should have at least the “copyright” line and a pointer to where the full notice is found.

one line to give the program’s name and an idea of what it does.

     Copyright (C) 19yy name of author
     This program is free software; you can redistribute it and/or
     modify it under the terms of the GNU General Public License
     as published by the Free Software Foundation; either version 2
     of the License, or (at your option) any later version.
     This program is distributed in the hope that it will be useful,
     but WITHOUT ANY WARRANTY; without even the implied warranty of
     MERCHANTABILITY or FITNESS FOR A PARTICULAR PURPOSE.  See the
     GNU General Public License for more details.
     You should have received a copy of the GNU General Public License along
     with this program; if not, write to the Free Software Foundation, Inc.,
     59 Temple Place, Suite 330, Boston, MA 02111-1307, USA.

Also add information on how to contact you by electronic and paper mail.

If the program is interactive, make it output a short notice like this when it starts in an interactive mode:

     Gnomovision version 69, Copyright (C) 19yy name of author
     Gnomovision comes with ABSOLUTELY NO WARRANTY; for details
     type `show w'.  This is free software, and you are welcome
     to redistribute it under certain conditions; type `show c'
     for details.

The hypothetical commands ‘show w’ and ‘show c’ should show the appropriate parts of the General Public License. Of course, the commands you use may be called something other than ‘show w’ and ‘show c’; they could even be mouse-clicks or menu items—whatever suits your program.

You should also get your employer (if you work as a programmer) or your school, if any, to sign a “copyright disclaimer” for the program, if necessary. Here is a sample; alter the names:

     Yoyodyne, Inc., hereby disclaims all copyright
     interest in the program `Gnomovision'
     (which makes passes at compilers) written
     by James Hacker.

signature of Ty Coon, 1 April 1989

     Ty Coon, President of Vice

This General Public License does not permit incorporating your program into proprietary programs. If your program is a subroutine library, you may consider it more useful to permit linking proprietary applications with the library. If this is what you want to do, use the GNU Library General Public License instead of this License.

Introduction

You are reading about GNU Emacs, the GNU incarnation of the advanced, self-documenting, customizable, extensible real-time display editor Emacs. (The ‘G’ in ‘GNU’ is not silent.)

We say that Emacs is a display editor because normally the text being edited is visible on the screen and is updated automatically as you type your commands. See Chapter 1 [Screen], page 15.

We call it a real-time editor because the display is updated very frequently, usually after each character or pair of characters you type. This minimizes the amount of information you must keep in your head as you edit. See Chapter 4 [Basic Editing], page 29.

We call Emacs advanced because it provides facilities that go beyond simple insertion and deletion: controlling subprocesses; automatic indentation of programs; viewing two or more files at once; editing formatted text; and dealing in terms of characters, words, lines, sentences, paragraphs, and pages, as well as expressions and comments in several different programming languages.

Self-documenting means that at any time you can type a special character, Control-h, to find out what your options are. You can also use it to find out what any command does, or to find all the commands that pertain to a topic. See Chapter 7 [Help], page 49.

Customizable means that you can change the definitions of Emacs commands in little ways. For example, if you use a programming language in which comments start with ‘<**’ and end with ‘**>’, you can tell the Emacs comment manipulation commands to use those strings (see Section 22.7 [Comments], page 237). Another sort of customization is rearrangement of the command set. For example, if you prefer the four basic cursor motion commands (up, down, left and right) on keys in a diamond pattern on the keyboard, you can rebind the keys that way. See Chapter 31 [Customization], page 383.

Extensible means that you can go beyond simple customization and write entirely new commands, programs in the Lisp language to be run by Emacs’s own Lisp interpreter. Emacs is an “on-line extensible” system, which means that it is divided into many functions that call each other, any of which can be redefined in the middle of an editing session. Almost any part of Emacs can be replaced without making a separate copy of all of Emacs. Most of the editing commands of Emacs are written in Lisp already; the few exceptions could have been written in Lisp but are written in C for efficiency. Although only a programmer can write an extension, anybody can use it afterward. If you want to learn Emacs Lisp programming, we recommend the Introduction to Emacs Lisp by Robert J. Chassell, also published by the Free Software Foundation.

When run under the X Window System, Emacs provides its own menus and convenient bindings to mouse buttons. But Emacs can provide many of the benefits of a window system on a text-only terminal. For instance, you can look at or edit several files at once, move text between files, and edit files while running shell commands.

1 The Organization of the Screen

On a text-only terminal, the Emacs display occupies the whole screen. On the X Window System, Emacs creates its own X windows to use. We use the term frame to mean an entire text-only screen or an entire X window used by Emacs. Emacs uses both kinds of frames in the same way to display your editing. Emacs normally starts out with just one frame, but you can create additional frames if you wish. See Chapter 17 [Frames], page 155.

When you start Emacs, the entire frame except for the first and last lines is devoted to the text you are editing. This area is called the window. The first line is a menu bar, and the last line is a special echo area or minibuffer window where prompts appear and where you can enter responses. See below for more information about these special lines.

You can subdivide the large text window horizontally or vertically into multiple text windows, each of which can be used for a different file (see Chapter 16 [Windows], page 149). In this manual, the word “window” always refers to the subdivisions of a frame within Emacs.

The window that the cursor is in is the selected window, in which editing takes place. Most Emacs commands implicitly apply to the text in the selected window (though mouse commands generally operate on whatever window you click them in, whether selected or not). The other windows display text for reference only, unless/until you select them. If you use multiple frames under the X Window System, then giving the input focus to a particular frame selects a window in that frame.

Each window’s last line is a mode line, which describes what is going on in that window. It appears in inverse video, if the terminal supports that, and its contents begin with ‘--:-- *scratch*’ when Emacs starts. The mode line displays status information such as what buffer is being displayed above it in the window, what major and minor modes are in use, and whether the buffer contains unsaved changes.

1.1 Point

Within Emacs, the terminal’s cursor shows the location at which editing commands will take effect. This location is called point. Many Emacs commands move point through the text, so that you can edit at different places in it. You can also place point by clicking mouse button 1.

While the cursor appears to point at a character, you should think of point as between two characters; it points before the character that appears under the cursor. For example, if your text looks like ‘frob’ with the cursor over the ‘b’, then point is between the ‘o’ and the ‘b’. If you insert the character ‘!’ at that position, the result is ‘fro!b’, with point between the ‘!’ and the ‘b’. Thus, the cursor remains over the ‘b’, as before.

Sometimes people speak of “the cursor” when they mean “point,” or speak of commands that move point as “cursor motion” commands.

Terminals have only one cursor, and when output is in progress it must appear where the typing is being done. This does not mean that point is moving. It is only that Emacs has no way to show you the location of point except when the terminal is idle.

If you are editing several files in Emacs, each in its own buffer, each buffer has its own point location. A buffer that is not currently displayed remembers where point is in case you display it again later.

When there are multiple windows in a frame, each window has its own point location. The cursor shows the location of point in the selected window. This also is how you can tell which window is selected. If the same buffer appears in more than one window, each window has its own position for point in that buffer.

When there are multiple frames, each frame can display one cursor. The cursor in the selected frame is solid; the cursor in other frames is a hollow box, and appears in the window that would be selected if you give the input focus to that frame.

The term ‘point’ comes from the character ‘.’, which was the command in TECO (the language in which the original Emacs was written) for accessing the value now called ‘point’.

1.2 The Echo Area

The line at the bottom of the frame (below the mode line) is the echo area. It is used to display small amounts of text for several purposes.

Echoing means displaying the characters that you type. Outside Emacs, the operating system normally echoes all your input. Emacs handles echoing differently.

Single-character commands do not echo in Emacs, and multi-character commands echo only if you pause while typing them. As soon as you pause for more than a second in the middle of a command, Emacs echoes all the characters of the command so far. This is to prompt you for the rest of the command. Once echoing has started, the rest of the command echoes immediately as you type it. This behavior is designed to give confident users fast response, while giving hesitant users maximum feedback. You can change this behavior by setting a variable (see Section 11.7 [Display Vars], page 81).

If a command cannot be executed, it may print an error message in the echo area. Error messages are accompanied by a beep or by flashing the screen. Also, any input you have typed ahead is thrown away when an error happens.

Some commands print informative messages in the echo area. These messages look much like error messages, but they are not announced with a beep and do not throw away input. Sometimes the message tells you what the command has done, when this is not obvious from looking at the text being edited. Sometimes the sole purpose of a command is to print a message giving you specific information—for example, C-x = prints a message describing the character position of point in the text and its current column in the window. Commands that take a long time often display messages ending in ‘...’ while they are working, and add ‘done’ at the end when they are finished.

Echo-area informative messages are saved in an editor buffer named ‘*Messages*’. (We have not explained buffers yet; see Chapter 15 [Buffers], page 141, for more information about them.) If you miss a message that appears briefly on the screen, you can switch to the ‘*Messages*’ buffer to see it again. (Successive progress messages are often collapsed into one in that buffer.)

The size of ‘*Messages*’ is limited to a certain number of lines. The variable message-log-max specifies how many lines. Once the buffer has that many lines, each line added at the end deletes one line from the beginning. See Section 31.2 [Variables], page 385, for how to set variables such as

message-log-max.

The echo area is also used to display the minibuffer, a window that is used for reading arguments to commands, such as the name of a file to be edited. When the minibuffer is in use, the echo area begins with a prompt string that usually ends with a colon; also, the cursor appears in that line because it is the selected window. You can always get out of the minibuffer by typing C-g. See Chapter 5 [Minibuffer], page 39.

1.3 The Mode Line

Each text window’s last line is a mode line, which describes what is going on in that window. When there is only one text window, the mode line appears right above the echo area; it is the next-to-last line on the frame. The mode line is in inverse video if the terminal supports that, and it starts and ends with dashes.

Normally, the mode line looks like this:

-cs:ch buf (major minor)--line--pos------ This gives information about the buffer being displayed in the window: the buffer’s name, what major and minor modes are in use, whether the buffer’s text has been changed, and how far down the buffer you are currently looking.

ch contains two stars ‘**’ if the text in the buffer has been edited (the buffer is “modified”), or ‘--’ if the buffer has not been edited. For a read-only buffer, it is ‘%*’ if the buffer is modified, and ‘%%’ otherwise.

buf is the name of the window’s buffer. In most cases this is the same as the name of a file you are editing. See Chapter 15 [Buffers], page 141.

The buffer displayed in the selected window (the window that the cursor is in) is also Emacs’s selected buffer, the one that editing takes place in.

When we speak of what some command does to “the buffer,” we are talking about the currently selected buffer.

line is ‘L’ followed by the current line number of point. This is present when Line Number mode is enabled (which it normally is). You can optionally display the current column number too, by turning on Column Number mode (which is not enabled by default because it is somewhat slower). See Section 11.5 [Optional Mode Line], page 80.

pos tells you whether there is additional text above the top of the window, or below the bottom. If your buffer is small and it is all visible in the window, pos is ‘All’. Otherwise, it is ‘Top’ if you are looking at the beginning of the buffer, ‘Bot’ if you are looking at the end of the buffer, or ‘nn%’, where nn is the percentage of the buffer above the top of the window.

major is the name of the major mode in effect in the buffer. At any time, each buffer is in one and only one of the possible major modes. The major modes available include Fundamental mode (the least specialized), Text mode, Lisp mode, C mode, Texinfo mode, and many others. See Chapter 19 [Major Modes], page 185, for details of how the modes differ and how to select one.

Some major modes display additional information after the major mode name. For example, Rmail buffers display the current message number and the total number of messages. Compilation buffers and Shell buffers display the status of the subprocess.

minor is a list of some of the minor modes that are turned on at the moment in the window’s chosen buffer. For example, ‘Fill’ means that Auto Fill mode is on. ‘Abbrev’ means that Word Abbrev mode is on. ‘Ovwrt’ means that Overwrite mode is on. See Section 31.1 [Minor Modes], page 383, for more information. ‘Narrow’ means that the buffer being displayed has editing restricted to only a portion of its text. This is not really a minor mode, but is like one. See Section 30.8 [Narrowing], page 374. ‘Def’ means that a keyboard macro is being defined. See Section 31.3 [Keyboard Macros], page 397.

In addition, if Emacs is currently inside a recursive editing level, square brackets (‘[...]’) appear around the parentheses that surround the modes. If Emacs is in one recursive editing level within another, double square brackets appear, and so on. Since recursive editing levels affect Emacs globally, not just one buffer, the square brackets appear in every window’s mode line or not in any of them. See Section 30.12 [Recursive Edit], page 377.

cs states the coding system used for the file you are editing. A dash indicates the default state of affairs: no code conversion, except for endof-line translation if the file contents call for that. ‘=’ means no conversion whatsoever. Nontrivial code conversions are represented by various letters— for example, ‘1’ refers to ISO Latin-1. See Section 18.6 [Coding Systems], page 175, for more information. If you are using an input method, cs starts with a string such as ‘i>’, where i identifies the input method. (Some input methods show ‘+’ or ‘@’ instead of ‘>’.) See Section 18.4 [Input Methods], page 173.

When you are using a character-only terminal (not a window system), cs uses three characters to describe, respectively, the coding system for keyboard input, the coding system for terminal output, and the coding system used for the file you are editing.

When multibyte characters are not enabled, cs does not appear at all. See Section 18.2 [Enabling Multibyte], page 171.

The colon after cs can change to another character in certain circumstances. Emacs uses newline to separate lines in the buffer. Some files use different conventions for separating lines: either carriage-return linefeed (the MS-DOS convention) or just carriage-return (the Macintosh convention). If the buffer’s file uses carriage-return linefeed, the colon changes to a backslash (‘\’). If the file uses just carriage-return, the colon indicator changes to a forward slash (‘/’).

See Section 11.5 [Optional Mode Line], page 80, for features that add other handy information to the mode line, such as the current column number of point, the current time, and whether new mail for you has arrived.

1.4 The Menu Bar

Each Emacs frame normally has a menu bar at the top which you can use to perform certain common operations. There’s no need to list them here, as you can more easily see for yourself.

When you are using a window system, you can use the mouse to choose a command from the menu bar. An arrow pointing right, after the menu item, indicates that the item leads to a subsidiary menu; ‘...’ at the end means that the command will read arguments from the keyboard before it actually does anything.

To view the full command name and documentation for a menu item, type C-h k, and then select the menu bar with the mouse in the usual way (see Section 7.1 [Key Help], page 51).

On text-only terminals with no mouse, you can use the menu bar by typing M-` or ^{⟨}\underline{\overline{F10}}^{⟩} (these run the command tmm-menubar). This command enters a mode in which you can select a menu item from the keyboard. A provisional choice appears in the echo area. You can use the left and right arrow keys to move through the menu to different choices. When you have found the choice you want, type ^{⟨}\underline{\overline{RET}}^{⟩} to select it.

Each menu item also has an assigned letter or digit which designates that item; it is usually the initial of some word in the item’s name. This letter or digit is separated from the item name by ‘=>’. You can type the item’s letter or digit to select the item.

Some of the commands in the menu bar have ordinary key bindings as well; if so, the menu lists one equivalent key binding in parentheses after the item itself.

2 Characters, Keys and Commands

This chapter explains the character sets used by Emacs for input commands and for the contents of files, and also explains the concepts of keys and commands, which are fundamental for understanding how Emacs interprets your keyboard and mouse input.

2.1 Kinds of User Input

GNU Emacs uses an extension of the ASCII character set for keyboard input; it also accepts non-character input events including function keys and mouse button actions.

ASCII consists of 128 character codes. Some of these codes are assigned graphic symbols such as ‘a’ and ‘=’; the rest are control characters, such as Control-a (usually written C-a for short). C-a gets its name from the fact that you type it by holding down the ^{⟨}\underline{\overline{CTRL}}^{⟩} key while pressing a.

Some ASCII control characters have special names, and most terminals have special keys you can type them with: for example, ^{⟨}\underline{\overline{RET}}^{⟩}, ^{⟨}\underline{\overline{TAB}}^{⟩}, ^{⟨}\underline{\overline{DEL}}^{⟩} and ^{⟨}\underline{\overline{ESC}}^{⟩}. The space character is usually referred to below as ^{⟨}\underline{\overline{SPC}}^{⟩}, even though strictly speaking it is a graphic character whose graphic happens to be blank. Some keyboards have a key labeled “linefeed” which is an alias for C-j.

Emacs extends the ASCII character set with thousands more printing characters (see Chapter 18 [International], page 171), additional control characters, and a few more modifiers that can be combined with any character.

On ASCII terminals, there are only 32 possible control characters. These are the control variants of letters and ‘@[]\^_’. In addition, the shift key is meaningless with control characters: C-a and C-A are the same character, and Emacs cannot distinguish them.

But the Emacs character set has room for control variants of all printing characters, and for distinguishing between C-a and C-A. X Windows makes it possible to enter all these characters. For example, C-- (that’s Control-Minus) and C-5 are meaningful Emacs commands under X.

Another Emacs character-set extension is additional modifier bits. Only one modifier bit is commonly used; it is called Meta. Every character has a Meta variant; examples include Meta-a (normally written M-a, for short), M-A (not the same character as M-a, but those two characters normally have the same meaning in Emacs), M-^{⟨}\underline{\overline{RET}}^{⟩}, and M-C-a. For reasons of tradition, we usually write C-M-a rather than M-C-a; logically speaking, the order in which the modifier keys ^{⟨}\underline{\overline{CTRL}}^{⟩} and ^{⟨}\underline{\overline{META}}^{⟩} are mentioned does not matter.

Some terminals have a ^{⟨}\underline{\overline{META}}^{⟩} key, and allow you to type Meta characters by holding this key down. Thus, Meta-a is typed by holding down ^{⟨}\underline{\overline{META}}^{⟩} and pressing a. The ^{⟨}\underline{\overline{META}}^{⟩} key works much like the ^{⟨}\underline{\overline{SHIFT}}^{⟩} key. Such a key is not always labeled ^{⟨}\underline{\overline{META}}^{⟩}, however, as this function is often a special option for a key with some other primary purpose.

If there is no ^{⟨}\underline{\overline{META}}^{⟩} key, you can still type Meta characters using twocharacter sequences starting with ^{⟨}\underline{\overline{ESC}}^{⟩}. Thus, to enter M-a, you could type ⟨\underline{\overline{ESC}}^{⟩} a. To enter C-M-a, you would type ^{⟨}\underline{\overline{ESC}}^{⟩} C-a. ^{⟨}\underline{\overline{ESC}}^{⟩} is allowed on terminals with ^{⟨}\underline{\overline{META}}^{⟩} keys, too, in case you have formed a habit of using it.

X Windows provides several other modifier keys that can be applied to any input character. These are called ^{⟨}\underline{\overline{SUPER}}^{⟩}, ^{⟨}\underline{\overline{HYPER}}^{⟩} and ^{⟨}\underline{\overline{ALT}}^{⟩}. We write ‘s-’, ‘H-’ and ‘A-’ to say that a character uses these modifiers. Thus, s-H-C-x is short for Super-Hyper-Control-x. Not all X terminals actually provide keys for these modifier flags—in fact, many terminals have a key labeled ⟨\underline{\overline{ALT}}^{⟩} which is really a ^{⟨}\underline{\overline{META}}^{⟩} key. The standard key bindings of Emacs do not include any characters with these modifiers. But you can assign them meanings of your own by customizing Emacs.

Keyboard input includes keyboard keys that are not characters at all: for example function keys and arrow keys. Mouse buttons are also outside the gamut of characters. You can modify these events with the modifier keys ⟨\underline{\overline{CTRL}}^{⟩}, ^{⟨}\underline{\overline{META}}^{⟩}, ^{⟨}\underline{\overline{SUPER}}^{⟩}, ^{⟨}\underline{\overline{HYPER}}^{⟩} and ^{⟨}\underline{\overline{ALT}}^{⟩}, just like keyboard characters.

Input characters and non-character inputs are collectively called input events. See section “Input Events” in The Emacs Lisp Reference Manual, for more information. If you are not doing Lisp programming, but simply want to redefine the meaning of some characters or non-character events, see Chapter 31 [Customization], page 383.

ASCII terminals cannot really send anything to the computer except ASCII characters. These terminals use a sequence of characters to represent each function key. But that is invisible to the Emacs user, because the keyboard input routines recognize these special sequences and convert them to function key events before any other part of Emacs gets to see them.

2.2 Keys

A key sequence (key, for short) is a sequence of input events that are meaningful as a unit—as “a single command.” Some Emacs command sequences are just one character or one event; for example, just C-f is enough to move forward one character. But Emacs also has commands that take two or more events to invoke.

If a sequence of events is enough to invoke a command, it is a complete key. Examples of complete keys include C-a, X, ^{⟨}\underline{\overline{RET}}^{⟩}, ^{⟨}\underline{\overline{NEXT}}^{⟩} (a function key), ^{⟨}\underline{\overline{DOWN}}^{⟩} (an arrow key), C-x C-f, and C-x 4 C-f. If it isn’t long enough to be complete, we call it a prefix key. The above examples show that C-x and C-x 4 are prefix keys. Every key sequence is either a complete key or a prefix key.

Most single characters constitute complete keys in the standard Emacs command bindings. A few of them are prefix keys. A prefix key combines with the following input event to make a longer key sequence, which may itself be complete or a prefix. For example, C-x is a prefix key, so C-x and the next input event combine to make a two-character key sequence. Most of these key sequences are complete keys, including C-x C-f and C-x b. A few, such as C-x 4 and C-x r, are themselves prefix keys that lead to threecharacter key sequences. There’s no limit to the length of a key sequence, but in practice people rarely use sequences longer than four events.

By contrast, you can’t add more events onto a complete key. For example, the two-character sequence C-f C-k is not a key, because the C-f is a complete key in itself. It’s impossible to give C-f C-k an independent meaning as a command. C-f C-k is two key sequences, not one.

All told, the prefix keys in Emacs are C-c, C-h, C-x, C-x ^{⟨}\underline{\overline{RET}}^{⟩}, C-x @, C-x

a, C-x n, C-x r, C-x v, C-x 4, C-x 5, C-x 6, ^{⟨}\underline{\overline{ESC}}^{⟩}, M-g and M-j. But this list

is not cast in concrete; it is just a matter of Emacs’s standard key bindings. If you customize Emacs, you can make new prefix keys, or eliminate these. See Section 31.4 [Key Bindings], page 400.

If you do make or eliminate prefix keys, that changes the set of possible key sequences. For example, if you redefine C-f as a prefix, C-f C-k automatically becomes a key (complete, unless you define it too as a prefix). Conversely, if you remove the prefix definition of C-x 4, then C-x 4 f (or C-x 4 anything) is no longer a key.

Typing the help character (C-h or ^{⟨}\underline{\overline{F1}}^{⟩}) after a prefix character displays a list of the commands starting with that prefix. There are a few prefix characters for which C-h does not work—for historical reasons, they have other meanings for C-h which are not easy to change. But ^{⟨}\underline{\overline{F1}}^{⟩} should work for all prefix characters.

2.3 Keys and Commands

This manual is full of passages that tell you what particular keys do. But Emacs does not assign meanings to keys directly. Instead, Emacs assigns meanings to named commands, and then gives keys their meanings by binding them to commands.

Every command has a name chosen by a programmer. The name is usually made of a few English words separated by dashes; for example, nextline or forward-word. A command also has a function definition which is a Lisp program; this is what makes the command do what it does. In Emacs Lisp, a command is actually a special kind of Lisp function; one which specifies how to read arguments for it and call it interactively. For more information on commands and functions, see section “What Is a Function” in The Emacs Lisp Reference Manual. (The definition we use in this manual is simplified slightly.)

The bindings between keys and commands are recorded in various tables called keymaps. See Section 31.4.1 [Keymaps], page 400.

When we say that “C-n moves down vertically one line” we are glossing over a distinction that is irrelevant in ordinary use but is vital in understanding how to customize Emacs. It is the command next-line that is programmed to move down vertically. C-n has this effect because it is bound to that command. If you rebind C-n to the command forward-word then C-n will move forward by words instead. Rebinding keys is a common method of customization.

In the rest of this manual, we usually ignore this subtlety to keep things simple. To give the information needed for customization, we state the name of the command which really does the work in parentheses after mentioning the key that runs it. For example, we will say that “The command C-n (next-line) moves point vertically down,” meaning that next-line is a command that moves vertically down and C-n is a key that is standardly bound to it.

While we are on the subject of information for customization only, it’s a good time to tell you about variables. Often the description of a command will say, “To change this, set the variable mumble-foo.” A variable is a name used to remember a value. Most of the variables documented in this manual exist just to facilitate customization: some command or other part of Emacs examines the variable and behaves differently according to the value that you set. Until you are interested in customizing, you can ignore the information about variables. When you are ready to be interested, read the basic information on variables, and then the information on individual variables will make sense. See Section 31.2 [Variables], page 385.

2.4 Character Set for Text

Text in Emacs buffers is a sequence of 8-bit bytes. Each byte can hold a single ASCII character. Both ASCII control characters (octal codes 000 through 037, and 0177) and ASCII printing characters (codes 040 through 0176) are allowed; however, non-ASCII control characters cannot appear in a buffer. The other modifier flags used in keyboard input, such as Meta, are not allowed in buffers either.

Some ASCII control characters serve special purposes in text, and have special names. For example, the newline character (octal code 012) is used in the buffer to end a line, and the tab character (octal code 011) is used for indenting to the next tab stop column (normally every 8 columns). See Section 11.6 [Text Display], page 80.

Non-ASCII printing characters can also appear in buffers. When multibyte characters are enabled, you can use any of the non-ASCII printing characters that Emacs supports. They have character codes starting at 256, octal 0400, and each one is represented as a sequence of two or more bytes. See Chapter 18 [International], page 171.

If you disable multibyte characters, then you can use only one alphabet of non-ASCII characters, but they all fit in one byte. They use codes 0200 through 0377. See Section 18.11 [Single-Byte European Support], page 183.

3 Entering and Exiting Emacs

The usual way to invoke Emacs is with the shell command ‘emacs’. Emacs clears the screen and then displays an initial help message and copyright notice. Some operating systems discard all type-ahead when Emacs starts up; they give Emacs no way to prevent this. Therefore, it is advisable to wait until Emacs clears the screen before typing your first editing command.

If you run Emacs from a shell window under the X Window System, run it in the background with ‘emacs&’. This way, Emacs does not tie up the shell window, so you can use that to run other shell commands while Emacs operates its own X windows. You can begin typing Emacs commands as soon as you direct your keyboard input to the Emacs frame.

When Emacs starts up, it makes a buffer named ‘*scratch*’. That’s the buffer you start out in. The ‘*scratch*’ buffer uses Lisp Interaction mode; you can use it to type Lisp expressions and evaluate them, or you can ignore that capability and simply doodle. (You can specify a different major mode for this buffer by setting the variable initial-major-mode in your init file. See Section 31.7 [Init File], page 411.)

It is possible to specify files to be visited, Lisp files to be loaded, and functions to be called, by giving Emacs arguments in the shell command line. See Appendix A [Command Arguments], page 433. But we don’t recommend doing this. The feature exists mainly for compatibility with other editors.

Many other editors are designed to be started afresh each time you want to edit. You edit one file and then exit the editor. The next time you want to edit either another file or the same one, you must run the editor again. With these editors, it makes sense to use a command-line argument to say which file to edit.

But starting a new Emacs each time you want to edit a different file does not make sense. For one thing, this would be annoyingly slow. For another, this would fail to take advantage of Emacs’s ability to visit more than one file in a single editing session. And it would lose the other accumulated context, such as registers, undo history, and the mark ring.

The recommended way to use GNU Emacs is to start it only once, just after you log in, and do all your editing in the same Emacs session. Each time you want to edit a different file, you visit it with the existing Emacs, which eventually comes to have many files in it ready for editing. Usually you do not kill the Emacs until you are about to log out. See Chapter 14 [Files], page 105, for more information on visiting more than one file.

3.1 Exiting Emacs

There are two commands for exiting Emacs because there are two kinds of exiting: suspending Emacs and killing Emacs.

Suspending means stopping Emacs temporarily and returning control to its parent process (usually a shell), allowing you to resume editing later in the same Emacs job, with the same buffers, same kill ring, same undo history, and so on. This is the usual way to exit.

Killing Emacs means destroying the Emacs job. You can run Emacs again later, but you will get a fresh Emacs; there is no way to resume the same editing session after it has been killed.

C-z Suspend Emacs (suspend-emacs) or iconify a frame (iconify-

          or-deiconify-frame).
C-x C-c Kill Emacs (save-buffers-kill-emacs).

To suspend Emacs, type C-z (suspend-emacs). This takes you back to the shell from which you invoked Emacs. You can resume Emacs with the shell command ‘%emacs’ in most common shells.

On systems that do not support suspending programs, C-z starts an inferior shell that communicates directly with the terminal. Emacs waits until you exit the subshell. (The way to do that is probably with C-d or ‘exit’, but it depends on which shell you use.) The only way on these systems to get back to the shell from which Emacs was run (to log out, for example) is to kill Emacs.

Suspending also fails if you run Emacs under a shell that doesn’t support suspending programs, even if the system itself does support it. In such a case, you can set the variable cannot-suspend to a non-nil value to force C-z to start an inferior shell. (One might also describe Emacs’s parent shell as “inferior” for failing to support job control properly, but that is a matter of taste.)

When Emacs communicates directly with an X server and creates its own dedicated X windows, C-z has a different meaning. Suspending an applications that uses its own X windows is not meaningful or useful. Instead, C-z runs the command iconify-or-deiconify-frame, which temporarily closes up the selected Emacs frame (see Chapter 17 [Frames], page 155). The way to get back to a shell window is with the window manager.

   To kill Emacs, type C-x C-c (save-buffers-kill-emacs). A two-

character key is used for this to make it harder to type. This command first offers to save any modified file-visiting buffers. If you do not save them all, it asks for reconfirmation with yes before killing Emacs, since any changes not saved will be lost forever. Also, if any subprocesses are still running, C-x C-c asks for confirmation about them, since killing Emacs will kill the subprocesses immediately.

There is no way to restart an Emacs session once you have killed it. You can, however, arrange for Emacs to record certain session information, such as which files are visited, when you kill it, so that the next time you restart Emacs it will try to visit the same files and so on. See Section 30.11 [Saving Emacs Sessions], page 377.

The operating system usually listens for certain special characters whose meaning is to kill or suspend the program you are running. This operating system feature is turned off while you are in Emacs. The meanings of C-z and C-x C-c as keys in Emacs were inspired by the use of C-z and C-c on several operating systems as the characters for stopping or killing a program, but that is their only relationship with the operating system. You can customize these keys to run any commands of your choice (see Section 31.4.1 [Keymaps], page 400).

28 GNU Emacs Manual

4 Basic Editing Commands

We now give the basics of how to enter text, make corrections, and save the text in a file. If this material is new to you, you might learn it more easily by running the Emacs learn-by-doing tutorial. To use the tutorial,

run Emacs and type Control-h t (help-with-tutorial).

To clear the screen and redisplay, type C-l (recenter).

4.1 Inserting Text

To insert printing characters into the text you are editing, just type them. This inserts the characters you type into the buffer at the cursor (that is, at point; see Section 1.1 [Point], page 15). The cursor moves forward, and any text after the cursor moves forward too. If the text in the buffer is ‘FOOBAR’, with the cursor before the ‘B’, then if you type XX, you get ‘FOOXXBAR’, with the cursor still before the ‘B’.

To delete text you have just inserted, use^{⟨}\underline{\overline{DEL}}^{⟩}. ^{⟨}\underline{\overline{DEL}}^{⟩} deletes the character before the cursor (not the one that the cursor is on top of or under; that is the character after the cursor). The cursor and all characters after it move backwards. Therefore, if you type a printing character and then type ^{⟨}\underline{\overline{DEL}}^{⟩}, they cancel out.

To end a line and start typing a new one, type ^{⟨}\underline{\overline{RET}}^{⟩}. This inserts a newline character in the buffer. If point is in the middle of a line, ^{⟨}\underline{\overline{RET}}^{⟩} splits the line. Typing ^{⟨}\underline{\overline{DEL}}^{⟩} when the cursor is at the beginning of a line deletes the preceding newline, thus joining the line with the preceding line.

Emacs can split lines automatically when they become too long, if you turn on a special minor mode called Auto Fill mode. See Section 21.5 [Filling], page 197, for how to use Auto Fill mode.

If you prefer to have text characters replace (overwrite) existing text rather than shove it to the right, you can enable Overwrite mode, a minor mode. See Section 31.1 [Minor Modes], page 383.

Direct insertion works for printing characters and ^{⟨}\underline{\overline{SPC}}^{⟩}, but other characters act as editing commands and do not insert themselves. If you need to insert a control character or a character whose code is above 200 octal, you must quote it by typing the character Control-q (quoted-insert) first. (This character’s name is normally written C-q for short.) There are two ways to use C-q:

• C-q followed by any non-graphic character (even C-g) inserts that character.

• C-q followed by a sequence of octal digits inserts the character with the specified octal character code. You can use any number of octal digits; any non-digit terminates the sequence. If the terminating character is ⟨\underline{\overline{RET}}^{⟩}, it serves only to terminate the sequence; any other non-digit is itself used as input after terminating the sequence. (The use of octal sequences is disabled in ordinary non-binary Overwrite mode, to give you a convenient way to insert a digit instead of overwriting with it.)

When multibyte characters are enabled, octal codes 0200 through 0377 are not valid as characters; if you specify a code in this range, C-q assumes that you intend to use some ISO Latin-n character set, and converts the specified code to the corresponding Emacs character code. See Section 18.2 [Enabling Multibyte], page 171. You select which ISO Latin character set though your choice of language environment (see Section 18.3 [Language Environments], page 172).

To use decimal or hexadecimal instead of octal, set the variable readquoted-char-radix to 10 or 16. If the radix is greater than 10, some letters starting with a serve as part of a character code, just like digits.

A numeric argument to C-q specifies how many copies of the quoted character should be inserted (see Section 4.10 [Arguments], page 37).

Customization information: ^{⟨}\underline{\overline{DEL}}^{⟩} in most modes runs the command delete-backward-char; ^{⟨}\underline{\overline{RET}}^{⟩} runs the command newline, and selfinserting printing characters run the command self-insert, which inserts whatever character was typed to invoke it. Some major modes rebind ^{⟨}\underline{\overline{DEL}}^{⟩} to other commands.

4.2 Changing the Location of Point

To do more than insert characters, you have to know how to move point (see Section 1.1 [Point], page 15). The simplest way to do this is with arrow keys, or by clicking the left mouse button where you want to move to.

There are also control and meta characters for cursor motion. Some are equivalent to the arrow keys (these date back to the days before terminals had arrow keys, and are usable on terminals which don’t have them). Others do more sophisticated things.

C-a Move to the beginning of the line (beginning-of-line).

C-e Move to the end of the line (end-of-line).

C-f Move forward one character (forward-char).

C-b Move backward one character (backward-char).

M-f Move forward one word (forward-word).

M-b Move backward one word (backward-word).

C-n Move down one line, vertically (next-line). This command attempts to keep the horizontal position unchanged, so if you start in the middle of one line, you end in the middle of the next. When on the last line of text, C-n creates a new line and moves onto it.

C-p Move up one line, vertically (previous-line).

M-r Move point to left margin, vertically centered in the window (move-to-window-line). Text does not move on the screen. A numeric argument says which screen line to place point on. It counts screen lines down from the top of the window (zero for the top line). A negative argument counts lines from the bottom (−1 for the bottom line).

M-< Move to the top of the buffer (beginning-of-buffer). With numeric argument n, move to n/10 of the way from the top. See Section 4.10 [Arguments], page 37, for more information on numeric arguments.

M-> Move to the end of the buffer (end-of-buffer).

M-x goto-char

Read a number n and move point to buffer position n. Position 1 is the beginning of the buffer.

M-x goto-line

Read a number n and move point to line number n. Line 1 is the beginning of the buffer.

C-x C-n Use the current column of point as the semipermanent goal column for C-n and C-p (set-goal-column). Henceforth, those commands always move to this column in each line moved into, or as close as possible given the contents of the line. This goal column remains in effect until canceled.

C-u C-x C-n

Cancel the goal column. Henceforth, C-n and C-p once again try to stick to a fixed horizontal position, as usual.

If you set the variable track-eol to a non-nil value, then C-n and C-p when at the end of the starting line move to the end of another line. Normally, track-eol is nil. See Section 31.2 [Variables], page 385, for how to set variables such as track-eol.

Normally, C-n on the last line of a buffer appends a newline to it. If the variable next-line-add-newlines is nil, then C-n gets an error instead (like C-p on the first line).

4.3 Erasing Text

⟨\underline{\overline{DEL}}^{⟩} Delete the character before point (delete-backward-char).

C-d Delete the character after point (delete-char).

C-k Kill to the end of the line (kill-line).

M-d Kill forward to the end of the next word (kill-word).

M-^{⟨}\underline{\overline{DEL}}^{⟩} Kill back to the beginning of the previous word (backward-kill-

          word).

You already know about the ^{⟨}\underline{\overline{DEL}}^{⟩} key which deletes the character before point (that is, before the cursor). Another key, Control-d (C-d for short), deletes the character after point (that is, the character that the cursor is on). This shifts the rest of the text on the line to the left. If you type C-d at the end of a line, it joins together that line and the next line.

To erase a larger amount of text, use the C-k key, which kills a line at a time. If you type C-k at the beginning or middle of a line, it kills all the text up to the end of the line. If you type C-k at the end of a line, it joins that line and the next line.

See Section 9.1 [Killing], page 63, for more flexible ways of killing text.

4.4 Undoing Changes

You can undo all the recent changes in the buffer text, up to a certain point. Each buffer records changes individually, and the undo command always applies to the current buffer. Usually each editing command makes a separate entry in the undo records, but some commands such as queryreplace make many entries, and very simple commands such as self-inserting characters are often grouped to make undoing less tedious.

C-x u Undo one batch of changes—usually, one command worth (undo).

C-_The same.

C-u C-x u Undo one batch of changes in the region.

The command C-x u or C-_is how you undo. The first time you give this command, it undoes the last change. Point moves back to where it was before the command that made the change.

Consecutive repetitions of C-_or C-x u undo earlier and earlier changes, back to the limit of the undo information available. If all recorded changes have already been undone, the undo command prints an error message and does nothing.

Any command other than an undo command breaks the sequence of undo commands. Starting from that moment, the previous undo commands become ordinary changes that you can undo. Thus, to redo changes you have undone, type C-f or any other command that will harmlessly break the sequence of undoing, then type more undo commands.

Ordinary undo applies to all changes made in the current buffer. You can also perform selective undo, limited to the current region. To do this, specify the region you want, then run the undo command with a prefix argument (the value does not matter): C-u C-x u or C-u C-_. This undoes the most recent change in the region. To undo further changes in the same region, repeat the undo command (no prefix argument is needed). In Transient Mark mode, any use of undo when there is an active region performs selective undo; you do not need a prefix argument.

If you notice that a buffer has been modified accidentally, the easiest way to recover is to type C-_repeatedly until the stars disappear from the front of the mode line. At this time, all the modifications you made have been canceled. Whenever an undo command makes the stars disappear from the mode line, it means that the buffer contents are the same as they were when the file was last read in or saved.

If you do not remember whether you changed the buffer deliberately, type C-_once. When you see the last change you made undone, you will see whether it was an intentional change. If it was an accident, leave it undone. If it was deliberate, redo the change as described above.

Not all buffers record undo information. Buffers whose names start with spaces don’t; these buffers are used internally by Emacs and its extensions to hold text that users don’t normally look at or edit.

You cannot undo mere cursor motion; only changes in the buffer contents save undo information. However, some cursor motion commands set the mark, so if you use these commands from time to time, you can move back to the neighborhoods you have moved through by popping the mark ring (see Section 8.5 [Mark Ring], page 61).

When the undo information for a buffer becomes too large, Emacs discards the oldest undo information from time to time (during garbage collection). You can specify how much undo information to keep by setting two variables: undo-limit and undo-strong-limit. Their values are expressed in units of bytes of space.

The variable undo-limit sets a soft limit: Emacs keeps undo data for enough commands to reach this size, and perhaps exceed it, but does not keep data for any earlier commands beyond that. Its default value is 20000. The variable undo-strong-limit sets a stricter limit: the command which pushes the size past this amount is itself forgotten. Its default value is 30000.

Regardless of the values of those variables, the most recent change is never discarded, so there is no danger that garbage collection occurring right after an unintentional large change might prevent you from undoing it.

The reason the undo command has two keys, C-x u and C-_, set up to run it is that it is worthy of a single-character key, but on some keyboards it is not obvious how to type C-_. C-x u is an alternative you can type straightforwardly on any terminal.

4.5 Files

The commands described above are sufficient for creating and altering text in an Emacs buffer; the more advanced Emacs commands just make things easier. But to keep any text permanently you must put it in a file. Files are named units of text which are stored by the operating system for you to retrieve later by name. To look at or use the contents of a file in any way, including editing the file with Emacs, you must specify the file name.

Consider a file named ‘/usr/rms/foo.c’. In Emacs, to begin editing this file, type

    C-x C-f /usr/rms/foo.c ^{⟨}\underline{\overline{RET}}^{⟩}

Here the file name is given as an argument to the command C-x C-f (findfile). That command uses the minibuffer to read the argument, and you type ^{⟨}\underline{\overline{RET}}^{⟩} to terminate the argument (see Chapter 5 [Minibuffer], page 39).

Emacs obeys the command by visiting the file: creating a buffer, copying the contents of the file into the buffer, and then displaying the buffer for you to edit. If you alter the text, you can save the new text in the file by typing C-x C-s (save-buffer). This makes the changes permanent by copying the altered buffer contents back into the file ‘/usr/rms/foo.c’. Until you save, the changes exist only inside Emacs, and the file ‘foo.c’ is unaltered.

To create a file, just visit the file with C-x C-f as if it already existed. This creates an empty buffer in which you can insert the text you want to put in the file. The file is actually created when you save this buffer with

C-x C-s.

Of course, there is a lot more to learn about using files. See Chapter 14 [Files], page 105.

4.6 Help

If you forget what a key does, you can find out with the Help character, which is C-h (or ^{⟨}\underline{\overline{F1}}^{⟩}, which is an alias for C-h). Type C-h k followed by the key you want to know about; for example, C-h k C-n tells you all about what C-n does. C-h is a prefix key; C-h k is just one of its subcommands (the command describe-key). The other subcommands of C-h provide different kinds of help. Type C-h twice to get a description of all the help facilities. See Chapter 7 [Help], page 49.

4.7 Blank Lines

Here are special commands and techniques for putting in and taking out blank lines.

C-o Insert one or more blank lines after the cursor (open-line).

C-x C-o Delete all but one of many consecutive blank lines (delete-

          blank-lines).

When you want to insert a new line of text before an existing line, you can do it by typing the new line of text, followed by ^{⟨}\underline{\overline{RET}}^{⟩}. However, it may be easier to see what you are doing if you first make a blank line and then insert the desired text into it. This is easy to do using the key C-o (open-line), which inserts a newline after point but leaves point in front of the newline. After C-o, type the text for the new line. C-o F O O has the same effect as F O O ^{⟨}\underline{\overline{RET}}^{⟩}, except for the final location of point.

You can make several blank lines by typing C-o several times, or by giving it a numeric argument to tell it how many blank lines to make. See Section 4.10 [Arguments], page 37, for how. If you have a fill prefix, then C-o command inserts the fill prefix on the new line, when you use it at the beginning of a line. See Section 21.5.3 [Fill Prefix], page 200.

The easy way to get rid of extra blank lines is with the command C-x C-o (delete-blank-lines). C-x C-o in a run of several blank lines deletes all but one of them. C-x C-o on a solitary blank line deletes that blank line. When point is on a nonblank line, C-x C-o deletes any blank lines following that nonblank line.

4.8 Continuation Lines

If you add too many characters to one line without breaking it with ^{⟨}\underline{\overline{RET}}^{⟩}, the line will grow to occupy two (or more) lines on the screen, with a ‘\’ at the extreme right margin of all but the last of them. The ‘\’ says that the following screen line is not really a distinct line in the text, but just the continuation of a line too long to fit the screen. Continuation is also called line wrapping.

Sometimes it is nice to have Emacs insert newlines automatically when a line gets too long. Continuation on the screen does not do that. Use Auto Fill mode (see Section 21.5 [Filling], page 197) if that’s what you want.

As an alternative to continuation, Emacs can display long lines by truncation. This means that all the characters that do not fit in the width of the screen or window do not appear at all. They remain in the buffer, temporarily invisible. ‘$’ is used in the last column instead of ‘\’ to inform you that truncation is in effect.

Truncation instead of continuation happens whenever horizontal scrolling is in use, and optionally in all side-by-side windows (see Chapter 16 [Windows], page 149). You can enable truncation for a particular buffer by setting the variable truncate-lines to non-nil in that buffer. (See Section 31.2 [Variables], page 385.) Altering the value of truncate-lines makes it local to the current buffer; until that time, the default value is in effect. The default is initially nil. See Section 31.2.4 [Locals], page 393.

See Section 11.7 [Display Vars], page 81, for additional variables that affect how text is displayed.

4.9 Cursor Position Information

Here are commands to get information about the size and position of parts of the buffer, and to count lines.

M-x what-page

Print page number of point, and line number within page.

M-x what-line

Print line number of point in the buffer.

M-x line-number-mode

Toggle automatic display of current line number.

M-= Print number of lines in the current region (count-lines-region). See Chapter 8 [Mark], page 57, for information about the region.

C-x = Print character code of character after point, character position of point, and column of point (what-cursor-position).

There are two commands for working with line numbers. M-x what-line computes the current line number and displays it in the echo area. To go to a given line by number, use M-x goto-line; it prompts you for the number. These line numbers count from one at the beginning of the buffer.

You can also see the current line number in the mode line; See Section 1.3 [Mode Line], page 17. If you narrow the buffer, then the line number in the mode line is relative to the accessible portion (see Section 30.8 [Narrowing], page 374). By contrast, what-line shows both the line number relative to the narrowed region and the line number relative to the whole buffer.

By contrast, M-x what-page counts pages from the beginning of the file, and counts lines within the page, printing both numbers. See Section 21.4 [Pages], page 196.

While on this subject, we might as well mention M-= (count-lines-region), which prints the number of lines in the region (see Chapter 8 [Mark], page 57). See Section 21.4 [Pages], page 196, for the command C-x l which counts the lines in the current page.

The command C-x = (what-cursor-position) can be used to find out the column that the cursor is in, and other miscellaneous information about point. It prints a line in the echo area that looks like this:

     Char: c (0143, 99, 0x63)  point=21044 of 26883(78%)  column 53

(In fact, this is the output produced when point is before the ‘column’ in the example.)

The four values after ‘Char:’ describe the character that follows point, first by showing it and then by giving its character code in octal, decimal and hex.

‘point=’ is followed by the position of point expressed as a character count. The front of the buffer counts as position 1, one character later as 2, and so on. The next, larger, number is the total number of characters in the buffer. Afterward in parentheses comes the position expressed as a percentage of the total size.

‘column’ is followed by the horizontal position of point, in columns from the left edge of the window.

If the buffer has been narrowed, making some of the text at the beginning and the end temporarily inaccessible, C-x = prints additional text describing the currently accessible range. For example, it might display this:

     Char: C (0103, 67, 0x43)  point=252 of 889(28%) <231 - 599>  column 0

where the two extra numbers give the smallest and largest character position that point is allowed to assume. The characters between those two positions are the accessible ones. See Section 30.8 [Narrowing], page 374.

If point is at the end of the buffer (or the end of the accessible part), C-x = omits any description of the character after point. The output might look like this:

     point=26957 of 26956(100%)  column 0

4.10 Numeric Arguments

In mathematics and computer usage, the word argument means “data provided to a function or operation.” You can give any Emacs command a numeric argument (also called a prefix argument). Some commands interpret the argument as a repetition count. For example, C-f with an argument of ten moves forward ten characters instead of one. With these commands, no argument is equivalent to an argument of one. Negative arguments tell most such commands to move or act in the opposite direction.

If your terminal keyboard has a ^{⟨}\underline{\overline{META}}^{⟩} key, the easiest way to specify a numeric argument is to type digits and/or a minus sign while holding down the ^{⟨}\underline{\overline{META}}^{⟩} key. For example,

    M-5 C-n

would move down five lines. The characters Meta-1, Meta-2, and so on, as well as Meta--, do this because they are keys bound to commands (digitargument and negative-argument) that are defined to contribute to an argument for the next command. Digits and - modified with Control, or Control and Meta, also specify numeric arguments.

Another way of specifying an argument is to use the C-u (universalargument) command followed by the digits of the argument. With C-u, you can type the argument digits without holding down modifier keys; C-u works on all terminals. To type a negative argument, type a minus sign after C-u. Just a minus sign without digits normally means −1.

C-u followed by a character which is neither a digit nor a minus sign has the special meaning of “multiply by four.” It multiplies the argument for the next command by four. C-u twice multiplies it by sixteen. Thus, C-u C-u C-f moves forward sixteen characters. This is a good way to move forward “fast,” since it moves about 1/5 of a line in the usual size screen. Other useful combinations are C-u C-n, C-u C-u C-n (move down a good fraction of a screen), C-u C-u C-o (make “a lot” of blank lines), and C-u C-k (kill four lines).

Some commands care only about whether there is an argument, and not about its value. For example, the command M-q (fill-paragraph) with no argument fills text; with an argument, it justifies the text as well. (See Section 21.5 [Filling], page 197, for more information on M-q.) Plain C-u is a handy way of providing an argument for such commands.

Some commands use the value of the argument as a repeat count, but do something peculiar when there is no argument. For example, the command C-k (kill-line) with argument n kills n lines, including their terminating newlines. But C-k with no argument is special: it kills the text up to the next newline, or, if point is right at the end of the line, it kills the newline itself. Thus, two C-k commands with no arguments can kill a nonblank line, just like C-k with an argument of one. (See Section 9.1 [Killing], page 63, for more information on C-k.)

A few commands treat a plain C-u differently from an ordinary argument. A few others may treat an argument of just a minus sign differently from an argument of −1. These unusual cases are described when they come up; they are always for reasons of convenience of use of the individual command.

You can use a numeric argument to insert multiple copies of a character. This is straightforward unless the character is a digit; for example, C-u 6 4 a inserts 64 copies of the character ‘a’. But this does not work for inserting digits; C-u 6 4 1 specifies an argument of 641, rather than inserting anything. To separate the digit to insert from the argument, type another C-u; for example, C-u 6 4 C-u 1 does insert 64 copies of the character ‘1’.

We use the term “prefix argument” as well as “numeric argument” to emphasize that you type the argument before the command, and to distinguish these arguments from minibuffer arguments that come after the command.

4.11 Repeating a Command

The command C-x z (repeat) provides another way to repeat an Emacs command many times. This command repeats the previous Emacs command, whatever that was. Repeating a command uses the same arguments that were used before; it does not read new arguments each time.

To repeat the command more than once, type additional z’s: each z repeats the command one more time. Repetition ends when you type a character other than z, or press a mouse button.

For example, suppose you type C-u 2 0 C-d to delete 20 characters. You can repeat that command (including its argument) three additional times, to delete a total of 80 characters, by typing C-x z z z. The first C-x z repeats the command once, and each subsequent z repeats it once again.

5 The Minibuffer

The minibuffer is the facility used by Emacs commands to read arguments more complicated than a single number. Minibuffer arguments can be file names, buffer names, Lisp function names, Emacs command names, Lisp expressions, and many other things, depending on the command reading the argument. You can use the usual Emacs editing commands in the minibuffer to edit the argument text.

When the minibuffer is in use, it appears in the echo area, and the terminal’s cursor moves there. The beginning of the minibuffer line displays a prompt which says what kind of input you should supply and how it will be used. Often this prompt is derived from the name of the command that the argument is for. The prompt normally ends with a colon.

Sometimes a default argument appears in parentheses after the colon; it too is part of the prompt. The default will be used as the argument value if you enter an empty argument (for example, just type ^{⟨}\underline{\overline{RET}}^{⟩}). For example, commands that read buffer names always show a default, which is the name of the buffer that will be used if you type just ^{⟨}\underline{\overline{RET}}^{⟩}.

The simplest way to enter a minibuffer argument is to type the text you want, terminated by ^{⟨}\underline{\overline{RET}}^{⟩} which exits the minibuffer. You can cancel the command that wants the argument, and get out of the minibuffer, by typing

C-g.

Since the minibuffer uses the screen space of the echo area, it can conflict with other ways Emacs customarily uses the echo area. Here is how Emacs handles such conflicts:

• If a command gets an error while you are in the minibuffer, this does not cancel the minibuffer. However, the echo area is needed for the error message and therefore the minibuffer itself is hidden for a while. It comes back after a few seconds, or as soon as you type anything.

• If in the minibuffer you use a command whose purpose is to print a message in the echo area, such as C-x =, the message is printed normally, and the minibuffer is hidden for a while. It comes back after a few seconds, or as soon as you type anything.

• Echoing of keystrokes does not take place while the minibuffer is in use.

5.1 Minibuffers for File Names

Sometimes the minibuffer starts out with text in it. For example, when you are supposed to give a file name, the minibuffer starts out containing the default directory, which ends with a slash. This is to inform you which directory the file will be found in if you do not specify a directory.

For example, the minibuffer might start out with these contents:

    Find File: /u2/emacs/src/

where ‘Find File: ’ is the prompt. Typing buffer.c specifies the file ‘/u2/emacs/src/buffer.c’. To find files in nearby directories, use ..; thus, if you type ../lisp/simple.el, you will get the file named ‘/u2/emacs/lisp/simple.el’. Alternatively, you can kill with M-^{⟨}\underline{\overline{DEL}}^{⟩} the directory names you don’t want (see Section 21.1 [Words], page 193).

If you don’t want any of the default, you can kill it with C-a C-k. But you don’t need to kill the default; you can simply ignore it. Insert an absolute file name, one starting with a slash or a tilde, after the default directory. For example, to specify the file ‘/etc/termcap’, just insert that name, giving these minibuffer contents:

    Find File: /u2/emacs/src//etc/termcap

Two slashes in a row are not normally meaningful in a file name, but they are allowed in GNU Emacs. They mean, “ignore everything before the second slash in the pair.” Thus, ‘/u2/emacs/src/’ is ignored in the example above, and you get the file ‘/etc/termcap’.

If you set insert-default-directory to nil, the default directory is not inserted in the minibuffer. This way, the minibuffer starts out empty. But the name you type, if relative, is still interpreted with respect to the same default directory.

5.2 Editing in the Minibuffer

The minibuffer is an Emacs buffer (albeit a peculiar one), and the usual Emacs commands are available for editing the text of an argument you are entering.

Since ^{⟨}\underline{\overline{RET}}^{⟩} in the minibuffer is defined to exit the minibuffer, you can’t use it to insert a newline in the minibuffer. To do that, type C-o or C-q C-j. (Recall that a newline is really the character control-J.)

The minibuffer has its own window which always has space on the screen but acts as if it were not there when the minibuffer is not in use. When the minibuffer is in use, its window is just like the others; you can switch to another window with C-x o, edit text in other windows and perhaps even visit more files, before returning to the minibuffer to submit the argument. You can kill text in another window, return to the minibuffer window, and then yank the text to use it in the argument. See Chapter 16 [Windows], page 149.

There are some restrictions on the use of the minibuffer window, however. You cannot switch buffers in it—the minibuffer and its window are permanently attached. Also, you cannot split or kill the minibuffer window. But you can make it taller in the normal fashion with C-x ^. If you enable Resize-Minibuffer mode, then the minibuffer window expands vertically as necessary to hold the text that you put in the minibuffer. Use M-x resize-minibuffer-mode to enable or disable this minor mode (see Section 31.1 [Minor Modes], page 383).

Scrolling works specially in the minibuffer window. When the minibuffer is just one line high, and it contains a long line of text that won’t fit on the screen, scrolling automatically maintains an overlap of a certain number of characters from one continuation line to the next. The variable minibufferscroll-overlap specifies how many characters of overlap; the default is 20.

If while in the minibuffer you issue a command that displays help text of any sort in another window, you can use the C-M-v command while in the minibuffer to scroll the help text. This lasts until you exit the minibuffer. This feature is especially useful if a completing minibuffer gives you a list of possible completions. See Section 16.3 [Other Window], page 150.

Emacs normally disallows most commands that use the minibuffer while the minibuffer is active. This rule is to prevent recursive minibuffers from confusing novice users. If you want to be able to use such commands in the

minibuffer, set the variable enable-recursive-minibuffers to a non-nil

value.

5.3 Completion

For certain kinds of arguments, you can use completion to enter the argument value. Completion means that you type part of the argument, then Emacs visibly fills in the rest, or as much as can be determined from the part you have typed.

When completion is available, certain keys—^{⟨}\underline{\overline{TAB}}^{⟩}, ^{⟨}\underline{\overline{RET}}^{⟩}, and ^{⟨}\underline{\overline{SPC}}^{⟩}—are rebound to complete the text present in the minibuffer into a longer string that it stands for, by matching it against a set of completion alternatives provided by the command reading the argument. ? is defined to display a list of possible completions of what you have inserted.

For example, when M-x uses the minibuffer to read the name of a command, it provides a list of all available Emacs command names to complete against. The completion keys match the text in the minibuffer against all the command names, find any additional name characters implied by the ones already present in the minibuffer, and add those characters to the ones you have given. This is what makes it possible to type M-x ins ^{⟨}\underline{\overline{SPC}}^{⟩} b ^{⟨}\underline{\overline{RET}}^{⟩} instead of M-x insert-buffer ^{⟨}\underline{\overline{RET}}^{⟩} (for example).

Case is normally significant in completion, because it is significant in most of the names that you can complete (buffer names, file names and command names). Thus, ‘fo’ does not complete to ‘Foo’. Completion does ignore case distinctions for certain arguments in which case does not matter.

5.3.1 Completion Example

A concrete example may help here. If you type M-x au ^{⟨}\underline{\overline{TAB}}^{⟩}, the ^{⟨}\underline{\overline{TAB}}^{⟩} looks for alternatives (in this case, command names) that start with ‘au’. There are several, including auto-fill-mode and auto-save-mode—but they are all the same as far as auto-, so the ‘au’ in the minibuffer changes

to ‘auto-’.

If you type ^{⟨}\underline{\overline{TAB}}^{⟩} again immediately, there are multiple possibilities for the very next character—it could be any of ‘cfilrs’—so no more characters are added; instead, ^{⟨}\underline{\overline{TAB}}^{⟩} displays a list of all possible completions in another window.

If you go on to type f ^{⟨}\underline{\overline{TAB}}^{⟩}, this ^{⟨}\underline{\overline{TAB}}^{⟩} sees ‘auto-f’. The only command name starting this way is auto-fill-mode, so completion fills in the rest of that. You now have ‘auto-fill-mode’ in the minibuffer after typing just au ⟨\underline{\overline{TAB}}^{⟩} f ^{⟨}\underline{\overline{TAB}}^{⟩}. Note that ^{⟨}\underline{\overline{TAB}}^{⟩} has this effect because in the minibuffer it is bound to the command minibuffer-complete when completion is available.

5.3.2 Completion Commands

Here is a list of the completion commands defined in the minibuffer when completion is available. ⟨\underline{\overline{TAB}}^{⟩} Complete the text in the minibuffer as much as possible

          (minibuffer-complete).

⟨\underline{\overline{SPC}}^{⟩} Complete the minibuffer text, but don’t go beyond one word

          (minibuffer-complete-word).

⟨\underline{\overline{RET}}^{⟩} Submit the text in the minibuffer as the argument, possibly completing first as described below (minibuffer-complete-and-

          exit).

? Print a list of all possible completions of the text in the minibuffer

          (minibuffer-list-completions).

⟨\underline{\overline{SPC}}^{⟩} completes much like ^{⟨}\underline{\overline{TAB}}^{⟩}, but never goes beyond the next hyphen or space. If you have ‘auto-f’ in the minibuffer and type ^{⟨}\underline{\overline{SPC}}^{⟩}, it finds that the completion is ‘auto-fill-mode’, but it stops completing after ‘fill-’. This gives ‘auto-fill-’. Another ^{⟨}\underline{\overline{SPC}}^{⟩} at this point completes all the way to ‘auto-fill-mode’. ^{⟨}\underline{\overline{SPC}}^{⟩} in the minibuffer when completion is available

runs the command minibuffer-complete-word.

Here are some commands you can use to choose a completion from a window that displays a list of completions:

Mouse-2 Clicking mouse button 2 on a completion in the list of possible completions chooses that completion (mouse-choose-completion). You normally use this command while point is in the minibuffer; but you must click in the list of completions, not in the minibuffer itself.

⟨\underline{\overline{PRIOR}}⟩

M-v Typing ^{⟨}\underline{\overline{PRIOR}}^{⟩} or ^{⟨}\underline{\overline{PAGE-UP}}^{⟩}, or M-v, while in the minibuffer, selects the window showing the completion list buffer (switchto-completions). This paves the way for using the commands below. (Selecting that window in the usual ways has the same effect, but this way is more convenient.)

⟨\underline{\overline{RET}}^{⟩} Typing ^{⟨}\underline{\overline{RET}}^{⟩} in the completion list buffer chooses the completion that point is in or next to (choose-completion). To use this command, you must first switch windows to the window that shows the list of completions.

⟨\underline{\overline{RIGHT}}^{⟩} Typing the right-arrow key ^{⟨}\underline{\overline{RIGHT}}^{⟩} in the completion list buffer moves point to the following completion (next-completion).

⟨\underline{\overline{LEFT}}^{⟩} Typing the left-arrow key ^{⟨}\underline{\overline{LEFT}}^{⟩} in the completion list buffer moves point toward the beginning of the buffer, to the previ-

      ous completion (previous-completion).

5.3.3 Strict Completion

There are three different ways that ^{⟨}\underline{\overline{RET}}^{⟩} can work in completing minibuffers, depending on how the argument will be used.

• Strict completion is used when it is meaningless to give any argument except one of the known alternatives. For example, when C-x k reads the name of a buffer to kill, it is meaningless to give anything but the name of an existing buffer. In strict completion, ^{⟨}\underline{\overline{RET}}^{⟩} refuses to exit if the text in the minibuffer does not complete to an exact match.

• Cautious completion is similar to strict completion, except that ^{⟨}\underline{\overline{RET}}^{⟩} exits only if the text was an exact match already, not needing completion. If the text is not an exact match, ^{⟨}\underline{\overline{RET}}^{⟩} does not exit, but it does complete the text. If it completes to an exact match, a second ^{⟨}\underline{\overline{RET}}^{⟩} will exit. Cautious completion is used for reading file names for files that must already exist.

• Permissive completion is used when any string whatever is meaningful, and the list of completion alternatives is just a guide. For example, when C-x C-f reads the name of a file to visit, any file name is allowed, in case you want to create a file. In permissive completion, ^{⟨}\underline{\overline{RET}}^{⟩} takes the text in the minibuffer exactly as given, without completing it.

The completion commands display a list of all possible completions in a window whenever there is more than one possibility for the very next character. Also, typing ? explicitly requests such a list. If the list of completions is long, you can scroll it with C-M-v (see Section 16.3 [Other Window], page 150).

5.3.4 Completion Options

When completion is done on file names, certain file names are usually ignored. The variable completion-ignored-extensions contains a list of strings; a file whose name ends in any of those strings is ignored as a possible completion. The standard value of this variable has several elements including ".o", ".elc", ".dvi" and "~". The effect is that, for example, ‘foo’ can complete to ‘foo.c’ even though ‘foo.o’ exists as well. However, if all the possible completions end in “ignored” strings, then they are not ignored. Ignored extensions do not apply to lists of completions—those always mention all possible completions.

Normally, a completion command that finds the next character is undetermined automatically displays a list of all possible completions. If the variable completion-auto-help is set to nil, this does not happen, and you must type ? to display the possible completions.

The complete library implements a more powerful kind of completion that can complete multiple words at a time. For example, it can complete the command name abbreviation p-b into print-buffer, because no other command starts with two words whose initials are ‘p’ and ‘b’. To use this library, put (load "complete") in your ‘~/.emacs’ file (see Section 31.7 [Init File], page 411).

Icomplete mode presents a constantly-updated display that tells you what completions are available for the text you’ve entered so far. The command to enable or disable this minor mode is M-x icomplete-mode.

5.4 Minibuffer History

Every argument that you enter with the minibuffer is saved on a minibuffer history list so that you can use it again later in another argument. Special commands load the text of an earlier argument in the minibuffer. They discard the old minibuffer contents, so you can think of them as moving through the history of previous arguments.

⟨\underline{\overline{up}}⟩

M-p Move to the next earlier argument string saved in the minibuffer

          history (previous-history-element).

⟨\underline{\overline{down}}⟩

M-n Move to the next later argument string saved in the minibuffer

          history (next-history-element).

M-r regexp ^{⟨}\underline{\overline{RET}}^{⟩} Move to an earlier saved argument in the minibuffer history

          that has a match for regexp (previous-matching-history-
          element).

M-s regexp ^{⟨}\underline{\overline{RET}}^{⟩} Move to a later saved argument in the minibuffer history that

          has a match for regexp (next-matching-history-element).

The simplest way to reuse the saved arguments in the history list is to move through the history list one element at a time. While in the minibuffer, use M-p or up-arrow (previous-history-element) to “move to” the next earlier minibuffer input, and use M-n or down-arrow (next-history-element) to “move to” the next later input.

The previous input that you fetch from the history entirely replaces the contents of the minibuffer. To use it as the argument, exit the minibuffer as usual with ^{⟨}\underline{\overline{RET}}^{⟩}. You can also edit the text before you reuse it; this does not change the history element that you “moved” to, but your new argument does go at the end of the history list in its own right.

For many minibuffer arguments there is a “default” value. In some cases, the minibuffer history commands know the default value. Then you can insert the default value into the minibuffer as text by using M-n to move “into the future” in the history. Eventually we hope to make this feature available whenever the minibuffer has a default value.

There are also commands to search forward or backward through the history; they search for history elements that match a regular expression that you specify with the minibuffer. M-r (previous-matching-history-element) searches older elements in the history, while M-s (next-matching-history-element) searches newer elements. By special dispensation, these commands can use the minibuffer to read their arguments even though you are already in the minibuffer when you issue them. As with incremental searching, an uppercase letter in the regular expression makes the search case-sensitive (see Section 12.6 [Search Case], page 92).

All uses of the minibuffer record your input on a history list, but there are separate history lists for different kinds of arguments. For example, there is a list for file names, used by all the commands that read file names. (As a special feature, this history list records the absolute file name, no more and no less, even if that is not how you entered the file name.)

There are several other very specific history lists, including one for command names read by M-x, one for buffer names, one for arguments of commands like query-replace, and one for compilation commands read by compile. Finally, there is one “miscellaneous” history list that most minibuffer arguments use.

The variable history-length specifies the maximum length of a minibuffer history list; once a list gets that long, the oldest element is deleted each time an element is added. If the value of history-length is t, though, there is no maximum length and elements are never deleted.

5.5 Repeating Minibuffer Commands

Every command that uses the minibuffer at least once is recorded on a special history list, together with the values of its arguments, so that you can repeat the entire command. In particular, every use of M-x is recorded there, since M-x uses the minibuffer to read the command name.

C-x ^{⟨}\underline{\overline{ESC}}^{⟩} ^{⟨}\underline{\overline{ESC}}^{⟩} Re-execute a recent minibuffer command (repeat-complex-

          command).
M-x list-command-history

Display the entire command history, showing all the commands C-x ^{⟨}\underline{\overline{ESC}}^{⟩} ^{⟨}\underline{\overline{ESC}}^{⟩} can repeat, most recent first.

C-x ^{⟨}\underline{\overline{ESC}}^{⟩} ^{⟨}\underline{\overline{ESC}}^{⟩} is used to re-execute a recent minibuffer-using command. With no argument, it repeats the last such command. A numeric argument specifies which command to repeat; one means the last one, and larger numbers specify earlier ones.

C-x ^{⟨}\underline{\overline{ESC}}^{⟩} ^{⟨}\underline{\overline{ESC}}^{⟩} works by turning the previous command into a Lisp expression and then entering a minibuffer initialized with the text for that expression. If you type just ^{⟨}\underline{\overline{RET}}^{⟩}, the command is repeated as before. You can also change the command by editing the Lisp expression. Whatever expression you finally submit is what will be executed. The repeated command is added to the front of the command history unless it is identical to the most recently executed command already there.

Even if you don’t understand Lisp syntax, it will probably be obvious which command is displayed for repetition. If you do not change the text, it will repeat exactly as before.

Once inside the minibuffer for C-x ^{⟨}\underline{\overline{ESC}}^{⟩} ^{⟨}\underline{\overline{ESC}}^{⟩}, you can use the minibuffer history commands (M-p, M-n, M-r, M-s; see Section 5.4 [Minibuffer History], page 44) to move through the history list of saved entire commands. After finding the desired previous command, you can edit its expression as usual and then resubmit it by typing ^{⟨}\underline{\overline{RET}}^{⟩} as usual.

The list of previous minibuffer-using commands is stored as a Lisp list in the variable command-history. Each element is a Lisp expression which describes one command and its arguments. Lisp programs can re-execute a command by calling eval with the command-history element.

6 Running Commands by Name

The Emacs commands that are used often or that must be quick to type are bound to keys—short sequences of characters—for convenient use. Other Emacs commands that do not need to be brief are not bound to keys; to run them, you must refer to them by name.

A command name is, by convention, made up of one or more words, separated by hyphens; for example, auto-fill-mode or manual-entry. The use of English words makes the command name easier to remember than a key made up of obscure characters, even though it is more characters to type.

The way to run a command by name is to start with M-x, type the command name, and finish it with ^{⟨}\underline{\overline{RET}}^{⟩}. M-x uses the minibuffer to read the command name. ^{⟨}\underline{\overline{RET}}^{⟩} exits the minibuffer and runs the command. The string ‘M-x’ appears at the beginning of the minibuffer as a prompt to remind you to enter the name of a command to be run. See Chapter 5 [Minibuffer], page 39, for full information on the features of the minibuffer.

You can use completion to enter the command name. For example, the command forward-char can be invoked by name by typing

    M-x forward-char ^{⟨}\underline{\overline{RET}}^{⟩}

or

M-x forw ^{⟨}\underline{\overline{TAB}}^{⟩} c ^{⟨}\underline{\overline{RET}}^{⟩}

Note that forward-char is the same command that you invoke with the key C-f. You can run any Emacs command by name using M-x, whether or not any keys are bound to it.

If you type C-g while the command name is being read, you cancel the M-x command and get out of the minibuffer, ending up at top level.

To pass a numeric argument to the command you are invoking with M-x, specify the numeric argument before the M-x. M-x passes the argument along to the command it runs. The argument value appears in the prompt while the command name is being read.

If the command you type has a key binding of its own, Emacs mentions this in the echo area, two seconds after the command finishes (if you don’t type anything else first). For example, if you type M-x forward-word, the message says that you can run the same command more easily by typing M-f. You can turn off these messages by setting suggest-key-bindings to

nil.

Normally, when describing in this manual a command that is run by name, we omit the ^{⟨}\underline{\overline{RET}}^{⟩} that is needed to terminate the name. Thus

we might speak of M-x auto-fill-mode rather than M-x auto-fill-mode

⟨\underline{\overline{RET}}^{⟩}. We mention the ^{⟨}\underline{\overline{RET}}^{⟩} only when there is a need to emphasize its presence, such as when we show the command together with following arguments.

48 GNU Emacs Manual

M-x works by running the command execute-extended-command, which is responsible for reading the name of another command and invoking it.

7 Help

Emacs provides extensive help features accessible through a single character, C-h. C-h is a prefix key that is used only for documentation-printing commands. The characters that you can type after C-h are called help options. One help option is C-h; that is how you ask for help about using C-h. To cancel, type C-g. The function key ^{⟨}\underline{\overline{F1}}^{⟩} is equivalent to C-h.

C-h C-h (help-for-help) displays a list of the possible help options, each with a brief description. Before you type a help option, you can use ^{⟨}\underline{\overline{SPC}}^{⟩} or ⟨\underline{\overline{DEL}}^{⟩} to scroll through the list.

C-h or ^{⟨}\underline{\overline{F1}}^{⟩} means “help” in various other contexts as well. For example, in the middle of query-replace, it describes the options available for how to operate on the current match. After a prefix key, it displays a list of the alternatives that can follow the prefix key. (A few prefix keys don’t support C-h, because they define other meanings for it, but they all support ^{⟨}\underline{\overline{F1}}^{⟩}.)

Most help buffers use a special major mode, Help mode, which lets you scroll conveniently with ^{⟨}\underline{\overline{SPC}}^{⟩} and ^{⟨}\underline{\overline{DEL}}^{⟩}.

Here is a summary of the defined help commands.

C-h a regexp ^{⟨}\underline{\overline{RET}}^{⟩} Display a list of commands whose names match regexp

          (apropos-command).

C-h b Display a table of all key bindings in effect now, in this order: minor mode bindings, major mode bindings, and global bindings

          (describe-bindings).

C-h c key Print the name of the command that key runs (describe-key-briefly). Here c stands for ‘character’. For more extensive information on key, use C-h k.

C-h f function ^{⟨}\underline{\overline{RET}}^{⟩} Display documentation on the Lisp function named function (describe-function). Since commands are Lisp functions, a command name may be used.

C-h h Display the ‘hello’ file, which shows examples of various character sets.

C-h i Run Info, the program for browsing documentation files (info). The complete Emacs manual is available on-line in Info.

C-h k key Display the name and documentation of the command that key

          runs (describe-key).

C-h l Display a description of the last 100 characters you typed (view-

          lossage).

C-h m Display documentation of the current major mode (describe-

          mode).

C-h n Display documentation of Emacs changes, most recent first (view-

emacs-news).

C-h p Find packages by topic keyword (finder-by-keyword).

C-h s Display current contents of the syntax table, plus an explanation of what they mean (describe-syntax). See Section 31.6 [Syntax], page 411.

C-h t Enter the Emacs interactive tutorial (help-with-tutorial).

C-h v var ^{⟨}\underline{\overline{RET}}^{⟩} Display the documentation of the Lisp variable var (describe-

variable).

C-h w command ^{⟨}\underline{\overline{RET}}^{⟩} Print which keys run the command named command (where-

is).

C-h C coding ^{⟨}\underline{\overline{RET}}^{⟩} Describe coding system coding (describe-coding-system).

C-h C ^{⟨}\underline{\overline{RET}}^{⟩} Describe the coding systems currently in use.

C-h I method ^{⟨}\underline{\overline{RET}}^{⟩} Describe an input method (describe-input-method).

C-h L language-env ^{⟨}\underline{\overline{RET}}^{⟩} Describe information on the character sets, coding systems and input methods used for language environment language-env

(describe-language-support).

C-h C-c Display the copying conditions for GNU Emacs.

C-h C-d Display information about getting new versions of GNU Emacs.

C-h C-f function ^{⟨}\underline{\overline{RET}}^{⟩} Enter Info and go to the node documenting the Emacs function

function (Info-goto-emacs-command-node).
C-h C-k key

Enter Info and go to the node where the key sequence key is

documented (Info-goto-emacs-key-command-node).

C-h C-p Display information about the GNU Project.

C-h ^{⟨}\underline{\overline{TAB}}^{⟩} symbol ^{⟨}\underline{\overline{RET}}^{⟩} Display the Info documentation on symbol symbol according to the programming language you are editing (info-lookup-

symbol).

7.1 Documentation for a Key

   The most basic C-h options are C-h c (describe-key-briefly) and

C-h k (describe-key). C-h c key prints in the echo area the name of the command that key is bound to. For example, C-h c C-f prints ‘forward-char’. Since command names are chosen to describe what the commands do, this is a good way to get a very brief description of what key does.

C-h k key is similar but gives more information: it displays the documentation string of the command as well as its name. This is too big for the echo area, so a window is used for the display.

C-h c and C-h k work for any sort of key sequences, including function keys and mouse events.

7.2 Help by Command or Variable Name

C-h f (describe-function) reads the name of a Lisp function using the minibuffer, then displays that function’s documentation string in a window. Since commands are Lisp functions, you can use this to get the documentation of a command that you know by name. For example,

    C-h f auto-fill-mode ^{⟨}\underline{\overline{RET}}^{⟩}

displays the documentation of auto-fill-mode. This is the only way to get the documentation of a command that is not bound to any key (one which you would normally run using M-x).

C-h f is also useful for Lisp functions that you are planning to use in a Lisp program. For example, if you have just written the expression (makevector len) and want to check that you are using make-vector properly, type C-h f make-vector ^{⟨}\underline{\overline{RET}}^{⟩}. Because C-h f allows all function names, not just command names, you may find that some of your favorite abbreviations that work in M-x don’t work in C-h f. An abbreviation may be unique among command names yet fail to be unique when other function names are allowed.

The function name for C-h f to describe has a default which is used if you type^{⟨}\underline{\overline{RET}}^{⟩} leaving the minibuffer empty. The default is the function called by the innermost Lisp expression in the buffer around point, provided that is a valid, defined Lisp function name. For example, if point is located following the text ‘(make-vector (car x)’, the innermost list containing point is the one that starts with ‘(make-vector’, so the default is to describe the function

make-vector.

C-h f is often useful just to verify that you have the right spelling for the function name. If C-h f mentions a name from the buffer as the default, that name must be defined as a Lisp function. If that is all you want to know, just type C-g to cancel the C-h f command, then go on editing.

C-h w command ^{⟨}\underline{\overline{RET}}^{⟩} tells you what keys are bound to command. It prints a list of the keys in the echo area. If it says the command is not on any key, you must use M-x to run it. C-h w runs the command where-is.

C-h v (describe-variable) is like C-h f but describes Lisp variables instead of Lisp functions. Its default is the Lisp symbol around or before point, but only if that is the name of a known Lisp variable. See Section 31.2 [Variables], page 385.

7.3 Apropos

A more sophisticated sort of question to ask is, “What are the commands for working with files?” To ask this question, type C-h a file ^{⟨}\underline{\overline{RET}}^{⟩}, which displays a list of all command names that contain ‘file’, including copyfile, find-file, and so on. With each command name appears a brief description of how to use the command, and what keys you can currently invoke it with. For example, it would say that you can invoke find-file by typing C-x C-f. The a in C-h a stands for ‘Apropos’; C-h a runs the command apropos-command. This command normally checks only commands (interactive functions); if you specify a prefix argument, it checks noninteractive functions as well.

Because C-h a looks only for functions whose names contain the string you specify, you must use ingenuity in choosing the string. If you are looking for commands for killing backwards and C-h a kill-backwards ^{⟨}\underline{\overline{RET}}^{⟩} doesn’t reveal any, don’t give up. Try just kill, or just backwards, or just back. Be persistent. Also note that you can use a regular expression as the argument, for more flexibility (see Section 12.5 [Regexps], page 88).

Here is a set of arguments to give to C-h a that covers many classes of Emacs commands, since there are strong conventions for naming the standard Emacs commands. By giving you a feel for the naming conventions, this set should also serve to aid you in developing a technique for picking apropos strings.

char, line, word, sentence, paragraph, region, page, sexp, list, defun, rect, buffer, frame, window, face, file, dir, register, mode, beginning, end, forward, backward, next, previous, up, down, search, goto, kill, delete, mark, insert, yank, fill, indent, case, change, set, what, list, find, view, describe, default. To list all user variables that match a regexp, use the command M-x

apropos-variable. This command shows only user variables and customization options by default; if you specify a prefix argument, it checks all variables.

To list all Lisp symbols that contain a match for a regexp, not just the ones that are defined as commands, use the command M-x apropos instead of C-h a. This command does not check key bindings by default; specify a numeric argument if you want it to check them.

The apropos-documentation command is like apropos except that it

searches documentation strings as well as symbol names for matches for the specified regular expression.

The apropos-value command is like apropos except that it searches symbols’ values for matches for the specified regular expression. This command does not check function definitions or property lists by default; specify a numeric argument if you want it to check them.

If the variable apropos-do-all is non-nil, the commands above all behave as if they had been given a prefix argument.

If you want more information about a function definition, variable or symbol property listed in the Apropos buffer, you can click on it with Mouse-2 or move there and type ^{⟨}\underline{\overline{RET}}^{⟩}.

7.4 Keyword Search for Lisp Libraries

The C-h p command lets you search the standard Emacs Lisp libraries by topic keywords. Here is a partial list of keywords you can use:

abbrev — abbreviation handling, typing shortcuts, macros. bib — support for the bibliography processor bib. c — C and C++ language support. calendar — calendar and time management support. comm — communications, networking, remote access to files. data — support for editing files of data. docs — support for Emacs documentation. emulations — emulations of other editors. extensions — emacs Lisp language extensions. faces — support for using faces (fonts and colors; see Section 17.13 [Faces], page 164). frames — support for Emacs frames and window systems. games — games, jokes and amusements. hardware — support for interfacing with exotic hardware. help — support for on-line help systems. hypermedia — support for links within text, or other media types. i18n — internationalization and alternate character-set support. internal — code for Emacs internals, build process, defaults. languages — specialized modes for editing programming languages. lisp — support for using Lisp (including Emacs Lisp). local — libraries local to your site. maint — maintenance aids for the Emacs development group. mail — modes for electronic-mail handling. matching — searching and matching. news — support for netnews reading and posting. non-text — support for editing files that are not ordinary text. oop — support for object-oriented programming. outlines — hierarchical outlining. processes — process, subshell, compilation, and job control support. terminals — support for terminal types. tex — support for the TEX formatter. tools — programming tools. unix — front-ends/assistants for, or emulators of, Unix features. vms — support code for VMS. wp — word processing.

7.5 Help for International Language Support

   You can use the command C-h L (describe-language-environment) to

find out the support for a specific language environment. See Section 18.3 [Language Environments], page 172. This tells you which languages this language environment is useful for, and lists the character sets, coding systems, and input methods that go with it. It also shows some sample text to illustrate scripts.

The command C-h h (view-hello-file) displays the file ‘etc/HELLO’, which shows how to say “hello” in many languages.

The command C-h I (describe-input-method) describes information about input methods—either a specified input method, or by default the input method in use. See Section 18.4 [Input Methods], page 173.

The command C-h C (describe-coding-system) describes information about coding systems—either a specified coding system, or the ones currently in use. See Section 18.6 [Coding Systems], page 175.

7.6 Help Mode Commands

Help buffers provide the commands of View mode (see Section 14.10 [Misc File Ops], page 137), plus a few special commands of their own. ⟨\underline{\overline{SPC}}^{⟩} Scroll forward. ⟨\underline{\overline{DEL}}^{⟩} Scroll backward. ⟨\underline{\overline{RET}}^{⟩} Follow a cross reference at point. ⟨\underline{\overline{TAB}}^{⟩} Move point forward to the next cross reference. S-^{⟨}\underline{\overline{TAB}}^{⟩} Move point back to the previous cross reference. Mouse-2 Follow a cross reference that you click on.

When a command name (see Chapter 6 [Running Commands by Name], page 47) or variable name (see Section 31.2 [Variables], page 385) appears in the documentation, it normally appears inside paired single-quotes. You can click on the name with Mouse-2, or move point there and type ^{⟨}\underline{\overline{RET}}^{⟩}, to view the documentation of that command or variable. Use C-c C-b to retrace your steps.

There are convenient commands for moving point to cross references in the help text. ^{⟨}\underline{\overline{TAB}}^{⟩} (help-next-ref) moves point down to the next cross reference. Use S-^{⟨}\underline{\overline{TAB}}^{⟩} to move point up to the previous cross reference

(help-previous-ref).

7.7 Other Help Commands

C-h i (info) runs the Info program, which is used for browsing through structured documentation files. The entire Emacs manual is available within Info. Eventually all the documentation of the GNU system will be available. Type h after entering Info to run a tutorial on using Info.

If you specify a numeric argument, C-h i prompts for the name of a documentation file. This way, you can browse a file which doesn’t have an entry in the top-level Info menu. It is also handy when you need to get to the documentation quickly, and you know the exact name of the file.

There are two special help commands for accessing Emacs documentation through Info. C-h C-f function ^{⟨}\underline{\overline{RET}}^{⟩} enters Info and goes straight to the documentation of the Emacs function function. C-h C-k key enters Info and goes straight to the documentation of the key key. These two

keys run the commands Info-goto-emacs-command-node and Info-goto-
emacs-key-command-node.

When editing a program, if you have an Info version of the manual for the programming language, you can use the command C-h C-i to refer to the manual documentation for a symbol (keyword, function or variable). The details of how this command works depend on the major mode.

If something surprising happens, and you are not sure what commands you typed, use C-h l (view-lossage). C-h l prints the last 100 command characters you typed in. If you see commands that you don’t know, you can use C-h c to find out what they do.

Emacs has numerous major modes, each of which redefines a few keys and makes a few other changes in how editing works. C-h m (describe-mode) prints documentation on the current major mode, which normally describes all the commands that are changed in this mode.

   C-h b (describe-bindings) and C-h s (describe-syntax) present other

information about the current Emacs mode. C-h b displays a list of all the key bindings now in effect; the local bindings defined by the current minor modes first, then the local bindings defined by the current major mode, and finally the global bindings (see Section 31.4 [Key Bindings], page 400). C-h s displays the contents of the syntax table, with explanations of each character’s syntax (see Section 31.6 [Syntax], page 411).

You can get a similar list for a particular prefix key by typing C-h after the prefix key. (There are a few prefix keys for which this does not work— those that provide their own bindings for C-h. One of these is ^{⟨}\underline{\overline{ESC}}^{⟩}, because ⟨\underline{\overline{ESC}}^{⟩} C-h is actually C-M-h, which marks a defun.)

The other C-h options display various files of useful information. C-h C-w displays the full details on the complete absence of warranty for GNU

Emacs. C-h n (view-emacs-news) displays the file ‘emacs/etc/NEWS’, which

contains documentation on Emacs changes arranged chronologically. C-h t (help-with-tutorial) displays the learn-by-doing Emacs tutorial. C-h C-c

(describe-copying) displays the file ‘emacs/etc/COPYING’, which tells you

the conditions you must obey in distributing copies of Emacs. C-h C-d

(describe-distribution) displays the file ‘emacs/etc/DISTRIB’, which

tells you how you can order a copy of the latest version of Emacs. C-h C-p (describe-project) displays general information about the GNU Project.

8 The Mark and the Region

Many Emacs commands operate on an arbitrary contiguous part of the current buffer. To specify the text for such a command to operate on, you set the mark at one end of it, and move point to the other end. The text between point and the mark is called the region. Emacs highlights the region whenever there is one, if you enable Transient Mark mode (see Section 8.2 [Transient Mark], page 58).

You can move point or the mark to adjust the boundaries of the region. It doesn’t matter which one is set first chronologically, or which one comes earlier in the text. Once the mark has been set, it remains where you put it until you set it again at another place. Each Emacs buffer has its own mark, so that when you return to a buffer that had been selected previously, it has the same mark it had before.

Many commands that insert text, such as C-y (yank) and M-x insert-buffer, position point and the mark at opposite ends of the inserted text, so that the region contains the text just inserted.

Aside from delimiting the region, the mark is also useful for remembering a spot that you may want to go back to. To make this feature more useful, each buffer remembers 16 previous locations of the mark in the mark ring.

8.1 Setting the Mark

Here are some commands for setting the mark:

C-^{⟨}\underline{\overline{SPC}}^{⟩} Set the mark where point is (set-mark-command).

C-@ The same.

C-x C-x Interchange mark and point (exchange-point-and-mark).
Drag-Mouse-1

Set point and the mark around the text you drag across.

Mouse-3 Set the mark where point is, then move point to where you click

          (mouse-save-then-kill).

For example, suppose you wish to convert part of the buffer to upper case, using the C-x C-u (upcase-region) command, which operates on the text in the region. You can first go to the beginning of the text to be capitalized, type C-^{⟨}\underline{\overline{SPC}}^{⟩} to put the mark there, move to the end, and then type C-x C-u. Or, you can set the mark at the end of the text, move to the beginning, and then type C-x C-u.

The most common way to set the mark is with the C-^{⟨}\underline{\overline{SPC}}^{⟩} command (set-mark-command). This sets the mark where point is. Then you can move point away, leaving the mark behind.

There are two ways to set the mark with the mouse. You can drag mouse button one across a range of text; that puts point where you release the mouse button, and sets the mark at the other end of that range. Or you can click mouse button three, which sets the mark at point (like C-^{⟨}\underline{\overline{SPC}}^{⟩}) and them moves point (like Mouse-1). Both of these methods copy the region into the kill ring in addition to setting the mark; that gives behavior consistent with other window-driven applications, but if you don’t want to modify the kill ring, you must use keyboard commands to set the mark. See Section 17.1 [Mouse Commands], page 155.

Ordinary terminals have only one cursor, so there is no way for Emacs to show you where the mark is located. You have to remember. The usual solution to this problem is to set the mark and then use it soon, before you forget where it is. Alternatively, you can see where the mark is with

the command C-x C-x (exchange-point-and-mark) which puts the mark

where point was and point where the mark was. The extent of the region is unchanged, but the cursor and point are now at the previous position of the mark. In Transient Mark mode, this command reactivates the mark.

C-x C-x is also useful when you are satisfied with the position of point but want to move the other end of the region (where the mark is); do C-x C-x to put point at that end of the region, and then move it. A second use of C-x C-x, if necessary, puts the mark at the new position with point back at its original position.

There is no such character as C-^{⟨}\underline{\overline{SPC}}^{⟩} in ASCII; when you type ^{⟨}\underline{\overline{SPC}}^{⟩} while holding down ^{⟨}\underline{\overline{CTRL}}^{⟩}, what you get on most ordinary terminals is the character C-@. This key is actually bound to set-mark-command. But unless you are unlucky enough to have a terminal where typing C-^{⟨}\underline{\overline{SPC}}^{⟩} does not produce C-@, you might as well think of this character as C-^{⟨}\underline{\overline{SPC}}^{⟩}. Under X, C-^{⟨}\underline{\overline{SPC}}^{⟩} is actually a distinct character, but its binding is still set-mark-

command.

8.2 Transient Mark Mode

Emacs can highlight the current region, using X Windows. But normally it does not. Why not?

Highlighting the region doesn’t work well ordinarily in Emacs, because once you have set a mark, there is always a region (in that buffer). And highlighting the region all the time would be a nuisance.

You can turn on region highlighting by enabling Transient Mark mode. This is a more rigid mode of operation in which the region “lasts” only temporarily, so you must set up a region for each command that uses one. In Transient Mark mode, most of the time there is no region; therefore, highlighting the region when it exists is convenient.

To enable Transient Mark mode, type M-x transient-mark-mode. This command toggles the mode, so you can repeat the command to turn off the mode.

Here are the details of Transient Mark mode:

• To set the mark, type C-^{⟨}\underline{\overline{SPC}}^{⟩} (set-mark-command). This makes the mark active; as you move point, you will see the region highlighting grow and shrink.

• The mouse commands for specifying the mark also make it active. So do keyboard commands whose purpose is to specify a region, including

    M-@, C-M-@, M-h, C-M-h, C-x C-p, and C-x h.

• When the mark is active, you can execute commands that operate on the region, such as killing, indenting, or writing to a file.

• Any change to the buffer, such as inserting or deleting a character, deactivates the mark. This means any subsequent command that operates on a region will get an error and refuse to operate. You can make the region active again by typing C-x C-x.

• Commands like M-> and C-s that “leave the mark behind” in addition to some other primary purpose do not activate the new mark. You can activate the new region by executing C-x C-x (exchange-point-and-

    mark).

• C-s when the mark is active does not alter the mark.

• Quitting with C-g deactivates the mark.

Highlighting of the region uses the region face; you can customize how the region is highlighted by changing this face. See Section 31.2.2.3 [Face Customization], page 390.

When multiple windows show the same buffer, they can have different regions, because they can have different values of point (though they all share one common mark position). Ordinarily, only the selected window highlights its region (see Chapter 16 [Windows], page 149). However, if the

variable highlight-nonselected-windows is non-nil, then each window

highlights its own region (provided that Transient Mark mode is enabled and the window’s buffer’s mark is active).

When Transient Mark mode is not enabled, every command that sets the mark also activates it, and nothing ever deactivates it.

If the variable mark-even-if-inactive is non-nil in Transient Mark mode, then commands can use the mark and the region even when it is inactive. Region highlighting appears and disappears just as it normally does in Transient Mark mode, but the mark doesn’t really go away when the highlighting disappears.

Transient Mark mode is also sometimes known as “Zmacs mode” because the Zmacs editor on the MIT Lisp Machine handled the mark in a similar way.

8.3 Operating on the Region

Once you have a region and the mark is active, here are some of the ways you can operate on the region:

• Kill it with C-w (see Section 9.1 [Killing], page 63).

• Save it in a register with C-x r s (see Chapter 10 [Registers], page 73).

• Save it in a buffer or a file (see Section 9.3 [Accumulating Text], page 68).

• Convert case with C-x C-l or C-x C-u (see Section 21.6 [Case], page 202).

• Indent it with C-x ^{⟨}\underline{\overline{TAB}}^{⟩} or C-M-\ (see Chapter 20 [Indentation], page 189).

• Fill it as text with M-x fill-region (see Section 21.5 [Filling], page 197).

• Print hardcopy with M-x print-region (see Section 30.4 [Hardcopy], page 369).

• Evaluate it as Lisp code with M-x eval-region (see Section 23.8 [Lisp Eval], page 278). Most commands that operate on the text in the region have the word

region in their names.

8.4 Commands to Mark Textual Objects

Here are the commands for placing point and the mark around a textual object such as a word, list, paragraph or page.

M-@ Set mark after end of next word (mark-word). This command and the following one do not move point.

C-M-@ Set mark after end of next Lisp expression (mark-sexp).

M-h Put region around current paragraph (mark-paragraph).

C-M-h Put region around current Lisp defun (mark-defun).

C-x h Put region around entire buffer (mark-whole-buffer).

C-x C-p Put region around current page (mark-page).

M-@ (mark-word) puts the mark at the end of the next word, while C-M-@ (mark-sexp) puts it at the end of the next Lisp expression. These commands handle arguments just like M-f and C-M-f.

Other commands set both point and mark, to delimit an object in the buffer. For example, M-h (mark-paragraph) moves point to the beginning of the paragraph that surrounds or follows point, and puts the mark at the end of that paragraph (see Section 21.3 [Paragraphs], page 195). It prepares the region so you can indent, case-convert, or kill a whole paragraph.

C-M-h (mark-defun) similarly puts point before and the mark after the current or following defun (see Section 22.4 [Defuns], page 224). C-x C-p (mark-page) puts point before the current page, and mark at the end (see Section 21.4 [Pages], page 196). The mark goes after the terminating page delimiter (to include it), while point goes after the preceding page delimiter (to exclude it). A numeric argument specifies a later page (if positive) or an earlier page (if negative) instead of the current page.

Finally, C-x h (mark-whole-buffer) sets up the entire buffer as the region, by putting point at the beginning and the mark at the end.

In Transient Mark mode, all of these commands activate the mark.

8.5 The Mark Ring

Aside from delimiting the region, the mark is also useful for remembering a spot that you may want to go back to. To make this feature more useful, each buffer remembers 16 previous locations of the mark, in the mark ring. Commands that set the mark also push the old mark onto this ring. To return to a marked location, use C-u C-^{⟨}\underline{\overline{SPC}}^{⟩} (or C-u C-@); this is the command set-mark-command given a numeric argument. It moves point to where the mark was, and restores the mark from the ring of former marks. Thus, repeated use of this command moves point to all of the old marks on the ring, one by one. The mark positions you move through in this way are not lost; they go to the end of the ring.

Each buffer has its own mark ring. All editing commands use the current buffer’s mark ring. In particular, C-u C-^{⟨}\underline{\overline{SPC}}^{⟩} always stays in the same buffer.

Many commands that can move long distances, such as M-< (beginningof-buffer), start by setting the mark and saving the old mark on the mark ring. This is to make it easier for you to move back later. Searches set the mark if they move point. You can tell when a command sets the mark because it displays ‘Mark Set’ in the echo area.

If you want to move back to the same place over and over, the mark ring may not be convenient enough. If so, you can record the position in a register for later retrieval (see Section 10.1 [RegPos], page 73).

The variable mark-ring-max specifies the maximum number of entries to keep in the mark ring. If that many entries exist and another one is pushed, the last one in the list is discarded. Repeating C-u C-^{⟨}\underline{\overline{SPC}}^{⟩} cycles through the positions currently in the ring.

The variable mark-ring holds the mark ring itself, as a list of marker objects, with the most recent first. This variable is local in every buffer.

8.6 The Global Mark Ring

In addition to the ordinary mark ring that belongs to each buffer, Emacs has a single global mark ring. It records a sequence of buffers in which you have recently set the mark, so you can go back to those buffers.

Setting the mark always makes an entry on the current buffer’s mark ring. If you have switched buffers since the previous mark setting, the new mark position makes an entry on the global mark ring also. The result is that the global mark ring records a sequence of buffers that you have been in, and, for each buffer, a place where you set the mark.

The command C-x C-^{⟨}\underline{\overline{SPC}}^{⟩} (pop-global-mark) jumps to the buffer and position of the latest entry in the global ring. It also rotates the ring, so that successive uses of C-x C-^{⟨}\underline{\overline{SPC}}^{⟩} take you to earlier and earlier buffers.

9 Killing and Moving Text

Killing means erasing text and copying it into the kill ring, from which it can be retrieved by yanking it. Some systems use the terms “cutting” and “pasting” for these operations.

The commonest way of moving or copying text within Emacs is to kill it and later yank it elsewhere in one or more places. This is very safe because Emacs remembers several recent kills, not just the last one. It is versatile, because the many commands for killing syntactic units can also be used for moving those units. But there are other ways of copying text for special purposes.

Emacs has only one kill ring for all buffers, so you can kill text in one buffer and yank it in another buffer.

9.1 Deletion and Killing

Most commands which erase text from the buffer save it in the kill ring so that you can move or copy it to other parts of the buffer. These commands are known as kill commands. The rest of the commands that erase text do not save it in the kill ring; they are known as delete commands. (This distinction is made only for erasure of text in the buffer.) If you do a kill or delete command by mistake, you can use the C-x u (undo) command to undo it (see Section 4.4 [Undo], page 32).

The delete commands include C-d (delete-char) and ^{⟨}\underline{\overline{DEL}}^{⟩} (deletebackward-char), which delete only one character at a time, and those commands that delete only spaces or newlines. Commands that can destroy significant amounts of nontrivial data generally kill. The commands’ names and individual descriptions use the words ‘kill’ and ‘delete’ to say which they do.

9.1.1 Deletion

C-d Delete next character (delete-char).

⟨\underline{\overline{DEL}}^{⟩} Delete previous character (delete-backward-char).

M-\ Delete spaces and tabs around point (delete-horizontal-

          space).

M-^{⟨}\underline{\overline{SPC}}^{⟩} Delete spaces and tabs around point, leaving one space (just-

          one-space).

C-x C-o Delete blank lines around the current line (delete-blank-lines).

M-^Join two lines by deleting the intervening newline, along with any indentation following it (delete-indentation).

The most basic delete commands are C-d (delete-char) and ^{⟨}\underline{\overline{DEL}}^{⟩} (delete-backward-char). C-d deletes the character after point, the one the cursor is “on top of.” This doesn’t move point. ^{⟨}\underline{\overline{DEL}}^{⟩} deletes the character before the cursor, and moves point back. You can delete newlines like any other characters in the buffer; deleting a newline joins two lines. Actually, C-d and ^{⟨}\underline{\overline{DEL}}^{⟩} aren’t always delete commands; when given arguments, they kill instead, since they can erase more than one character this way.

The other delete commands are those which delete only whitespace characters: spaces, tabs and newlines. M-\ (delete-horizontal-space) deletes all the spaces and tab characters before and after point. M-^{⟨}\underline{\overline{SPC}}^{⟩} (just-one-space) does likewise but leaves a single space after point, regardless of the number of spaces that existed previously (even zero).

C-x C-o (delete-blank-lines) deletes all blank lines after the current line. If the current line is blank, it deletes all blank lines preceding the current line as well (leaving one blank line, the current line).

M-^(delete-indentation) joins the current line and the previous line, by deleting a newline and all surrounding spaces, usually leaving a single space. See Chapter 20 [Indentation], page 189.

9.1.2 Killing by Lines

C-k Kill rest of line or one or more lines (kill-line).

The simplest kill command is C-k. If given at the beginning of a line, it kills all the text on the line, leaving it blank. When used on a blank line, it kills the whole line including its newline. To kill an entire non-blank line, go to the beginning and type C-k twice.

More generally, C-k kills from point up to the end of the line, unless it is at the end of a line. In that case it kills the newline following point, thus merging the next line into the current one. Spaces and tabs that you can’t see at the end of the line are ignored when deciding which case applies, so if point appears to be at the end of the line, you can be sure C-k will kill the newline.

When C-k is given a positive argument, it kills that many lines and the newlines that follow them (however, text on the current line before point is spared). With a negative argument −n, it kills n lines preceding the current line (together with the text on the current line before point). Thus, C-u - 2 C-k at the front of a line kills the two previous lines.

C-k with an argument of zero kills the text before point on the current line.

If the variable kill-whole-line is non-nil, C-k at the very beginning of a line kills the entire line including the following newline. This variable is normally nil.

9.1.3 Other Kill Commands

C-w Kill region (from point to the mark) (kill-region).

M-d Kill word (kill-word). See Section 21.1 [Words], page 193.

M-^{⟨}\underline{\overline{DEL}}^{⟩} Kill word backwards (backward-kill-word).

C-x ^{⟨}\underline{\overline{DEL}}^{⟩} Kill back to beginning of sentence (backward-kill-sentence). See Section 21.2 [Sentences], page 194.

M-k Kill to end of sentence (kill-sentence).

C-M-k Kill sexp (kill-sexp). See Section 22.2 [Lists], page 222.

M-z char Kill through the next occurrence of char (zap-to-char).

A kill command which is very general is C-w (kill-region), which kills everything between point and the mark. With this command, you can kill any contiguous sequence of characters, if you first set the region around them.

A convenient way of killing is combined with searching: M-z (zap-to-char) reads a character and kills from point up to (and including) the next occurrence of that character in the buffer. A numeric argument acts as a repeat count. A negative argument means to search backward and kill text before point.

Other syntactic units can be killed: words, with M-^{⟨}\underline{\overline{DEL}}^{⟩} and M-d (see Section 21.1 [Words], page 193); sexps, with C-M-k (see Section 22.2 [Lists], page 222); and sentences, with C-x ^{⟨}\underline{\overline{DEL}}^{⟩} and M-k (see Section 21.2 [Sentences], page 194).

You can use kill commands in read-only buffers. They don’t actually change the buffer, and they beep to warn you of that, but they do copy the text you tried to kill into the kill ring, so you can yank it into other buffers. Most of the kill commands move point across the text they copy in this way, so that successive kill commands build up a single kill ring entry as usual.

9.2 Yanking

Yanking means reinserting text previously killed. This is what some systems call “pasting.” The usual way to move or copy text is to kill it and then yank it elsewhere one or more times.

C-y Yank last killed text (yank).

M-y Replace text just yanked with an earlier batch of killed text (yank-

          pop).

M-w Save region as last killed text without actually killing it (kill-

          ring-save).

C-M-w Append next kill to last batch of killed text (append-next-kill).

9.2.1 The Kill Ring

All killed text is recorded in the kill ring, a list of blocks of text that have been killed. There is only one kill ring, shared by all buffers, so you can kill text in one buffer and yank it in another buffer. This is the usual way to move text from one file to another. (See Section 9.3 [Accumulating Text], page 68, for some other ways.)

The command C-y (yank) reinserts the text of the most recent kill. It leaves the cursor at the end of the text. It sets the mark at the beginning of the text. See Chapter 8 [Mark], page 57.

C-u C-y leaves the cursor in front of the text, and sets the mark after it. This happens only if the argument is specified with just a C-u, precisely. Any other sort of argument, including C-u and digits, specifies an earlier kill to yank (see Section 9.2.3 [Earlier Kills], page 67).

To copy a block of text, you can use M-w (kill-ring-save), which copies the region into the kill ring without removing it from the buffer. This is approximately equivalent to C-w followed by C-x u, except that M-w does not alter the undo history and does not temporarily change the screen.

9.2.2 Appending Kills

Normally, each kill command pushes a new entry onto the kill ring. However, two or more kill commands in a row combine their text into a single entry, so that a single C-y yanks all the text as a unit, just as it was before it was killed.

Thus, if you want to yank text as a unit, you need not kill all of it with one command; you can keep killing line after line, or word after word, until you have killed it all, and you can still get it all back at once.

Commands that kill forward from point add onto the end of the previous killed text. Commands that kill backward from point add text onto the beginning. This way, any sequence of mixed forward and backward kill commands puts all the killed text into one entry without rearrangement. Numeric arguments do not break the sequence of appending kills. For example, suppose the buffer contains this text:

    This is a line ⋆of sample text.

with point shown by ⋆. If you type M-d M-^{⟨}\underline{\overline{DEL}}^{⟩} M-d M-^{⟨}\underline{\overline{DEL}}^{⟩}, killing alternately forward and backward, you end up with ‘a line of sample’ as one entry in the kill ring, and ‘This is text.’ in the buffer. (Note the double space, which you can clean up with M-^{⟨}\underline{\overline{SPC}}^{⟩} or M-q.)

Another way to kill the same text is to move back two words with M-b M-b, then kill all four words forward with C-u M-d. This produces exactly the same results in the buffer and in the kill ring. M-f M-f C-u M-^{⟨}\underline{\overline{DEL}}^{⟩} kills the same text, all going backward; once again, the result is the same. The text in the kill ring entry always has the same order that it had in the buffer before you killed it.

If a kill command is separated from the last kill command by other commands (not just numeric arguments), it starts a new entry on the kill ring. But you can force it to append by first typing the command C-M-w (appendnext-kill) right before it. The C-M-w tells the following command, if it is a kill command, to append the text it kills to the last killed text, instead of starting a new entry. With C-M-w, you can kill several separated pieces of text and accumulate them to be yanked back in one place.

A kill command following M-w does not append to the text that M-w copied into the kill ring.

9.2.3 Yanking Earlier Kills

To recover killed text that is no longer the most recent kill, use the M-y command (yank-pop). It takes the text previously yanked and replaces it with the text from an earlier kill. So, to recover the text of the next-to-the-last kill, first use C-y to yank the last kill, and then use M-y to replace it with the previous kill. M-y is allowed only after a C-y or another M-y.

You can understand M-y in terms of a “last yank” pointer which points at an entry in the kill ring. Each time you kill, the “last yank” pointer moves to the newly made entry at the front of the ring. C-y yanks the entry which the “last yank” pointer points to. M-y moves the “last yank” pointer to a different entry, and the text in the buffer changes to match. Enough M-y commands can move the pointer to any entry in the ring, so you can get any entry into the buffer. Eventually the pointer reaches the end of the ring; the next M-y moves it to the first entry again.

M-y moves the “last yank” pointer around the ring, but it does not change the order of the entries in the ring, which always runs from the most recent kill at the front to the oldest one still remembered.

M-y can take a numeric argument, which tells it how many entries to advance the “last yank” pointer by. A negative argument moves the pointer toward the front of the ring; from the front of the ring, it moves “around” to the last entry and continues forward from there.

Once the text you are looking for is brought into the buffer, you can stop doing M-y commands and it will stay there. It’s just a copy of the kill ring entry, so editing it in the buffer does not change what’s in the ring. As long as no new killing is done, the “last yank” pointer remains at the same place in the kill ring, so repeating C-y will yank another copy of the same previous kill.

If you know how many M-y commands it would take to find the text you want, you can yank that text in one step using C-y with a numeric argument. C-y with an argument restores the text the specified number of entries back in the kill ring. Thus, C-u 2 C-y gets the next-to-the-last block of killed text. It is equivalent to C-y M-y. C-y with a numeric argument starts counting from the “last yank” pointer, and sets the “last yank” pointer to the entry that it yanks.

The length of the kill ring is controlled by the variable kill-ring-max; no more than that many blocks of killed text are saved.

The actual contents of the kill ring are stored in a variable named killring; you can view the entire contents of the kill ring with the command

C-h v kill-ring.

9.3 Accumulating Text

Usually we copy or move text by killing it and yanking it, but there are other methods convenient for copying one block of text in many places, or for copying many scattered blocks of text into one place. To copy one block to many places, store it in a register (see Chapter 10 [Registers], page 73). Here we describe the commands to accumulate scattered pieces of text into a buffer or into a file.

M-x append-to-buffer

Append region to contents of specified buffer.

M-x prepend-to-buffer

Prepend region to contents of specified buffer.

M-x copy-to-buffer

Copy region into specified buffer, deleting that buffer’s old contents.

M-x insert-buffer

Insert contents of specified buffer into current buffer at point.

M-x append-to-file

Append region to contents of specified file, at the end.

To accumulate text into a buffer, use M-x append-to-buffer. This reads a buffer name, then inserts a copy of the region into the buffer specified. If you specify a nonexistent buffer, append-to-buffer creates the buffer. The text is inserted wherever point is in that buffer. If you have been using the buffer for editing, the copied text goes into the middle of the text of the buffer, wherever point happens to be in it.

Point in that buffer is left at the end of the copied text, so successive uses of append-to-buffer accumulate the text in the specified buffer in the same order as they were copied. Strictly speaking, append-to-buffer does not always append to the text already in the buffer—it appends only if point in that buffer is at the end. However, if append-to-buffer is the only command you use to alter a buffer, then point is always at the end.

   M-x prepend-to-buffer is just like append-to-buffer except that point

in the other buffer is left before the copied text, so successive prependings add text in reverse order. M-x copy-to-buffer is similar except that any existing text in the other buffer is deleted, so the buffer is left containing just the text newly copied into it.

To retrieve the accumulated text from another buffer, use the command M-x insert-buffer; this too takes buffername as an argument. It inserts a copy of the text in buffer buffername into the selected buffer. You can alternatively select the other buffer for editing, then optionally move text from it by killing. See Chapter 15 [Buffers], page 141, for background information on buffers.

Instead of accumulating text within Emacs, in a buffer, you can append text directly into a file with M-x append-to-file, which takes filename as an argument. It adds the text of the region to the end of the specified file. The file is changed immediately on disk.

You should use append-to-file only with files that are not being visited in Emacs. Using it on a file that you are editing in Emacs would change the file behind Emacs’s back, which can lead to losing some of your editing.

9.4 Rectangles

The rectangle commands operate on rectangular areas of the text: all the characters between a certain pair of columns, in a certain range of lines. Commands are provided to kill rectangles, yank killed rectangles, clear them out, fill them with blanks or text, or delete them. Rectangle commands are useful with text in multicolumn formats, and for changing text into or out of such formats.

When you must specify a rectangle for a command to work on, you do it by putting the mark at one corner and point at the opposite corner. The rectangle thus specified is called the region-rectangle because you control it in about the same way the region is controlled. But remember that a given combination of point and mark values can be interpreted either as a region or as a rectangle, depending on the command that uses them.

If point and the mark are in the same column, the rectangle they delimit is empty. If they are in the same line, the rectangle is one line high. This asymmetry between lines and columns comes about because point (and likewise the mark) is between two columns, but within a line.

C-x r k Kill the text of the region-rectangle, saving its contents as the “last killed rectangle” (kill-rectangle).

C-x r d Delete the text of the region-rectangle (delete-rectangle).

C-x r y Yank the last killed rectangle with its upper left corner at point

          (yank-rectangle).

C-x r o Insert blank space to fill the space of the region-rectangle (openrectangle). This pushes the previous contents of the regionrectangle rightward.

M-x clear-rectangle

Clear the region-rectangle by replacing its contents with spaces.

M-x delete-whitespace-rectangle

Delete whitespace in each of the lines on the specified rectangle, starting from the left edge column of the rectangle.

C-x r t ^{⟨}\underline{\overline{RET}}^{⟩} string ^{⟨}\underline{\overline{RET}}^{⟩} Insert string on each line of the region-rectangle (string-

          rectangle).

The rectangle operations fall into two classes: commands deleting and inserting rectangles, and commands for blank rectangles.

There are two ways to get rid of the text in a rectangle: you can discard the text (delete it) or save it as the “last killed” rectangle. The commands

for these two ways are C-x r d (delete-rectangle) and C-x r k (kill-

rectangle). In either case, the portion of each line that falls inside the rectangle’s boundaries is deleted, causing following text (if any) on the line to move left into the gap.

Note that “killing” a rectangle is not killing in the usual sense; the rectangle is not stored in the kill ring, but in a special place that can only record the most recent rectangle killed. This is because yanking a rectangle is so different from yanking linear text that different yank commands have to be used and yank-popping is hard to make sense of.

To yank the last killed rectangle, type C-x r y (yank-rectangle). Yanking a rectangle is the opposite of killing one. Point specifies where to put the rectangle’s upper left corner. The rectangle’s first line is inserted there, the rectangle’s second line is inserted at a position one line vertically down, and so on. The number of lines affected is determined by the height of the saved rectangle.

You can convert single-column lists into double-column lists using rectangle killing and yanking; kill the second half of the list as a rectangle and then yank it beside the first line of the list. See Section 30.9 [Two-Column], page 375, for another way to edit multi-column text.

You can also copy rectangles into and out of registers with C-x r r r and C-x r i r. See Section 10.3 [Rectangle Registers], page 74.

There are two commands you can use for making blank rectangles: M-x clear-rectangle which blanks out existing text, and C-x r o (openrectangle) which inserts a blank rectangle. Clearing a rectangle is equivalent to deleting it and then inserting a blank rectangle of the same size.

   The command M-x delete-whitespace-rectangle deletes horizontal

whitespace starting from a particular column. This applies to each of the lines in the rectangle, and the column is specified by the left edge of the rectangle. The right edge of the rectangle does not make any difference to this command.

The command C-x r t (M-x string-rectangle) replaces the rectangle with a specified string (inserted once on each line). The string’s width need not be the same as the width of the rectangle. if the string’s width is less, the Chapter 9: Killing and Moving Text 71 text after the rectangle shifts left; if the stringis wider than the rectangle, the text after the rectangle shifts right.

72 GNU Emacs Manual

10 Registers

Emacs registers are places you can save text or positions for later use. Once you save text or a rectangle in a register, you can copy it into the buffer once or many times; you can move point to a position saved in a register once or many times.

Each register has a name which is a single character. A register can store a piece of text, a rectangle, a position, a window configuration, or a file name, but only one thing at any given time. Whatever you store in a register remains there until you store something else in that register. To see what a register r contains, use M-x view-register.

M-x view-register ^{⟨}\underline{\overline{RET}}^{⟩} r

Display a description of what register r contains.

10.1 Saving Positions in Registers

Saving a position records a place in a buffer so that you can move back there later. Moving to a saved position switches to that buffer and moves point to that place in it.

C-x r ^{⟨}\underline{\overline{SPC}}^{⟩} r Save position of point in register r (point-to-register).

C-x r j r Jump to the position saved in register r (jump-to-register).

To save the current position of point in a register, choose a name r and type C-x r ^{⟨}\underline{\overline{SPC}}^{⟩} r. The register r retains the position thus saved until you store something else in that register.

The command C-x r j r moves point to the position recorded in register r. The register is not affected; it continues to record the same position. You can jump to the saved position any number of times.

If you use C-x r j to go to a saved position, but the buffer it was saved from has been killed, C-x r j tries to create the buffer again by visiting the same file. Of course, this works only for buffers that were visiting files.

10.2 Saving Text in Registers

When you want to insert a copy of the same piece of text several times, it may be inconvenient to yank it from the kill ring, since each subsequent kill moves that entry further down the ring. An alternative is to store the text in a register and later retrieve it.

C-x r s r Copy region into register r (copy-to-register).

C-x r i r Insert text from register r (insert-register).

C-x r s r stores a copy of the text of the region into the register named r. Given a numeric argument, C-x r s r deletes the text from the buffer as well.

C-x r i r inserts in the buffer the text from register r. Normally it leaves point before the text and places the mark after, but with a numeric argument (C-u) it puts point after the text and the mark before.

10.3 Saving Rectangles in Registers

A register can contain a rectangle instead of linear text. The rectangle is represented as a list of strings. See Section 9.4 [Rectangles], page 69, for basic information on how to specify a rectangle in the buffer.

C-x r r r Copy the region-rectangle into register r (copy-rectangle-to-register). With numeric argument, delete it as well.

C-x r i r Insert the rectangle stored in register r (if it contains a rectangle)

          (insert-register).

The C-x r i r command inserts a text string if the register contains one, and inserts a rectangle if the register contains one.

See also the command sort-columns, which you can think of as sorting a rectangle. See Section 30.7 [Sorting], page 372.

10.4 Saving Window Configurations in Registers

You can save the window configuration of the selected frame in a register, or even the configuration of all windows in all frames, and restore the configuration later. C-x r w r Save the state of the selected frame’s windows in register r

          (window-configuration-to-register).

C-x r f r Save the state of all frames, including all their windows, in register

          r (frame-configuration-to-register).

Use C-x r j r to restore a window or frame configuration. This is the same command used to restore a cursor position. When you restore a frame configuration, any existing frames not included in the configuration become invisible. If you wish to delete these frames instead, use C-u C-x r j r.

10.5 Keeping Numbers in Registers

There are commands to store a number in a register, to insert the number in the buffer in decimal, and to increment it. These commands can be useful in keyboard macros (see Section 31.3 [Keyboard Macros], page 397). C-u number C-x r n reg

Store number into register reg (number-to-register).

C-u number C-x r + reg Increment the number in register reg by number (increment-

          register).
C-x r g reg

Insert the number from register reg into the buffer.

C-x r g is the same command used to insert any other sort of register contents into the buffer.

10.6 Keeping File Names in Registers

If you visit certain file names frequently, you can visit them more conveniently if you put their names in registers. Here’s the Lisp code used to put a file name in a register:

     (set-register ?r '(file . name))

For example,

     (set-register ?z '(file . "/gd/gnu/emacs/19.0/src/ChangeLog"))

puts the file name shown in register ‘z’.

To visit the file whose name is in register r, type C-x r j r. (This is the same command used to jump to a position or restore a frame configuration.)

10.7 Bookmarks

Bookmarks are somewhat like registers in that they record positions you can jump to. Unlike registers, they have long names, and they persist automatically from one Emacs session to the next. The prototypical use of bookmarks is to record “where you were reading” in various files.

C-x r m ^{⟨}\underline{\overline{RET}}^{⟩}

Set the bookmark for the visited file, at point.

C-x r m bookmark ^{⟨}\underline{\overline{RET}}^{⟩} Set the bookmark named bookmark at point (bookmark-set).

C-x r b bookmark ^{⟨}\underline{\overline{RET}}^{⟩} Jump to the bookmark named bookmark (bookmark-jump).

C-x r l List all bookmarks (list-bookmarks).
M-x bookmark-save

Save all the current bookmark values in the default bookmark file.

The prototypical use for bookmarks is to record one current position in each of several files. So the command C-x r m, which sets a bookmark, uses the visited file name as the default for the bookmark name. If you name each bookmark after the file it points to, then you can conveniently revisit any of those files with C-x r b, and move to the position of the bookmark at the same time.

To display a list of all your bookmarks in a separate buffer, type C-x r l (list-bookmarks). If you switch to that buffer, you can use it to edit your bookmark definitions or annotate the bookmarks. Type C-h m in that buffer for more information about its special editing commands.

When you kill Emacs, Emacs offers to save your bookmark values in your default bookmark file, ‘~/.emacs.bmk’, if you have changed any bookmark values. You can also save the bookmarks at any time with the M-x bookmark-save command. The bookmark commands load your default bookmark file automatically. This saving and loading is how bookmarks persist from one Emacs session to the next.

If you set the variable bookmark-save-flag to 1, then each command that sets a bookmark will also save your bookmarks; this way, you don’t lose any bookmark values even if Emacs crashes. (The value, if a number, says how many bookmark modifications should go by between saving.)

Bookmark position values are saved with surrounding context, so that bookmark-jump can find the proper position even if the file is modified slightly. The variable bookmark-search-size says how many characters of context to record, on each side of the bookmark’s position.

Here are some additional commands for working with bookmarks:

M-x bookmark-load ^{⟨}\underline{\overline{RET}}^{⟩} filename ^{⟨}\underline{\overline{RET}}^{⟩} Load a file named filename that contains a list of bookmark values. You can use this command, as well as bookmark-write, to work with other files of bookmark values in addition to your default bookmark file.

M-x bookmark-write ^{⟨}\underline{\overline{RET}}^{⟩} filename ^{⟨}\underline{\overline{RET}}^{⟩}

Save all the current bookmark values in the file filename.

M-x bookmark-delete ^{⟨}\underline{\overline{RET}}^{⟩} bookmark ^{⟨}\underline{\overline{RET}}^{⟩}

Delete the bookmark named bookmark.

M-x bookmark-insert-location ^{⟨}\underline{\overline{RET}}^{⟩} bookmark ^{⟨}\underline{\overline{RET}}^{⟩}

Insert in the buffer the name of the file that bookmark bookmark points to.

M-x bookmark-insert ^{⟨}\underline{\overline{RET}}^{⟩} bookmark ^{⟨}\underline{\overline{RET}}^{⟩}

Insert in the buffer the contents of the file that bookmark bookmark points to.

11 Controlling the Display

Since only part of a large buffer fits in the window, Emacs tries to show a part that is likely to be interesting. Display-control commands allow you to specify which part of the text you want to see, and how to display it.

11.1 Scrolling

If a buffer contains text that is too large to fit entirely within a window that is displaying the buffer, Emacs shows a contiguous portion of the text. The portion shown always contains point.

Scrolling means moving text up or down in the window so that different parts of the text are visible. Scrolling forward means that text moves up, and new text appears at the bottom. Scrolling backward moves text down and new text appears at the top.

Scrolling happens automatically if you move point past the bottom or top of the window. You can also explicitly request scrolling with the commands in this section. C-l Clear screen and redisplay, scrolling the selected window to center

point vertically within it (recenter).

C-v Scroll forward (a windowful or a specified number of lines)

          (scroll-up).

⟨\underline{\overline{NEXT}}^{⟩} Likewise, scroll forward. M-v Scroll backward (scroll-down). ⟨\underline{\overline{PRIOR}}^{⟩} Likewise, scroll backward. arg C-l Scroll so point is on line arg (recenter). C-M-l Scroll heuristically to bring useful information onto the screen

          (reposition-window).

The most basic scrolling command is C-l (recenter) with no argument. It clears the entire screen and redisplays all windows. In addition, it scrolls the selected window so that point is halfway down from the top of the window.

The scrolling commands C-v and M-v let you move all the text in the window up or down a few lines. C-v (scroll-up) with an argument shows you that many more lines at the bottom of the window, moving the text and point up together as C-l might. C-v with a negative argument shows you more lines at the top of the window. M-v (scroll-down) is like C-v, but moves in the opposite direction. The function keys ^{⟨}\underline{\overline{NEXT}}^{⟩} and ^{⟨}\underline{\overline{PRIOR}}^{⟩} are equivalent to C-v and M-v.

The names of scroll commands are based on the direction that the text moves in the window. Thus, the command to scroll forward is called scrollup because it moves the text upward on the screen.

To read the buffer a windowful at a time, use C-v with no argument. It takes the last two lines at the bottom of the window and puts them at the top, followed by nearly a whole windowful of lines not previously visible. If point was in the text scrolled off the top, it moves to the new top of the window. M-v with no argument moves backward with overlap similarly. The number of lines of overlap across a C-v or M-v is controlled by the variable

next-screen-context-lines; by default, it is 2.

Some users like the full-screen scroll commands to keep point at the same screen position. To enable this behavior, set the variable scroll-preserve-screen-position to a non-nil value. This mode is convenient for browsing through a file by scrolling by screenfuls; if you come back to the screen where you started, point goes back to its starting value. However, this mode is inconvenient when you move to the next screen in order to move point to the text there.

Another way to do scrolling is with C-l with a numeric argument. C-l does not clear the screen when given an argument; it only scrolls the selected window. With a positive argument n, it repositions text to put point n lines down from the top. An argument of zero puts point on the very top line. Point does not move with respect to the text; rather, the text and point move rigidly on the screen. C-l with a negative argument puts point that many lines from the bottom of the window. For example, C-u - 1 C-l puts point on the bottom line, and C-u - 5 C-l puts it five lines from the bottom. Just C-u as argument, as in C-u C-l, scrolls point to the center of the selected window.

The C-M-l command (reposition-window) scrolls the current window heuristically in a way designed to get useful information onto the screen. For example, in a Lisp file, this command tries to get the entire current defun onto the screen if possible.

Scrolling happens automatically if point has moved out of the visible portion of the text when it is time to display. Normally, automatic scrolling centers point vertically within the window. However, if you set scrollconservatively to a small number n, then if you move point just a little off the screen—less than n lines—then Emacs scrolls the text just far enough to bring point back on screen. By default, scroll-conservatively is 0.

The variable scroll-margin restricts how close point can come to the top or bottom of a window. Its value is a number of screen lines; if point comes within that many lines of the top or bottom of the window, Emacs recenters the window. By default, scroll-margin is 0.

11.2 Horizontal Scrolling

Horizontal scrolling means shifting all the lines sideways within a window—so that some of the text near the left margin is not displayed at all.

C-x < Scroll text in current window to the left (scroll-left).

C-x > Scroll to the right (scroll-right).

When a window has been scrolled horizontally, text lines are truncated rather than continued (see Section 4.8 [Continuation Lines], page 35), with a ‘$’ appearing in the first column when there is text truncated to the left, and in the last column when there is text truncated to the right.

The command C-x < (scroll-left) scrolls the selected window to the left by n columns with argument n. This moves part of the beginning of each line off the left edge of the window. With no argument, it scrolls by almost the full width of the window (two columns less, to be precise).

C-x > (scroll-right) scrolls similarly to the right. The window cannot be scrolled any farther to the right once it is displayed normally (with each line starting at the window’s left margin); attempting to do so has no effect. This means that you don’t have to calculate the argument precisely for C-x >; any sufficiently large argument will restore the normal display.

You can request automatic horizontal scrolling by enabling Hscroll mode. When this mode is enabled, Emacs scrolls a window horizontally whenever that is necessary to keep point visible and not too far from the left or right edge. The command to enable or disable this mode is M-x hscroll-mode.

11.3 Follow Mode

Follow mode is a minor mode that makes two windows showing the same buffer scroll as one tall “virtual window.” To use Follow mode, go to a frame with just one window, split it into two side-by-side windows using C-x 3, and then type M-x follow-mode. From then on, you can edit the buffer in either of the two windows, or scroll either one; the other window follows it.

To turn off Follow mode, type M-x follow-mode a second time.

11.4 Selective Display

Emacs has the ability to hide lines indented more than a certain number of columns (you specify how many columns). You can use this to get an overview of a part of a program.

To hide lines, type C-x $ (set-selective-display) with a numeric argument n. Then lines with at least n columns of indentation disappear from the screen. The only indication of their presence is that three dots (‘...’) appear at the end of each visible line that is followed by one or more hidden ones.

The commands C-n and C-p move across the hidden lines as if they were not there.

The hidden lines are still present in the buffer, and most editing commands see them as usual, so you may find point in the middle of the hidden text. When this happens, the cursor appears at the end of the previous line, after the three dots. If point is at the end of the visible line, before the newline that ends it, the cursor appears before the three dots.

To make all lines visible again, type C-x $ with no argument.

If you set the variable selective-display-ellipses to nil, the three dots do not appear at the end of a line that precedes hidden lines. Then there is no visible indication of the hidden lines. This variable becomes local automatically when set.

11.5 Optional Mode Line Features

The current line number of point appears in the mode line when Line Number mode is enabled. Use the command M-x line-number-mode to turn this mode on and off; normally it is on. The line number appears before the buffer percentage pos, with the letter ‘L’ to indicate what it is. See Section 31.1 [Minor Modes], page 383, for more information about minor modes and about how to use this command.

If the buffer is very large (larger than the value of line-number-display-limit), then the line number doesn’t appear. Emacs doesn’t compute the line number when the buffer is large, because that would be too slow. If you have narrowed the buffer (see Section 30.8 [Narrowing], page 374), the displayed line number is relative to the accessible portion of the buffer.

You can also display the current column number by turning on Column Number mode. It displays the current column number preceded by the letter ‘C’. Type M-x column-number-mode to toggle this mode.

Emacs can optionally display the time and system load in all mode lines. To enable this feature, type M-x display-time. The information added to the mode line usually appears after the buffer name, before the mode names and their parentheses. It looks like this:

hh:mmpm l.ll Here hh and mm are the hour and minute, followed always by ‘am’ or ‘pm’. l.ll is the average number of running processes in the whole system recently. (Some fields may be missing if your operating system cannot support them.) If you prefer time display in 24-hour format, set the variable display-time-

24hr-format to t.

The word ‘Mail’ appears after the load level if there is mail for you that you have not read yet.

11.6 How Text Is Displayed

ASCII printing characters (octal codes 040 through 0176) in Emacs buffers are displayed with their graphics. So are non-ASCII multibyte printing characters (octal codes above 0400).

Some ASCII control characters are displayed in special ways. The newline character (octal code 012) is displayed by starting a new line. The tab character (octal code 011) is displayed by moving to the next tab stop column (normally every 8 columns).

Other ASCII control characters are normally displayed as a caret (‘^’) followed by the non-control version of the character; thus, control-A is displayed as ‘^A’.

Non-ASCII characters 0200 through 0377 are displayed with octal escape sequences; thus, character code 0243 (octal) is displayed as ‘\243’. However, if you enable European display, most of these characters become non-ASCII printing characters, and are displayed using their graphics (assuming your terminal supports them). See Section 18.11 [Single-Byte European Support], page 183.

11.7 Variables Controlling Display

This section contains information for customization only. Beginning users should skip it.

The variable mode-line-inverse-video controls whether the mode line is displayed in inverse video (assuming the terminal supports it); nil means don’t do so. See Section 1.3 [Mode Line], page 17. If you specify the fore-

ground color for the modeline face, and mode-line-inverse-video is non-

nil, then the default background color for that face is the usual foreground color. See Section 17.13 [Faces], page 164.

If the variable inverse-video is non-nil, Emacs attempts to invert all the lines of the display from what they normally are.

If the variable visible-bell is non-nil, Emacs attempts to make the whole screen blink when it would normally make an audible bell sound. This variable has no effect if your terminal does not have a way to make the screen blink.

When you reenter Emacs after suspending, Emacs normally clears the screen and redraws the entire display. On some terminals with more than one page of memory, it is possible to arrange the termcap entry so that the ‘ti’ and ‘te’ strings (output to the terminal when Emacs is entered and exited, respectively) switch between pages of memory so as to use one page for Emacs and another page for other output. Then you might want to set the variable no-redraw-on-reenter non-nil; this tells Emacs to assume, when resumed, that the screen page it is using still contains what Emacs last wrote there.

The variable echo-keystrokes controls the echoing of multi-character keys; its value is the number of seconds of pause required to cause echoing to start, or zero meaning don’t echo at all. See Section 1.2 [Echo Area], page 16.

If the variable ctl-arrow is nil, control characters in the buffer are displayed with octal escape sequences, except for newline and tab. Altering the value of ctl-arrow makes it local to the current buffer; until that time, the default value is in effect. The default is initially t. See section “Display Tables” in The Emacs Lisp Reference Manual.

Normally, a tab character in the buffer is displayed as whitespace which extends to the next display tab stop position, and display tab stops come at intervals equal to eight spaces. The number of spaces per tab is controlled by the variable tab-width, which is made local by changing it, just like ctlarrow. Note that how the tab character in the buffer is displayed has nothing to do with the definition of ^{⟨}\underline{\overline{TAB}}^{⟩} as a command. The variable tab-width must have an integer value between 1 and 1000, inclusive.

If the variable truncate-lines is non-nil, then each line of text gets just one screen line for display; if the text line is too long, display shows only the part that fits. If truncate-lines is nil, then long text lines display as more than one screen line, enough to show the whole text of the line. See Section 4.8 [Continuation Lines], page 35. Altering the value of truncatelines makes it local to the current buffer; until that time, the default value is in effect. The default is initially nil.

   If the variable truncate-partial-width-windows is non-nil, it forces

truncation rather than continuation in any window less than the full width of the screen or frame, regardless of the value of truncate-lines. For information about side-by-side windows, see Section 16.2 [Split Window], page 150. See also section “Display” in The Emacs Lisp Reference Manual.

The variable baud-rate holds the output speed of the terminal, as far as Emacs knows. Setting this variable does not change the speed of actual data transmission, but the value is used for calculations such as padding. It also affects decisions about whether to scroll part of the screen or redraw it instead—even when using a window system. (We designed it this way, despite the fact that a window system has no true “output speed,” to give you a way to tune these decisions.)

You can customize the way any particular character code is displayed by means of a display table. See section “Display Tables” in The Emacs Lisp Reference Manual.

12 Searching and Replacement

Like other editors, Emacs has commands for searching for occurrences of a string. The principal search command is unusual in that it is incremental; it begins to search before you have finished typing the search string. There are also nonincremental search commands more like those of other editors.

Besides the usual replace-string command that finds all occurrences of one string and replaces them with another, Emacs has a fancy replacement command called query-replace which asks interactively which occurrences to replace.

12.1 Incremental Search

An incremental search begins searching as soon as you type the first character of the search string. As you type in the search string, Emacs shows you where the string (as you have typed it so far) would be found. When you have typed enough characters to identify the place you want, you can stop. Depending on what you plan to do next, you may or may not need to terminate the search explicitly with ^{⟨}\underline{\overline{RET}}^{⟩}. C-s Incremental search forward (isearch-forward). C-r Incremental search backward (isearch-backward).

C-s starts an incremental search. C-s reads characters from the keyboard and positions the cursor at the first occurrence of the characters that you have typed. If you type C-s and then F, the cursor moves right after the first ‘F’. Type an O, and see the cursor move to after the first ‘FO’. After another O, the cursor is after the first ‘FOO’ after the place where you started the search. At each step, the buffer text that matches the search string is highlighted, if the terminal can do that; at each step, the current search string is updated in the echo area.

If you make a mistake in typing the search string, you can cancel characters with ^{⟨}\underline{\overline{DEL}}^{⟩}. Each ^{⟨}\underline{\overline{DEL}}^{⟩} cancels the last character of search string. This does not happen until Emacs is ready to read another input character; first it must either find, or fail to find, the character you want to erase. If you do not want to wait for this to happen, use C-g as described below.

When you are satisfied with the place you have reached, you can type ⟨\underline{\overline{RET}}^{⟩}, which stops searching, leaving the cursor where the search brought it. Also, any command not specially meaningful in searches stops the searching and is then executed. Thus, typing C-a would exit the search and then move to the beginning of the line. ^{⟨}\underline{\overline{RET}}^{⟩} is necessary only if the next command you want to type is a printing character, ^{⟨}\underline{\overline{DEL}}^{⟩}, ^{⟨}\underline{\overline{RET}}^{⟩}, or another control character that is special within searches (C-q, C-w, C-r, C-s, C-y, M-y, M-r, or M-s).

Sometimes you search for ‘FOO’ and find it, but not the one you expected to find. There was a second ‘FOO’ that you forgot about, before the one you were aiming for. In this event, type another C-s to move to the next occurrence of the search string. This can be done any number of times. If you overshoot, you can cancel some C-s characters with ^{⟨}\underline{\overline{DEL}}^{⟩}.

After you exit a search, you can search for the same string again by typing just C-s C-s: the first C-s is the key that invokes incremental search, and the second C-s means “search again.”

To reuse earlier search strings, use the search ring. The commands M-p and M-n move through the ring to pick a search string to reuse. These commands leave the selected search ring element in the minibuffer, where you can edit it. Type C-s or C-r to terminate editing the string and search for it.

If your string is not found at all, the echo area says ‘Failing I-Search’. The cursor is after the place where Emacs found as much of your string as it could. Thus, if you search for ‘FOOT’, and there is no ‘FOOT’, you might see the cursor after the ‘FOO’ in ‘FOOL’. At this point there are several things you can do. If your string was mistyped, you can rub some of it out and correct it. If you like the place you have found, you can type ^{⟨}\underline{\overline{RET}}^{⟩} or some other Emacs command to “accept what the search offered.” Or you can type C-g, which removes from the search string the characters that could not be found (the ‘T’ in ‘FOOT’), leaving those that were found (the ‘FOO’ in ‘FOOT’). A second C-g at that point cancels the search entirely, returning point to where it was when the search started.

An upper-case letter in the search string makes the search case-sensitive. If you delete the upper-case character from the search string, it ceases to have this effect. See Section 12.6 [Search Case], page 92.

If a search is failing and you ask to repeat it by typing another C-s, it starts again from the beginning of the buffer. Repeating a failing reverse search with C-r starts again from the end. This is called wrapping around. ‘Wrapped’ appears in the search prompt once this has happened. If you keep on going past the original starting point of the search, it changes to ‘Overwrapped’, which means that you are revisiting matches that you have already seen.

The C-g “quit” character does special things during searches; just what it does depends on the status of the search. If the search has found what you specified and is waiting for input, C-g cancels the entire search. The cursor moves back to where you started the search. If C-g is typed when there are characters in the search string that have not been found—because Emacs is still searching for them, or because it has failed to find them—then the search string characters which have not been found are discarded from the search string. With them gone, the search is now successful and waiting for more input, so a second C-g will cancel the entire search.

To search for a newline, type C-j. To search for another control character, such as control-S or carriage return, you must quote it by typing C-q first. This function of C-q is analogous to its use for insertion (see Section 4.1 [Inserting Text], page 29): it causes the following character to be treated the way any “ordinary” character is treated in the same context. You can also specify a character by its octal code: enter C-q followed by a sequence of octal digits.

You can change to searching backwards with C-r. If a search fails because the place you started was too late in the file, you should do this. Repeated C-r keeps looking for more occurrences backwards. A C-s starts going forwards again. C-r in a search can be canceled with ^{⟨}\underline{\overline{DEL}}^{⟩}.

If you know initially that you want to search backwards, you can use C-r instead of C-s to start the search, because C-r as a key runs a command (isearch-backward) to search backward. A backward search finds matches that are entirely before the starting point, just as a forward search finds matches that begin after it.

The characters C-y and C-w can be used in incremental search to grab text from the buffer into the search string. This makes it convenient to search for another occurrence of text at point. C-w copies the word after point as part of the search string, advancing point over that word. Another C-s to repeat the search will then search for a string including that word. C-y is similar to C-w but copies all the rest of the current line into the search string. Both C-y and C-w convert the text they copy to lower case if the search is currently not case-sensitive; this is so the search remains case-insensitive.

The character M-y copies text from the kill ring into the search string. It uses the same text that C-y as a command would yank. See Section 9.2 [Yanking], page 65.

When you exit the incremental search, it sets the mark to where point was, before the search. That is convenient for moving back there. In Transient Mark mode, incremental search sets the mark without activating it, and does so only if the mark is not already active.

To customize the special characters that incremental search understands, alter their bindings in the keymap isearch-mode-map. For a list of bindings, look at the documentation of isearch-mode with C-h f isearch-mode ⟨\underline{\overline{RET}}^{⟩}.

12.1.1 Slow Terminal Incremental Search

Incremental search on a slow terminal uses a modified style of display that is designed to take less time. Instead of redisplaying the buffer at each place the search gets to, it creates a new single-line window and uses that to display the line that the search has found. The single-line window comes into play as soon as point gets outside of the text that is already on the screen.

When you terminate the search, the single-line window is removed. Then Emacs redisplays the window in which the search was done, to show its new position of point.

The slow terminal style of display is used when the terminal baud rate is less than or equal to the value of the variable search-slow-speed, initially 1200.

The number of lines to use in slow terminal search display is controlled by the variable search-slow-window-lines. Its normal value is 1.

12.2 Nonincremental Search

Emacs also has conventional nonincremental search commands, which require you to type the entire search string before searching begins.

C-s ^{⟨}\underline{\overline{RET}}^{⟩} string ^{⟨}\underline{\overline{RET}}^{⟩} Search for string.

C-r ^{⟨}\underline{\overline{RET}}^{⟩} string ^{⟨}\underline{\overline{RET}}^{⟩} Search backward for string.

To do a nonincremental search, first type C-s ^{⟨}\underline{\overline{RET}}^{⟩}. This enters the minibuffer to read the search string; terminate the string with ^{⟨}\underline{\overline{RET}}^{⟩}, and then the search takes place. If the string is not found, the search command gets an error.

The way C-s ^{⟨}\underline{\overline{RET}}^{⟩} works is that the C-s invokes incremental search, which is specially programmed to invoke nonincremental search if the argument you give it is empty. (Such an empty argument would otherwise be useless.) C-r ^{⟨}\underline{\overline{RET}}^{⟩} also works this way.

However, nonincremental searches performed using C-s ^{⟨}\underline{\overline{RET}}^{⟩} do not call search-forward right away. The first thing done is to see if the next character is C-w, which requests a word search.

Forward and backward nonincremental searches are implemented by the

commands search-forward and search-backward. These commands may

be bound to keys in the usual manner. The feature that you can get to them via the incremental search commands exists for historical reasons, and to avoid the need to find suitable key sequences for them.

12.3 Word Search

Word search searches for a sequence of words without regard to how the words are separated. More precisely, you type a string of many words, using single spaces to separate them, and the string can be found even if there are multiple spaces, newlines or other punctuation between the words.

Word search is useful for editing a printed document made with a text formatter. If you edit while looking at the printed, formatted version, you can’t tell where the line breaks are in the source file. With word search, you can search without having to know them.

C-s ^{⟨}\underline{\overline{RET}}^{⟩} C-w words ^{⟨}\underline{\overline{RET}}^{⟩} Search for words, ignoring details of punctuation.

C-r ^{⟨}\underline{\overline{RET}}^{⟩} C-w words ^{⟨}\underline{\overline{RET}}^{⟩} Search backward for words, ignoring details of punctuation.

Word search is a special case of nonincremental search and is invoked with C-s ^{⟨}\underline{\overline{RET}}^{⟩} C-w. This is followed by the search string, which must always be terminated with^{⟨}\underline{\overline{RET}}^{⟩}. Being nonincremental, this search does not start until the argument is terminated. It works by constructing a regular expression and searching for that; see Section 12.4 [Regexp Search], page 87.

Use C-r ^{⟨}\underline{\overline{RET}}^{⟩} C-w to do backward word search.

Forward and backward word searches are implemented by the commands

word-search-forward and word-search-backward. These commands may

be bound to keys in the usual manner. The feature that you can get to them via the incremental search commands exists for historical reasons, and to avoid the need to find suitable key sequences for them.

12.4 Regular Expression Search

A regular expression (regexp, for short) is a pattern that denotes a class of alternative strings to match, possibly infinitely many. In GNU Emacs, you can search for the next match for a regexp either incrementally or not.

Incremental search for a regexp is done by typing C-M-s (isearchforward-regexp). This command reads a search string incrementally just like C-s, but it treats the search string as a regexp rather than looking for an exact match against the text in the buffer. Each time you add text to the search string, you make the regexp longer, and the new regexp is searched for. Invoking C-s with a prefix argument (its value does not matter) is another way to do a forward incremental regexp search. To search backward

for a regexp, use C-M-r (isearch-backward-regexp), or C-r with a prefix

argument.

All of the control characters that do special things within an ordinary incremental search have the same function in incremental regexp search. Typing C-s or C-r immediately after starting the search retrieves the last incremental search regexp used; that is to say, incremental regexp and nonregexp searches have independent defaults. They also have separate search rings that you can access with M-p and M-n.

If you type ^{⟨}\underline{\overline{SPC}}^{⟩} in incremental regexp search, it matches any sequence of whitespace characters, including newlines. If you want to match just a space, type C-q ^{⟨}\underline{\overline{SPC}}^{⟩}.

Note that adding characters to the regexp in an incremental regexp search can make the cursor move back and start again. For example, if you have searched for ‘foo’ and you add ‘\|bar’, the cursor backs up in case the first ‘bar’ precedes the first ‘foo’.

Nonincremental search for a regexp is done by the functions re-search-forward and re-search-backward. You can invoke these with M-x, or bind them to keys, or invoke them by way of incremental regexp search with C-M-s ^{⟨}\underline{\overline{RET}}^{⟩} and C-M-r ^{⟨}\underline{\overline{RET}}^{⟩}.

If you use the incremental regexp search commands with a prefix argument, they perform ordinary string search, like isearch-forward and isearch-backward. See Section 12.1 [Incremental Search], page 83.

12.5 Syntax of Regular Expressions

Regular expressions have a syntax in which a few characters are special constructs and the rest are ordinary. An ordinary character is a simple regular expression which matches that same character and nothing else. The special characters are ‘$’, ‘^’, ‘.’, ‘*’, ‘+’, ‘?’, ‘[’, ‘]’ and ‘\’. Any other character appearing in a regular expression is ordinary, unless a ‘\’ precedes it.

For example, ‘f’ is not a special character, so it is ordinary, and therefore ‘f’ is a regular expression that matches the string ‘f’ and no other string. (It does not match the string ‘ff’.) Likewise, ‘o’ is a regular expression that matches only ‘o’. (When case distinctions are being ignored, these regexps also match ‘F’ and ‘O’, but we consider this a generalization of “the same string,” rather than an exception.)

Any two regular expressions a and b can be concatenated. The result is a regular expression which matches a string if a matches some amount of the beginning of that string and b matches the rest of the string.

As a simple example, we can concatenate the regular expressions ‘f’ and ‘o’ to get the regular expression ‘fo’, which matches only the string ‘fo’. Still trivial. To do something nontrivial, you need to use one of the special characters. Here is a list of them. . (Period) is a special character that matches any single character except a

newline. Using concatenation, we can make regular expressions like ‘a.b’, which matches any three-character string that begins with ‘a’ and ends with ‘b’.

* is not a construct by itself; it is a postfix operator that means to match the preceding regular expression repetitively as many times as possible. Thus, ‘o*’ matches any number of ‘o’s (including no ‘o’s). ‘*’ always applies to the smallest possible preceding expression. Thus, ‘fo*’ has a repeating ‘o’, not a repeating ‘fo’. It matches ‘f’, ‘fo’, ‘foo’, and so on. The matcher processes a ‘*’ construct by matching, immediately, as many repetitions as can be found. Then it continues with the rest of the pattern. If that fails, backtracking occurs, discarding some of the matches of the ‘*’-modified construct in case that makes it possible to match the rest of the pattern. For example, in matching ‘ca*ar’ against the string ‘caaar’, the ‘a*’ first tries to match all three ‘a’s; but the rest of the pattern is ‘ar’ and there is only ‘r’ left to match, so this try fails. The next alternative is for ‘a*’ to match only two ‘a’s. With this choice, the rest of the regexp matches successfully.

+ is a postfix operator, similar to ‘*’ except that it must match the preceding expression at least once. So, for example, ‘ca+r’ matches the strings ‘car’ and ‘caaaar’ but not the string ‘cr’, whereas ‘ca*r’ matches all three strings.

? is a postfix operator, similar to ‘*’ except that it can match the preceding expression either once or not at all. For example, ‘ca?r’ matches ‘car’ or ‘cr’; nothing else.

[ ... ] is a character set, which begins with ‘[’ and is terminated by ‘]’. In the simplest case, the characters between the two brackets are what this set can match. Thus, ‘[ad]’ matches either one ‘a’ or one ‘d’, and ‘[ad]*’ matches any string composed of just ‘a’s and ‘d’s (including the empty string), from which it follows that ‘c[ad]*r’ matches ‘cr’, ‘car’, ‘cdr’, ‘caddaar’, etc. You can also include character ranges in a character set, by writing the starting and ending characters with a ‘-’ between them. Thus, ‘[a-z]’ matches any lower-case ASCII letter. Ranges may be intermixed freely with individual characters, as in ‘[a-z$%.]’, which matches any lower-case ASCII letter or ‘$’, ‘%’ or period. Note that the usual regexp special characters are not special inside a character set. A completely different set of special characters exists inside character sets: ‘]’, ‘-’ and ‘^’. To include a ‘]’ in a character set, you must make it the first character. For example, ‘[]a]’ matches ‘]’ or ‘a’. To include a ‘-’, write ‘-’ as the first or last character of the set, or put it after a range. Thus, ‘[]-]’ matches both ‘]’ and ‘-’. To include ‘^’ in a set, put it anywhere but at the beginning of the set. When you use a range in case-insensitive search, you should write both ends of the range in upper case, or both in lower case, or both should be non-letters. The behavior of a mixedcase range such as ‘A-z’ is somewhat ill-defined, and it may change in future Emacs versions.

[^... ] ‘[^’ begins a complemented character set, which matches any character except the ones specified. Thus, ‘[^a-z0-9A-Z]’ matches all characters except letters and digits. ‘^’ is not special in a character set unless it is the first character. The character following the ‘^’ is treated as if it were first (in other words, ‘-’ and ‘]’ are not special there).

A complemented character set can match a newline, unless newline is mentioned as one of the characters not to match. This is in contrast to the handling of regexps in programs such as grep.

^is a special character that matches the empty string, but only at the beginning of a line in the text being matched. Otherwise it fails to match anything. Thus, ‘^foo’ matches a ‘foo’ that occurs at the beginning of a line.

$ is similar to ‘^’ but matches only at the end of a line. Thus, ‘x+$’ matches a string of one ‘x’ or more at the end of a line.

\ has two functions: it quotes the special characters (including ‘\’), and it introduces additional special constructs. Because ‘\’ quotes special characters, ‘\$’ is a regular expression that matches only ‘$’, and ‘\[’ is a regular expression that matches only ‘[’, and so on.

Note: for historical compatibility, special characters are treated as ordinary ones if they are in contexts where their special meanings make no sense. For example, ‘*foo’ treats ‘*’ as ordinary since there is no preceding expression on which the ‘*’ can act. It is poor practice to depend on this behavior; it is better to quote the special character anyway, regardless of where it appears.

For the most part, ‘\’ followed by any character matches only that character. However, there are several exceptions: two-character sequences starting with ‘\’ that have special meanings. The second character in the sequence is always an ordinary character when used on its own. Here is a table of ‘\’ constructs.

\| specifies an alternative. Two regular expressions a and b with ‘\|’ in between form an expression that matches some text if either a matches it or b matches it. It works by trying to match a, and if that fails, by trying to match b. Thus, ‘foo\|bar’ matches either ‘foo’ or ‘bar’ but no other string. ‘\|’ applies to the largest possible surrounding expressions. Only a surrounding ‘\( ... \)’ grouping can limit the grouping power of ‘\|’. Full backtracking capability exists to handle multiple uses of ‘\|’.

\( ... \) is a grouping construct that serves three purposes:

1. To enclose a set of ‘\|’ alternatives for other operations. Thus, ‘\(foo\|bar\)x’ matches either ‘foox’ or ‘barx’.

2. To enclose a complicated expression for the postfix operators ‘*’, ‘+’ and ‘?’ to operate on. Thus, ‘ba\(na\)*’ matches ‘bananana’, etc., with any (zero or more) number of ‘na’ strings.

3. To record a matched substring for future reference. This last application is not a consequence of the idea of a parenthetical grouping; it is a separate feature that is assigned as a second meaning to the same ‘\( ... \)’ construct. In practice there is no conflict between the two meanings.

\d matches the same text that matched the dth occurrence of a ‘\( ... \)’ construct. After the end of a ‘\( ... \)’ construct, the matcher remembers the beginning and end of the text matched by that construct. Then, later on in the regular expression, you can use ‘\’ followed by the digit d to mean “match the same text matched the dth time by the ‘\( ... \)’ construct.” The strings matching the first nine ‘\( ... \)’ constructs appearing in a regular expression are assigned numbers 1 through 9 in the order that the open-parentheses appear in the regular expression. So you can use ‘\1’ through ‘\9’ to refer to the text matched by the corresponding ‘\( ... \)’ constructs. For example, ‘\(.*\)\1’ matches any newline-free string that is composed of two identical halves. The ‘\(.*\)’ matches the first half, which may be anything, but the ‘\1’ that follows must match the same exact text. If a particular ‘\( ... \)’ construct matches more than once (which can easily happen if it is followed by ‘*’), only the last match is recorded.

\` matches the empty string, but only at the beginning of the buffer or string being matched against.

\' matches the empty string, but only at the end of the buffer or string being matched against.

\= matches the empty string, but only at point. \b matches the empty string, but only at the beginning or end of

a word. Thus, ‘\bfoo\b’ matches any occurrence of ‘foo’ as a separate word. ‘\bballs?\b’ matches ‘ball’ or ‘balls’ as a separate word. ‘\b’ matches at the beginning or end of the buffer regardless of what text appears next to it.

\B matches the empty string, but not at the beginning or end of a word.

\< matches the empty string, but only at the beginning of a word. ‘\<’ matches at the beginning of the buffer only if a wordconstituent character follows.

\> matches the empty string, but only at the end of a word. ‘\>’ matches at the end of the buffer only if the contents end with a word-constituent character.

\w matches any word-constituent character. The syntax table determines which characters these are. See Section 31.6 [Syntax], page 411.

\W matches any character that is not a word-constituent.

\sc matches any character whose syntax is c. Here c is a character that represents a syntax code: thus, ‘w’ for word constituent, ‘-’ for whitespace, ‘(’ for open parenthesis, etc. Represent a character of whitespace (which can be a newline) by either ‘-’ or a space character.

\Sc matches any character whose syntax is not c.

The constructs that pertain to words and syntax are controlled by the setting of the syntax table (see Section 31.6 [Syntax], page 411).

Here is a complicated regexp, used by Emacs to recognize the end of a sentence together with any whitespace that follows. It is given in Lisp syntax to enable you to distinguish the spaces from the tab characters. In Lisp syntax, the string constant begins and ends with a double-quote. ‘\"’ stands for a double-quote as part of the regexp, ‘\\’ for a backslash as part of the regexp, ‘\t’ for a tab and ‘\n’ for a newline.

    "[.?!][]\"')]*\\($\\|\t\\|  \\)[ \t\n]*"

This contains four parts in succession: a character set matching period, ‘?’, or ‘!’; a character set matching close-brackets, quotes, or parentheses, repeated any number of times; an alternative in backslash-parentheses that matches end-of-line, a tab, or two spaces; and a character set matching whitespace characters, repeated any number of times.

To enter the same regexp interactively, you would type ^{⟨}\underline{\overline{TAB}}^{⟩} to enter a tab, and C-j to enter a newline. You would also type single backslashes as themselves, instead of doubling them for Lisp syntax.

12.6 Searching and Case

Incremental searches in Emacs normally ignore the case of the text they are searching through, if you specify the text in lower case. Thus, if you specify searching for ‘foo’, then ‘Foo’ and ‘foo’ are also considered a match. Regexps, and in particular character sets, are included: ‘[ab]’ would match ‘a’ or ‘A’ or ‘b’ or ‘B’.

An upper-case letter anywhere in the incremental search string makes the search case-sensitive. Thus, searching for ‘Foo’ does not find ‘foo’ or ‘FOO’. This applies to regular expression search as well as to string search. The effect ceases if you delete the upper-case letter from the search string.

If you set the variable case-fold-search to nil, then all letters must match exactly, including case. This is a per-buffer variable; altering the variable affects only the current buffer, but there is a default value which you can change as well. See Section 31.2.4 [Locals], page 393. This variable applies to nonincremental searches also, including those performed by the replace commands (see Section 12.7 [Replace], page 93) and the minibuffer history matching commands (see Section 5.4 [Minibuffer History], page 44).

12.7 Replacement Commands

Global search-and-replace operations are not needed as often in Emacs as they are in other editors^{1}, but they are available. In addition to the simple M-x replace-string command which is like that found in most editors, there is a M-x query-replace command which asks you, for each occurrence of the pattern, whether to replace it.

The replace commands normally operate on the text from point to the end of the buffer; however, in Transient Mark mode, when the mark is active, they operate on the region. The replace commands all replace one string (or regexp) with one replacement string. It is possible to perform several replacements in parallel using the command expand-region-abbrevs (see Section 24.3 [Expanding Abbrevs], page 282).

12.7.1 Unconditional Replacement

M-x replace-string ^{⟨}\underline{\overline{RET}}^{⟩} string ^{⟨}\underline{\overline{RET}}^{⟩} newstring ^{⟨}\underline{\overline{RET}}^{⟩} Replace every occurrence of string with newstring.

M-x replace-regexp ^{⟨}\underline{\overline{RET}}^{⟩} regexp ^{⟨}\underline{\overline{RET}}^{⟩} newstring ^{⟨}\underline{\overline{RET}}^{⟩} Replace every match for regexp with newstring.

To replace every instance of ‘foo’ after point with ‘bar’, use the command M-x replace-string with the two arguments ‘foo’ and ‘bar’. Replacement happens only in the text after point, so if you want to cover the whole buffer you must go to the beginning first. All occurrences up to the end of the buffer are replaced; to limit replacement to part of the buffer, narrow to that part of the buffer before doing the replacement (see Section 30.8 [Narrowing], page 374). In Transient Mark mode, when the region is active, replacement is limited to the region (see Section 8.2 [Transient Mark], page 58).

When replace-string exits, it leaves point at the last occurrence replaced. It sets the mark to the prior position of point (where the replacestring command was issued); use C-u C-^{⟨}\underline{\overline{SPC}}^{⟩} to move back there.

A numeric argument restricts replacement to matches that are surrounded by word boundaries. The argument’s value doesn’t matter.

1 In some editors, search-and-replace operations are the only convenient way to make a single change in the text.

12.7.2 Regexp Replacement

The M-x replace-string command replaces exact matches for a single string. The similar command M-x replace-regexp replaces any match for a specified pattern.

In replace-regexp, the newstring need not be constant: it can refer to all or part of what is matched by the regexp. ‘\&’ in newstring stands for the entire match being replaced. ‘\d’ in newstring, where d is a digit, stands for whatever matched the dth parenthesized grouping in regexp. To include a ‘\’ in the text to replace with, you must enter ‘\\’. For example,

    M-x replace-regexp ^{⟨}\underline{\overline{RET}}^{⟩} c[ad]+r ^{⟨}\underline{\overline{RET}}^{⟩} \&-safe ^{⟨}\underline{\overline{RET}}^{⟩}

replaces (for example) ‘cadr’ with ‘cadr-safe’ and ‘cddr’ with ‘cddr-safe’.

    M-x replace-regexp ^{⟨}\underline{\overline{RET}}^{⟩} \(c[ad]+r\)-safe ^{⟨}\underline{\overline{RET}}^{⟩} \1 ^{⟨}\underline{\overline{RET}}^{⟩}

performs the inverse transformation.

12.7.3 Replace Commands and Case

If the arguments to a replace command are in lower case, it preserves case when it makes a replacement. Thus, the command

    M-x replace-string ^{⟨}\underline{\overline{RET}}^{⟩} foo ^{⟨}\underline{\overline{RET}}^{⟩} bar ^{⟨}\underline{\overline{RET}}^{⟩}

replaces a lower case ‘foo’ with a lower case ‘bar’, an all-caps ‘FOO’ with ‘BAR’, and a capitalized ‘Foo’ with ‘Bar’. (These three alternatives—lower case, all caps, and capitalized, are the only ones that replace-string can distinguish.)

If upper-case letters are used in the second argument, they remain upper case every time that argument is inserted. If upper-case letters are used in the first argument, the second argument is always substituted exactly as given, with no case conversion. Likewise, if the variable case-replace is set to nil, replacement is done without case conversion. If case-fold-search is set to nil, case is significant in matching occurrences of ‘foo’ to replace; this also inhibits case conversion of the replacement string.

12.7.4 Query Replace

M-% string ^{⟨}\underline{\overline{RET}}^{⟩} newstring ^{⟨}\underline{\overline{RET}}^{⟩} M-x query-replace ^{⟨}\underline{\overline{RET}}^{⟩} string ^{⟨}\underline{\overline{RET}}^{⟩} newstring ^{⟨}\underline{\overline{RET}}^{⟩}

Replace some occurrences of string with newstring.

C-M-% regexp ^{⟨}\underline{\overline{RET}}^{⟩} newstring ^{⟨}\underline{\overline{RET}}^{⟩} M-x query-replace-regexp ^{⟨}\underline{\overline{RET}}^{⟩} regexp ^{⟨}\underline{\overline{RET}}^{⟩} newstring ^{⟨}\underline{\overline{RET}}^{⟩}

Replace some matches for regexp with newstring.

If you want to change only some of the occurrences of ‘foo’ to ‘bar’, not all of them, then you cannot use an ordinary replace-string. Instead, use M-% (query-replace). This command finds occurrences of ‘foo’ one by one, displays each occurrence and asks you whether to replace it. A numeric argument to query-replace tells it to consider only occurrences that are bounded by word-delimiter characters. This preserves case, just like replace-string, provided case-replace is non-nil, as it normally is.

Aside from querying, query-replace works just like replace-string,

and query-replace-regexp works just like replace-regexp. This com-

mand is run by C-M-%.

The things you can type when you are shown an occurrence of string or a match for regexp are:

⟨\underline{\overline{SPC}}^{⟩} to replace the occurrence with newstring.

⟨\underline{\overline{DEL}}^{⟩} to skip to the next occurrence without replacing this one.

, (Comma)

to replace this occurrence and display the result. You are then asked for another input character to say what to do next. Since the replacement has already been made, ^{⟨}\underline{\overline{DEL}}^{⟩} and ^{⟨}\underline{\overline{SPC}}^{⟩} are equivalent in this situation; both move to the next occurrence. You can type C-r at this point (see below) to alter the replaced text. You can also type C-x u to undo the replacement; this exits the query-replace, so if you want to do further replacement you must use C-x ^{⟨}\underline{\overline{ESC}}^{⟩} ^{⟨}\underline{\overline{ESC}}^{⟩} ^{⟨}\underline{\overline{RET}}^{⟩} to restart (see Section 5.5 [Repetition], page 46).

⟨\underline{\overline{RET}}^{⟩} to exit without doing any more replacements.

. (Period) to replace this occurrence and then exit without searching for more occurrences.

! to replace all remaining occurrences without asking again.

^to go back to the position of the previous occurrence (or what used to be an occurrence), in case you changed it by mistake. This works by popping the mark ring. Only one ^in a row is meaningful, because only one previous replacement position is

kept during query-replace.

C-r to enter a recursive editing level, in case the occurrence needs to be edited rather than just replaced with newstring. When you are done, exit the recursive editing level with C-M-c to proceed to the next occurrence. See Section 30.12 [Recursive Edit], page 377.

C-w to delete the occurrence, and then enter a recursive editing level as in C-r. Use the recursive edit to insert text to replace the deleted occurrence of string. When done, exit the recursive editing level with C-M-c to proceed to the next occurrence.

C-l to redisplay the screen. Then you must type another character to specify what to do with this occurrence.

C-h to display a message summarizing these options. Then you must type another character to specify what to do with this occurrence.

Some other characters are aliases for the ones listed above: y, n and q are equivalent to ^{⟨}\underline{\overline{SPC}}^{⟩}, ^{⟨}\underline{\overline{DEL}}^{⟩} and ^{⟨}\underline{\overline{RET}}^{⟩}.

Aside from this, any other character exits the query-replace, and is then reread as part of a key sequence. Thus, if you type C-k, it exits the query-replace and then kills to end of line.

To restart a query-replace once it is exited, use C-x ^{⟨}\underline{\overline{ESC}}^{⟩} ^{⟨}\underline{\overline{ESC}}^{⟩}, which repeats the query-replace because it used the minibuffer to read its arguments. See Section 5.5 [Repetition], page 46.

See also Section 28.9 [Transforming File Names], page 328, for Dired commands to rename, copy, or link files by replacing regexp matches in file names.

12.8 Other Search-and-Loop Commands

Here are some other commands that find matches for a regular expression. They all operate from point to the end of the buffer. M-x occur ^{⟨}\underline{\overline{RET}}^{⟩} regexp ^{⟨}\underline{\overline{RET}}^{⟩}

Display a list showing each line in the buffer that contains a match for regexp. A numeric argument specifies the number of context lines to print before and after each matching line; the default is none. To limit the search to part of the buffer, narrow to that part (see Section 30.8 [Narrowing], page 374). If regexp contains upper-case letters, then the matching is done in a case-sensitive way. Otherwise, case-fold-search controls whether case is ignored. The buffer ‘*Occur*’ containing the output serves as a menu for finding the occurrences in their original context. Click Mouse-2 on an occurrence listed in ‘*Occur*’, or position point there and type ^{⟨}\underline{\overline{RET}}^{⟩}; this switches to the buffer that was searched and moves point to the original of the chosen occurrence.

M-x list-matching-lines

Synonym for M-x occur.

M-x count-matches ^{⟨}\underline{\overline{RET}}^{⟩} regexp ^{⟨}\underline{\overline{RET}}^{⟩}

Print the number of matches for regexp after point.

M-x flush-lines ^{⟨}\underline{\overline{RET}}^{⟩} regexp ^{⟨}\underline{\overline{RET}}^{⟩} Delete each line that follows point and contains a match for regexp.

M-x keep-lines ^{⟨}\underline{\overline{RET}}^{⟩} regexp ^{⟨}\underline{\overline{RET}}^{⟩} Delete each line that follows point and does not contain a match for regexp.

In addition, you can use grep from Emacs to search a collection of files for matches for a regular expression, then visit the matches either sequentially or in arbitrary order. See Section 23.2 [Grep Searching], page 270.

98 GNU Emacs Manual

13 Commands for Fixing Typos

In this chapter we describe the commands that are especially useful for the times when you catch a mistake in your text just after you have made it, or change your mind while composing text on the fly.

The most fundamental command for correcting erroneous editing is the undo command, C-x u or C-_. This command undoes a single command (usually), a part of a command (in the case of query-replace), or several consecutive self-inserting characters. Consecutive repetitions of C-_or C-x u undo earlier and earlier changes, back to the limit of the undo information available. See Section 4.4 [Undo], page 32, for for more information.

13.1 Killing Your Mistakes

⟨\underline{\overline{DEL}}^{⟩} Delete last character (delete-backward-char).

M-^{⟨}\underline{\overline{DEL}}^{⟩} Kill last word (backward-kill-word).

C-x ^{⟨}\underline{\overline{DEL}}^{⟩} Kill to beginning of sentence (backward-kill-sentence).

The ^{⟨}\underline{\overline{DEL}}^{⟩} character (delete-backward-char) is the most important correction command. It deletes the character before point. When ^{⟨}\underline{\overline{DEL}}^{⟩} follows a self-inserting character command, you can think of it as canceling that command. However, avoid the mistake of thinking of ^{⟨}\underline{\overline{DEL}}^{⟩} as a general way to cancel a command!

When your mistake is longer than a couple of characters, it might be more convenient to use M-^{⟨}\underline{\overline{DEL}}^{⟩} or C-x ^{⟨}\underline{\overline{DEL}}^{⟩}. M-^{⟨}\underline{\overline{DEL}}^{⟩} kills back to the start of the last word, and C-x ^{⟨}\underline{\overline{DEL}}^{⟩} kills back to the start of the last sentence. C-x ⟨\underline{\overline{DEL}}^{⟩} is particularly useful when you change your mind about the phrasing of the text you are writing. M-^{⟨}\underline{\overline{DEL}}^{⟩} and C-x ^{⟨}\underline{\overline{DEL}}^{⟩} save the killed text for C-y and M-y to retrieve. See Section 9.2 [Yanking], page 65.

M-^{⟨}\underline{\overline{DEL}}^{⟩} is often useful even when you have typed only a few characters wrong, if you know you are confused in your typing and aren’t sure exactly what you typed. At such a time, you cannot correct with ^{⟨}\underline{\overline{DEL}}^{⟩} except by looking at the screen to see what you did. Often it requires less thought to kill the whole word and start again.

13.2 Transposing Text

C-t Transpose two characters (transpose-chars).

M-t Transpose two words (transpose-words).

C-M-t Transpose two balanced expressions (transpose-sexps).

C-x C-t Transpose two lines (transpose-lines).

The common error of transposing two characters can be fixed, when they are adjacent, with the C-t command (transpose-chars). Normally, C-t transposes the two characters on either side of point. When given at the end of a line, rather than transposing the last character of the line with the newline, which would be useless, C-t transposes the last two characters on the line. So, if you catch your transposition error right away, you can fix it with just a C-t. If you don’t catch it so fast, you must move the cursor back to between the two transposed characters. If you transposed a space with the last character of the word before it, the word motion commands are a good way of getting there. Otherwise, a reverse search (C-r) is often the best way. See Chapter 12 [Search], page 83.

M-t (transpose-words) transposes the word before point with the word after point. It moves point forward over a word, dragging the word preceding or containing point forward as well. The punctuation characters between the words do not move. For example, ‘FOO, BAR’ transposes into ‘BAR, FOO’ rather than ‘BAR FOO,’.

C-M-t (transpose-sexps) is a similar command for transposing two expressions (see Section 22.2 [Lists], page 222), and C-x C-t (transposelines) exchanges lines. They work like M-t except in determining the division of the text into syntactic units.

A numeric argument to a transpose command serves as a repeat count: it tells the transpose command to move the character (word, sexp, line) before or containing point across several other characters (words, sexps, lines). For example, C-u 3 C-t moves the character before point forward across three other characters. It would change ‘f⋆oobar’ into ‘oobf⋆ar’. This is equivalent to repeating C-t three times. C-u - 4 M-t moves the word before point backward across four words. C-u - C-M-t would cancel the effect of plain

C-M-t.

A numeric argument of zero is assigned a special meaning (because otherwise a command with a repeat count of zero would do nothing): to transpose the character (word, sexp, line) ending after point with the one ending after the mark.

13.3 Case Conversion

M-- M-l Convert last word to lower case. Note Meta-- is Meta-minus.

M-- M-u Convert last word to all upper case.

M-- M-c Convert last word to lower case with capital initial.

A very common error is to type words in the wrong case. Because of this, the word case-conversion commands M-l, M-u and M-c have a special feature when used with a negative argument: they do not move the cursor. As soon as you see you have mistyped the last word, you can simply case-convert it and go on typing. See Section 21.6 [Case], page 202.

13.4 Checking and Correcting Spelling

This section describes the commands to check the spelling of a single word or of a portion of a buffer. These commands work with the spelling checker program Ispell, which is not part of Emacs.

M-x flyspell-mode

Enable Flyspell mode, which highlights all misspelled words.

M-$ Check and correct spelling of the word at point (ispell-word).

M-^{⟨}\underline{\overline{TAB}}^{⟩} Complete the word before point based on the spelling dictionary

          (ispell-complete-word).
M-x ispell-buffer

Check and correct spelling of each word in the buffer.

M-x ispell-region

Check and correct spelling of each word in the region.

M-x ispell-message

Check and correct spelling of each word in a draft mail message, excluding cited material.

M-x ispell-change-dictionary ^{⟨}\underline{\overline{RET}}^{⟩} dict ^{⟨}\underline{\overline{RET}}^{⟩}

Restart the Ispell process, using dict as the dictionary.

M-x ispell-kill-ispell

Kill the Ispell subprocess.

Flyspell mode is a fully-automatic way to check spelling as you edit in Emacs. It operates by checking words as you change or insert them. When it finds a word that it does not recognize, it highlights that word. This does not interfere with your editing, but when you see the highlighted word, you can move to it and fix it. Type M-x flyspell-mode to enable or disable this mode in the current buffer.

When Flyspell mode highlights a word as misspelled, you can click on it with Mouse-2 to display a menu of possible corrections and actions. You can also correct the word by editing it manually in any way you like.

The other Emacs spell-checking features check or look up words when you give an explicit command to do so. Checking all or part of the buffer is useful when you have text that was written outside of this Emacs session and might contain any number of misspellings.

To check the spelling of the word around or next to point, and optionally correct it as well, use the command M-$ (ispell-word). If the word is not correct, the command offers you various alternatives for what to do about it.

To check the entire current buffer, use M-x ispell-buffer. Use M-x ispell-region to check just the current region. To check spelling in an email message you are writing, use M-x ispell-message; that checks the whole buffer, but does not check material that is indented or appears to be cited from other messages.

Each time these commands encounter an incorrect word, they ask you what to do. They display a list of alternatives, usually including several “near-misses”—words that are close to the word being checked. Then you must type a character. Here are the valid responses:

⟨\underline{\overline{SPC}}^{⟩} Skip this word—continue to consider it incorrect, but don’t change it here.

r new ^{⟨}\underline{\overline{RET}}^{⟩} Replace the word (just this time) with new.

R new ^{⟨}\underline{\overline{RET}}^{⟩} Replace the word with new, and do a query-replace so you can replace it elsewhere in the buffer if you wish.

digit Replace the word (just this time) with one of the displayed nearmisses. Each near-miss is listed with a digit; type that digit to select it.

a Accept the incorrect word—treat it as correct, but only in this editing session.

A Accept the incorrect word—treat it as correct, but only in this editing session and for this buffer.

i Insert this word in your private dictionary file so that Ispell will consider it correct it from now on, even in future sessions.

u Insert the lower-case version of this word in your private dictionary file.

m Like i, but you can also specify dictionary completion information.

l word ^{⟨}\underline{\overline{RET}}^{⟩} Look in the dictionary for words that match word. These words become the new list of “near-misses”; you can select one of them to replace with by typing a digit. You can use ‘*’ in word as a wildcard.

C-g Quit interactive spell checking. You can restart it again afterward

          with C-u M-$.

X Same as C-g.

x Quit interactive spell checking and move point back to where it was when you started spell checking.

q Quit interactive spell checking and kill the Ispell subprocess.

C-l Refresh the screen.

C-z This key has its normal command meaning (suspend Emacs or iconify this frame).

The command ispell-complete-word, which is bound to the key M-^{⟨}\underline{\overline{TAB}}^{⟩} in Text mode and related modes, shows a list of completions based on spelling correction. Insert the beginning of a word, and then type M-^{⟨}\underline{\overline{TAB}}^{⟩}; the command displays a completion list window. To choose one of the completions listed, click Mouse-2 on it, or move the cursor there in the completions window and type ^{⟨}\underline{\overline{RET}}^{⟩}. See Section 21.7 [Text Mode], page 203.

Once started, the Ispell subprocess continues to run (waiting for something to do), so that subsequent spell checking commands complete more quickly. If you want to get rid of the Ispell process, use M-x ispell-kill-ispell. This is not usually necessary, since the process uses no time except when you do spelling correction.

Ispell uses two dictionaries: the standard dictionary and your private dictionary. The variable ispell-dictionary specifies the file name of the standard dictionary to use. A value of nil says to use the default dictionary. The command M-x ispell-change-dictionary sets this variable and then restarts the Ispell subprocess, so that it will use a different dictionary.

104 GNU Emacs Manual

14 File Handling

The operating system stores data permanently in named files. So most of the text you edit with Emacs comes from a file and is ultimately stored in a file.

To edit a file, you must tell Emacs to read the file and prepare a buffer containing a copy of the file’s text. This is called visiting the file. Editing commands apply directly to text in the buffer; that is, to the copy inside Emacs. Your changes appear in the file itself only when you save the buffer back into the file.

In addition to visiting and saving files, Emacs can delete, copy, rename, and append to files, keep multiple versions of them, and operate on file directories.

14.1 File Names

Most Emacs commands that operate on a file require you to specify the file name. (Saving and reverting are exceptions; the buffer knows which file name to use for them.) You enter the file name using the minibuffer (see Chapter 5 [Minibuffer], page 39). Completion is available, to make it easier to specify long file names. See Section 5.3 [Completion], page 41.

For most operations, there is a default file name which is used if you type just ^{⟨}\underline{\overline{RET}}^{⟩} to enter an empty argument. Normally the default file name is the name of the file visited in the current buffer; this makes it easy to operate on that file with any of the Emacs file commands.

Each buffer has a default directory, normally the same as the directory of the file visited in that buffer. When you enter a file name without a directory, the default directory is used. If you specify a directory in a relative fashion, with a name that does not start with a slash, it is interpreted with respect to the default directory. The default directory is kept in the variable defaultdirectory, which has a separate value in every buffer.

For example, if the default file name is ‘/u/rms/gnu/gnu.tasks’ then the default directory is ‘/u/rms/gnu/’. If you type just ‘foo’, which does not specify a directory, it is short for ‘/u/rms/gnu/foo’. ‘../.login’ would stand for ‘/u/rms/.login’. ‘new/foo’ would stand for the file name

‘/u/rms/gnu/new/foo’.

The command M-x pwd prints the current buffer’s default directory, and the command M-x cd sets it (to a value read using the minibuffer). A buffer’s default directory changes only when the cd command is used. A file-visiting buffer’s default directory is initialized to the directory of the file that is visited there. If you create a buffer with C-x b, its default directory is copied from that of the buffer that was current at the time.

The default directory actually appears in the minibuffer when the minibuffer becomes active to read a file name. This serves two purposes: it shows you what the default is, so that you can type a relative file name and know with certainty what it will mean, and it allows you to edit the default to specify a different directory. This insertion of the default directory is inhibited if the variable insert-default-directory is set to nil.

Note that it is legitimate to type an absolute file name after you enter the minibuffer, ignoring the presence of the default directory name as part of the text. The final minibuffer contents may look invalid, but that is not so. For example, if the minibuffer starts out with ‘/usr/tmp/’ and you add ‘/x1/rms/foo’, you get ‘/usr/tmp//x1/rms/foo’; but Emacs ignores everything through the first slash in the double slash; the result is ‘/x1/rms/foo’. See Section 5.1 [Minibuffer File], page 39.

‘$’ in a file name is used to substitute environment variables. For example, if you have used the shell command ‘export FOO=rms/hacks’ to set up an environment variable named FOO, then you can use ‘/u/$FOO/test.c’ or

‘/u/${FOO}/test.c’ as an abbreviation for ‘/u/rms/hacks/test.c’. The

environment variable name consists of all the alphanumeric characters after the ‘$’; alternatively, it may be enclosed in braces after the ‘$’. Note that shell commands to set environment variables affect Emacs only if done before Emacs is started.

To access a file with ‘$’ in its name, type ‘$$’. This pair is converted to a single ‘$’ at the same time as variable substitution is performed for single ‘$’. Alternatively, quote the whole file name with ‘/:’ (see Section 14.13 [Quoted File Names], page 139).

The Lisp function that performs the substitution is called substitutein-file-name. The substitution is performed only on file names read as such using the minibuffer.

You can include non-ASCII characters in file names if you set the variable file-name-coding-system to a non-nil value. See Section 18.8 [Specify Coding], page 179.

14.2 Visiting Files

C-x C-f Visit a file (find-file).

C-x C-r Visit a file for viewing, without allowing changes to it (find-

          file-read-only).

C-x C-v Visit a different file instead of the one visited last (find-

          alternate-file).

C-x 4 f Visit a file, in another window (find-file-other-window). Don’t alter what is displayed in the selected window.

C-x 5 f Visit a file, in a new frame (find-file-other-frame). Don’t alter what is displayed in the selected frame.

M-x find-file-literally

Visit a file with no conversion of the contents.

Visiting a file means copying its contents into an Emacs buffer so you can edit them. Emacs makes a new buffer for each file that you visit. We say that this buffer is visiting the file that it was created to hold. Emacs constructs the buffer name from the file name by throwing away the directory, keeping just the name proper. For example, a file named ‘/usr/rms/emacs.tex’ would get a buffer named ‘emacs.tex’. If there is already a buffer with that name, a unique name is constructed by appending ‘<2>’, ‘<3>’, or so on, using the lowest number that makes a name that is not already in use.

Each window’s mode line shows the name of the buffer that is being displayed in that window, so you can always tell what buffer you are editing.

The changes you make with editing commands are made in the Emacs buffer. They do not take effect in the file that you visited, or any place permanent, until you save the buffer. Saving the buffer means that Emacs writes the current contents of the buffer into its visited file. See Section 14.3 [Saving], page 109.

If a buffer contains changes that have not been saved, we say the buffer is modified. This is important because it implies that some changes will be lost if the buffer is not saved. The mode line displays two stars near the left margin to indicate that the buffer is modified.

To visit a file, use the command C-x C-f (find-file). Follow the command with the name of the file you wish to visit, terminated by a ^{⟨}\underline{\overline{RET}}^{⟩}.

The file name is read using the minibuffer (see Chapter 5 [Minibuffer], page 39), with defaulting and completion in the standard manner (see Section 14.1 [File Names], page 105). While in the minibuffer, you can abort

C-x C-f by typing C-g.

Your confirmation that C-x C-f has completed successfully is the appearance of new text on the screen and a new buffer name in the mode line. If the specified file does not exist and could not be created, or cannot be read, then you get an error, with an error message displayed in the echo area.

If you visit a file that is already in Emacs, C-x C-f does not make another copy. It selects the existing buffer containing that file. However, before doing so, it checks that the file itself has not changed since you visited or saved it last. If the file has changed, a warning message is printed. See Section 14.3.2 [Simultaneous Editing], page 113.

What if you want to create a new file? Just visit it. Emacs prints ‘(New File)’ in the echo area, but in other respects behaves as if you had visited an existing empty file. If you make any changes and save them, the file is created.

If the file you specify is actually a directory, C-x C-f invokes Dired, the Emacs directory browser, so that you can “edit” the contents of the directory (see Chapter 28 [Dired], page 321). Dired is a convenient way to delete, look at, or operate on the files in the directory. However, if the variable findfile-run-dired is nil, then it is an error to try to visit a directory.

If you visit a file that the operating system won’t let you modify, Emacs makes the buffer read-only, so that you won’t go ahead and make changes that you’ll have trouble saving afterward. You can make the buffer writable with C-x C-q (vc-toggle-read-only). See Section 15.3 [Misc Buffer], page 143.

Occasionally you might want to visit a file as read-only in order to protect yourself from entering changes accidentally; do so by visiting the file with

the command C-x C-r (find-file-read-only).

If you visit a nonexistent file unintentionally (because you typed the wrong file name), use the C-x C-v command (find-alternate-file) to visit the file you really wanted. C-x C-v is similar to C-x C-f, but it kills the current buffer (after first offering to save it if it is modified). When it reads the file name to visit, it inserts the entire default file name in the buffer, with point just after the directory part; this is convenient if you made a slight error in typing the name.

If you find a file which exists but cannot be read, C-x C-f signals an error.

   C-x 4 f (find-file-other-window) is like C-x C-f except that the buffer

containing the specified file is selected in another window. The window that was selected before C-x 4 f continues to show the same buffer it was already showing. If this command is used when only one window is being displayed, that window is split in two, with one window showing the same buffer as before, and the other one showing the newly requested file. See Chapter 16 [Windows], page 149.

C-x 5 f (find-file-other-frame) is similar, but opens a new frame, or makes visible any existing frame showing the file you seek. This feature is available only when you are using a window system. See Chapter 17 [Frames], page 155.

If you wish to edit a file as a sequence of characters with no special encoding or conversion, use the M-x find-file-literally command. It visits a file, like C-x C-f, but does not do format conversion (see Section 21.11 [Formatted Text], page 213), character code conversion (see Section 18.6 [Coding Systems], page 175), or automatic uncompression (see Section 14.11 [Compressed Files], page 138). If you already have visited the same file in the usual (non-literal) manner, this command asks you whether to visit it literally instead.

Two special hook variables allow extensions to modify the operation of visiting files. Visiting a file that does not exist runs the functions in the list find-file-not-found-hooks; this variable holds a list of functions, and the functions are called one by one until one of them returns non-nil. Any visiting of a file, whether extant or not, expects find-file-hooks to contain a list of functions and calls them all, one by one. In both cases the functions receive no arguments. Of these two variables, find-file-not-found-hooks takes effect first. These variables are not normal hooks, and their names end in ‘-hooks’ rather than ‘-hook’ to indicate that fact. See Section 31.2.3 [Hooks], page 392.

There are several ways to specify automatically the major mode for editing the file (see Section 19.1 [Choosing Modes], page 185), and to specify local variables defined for that file (see Section 31.2.5 [File Variables], page 395).

14.3 Saving Files

Saving a buffer in Emacs means writing its contents back into the file that was visited in the buffer.

C-x C-s Save the current buffer in its visited file (save-buffer).

C-x s Save any or all buffers in their visited files (save-some-buffers).

M-~ Forget that the current buffer has been changed (not-modified).

C-x C-w Save the current buffer in a specified file (write-file).

M-x set-visited-file-name

Change file the name under which the current buffer will be saved.

When you wish to save the file and make your changes permanent, type C-x C-s (save-buffer). After saving is finished, C-x C-s displays a message like this:

    Wrote /u/rms/gnu/gnu.tasks

If the selected buffer is not modified (no changes have been made in it since the buffer was created or last saved), saving is not really done, because it would have no effect. Instead, C-x C-s displays a message like this in the echo area:

    (No changes need to be saved)

The command C-x s (save-some-buffers) offers to save any or all modified buffers. It asks you what to do with each buffer. The possible responses are analogous to those of query-replace:

y Save this buffer and ask about the rest of the buffers.

n Don’t save this buffer, but ask about the rest of the buffers.

! Save this buffer and all the rest with no more questions.

⟨\underline{\overline{RET}}^{⟩} Terminate save-some-buffers without any more saving.

. Save this buffer, then exit save-some-buffers without even asking about other buffers.

C-r View the buffer that you are currently being asked about. When you exit View mode, you get back to save-some-buffers, which asks the question again.

C-h Display a help message about these options.

C-x C-c, the key sequence to exit Emacs, invokes save-some-buffers and therefore asks the same questions.

If you have changed a buffer but you do not want to save the changes, you should take some action to prevent it. Otherwise, each time you use C-x s or C-x C-c, you are liable to save this buffer by mistake. One thing you can do is type M-~ (not-modified), which clears out the indication that the buffer is modified. If you do this, none of the save commands will believe that the buffer needs to be saved. (‘~’ is often used as a mathematical symbol for ‘not’; thus M-~ is ‘not’, metafied.) You could also use set-visited-file-name (see below) to mark the buffer as visiting a different file name, one which is not in use for anything important. Alternatively, you can cancel all the changes made since the file was visited or saved, by reading the text from the file again. This is called reverting. See Section 14.4 [Reverting], page 114. You could also undo all the changes by repeating the undo command C-x u until you have undone all the changes; but reverting is easier.

M-x set-visited-file-name alters the name of the file that the current buffer is visiting. It reads the new file name using the minibuffer. Then it specifies the visited file name and changes the buffer name correspondingly (as long as the new name is not in use). set-visited-file-name does not save the buffer in the newly visited file; it just alters the records inside Emacs in case you do save later. It also marks the buffer as “modified” so that C-x C-s in that buffer will save.

If you wish to mark the buffer as visiting a different file and save it right away, use C-x C-w (write-file). It is precisely equivalent to set-visited-file-name followed by C-x C-s. C-x C-s used on a buffer that is not visiting a file has the same effect as C-x C-w; that is, it reads a file name, marks the buffer as visiting that file, and saves it there. The default file name in a buffer that is not visiting a file is made by combining the buffer name with the buffer’s default directory.

If the new file name implies a major mode, then C-x C-w switches to that major mode, in most cases. The command set-visited-file-name also does this. See Section 19.1 [Choosing Modes], page 185.

If Emacs is about to save a file and sees that the date of the latest version on disk does not match what Emacs last read or wrote, Emacs notifies you of this fact, because it probably indicates a problem caused by simultaneous editing and requires your immediate attention. See Section 14.3.2 [Simultaneous Editing], page 113.

If the variable require-final-newline is non-nil, Emacs puts a newline at the end of any file that doesn’t already end in one, every time a file is saved or written. The default is nil.

14.3.1 Backup Files

On most operating systems, rewriting a file automatically destroys all record of what the file used to contain. Thus, saving a file from Emacs throws away the old contents of the file—or it would, except that Emacs carefully copies the old contents to another file, called the backup file, before actually saving. (This assumes that the variable make-backup-files is nonnil. Backup files are not written if this variable is nil.) Emacs does not normally make backup files for files in ‘/tmp’.

At your option, Emacs can keep either a single backup file or a series of numbered backup files for each file that you edit.

Emacs makes a backup for a file only the first time the file is saved from one buffer. No matter how many times you save a file, its backup file continues to contain the contents from before the file was visited. Normally this means that the backup file contains the contents from before the current editing session; however, if you kill the buffer and then visit the file again, a new backup file will be made by the next save.

You can also explicitly request making another backup file from a buffer even though it has already been saved at least once. If you save the buffer with C-u C-x C-s, the version thus saved will be made into a backup file if you save the buffer again. C-u C-u C-x C-s saves the buffer, but first makes the previous file contents into a new backup file. C-u C-u C-u C-x C-s does both things: it makes a backup from the previous contents, and arranges to make another from the newly saved contents, if you save again.

14.3.1.1 Single or Numbered Backups

If you choose to have a single backup file (this is the default), the backup file’s name is constructed by appending ‘~’ to the file name being edited; thus, the backup file for ‘eval.c’ would be ‘eval.c~’.

If you choose to have a series of numbered backup files, backup file names are made by appending ‘.~’, the number, and another ‘~’ to the original file name. Thus, the backup files of ‘eval.c’ would be called ‘eval.c.~1~’, ‘eval.c.~2~’, and so on, through names like ‘eval.c.~259~’ and beyond.

If protection stops you from writing backup files under the usual names, the backup file is written as ‘%backup%~’ in your home directory. Only one such file can exist, so only the most recently made such backup is available.

The choice of single backup or numbered backups is controlled by the variable version-control. Its possible values are

t Make numbered backups.

nil Make numbered backups for files that have numbered backups already. Otherwise, make single backups.

never Do not in any case make numbered backups; always make single backups.

You can set version-control locally in an individual buffer to control the making of backups for that buffer’s file. For example, Rmail mode locally sets version-control to never to make sure that there is only one backup for an Rmail file. See Section 31.2.4 [Locals], page 393.

If you set the environment variable VERSION_CONTROL, to tell various GNU utilities what to do with backup files, Emacs also obeys the environment variable by setting the Lisp variable version-control accordingly at startup. If the environment variable’s value is ‘t’ or ‘numbered’, then version-control becomes t; if the value is ‘nil’ or ‘existing’, then

version-control becomes nil; if it is ‘never’ or ‘simple’, then version-
control becomes never.

For files managed by a version control system (see Section 14.7 [Version Control], page 118), the variable vc-make-backup-files determines whether to make backup files. By default, it is nil, since backup files are redundant when you store all the previous versions in a version control system. See Section 14.7.9.2 [VC Workfile Handling], page 134.

14.3.1.2 Automatic Deletion of Backups

To prevent unlimited consumption of disk space, Emacs can delete numbered backup versions automatically. Generally Emacs keeps the first few backups and the latest few backups, deleting any in between. This happens every time a new backup is made.

   The two variables kept-old-versions and kept-new-versions control

this deletion. Their values are, respectively the number of oldest (lowestnumbered) backups to keep and the number of newest (highest-numbered) ones to keep, each time a new backup is made. Recall that these values are used just after a new backup version is made; that newly made backup is included in the count in kept-new-versions. By default, both variables are 2.

If delete-old-versions is non-nil, the excess middle versions are deleted without a murmur. If it is nil, the default, then you are asked whether the excess middle versions should really be deleted.

Dired’s . (Period) command can also be used to delete old versions. See Section 28.3 [Dired Deletion], page 321.

14.3.1.3 Copying vs. Renaming

Backup files can be made by copying the old file or by renaming it. This makes a difference when the old file has multiple names. If the old file is renamed into the backup file, then the alternate names become names for the backup file. If the old file is copied instead, then the alternate names remain names for the file that you are editing, and the contents accessed by those names will be the new contents.

The method of making a backup file may also affect the file’s owner and group. If copying is used, these do not change. If renaming is used, you become the file’s owner, and the file’s group becomes the default (different operating systems have different defaults for the group).

Having the owner change is usually a good idea, because then the owner always shows who last edited the file. Also, the owners of the backups show who produced those versions. Occasionally there is a file whose owner should not change; it is a good idea for such files to contain local variable lists to set backup-by-copying-when-mismatch locally (see Section 31.2.5 [File Variables], page 395).

The choice of renaming or copying is controlled by three variables. Renaming is the default choice. If the variable backup-by-copying is nonnil, copying is used. Otherwise, if the variable backup-by-copying-when-linked is non-nil, then copying is used for files that have multiple names, but renaming may still be used when the file being edited has only one

name. If the variable backup-by-copying-when-mismatch is non-nil, then

copying is used if renaming would cause the file’s owner or group to change.

backup-by-copying-when-mismatch is t by default if you start Emacs as

the superuser.

When a file is managed with a version control system (see Section 14.7 [Version Control], page 118), Emacs does not normally make backups in the usual way for that file. But check-in and check-out are similar in some ways to making backups. One unfortunate similarity is that these operations typically break hard links, disconnecting the file name you visited from any alternate names for the same file. This has nothing to do with Emacs—the version control system does it.

14.3.2 Protection against Simultaneous Editing

Simultaneous editing occurs when two users visit the same file, both make changes, and then both save them. If nobody were informed that this was happening, whichever user saved first would later find that his changes were lost.

On some systems, Emacs notices immediately when the second user starts to change the file, and issues an immediate warning. On all systems, Emacs checks when you save the file, and warns if you are about to overwrite another user’s changes. You can prevent loss of the other user’s work by taking the proper corrective action instead of saving the file.

When you make the first modification in an Emacs buffer that is visiting a file, Emacs records that the file is locked by you. (It does this by creating a symbolic link in the same directory with a different name.) Emacs removes the lock when you save the changes. The idea is that the file is locked whenever an Emacs buffer visiting it has unsaved changes.

If you begin to modify the buffer while the visited file is locked by someone else, this constitutes a collision. When Emacs detects a collision, it asks you what to do, by calling the Lisp function ask-user-about-lock. You can redefine this function for the sake of customization. The standard definition of this function asks you a question and accepts three possible answers:

s Steal the lock. Whoever was already changing the file loses the lock, and you gain the lock.

p Proceed. Go ahead and edit the file despite its being locked by someone else.

q Quit. This causes an error (file-locked) and the modification you were trying to make in the buffer does not actually take place.

Note that locking works on the basis of a file name; if a file has multiple names, Emacs does not realize that the two names are the same file and cannot prevent two users from editing it simultaneously under different names. However, basing locking on names means that Emacs can interlock the editing of new files that will not really exist until they are saved.

Some systems are not configured to allow Emacs to make locks, and there are cases where lock files cannot be written. In these cases, Emacs cannot detect trouble in advance, but it still can detect the collision when you try to save a file and overwrite someone else’s changes.

If Emacs or the operating system crashes, this may leave behind lock files which are stale. So you may occasionally get warnings about spurious collisions. When you determine that the collision is spurious, just use p to tell Emacs to go ahead anyway.

Every time Emacs saves a buffer, it first checks the last-modification date of the existing file on disk to verify that it has not changed since the file was last visited or saved. If the date does not match, it implies that changes were made in the file in some other way, and these changes are about to be lost if Emacs actually does save. To prevent this, Emacs prints a warning message and asks for confirmation before saving. Occasionally you will know why the file was changed and know that it does not matter; then you can answer yes and proceed. Otherwise, you should cancel the save with C-g and investigate the situation.

The first thing you should do when notified that simultaneous editing has already taken place is to list the directory with C-u C-x C-d (see Section 14.8 [Directories], page 136). This shows the file’s current author. You should attempt to contact him to warn him not to continue editing. Often the next step is to save the contents of your Emacs buffer under a different name, and use diff to compare the two files.

14.4 Reverting a Buffer

If you have made extensive changes to a file and then change your mind about them, you can get rid of them by reading in the previous version of the file. To do this, use M-x revert-buffer, which operates on the current buffer. Since reverting a buffer unintentionally could lose a lot of work, you must confirm this command with yes.

revert-buffer keeps point at the same distance (measured in characters) from the beginning of the file. If the file was edited only slightly, you will be at approximately the same piece of text after reverting as before. If you have made drastic changes, the same value of point in the old file may address a totally different piece of text.

Reverting marks the buffer as “not modified” until another change is made.

Some kinds of buffers whose contents reflect data bases other than files, such as Dired buffers, can also be reverted. For them, reverting means recalculating their contents from the appropriate data base. Buffers created explicitly with C-x b cannot be reverted; revert-buffer reports an error when asked to do so.

When you edit a file that changes automatically and frequently—for example, a log of output from a process that continues to run—it may be useful for Emacs to revert the file without querying you, whenever you visit the file again with C-x C-f.

To request this behavior, set the variable revert-without-query to a list of regular expressions. When a file name matches one of these regular expressions, find-file and revert-buffer will revert it automatically if it has changed—provided the buffer itself is not modified. (If you have edited the text, it would be wrong to discard your changes.)

14.5 Auto-Saving: Protection Against Disasters

Emacs saves all the visited files from time to time (based on counting your keystrokes) without being asked. This is called auto-saving. It prevents you from losing more than a limited amount of work if the system crashes.

When Emacs determines that it is time for auto-saving, each buffer is considered, and is auto-saved if auto-saving is turned on for it and it has been changed since the last time it was auto-saved. The message ‘Auto-saving...’ is displayed in the echo area during auto-saving, if any files are actually auto-saved. Errors occurring during auto-saving are caught so that they do not interfere with the execution of commands you have been typing.

14.5.1 Auto-Save Files

Auto-saving does not normally save in the files that you visited, because it can be very undesirable to save a program that is in an inconsistent state when you have made half of a planned change. Instead, auto-saving is done in a different file called the auto-save file, and the visited file is changed only when you request saving explicitly (such as with C-x C-s).

Normally, the auto-save file name is made by appending ‘#’ to the front and rear of the visited file name. Thus, a buffer visiting file ‘foo.c’ is auto-saved in a file ‘#foo.c#’. Most buffers that are not visiting files are auto-saved only if you request it explicitly; when they are auto-saved, the auto-save file name is made by appending ‘#%’ to the front and ‘#’ to the rear of buffer name. For example, the ‘*mail*’ buffer in which you compose messages to be sent is auto-saved in a file named ‘#%*mail*#’. Auto-save file names are made this way unless you reprogram parts of Emacs to do something different (the functions make-auto-save-file-name and autosave-file-name-p). The file name to be used for auto-saving in a buffer is calculated when auto-saving is turned on in that buffer.

When you delete a substantial part of the text in a large buffer, auto save turns off temporarily in that buffer. This is because if you deleted the text unintentionally, you might find the auto-save file more useful if it contains the deleted text. To reenable auto-saving after this happens, save the buffer

with C-x C-s, or use C-u 1 M-x auto-save.

If you want auto-saving to be done in the visited file, set the variable auto-save-visited-file-name to be non-nil. In this mode, there is really no difference between auto-saving and explicit saving.

A buffer’s auto-save file is deleted when you save the buffer in its visited file. To inhibit this, set the variable delete-auto-save-files to nil. Changing the visited file name with C-x C-w or set-visited-file-name renames any auto-save file to go with the new visited name.

14.5.2 Controlling Auto-Saving

Each time you visit a file, auto-saving is turned on for that file’s buffer if the variable auto-save-default is non-nil (but not in batch mode; see Chapter 3 [Entering Emacs], page 25). The default for this variable is t, so auto-saving is the usual practice for file-visiting buffers. Auto-saving can be turned on or off for any existing buffer with the command M-x

auto-save-mode. Like other minor mode commands, M-x auto-save-mode

turns auto-saving on with a positive argument, off with a zero or negative argument; with no argument, it toggles.

Emacs does auto-saving periodically based on counting how many characters you have typed since the last time auto-saving was done. The variable auto-save-interval specifies how many characters there are between autosaves. By default, it is 300.

Auto-saving also takes place when you stop typing for a while. The variable auto-save-timeout says how many seconds Emacs should wait before it does an auto save (and perhaps also a garbage collection). (The actual time period is longer if the current buffer is long; this is a heuristic which aims to keep out of your way when you are editing long buffers, in which auto-save takes an appreciable amount of time.) Auto-saving during idle periods accomplishes two things: first, it makes sure all your work is saved if you go away from the terminal for a while; second, it may avoid some auto-saving while you are actually typing.

Emacs also does auto-saving whenever it gets a fatal error. This includes killing the Emacs job with a shell command such as ‘kill %emacs’, or disconnecting a phone line or network connection.

You can request an auto-save explicitly with the command M-x

do-auto-save.

14.5.3 Recovering Data from Auto-Saves

You can use the contents of an auto-save file to recover from a loss of data with the command M-x recover-file ^{⟨}\underline{\overline{RET}}^{⟩} file ^{⟨}\underline{\overline{RET}}^{⟩}. This visits file and then (after your confirmation) restores the contents from its auto-save file ‘#file#’. You can then save with C-x C-s to put the recovered text into file itself. For example, to recover file ‘foo.c’ from its auto-save file ‘#foo.c#’, do:

    M-x recover-file ^{⟨}\underline{\overline{RET}}^{⟩} foo.c ^{⟨}\underline{\overline{RET}}^{⟩}

yes ^{⟨}\underline{\overline{RET}}^{⟩}

    C-x C-s

Before asking for confirmation, M-x recover-file displays a directory listing describing the specified file and the auto-save file, so you can compare their sizes and dates. If the auto-save file is older, M-x recover-file does not offer to read it.

If Emacs or the computer crashes, you can recover all the files you were editing from their auto save files with the command M-x recover-session. This first shows you a list of recorded interrupted sessions. Move point to the one you choose, and type C-c C-c.

Then recover-session asks about each of the files that were being edited during that session, asking whether to recover that file. If you answer y, it calls recover-file, which works in its normal fashion. It shows the dates of the original file and its auto-save file, and asks once again whether to recover that file.

When recover-session is done, the files you’ve chosen to recover are present in Emacs buffers. You should then save them. Only this—saving them—updates the files themselves.

Interrupted sessions are recorded for later recovery in files named ‘~/.saves-pid-hostname’. The ‘~/.saves’ portion of these names comes from the value of auto-save-list-file-prefix. You can arrange to record sessions in a different place by setting that variable in your ‘.emacs’ file, but you’ll have to redefine recover-session as well to make it look in the new

place. If you set auto-save-list-file-prefix to nil in your ‘.emacs’

file, sessions are not recorded for recovery.

14.6 File Name Aliases

Symbolic links and hard links both make it possible for several file names to refer to the same file. Hard links are alternate names that refer directly to the file; all the names are equally valid, and no one of them is preferred. By contrast, a symbolic link is a kind of defined alias: when ‘foo’ is a symbolic link to ‘bar’, you can use either name to refer to the file, but ‘bar’ is the real name, while ‘foo’ is just an alias. More complex cases occur when symbolic links point to directories.

If you visit two names for the same file, normally Emacs makes two different buffers, but it warns you about the situation.

If you wish to avoid visiting the same file in two buffers under differ-

ent names, set the variable find-file-existing-other-name to a non-nil

value. Then find-file uses the existing buffer visiting the file, no matter which of the file’s names you specify.

If the variable find-file-visit-truename is non-nil, then the file name recorded for a buffer is the file’s truename (made by replacing all symbolic links with their target names), rather than the name you specify.

Setting find-file-visit-truename also implies the effect of find-file-
existing-other-name.

14.7 Version Control

Version control systems are packages that can record multiple versions of a source file, usually storing the unchanged parts of the file just once. Version control systems also record history information such as the creation time of each version, who created it, and a description of what was changed in that version.

The Emacs version control interface is called VC. Its commands work with three version control systems—RCS, CVS and SCCS. The GNU project recommends RCS and CVS, which are free software and available from the Free Software Foundation.

14.7.1 Introduction to Version Control

VC allows you to use a version control system from within Emacs, integrating the version control operations smoothly with editing. VC provides a uniform interface to version control, so that regardless of which version control system is in use, you can use it the same way.

This section provides a general overview of version control, and describes the version control systems that VC supports. You can skip this section if you are already familiar with the version control system you want to use.

14.7.1.1 Supported Version Control Systems

VC currently works with three different version control systems or “back ends”: RCS, CVS, and SCCS.

RCS is a free version control system that is available from the Free Software Foundation. It is perhaps the most mature of the supported back ends, and the VC commands are conceptually closest to RCS. Almost everything you can do with RCS can be done through VC.

CVS is built on top of RCS, and extends the features of RCS, allowing for more sophisticated release management, and concurrent multi-user development. VC supports basic editing operations under CVS, but for some less common tasks you still need to call CVS from the command line. Note also that before using CVS you must set up a repository, which is a subject too complex to treat here.

SCCS is a proprietary but widely used version control system. In terms of capabilities, it is the weakest of the three that VC supports. VC compensates for certain features missing in SCCS (snapshots, for example) by implementing them itself, but some other VC features, such as multiple branches, are not available with SCCS. You should use SCCS only if for some reason you cannot use RCS.

14.7.1.2 Concepts of Version Control

When a file is under version control, we also say that it is registered in the version control system. Each registered file has a corresponding master file which represents the file’s present state plus its change history—enough to reconstruct the current version or any earlier version. Usually the master file also records a log entry for each version, describing in words what was changed in that version.

The file that is maintained under version control is sometimes called the work file corresponding to its master file. You edit the work file and make changes in it, as you would with an ordinary file. (With SCCS and RCS, you must lock the file before you start to edit it.) After you are done with a set of changes, you check the file in, which records the changes in the master file, along with a log entry for them.

With CVS, there are usually multiple work files corresponding to a single master file—often each user has his own copy. It is also possible to use RCS in this way, but this is not the usual way to use RCS.

A version control system typically has some mechanism to coordinate between users who want to change the same file. One method is locking (analogous to the locking that Emacs uses to detect simultaneous editing of a file, but distinct from it). The other method is to merge your changes with other people’s changes when you check them in.

With version control locking, work files are normally read-only so that you cannot change them. You ask the version control system to make a work file writable for you by locking it; only one user can do this at any given time. When you check in your changes, that unlocks the file, making the work file read-only again. This allows other users to lock the file to make further changes. SCCS always uses locking, and RCS normally does.

The other alternative for RCS is to let each user modify the work file at any time. In this mode, locking is not required, but it is permitted; check-in is still the way to record a new version.

CVS normally allows each user to modify his own copy of the work file at any time, but requires merging with changes from other users at checkin time. However, CVS can also be set up to require locking. (see Section 14.7.9.1 [Backend Options], page 134).

14.7.2 Version Control and the Mode Line

When you visit a file that is under version control, Emacs indicates this on the mode line. For example, ‘RCS-1.3’ says that RCS is used for that file, and the current version is 1.3.

The character between the back-end name and the version number indicates the version control status of the file. ‘-’ means that the work file is not locked (if locking is in use), or not modified (if locking is not in use). ‘:’ indicates that the file is locked, or that it is modified. If the file is locked by some other user (for instance, ‘jim’), that is displayed as ‘RCS:jim:1.3’.

14.7.3 Basic Editing under Version Control

The principal VC command is an all-purpose command that performs either locking or check-in, depending on the situation.

C-x C-q

C-x v v Perform the next logical version control operation on this file.

Strictly speaking, the command for this job is vc-next-action, bound to C-x v v. However, the normal meaning of C-x C-q is to make a readonly buffer writable, or vice versa; we have extended it to do the same job properly for files managed by version control, by performing the appropriate version control operations. When you type C-x C-q on a registered file, it acts like C-x v v.

The precise action of this command depends on the state of the file, and whether the version control system uses locking or not. SCCS and RCS normally use locking; CVS normally does not use locking.

14.7.3.1 Basic Version Control with Locking

If locking is used for the file (as with SCCS, and RCS in its default mode), C-x C-q can either lock a file or check it in:

• If the file is not locked, C-x C-q locks it, and makes it writable so that you can change it.

• If the file is locked by you, and contains changes, C-x C-q checks in the changes. In order to do this, it first reads the log entry for the new version. See Section 14.7.3.3 [Log Buffer], page 122.

• If the file is locked by you, but you have not changed it since you locked it, C-x C-q releases the lock and makes the file read-only again.

• If the file is locked by some other user, C-x C-q asks you whether you want to “steal the lock” from that user. If you say yes, the file becomes locked by you, but a message is sent to the person who had formerly locked the file, to inform him of what has happened.

These rules also apply when you use CVS in locking mode, except that there is no such thing as stealing a lock.

14.7.3.2 Basic Version Control without Locking

When there is no locking–the default for CVS–work files are always writable; you do not need to do anything before you begin to edit a file. The status indicator on the mode line is ‘-’ if the file is unmodified; it flips to ‘:’ as soon as you save any changes in the work file.

Here is what C-x C-q does when using CVS:

• If some other user has checked in changes into the master file, Emacs asks you whether you want to merge those changes into your own work file (see Section 14.7.6.3 [Merging], page 127). You must do this before you can check in your own changes.

• If there are no new changes in the master file, but you have made modifications in your work file, C-x C-q checks in your changes. In order to do this, it first reads the log entry for the new version. See Section 14.7.3.3 [Log Buffer], page 122.

• If the file is not modified, the C-x C-q does nothing.

These rules also apply when you use RCS in the mode that does not require locking, except that automatic merging of changes from the master file is not implemented. Unfortunately, this means that nothing informs you if another user has checked in changes in the same file since you began editing it, and when this happens, his changes will be effectively removed when you check in your version (though they will remain in the master file, so they will not be entirely lost). You must therefore verify the current version is unchanged, before you check in your changes. We hope to eliminate this risk and provide automatic merging with RCS in a future Emacs version.

In addition, locking is possible with RCS even in this mode, although it is not required; C-x C-q with an unmodified file locks the file, just as it does with RCS in its normal (locking) mode.

14.7.3.3 Features of the Log Entry Buffer

When you check in changes, C-x C-q first reads a log entry. It pops up a buffer called ‘*VC-Log*’ for you to enter the log entry. When you are finished, type C-c C-c in the ‘*VC-Log*’ buffer. That is when check-in really happens.

To abort check-in, just don’t type C-c C-c in that buffer. You can switch buffers and do other editing. As long as you don’t try to check in another file, the entry you were editing remains in the ‘*VC-Log*’ buffer, and you can go back to that buffer at any time to complete the check-in.

If you change several source files for the same reason, it is often convenient to specify the same log entry for many of the files. To do this, use the history of previous log entries. The commands M-n, M-p, M-s and M-r for doing this work just like the minibuffer history commands (except that these versions are used outside the minibuffer).

Each time you check in a file, the log entry buffer is put into VC Log mode, which involves running two hooks: text-mode-hook and vc-log-mode-hook. See Section 31.2.3 [Hooks], page 392.

14.7.4 Examining And Comparing Old Versions

One of the convenient features of version control is the ability to examine any version of a file, or compare two versions.

C-x v ~ version ^{⟨}\underline{\overline{RET}}^{⟩} Examine version version of the visited file, in a buffer of its own.

C-x v = Compare the current buffer contents with the latest checked-in version of the file.

C-u C-x v = file ^{⟨}\underline{\overline{RET}}^{⟩} oldvers ^{⟨}\underline{\overline{RET}}^{⟩} newvers ^{⟨}\underline{\overline{RET}}^{⟩} Compare the specified two versions of file.

C-x v g Display the result of the CVS annotate command using colors.

To examine an old version in toto, visit the file and then type C-x v ~ version ^{⟨}\underline{\overline{RET}}^{⟩} (vc-version-other-window). This puts the text of version version in a file named ‘filename.~version~’, and visits it in its own buffer in a separate window. (In RCS, you can also select an old version and create a branch from it. See Section 14.7.6 [Branches], page 126.)

But usually it is more convenient to compare two versions of the file, with the command C-x v = (vc-diff). Plain C-x v = compares the current buffer contents (saving them in the file if necessary) with the last checked-in version of the file. C-u C-x v =, with a numeric argument, reads a file name and two version numbers, then compares those versions of the specified file.

If you supply a directory name instead of the name of a registered file, this command compares the two specified versions of all registered files in that directory and its subdirectories.

You can specify a checked-in version by its number; an empty input specifies the current contents of the work file (which may be different from all the checked-in versions). You can also specify a snapshot name (see Section 14.7.7 [Snapshots], page 129) instead of one or both version numbers.

This command works by running the diff utility, getting the options from the variable diff-switches. It displays the output in a special buffer in another window. Unlike the M-x diff command, C-x v = does not try to locate the changes in the old and new versions. This is because normally one or both versions do not exist as files when you compare them; they exist only in the records of the master file. See Section 14.9 [Comparing Files], page 137, for more information about M-x diff.

For CVS-controlled files, you can display the result of the CVS annotate command, using colors to enhance the visual appearance. Use the command M-x vc-annotate to do this. Red means new, blue means old, and intermediate colors indicate intermediate ages. A prefix argument n specifies a stretch factor for the time scale; it makes each color cover a period n times as long.

14.7.5 The Secondary Commands of VC

This section explains the secondary commands of VC; those that you might use once a day.

14.7.5.1 Registering a File for Version Control

You can put any file under version control by simply visiting it, and then

typing C-x v i (vc-register).

C-x v i Register the visited file for version control.

To register the file, Emacs must choose which version control system to use for it. You can specify your choice explicitly by setting vc-default-back-end to RCS, CVS or SCCS. Otherwise, if there is a subdirectory named ‘RCS’, ‘SCCS’, or ‘CVS’, Emacs uses the corresponding version control system. In the absence of any specification, the default choice is RCS if RCS is installed, otherwise SCCS.

If locking is in use,C-x v i leaves the file unlocked and read-only. Type C-x C-q if you wish to start editing it. After registering a file with CVS, you must subsequently commit the initial version by typing C-x C-q.

The initial version number for a newly registered file is 1.1, by default. You can specify a different default by setting the variable vc-default-init-version, or you can give C-x v i a numeric argument; then it reads the initial version number for this particular file using the minibuffer.

If vc-initial-comment is non-nil, C-x v i reads an initial comment to describe the purpose of this source file. Reading the initial comment works like reading a log entry (see Section 14.7.3.3 [Log Buffer], page 122).

14.7.5.2 VC Status Commands

C-x v l Display version control state and change history.

To view the detailed version control status and history of a file, type C-x v l (vc-print-log). It displays the history of changes to the current file, including the text of the log entries. The output appears in a separate window.

14.7.5.3 Undoing Version Control Actions

C-x v u Revert the buffer and the file to the last checked-in version.

C-x v c Remove the last-entered change from the master for the visited file. This undoes your last check-in.

If you want to discard your current set of changes and revert to the last version checked in, use C-x v u (vc-revert-buffer). This leaves the file unlocked; if locking is in use, you must first lock the file again before you change it again. C-x v u requires confirmation, unless it sees that you haven’t made any changes since the last checked-in version.

C-x v u is also the command to unlock a file if you lock it and then decide not to change it.

To cancel a change that you already checked in, use C-x v c (vc-cancel-version). This command discards all record of the most recent checked-in version. C-x v c also offers to revert your work file and buffer to the previous version (the one that precedes the version that is deleted).

If you answer no, VC keeps your changes in the buffer, and locks the file. The no-revert option is useful when you have checked in a change and then discover a trivial error in it; you can cancel the erroneous check-in, fix the error, and check the file in again.

When C-x v c does not revert the buffer, it unexpands all version control headers in the buffer instead (see Section 14.7.8.3 [Version Headers], page 132). This is because the buffer no longer corresponds to any existing version. If you check it in again, the check-in process will expand the headers properly for the new version number.

However, it is impossible to unexpand the RCS ‘$Log$’ header automatically. If you use that header feature, you have to unexpand it by hand—by deleting the entry for the version that you just canceled.

Be careful when invoking C-x v c, as it is easy to lose a lot of work with it. To help you be careful, this command always requires confirmation with yes. Note also that this command is disabled under CVS, because canceling versions is very dangerous and discouraged with CVS.

14.7.5.4 Dired under VC

When you are working on a large program, it is often useful to find out which files have changed within an entire directory tree, or to view the status of all files under version control at once, and to perform version control operations on collections of files. You can use the command C-x v d (vcdirectory) to make a directory listing that includes only files relevant for version control.

C-x v d creates a buffer which uses VC Dired Mode. This looks much like an ordinary Dired buffer (see Chapter 28 [Dired], page 321); however, normally it shows only the noteworthy files (those locked or not up-to-date). This is called terse display. If you set the variable vc-dired-terse-display to nil, then VC Dired shows all relevant files—those managed under version control, plus all subdirectories (full display). The command v t in a VC Dired buffer toggles between terse display and full display (see Section 14.7.5.5 [VC Dired Commands], page 126).

By default, VC Dired produces a recursive listing of noteworthy or relevant files at or below the given directory. You can change this by setting the variable vc-dired-recurse to nil; then VC Dired shows only the files in the given directory.

The line for an individual file shows the version control state in the place of the hard link count, owner, group, and size of the file. If the file is unmodified, in sync with the master file, the version control state shown is blank. Otherwise it consists of text in parentheses. Under RCS and SCCS, the name of the user locking the file is shown; under CVS, an abbreviated version of the ‘cvs status’ output is used. Here is an example using RCS:

       /home/jim/project:
       -rw-r--r-- (jim)      Apr  2 23:39 file1
       -r--r--r--            Apr  5 20:21 file2

The files ‘file1’ and ‘file2’ are under version control, ‘file1’ is locked by user jim, and ‘file2’ is unlocked.

Here is an example using CVS:

       /home/joe/develop:
       -rw-r--r-- (modified) Aug  2  1997 file1.c
       -rw-r--r--            Apr  4 20:09 file2.c
       -rw-r--r-- (merge)    Sep 13  1996 file3.c

Here ‘file1.c’ is modified with respect to the repository, and ‘file2.c’ is not. ‘file3.c’ is modified, but other changes have also been checked in to the repository—you need to merge them with the work file before you can check it in.

When VC Dired displays subdirectories (in the “full” display mode), it omits some that should never contain any files under version control. By default, this includes Version Control subdirectories such as ‘RCS’ and ‘CVS’; you can customize this by setting the variable vc-directory-exclusion-

list.

You can fine-tune VC Dired’s format by typing C-u x v d—as in ordinary Dired, that allows you to specify additional switches for the ‘ls’ command.

14.7.5.5 VC Dired Commands

All the usual Dired commands work normally in VC Dired mode, except for v, which is redefined as the version control prefix. You can invoke VC commands such as vc-diff and vc-print-log by typing v =, or v l, and so on. Most of these commands apply to the file name on the current line.

The command v v (vc-next-action) operates on all the marked files, so that you can lock or check in several files at once. If it operates on more than one file, it handles each file according to its current state; thus, it might lock one file, but check in another file. This could be confusing; it is up to you to avoid confusing behavior by marking a set of files that are in a similar state.

If any files call for check-in, v v reads a single log entry, then uses it for all the files being checked in. This is convenient for registering or checking in several files at once, as part of the same change.

You can toggle between terse display (only locked files, or files not up-to-date) and full display at any time by typing v t vc-dired-toggle-terse-mode. There is also a special command * l (vc-dired-mark-locked), which marks all files currently locked (or, with CVS, all files not up-to-date). Thus, typing * l t k is another way to delete from the buffer all files except those currently locked.

14.7.6 Multiple Branches of a File

One use of version control is to maintain multiple “current” versions of a file. For example, you might have different versions of a program in which you are gradually adding various unfinished new features. Each such independent line of development is called a branch. VC allows you to create branches, switch between different branches, and merge changes from one branch to another. Please note, however, that branches are only supported for RCS at the moment.

A file’s main line of development is usually called the trunk. The versions on the trunk are normally numbered 1.1, 1.2, 1.3, etc. At any such version, you can start an independent branch. A branch starting at version 1.2 would have version number 1.2.1.1, and consecutive versions on this branch would have numbers 1.2.1.2, 1.2.1.3, 1.2.1.4, and so on. If there is a second branch also starting at version 1.2, it would consist of versions 1.2.2.1, 1.2.2.2, 1.2.2.3, etc.

If you omit the final component of a version number, that is called a branch number. It refers to the highest existing version on that branch—the head version of that branch. The branches in the example above have branch numbers 1.2.1 and 1.2.2.

14.7.6.1 Switching between Branches

To switch between branches, type C-u C-x C-q and specify the version number you want to select. This version is then visited unlocked (writeprotected), so you can examine it before locking it. Switching branches in this way is allowed only when the file is not locked.

You can omit the minor version number, thus giving only the branch number; this takes you to the head version on the chosen branch. If you only type ^{⟨}\underline{\overline{RET}}^{⟩}, Emacs goes to the highest version on the trunk.

After you have switched to any branch (including the main branch), you stay on it for subsequent VC commands, until you explicitly select some other branch.

14.7.6.2 Creating New Branches

To create a new branch from a head version (one that is the latest in the branch that contains it), first select that version if necessary, lock it with C-x C-q, and make whatever changes you want. Then, when you check in the changes, use C-u C-x C-q. This lets you specify the version number for the new version. You should specify a suitable branch number for a branch starting at the current version. For example, if the current version is 2.5, the branch number should be 2.5.1, 2.5.2, and so on, depending on the number of existing branches at that point.

To create a new branch at an older version (one that is no longer the head of a branch), first select that version, then lock it with C-x C-q. You’ll be asked to confirm, when you lock the old version, that you really mean to create a new branch—if you say no, you’ll be offered a chance to lock the latest version instead.

Then make your changes and type C-x C-q again to check in a new version. This automatically creates a new branch starting from the selected version. You need not specially request a new branch, because that’s the only way to add a new version at a point that is not the head of a branch.

After the branch is created, you “stay” on it. That means that subsequent check-ins create new versions on that branch. To leave the branch, you must explicitly select a different version with C-u C-x C-q. To transfer changes from one branch to another, use the merge command, described in the next section.

14.7.6.3 Merging Branches

When you have finished the changes on a certain branch, you will often want to incorporate them into the file’s main line of development (the trunk). This is not a trivial operation, because development might also have proceeded on the trunk, so that you must merge the changes into a file that has already been changed otherwise. VC allows you to do this (and other things) with the vc-merge command.

C-x v m (vc-merge)

Merge changes into the work file.

C-x v m (vc-merge) takes a set of changes and merges it into the current version of the work file. It first asks you for a branch number or a pair of version numbers in the minibuffer. Then it finds the changes from that branch, or between the two versions you specified, and merges them into the current version of the current file.

As an example, suppose that you have finished a certain feature on branch 1.3.1. In the meantime, development on the trunk has proceeded to version 1.5. To merge the changes from the branch to the trunk, first go to the head version of the trunk, by typing C-u C-x C-q RET. Version 1.5 is now current. If locking is used for the file, type C-x C-q to lock version 1.5 so that you can change it. Next, type C-x v m 1.3.1 RET. This takes the entire set of changes on branch 1.3.1 (relative to version 1.3, where the branch started, up to the last version on the branch) and merges it into the current version of the work file. You can now check in the changed file, thus creating version 1.6 containing the changes from the branch.

It is possible to do further editing after merging the branch, before the next check-in. But it is usually wiser to check in the merged version, then lock it and make the further changes. This will keep a better record of the history of changes.

When you merge changes into a file that has itself been modified, the changes might overlap. We call this situation a conflict, and reconciling the conflicting changes is called resolving a conflict.

Whenever conflicts occur during merging, VC detects them, tells you about them in the echo area, and asks whether you want help in merging. If you say yes, it starts an Ediff session (see section “Ediff” in The Ediff Manual).

If you say no, the conflicting changes are both inserted into the file, surrounded by conflict markers. The example below shows how a conflict region looks; the file is called ‘name’ and the current master file version with user B’s changes in it is 1.11.

     <<<<<<< name

User A’s version

     =======

User B’s version

     >>>>>>> 1.11

Then you can resolve the conflicts by editing the file manually. Or you can type M-x vc-resolve-conflicts after visiting the file. This starts an Ediff session, as described above.

14.7.6.4 Multi-User Branching

It is often useful for multiple developers to work simultaneously on different branches of a file. CVS allows this by default; for RCS, it is possible if you create multiple source directories. Each source directory should have a link named ‘RCS’ which points to a common directory of RCS master files. Then each source directory can have its own choice of selected versions, but all share the same common RCS records.

This technique works reliably and automatically, provided that the source files contain RCS version headers (see Section 14.7.8.3 [Version Headers], page 132). The headers enable Emacs to be sure, at all times, which version number is present in the work file.

If the files do not have version headers, you must instead tell Emacs explicitly in each session which branch you are working on. To do this, first find the file, then type C-u C-x C-q and specify the correct branch number. This ensures that Emacs knows which branch it is using during this particular editing session.

14.7.7 Snapshots

A snapshot is a named set of file versions (one for each registered file) that you can treat as a unit. One important kind of snapshot is a release, a (theoretically) stable version of the system that is ready for distribution to users.

14.7.7.1 Making and Using Snapshots

There are two basic commands for snapshots; one makes a snapshot with a given name, the other retrieves a named snapshot.

C-x v s name ^{⟨}\underline{\overline{RET}}^{⟩} Define the last saved versions of every registered file in or under the current directory as a snapshot named name (vc-create-

          snapshot).

C-x v r name ^{⟨}\underline{\overline{RET}}^{⟩} For all registered files at or below the current directory level, select whatever versions correspond to the snapshot name (vc-

          retrieve-snapshot).

This command reports an error if any files are locked at or below the current directory, without changing anything; this is to avoid overwriting work in progress.

A snapshot uses a very small amount of resources—just enough to record the list of file names and which version belongs to the snapshot. Thus, you need not hesitate to create snapshots whenever they are useful.

You can give a snapshot name as an argument to C-x v = or C-x v ~ (see Section 14.7.4 [Old Versions], page 122). Thus, you can use it to compare a snapshot against the current files, or two snapshots against each other, or a snapshot against a named version.

14.7.7.2 Snapshot Caveats

VC’s snapshot facilities are modeled on RCS’s named-configuration support. They use RCS’s native facilities for this, so under VC snapshots made using RCS are visible even when you bypass VC.

For SCCS, VC implements snapshots itself. The files it uses contain name/file/version-number triples. These snapshots are visible only through VC.

A snapshot is a set of checked-in versions. So make sure that all the files are checked in and not locked when you make a snapshot.

File renaming and deletion can create some difficulties with snapshots. This is not a VC-specific problem, but a general design issue in version control systems that no one has solved very well yet.

If you rename a registered file, you need to rename its master along with it (the command vc-rename-file does this automatically). If you are using SCCS, you must also update the records of the snapshot, to mention the file by its new name (vc-rename-file does this, too). An old snapshot that refers to a master file that no longer exists under the recorded name is invalid; VC can no longer retrieve it. It would be beyond the scope of this manual to explain enough about RCS and SCCS to explain how to update the snapshots by hand.

Using vc-rename-file makes the snapshot remain valid for retrieval, but it does not solve all problems. For example, some of the files in the program probably refer to others by name. At the very least, the makefile probably mentions the file that you renamed. If you retrieve an old snapshot, the renamed file is retrieved under its new name, which is not the name that the makefile expects. So the program won’t really work as retrieved.

14.7.8 Miscellaneous Commands and Features of VC

This section explains the less-frequently-used features of VC.

14.7.8.1 Change Logs and VC

If you use RCS or CVS for a program and also maintain a change log file for it (see Section 22.12 [Change Log], page 242), you can generate change log entries automatically from the version control log entries:

C-x v a Visit the current directory’s change log file and, for registered files in that directory, create new entries for versions checked in since the most recent entry in the change log file. (vc-update-

          change-log).

This command works with RCS or CVS only, not with SCCS.

C-u C-x v a

As above, but only find entries for the current buffer’s file.

M-1 C-x v a

As above, but find entries for all the currently visited files that are maintained with version control. This works only with RCS, and it puts all entries in the log for the default directory, which may not be appropriate.

For example, suppose the first line of ‘ChangeLog’ is dated 10 April 1992, and that the only check-in since then was by Nathaniel Bowditch to

‘rcs2log’ on 8 May 1992 with log text ‘Ignore log messages that start

with `#'.’. Then C-x v a visits ‘ChangeLog’ and inserts text like this:

     Fri May  8 21:45:00 1992  Nathaniel Bowditch  <nat@apn.org>
             * rcs2log: Ignore log messages that start with `#'.

You can then edit the new change log entry further as you wish.

Normally, the log entry for file ‘foo’ is displayed as ‘* foo: text of log entry’. The ‘:’ after ‘foo’ is omitted if the text of the log entry starts with ‘(functionname): ’. For example, if the log entry for

‘vc.el’ is ‘(vc-do-command): Check call-process status.’, then the text

in ‘ChangeLog’ looks like this:

     Wed May  6 10:53:00 1992  Nathaniel Bowditch  <nat@apn.org>
             * vc.el (vc-do-command): Check call-process status.

When C-x v a adds several change log entries at once, it groups related log entries together if they all are checked in by the same author at nearly the same time. If the log entries for several such files all have the same text, it coalesces them into a single entry. For example, suppose the most recent check-ins have the following log entries:

• For ‘vc.texinfo’: ‘Fix expansion typos.’
• For ‘vc.el’: ‘Don't call expand-file-name.’
• For ‘vc-hooks.el’: ‘Don't call expand-file-name.’

They appear like this in ‘ChangeLog’:

     Wed Apr  1 08:57:59 1992  Nathaniel Bowditch  <nat@apn.org>
             * vc.texinfo: Fix expansion typos.
             * vc.el, vc-hooks.el: Don't call expand-file-name.

Normally, C-x v a separates log entries by a blank line, but you can mark several related log entries to be clumped together (without an intervening blank line) by starting the text of each related log entry with a label of the form ‘{clumpname} ’. The label itself is not copied to ‘ChangeLog’. For example, suppose the log entries are:

• For ‘vc.texinfo’: ‘{expand} Fix expansion typos.’
• For ‘vc.el’: ‘{expand} Don't call expand-file-name.’
• For ‘vc-hooks.el’: ‘{expand} Don't call expand-file-name.’

Then the text in ‘ChangeLog’ looks like this:

     Wed Apr  1 08:57:59 1992  Nathaniel Bowditch  <nat@apn.org>
             * vc.texinfo: Fix expansion typos.
             * vc.el, vc-hooks.el: Don't call expand-file-name.

A log entry whose text begins with ‘#’ is not copied to ‘ChangeLog’. For example, if you merely fix some misspellings in comments, you can log the change with an entry beginning with ‘#’ to avoid putting such trivia into

‘ChangeLog’.

14.7.8.2 Renaming VC Work Files and Master Files

When you rename a registered file, you must also rename its master file correspondingly to get proper results. Use vc-rename-file to rename the source file as you specify, and rename its master file accordingly. It also updates any snapshots (see Section 14.7.7 [Snapshots], page 129) that mention the file, so that they use the new name; despite this, the snapshot thus modified may not completely work (see Section 14.7.7.2 [Snapshot Caveats], page 130).

You cannot use vc-rename-file on a file that is locked by someone else.

14.7.8.3 Inserting Version Control Headers

Sometimes it is convenient to put version identification strings directly into working files. Certain special strings called version headers are replaced in each successive version by the number of that version.

If you are using RCS, and version headers are present in your working files, Emacs can use them to determine the current version and the locking state of the files. This is more reliable than referring to the master files, which is done when there are no version headers. Note that in a multi-branch environment, version headers are necessary to make VC behave correctly (see Section 14.7.6.4 [Multi-User Branching], page 129).

Searching for version headers is controlled by the variable vc-consult-headers. If it is non-nil, Emacs searches for headers to determine the version number you are editing. Setting it to nil disables this feature.

You can use the C-x v h command (vc-insert-headers) to insert a suitable header string.

C-x v h Insert headers in a file for use with your version-control system.

The default header string is ‘$Id$’ for RCS and ‘%W%’ for SCCS. You can specify other headers to insert by setting the variable vc-header-alist. Its value is a list of elements of the form (program . string) where program is RCS or SCCS and string is the string to use.

Instead of a single string, you can specify a list of strings; then each string in the list is inserted as a separate header on a line of its own.

It is often necessary to use “superfluous” backslashes when writing the strings that you put in this variable. This is to prevent the string in the constant from being interpreted as a header itself if the Emacs Lisp file containing it is maintained with version control.

Each header is inserted surrounded by tabs, inside comment delimiters, on a new line at the start of the buffer. Normally the ordinary comment start and comment end strings of the current mode are used, but for certain modes, there are special comment delimiters for this purpose; the variable vc-comment-alist specifies them. Each element of this list has the form (mode starter ender).

The variable vc-static-header-alist specifies further strings to add based on the name of the buffer. Its value should be a list of elements of the form (regexp . format). Whenever regexp matches the buffer name, format is inserted as part of the header. A header line is inserted for each element that matches the buffer name, and for each string specified by vcheader-alist. The header line is made by processing the string from vcheader-alist with the format taken from the element. The default value

for vc-static-header-alist is as follows:
    (("\\.c$" .
      "\n#ifndef lint\nstatic char vcid[] = \"\%s\";\n\
    #endif /* lint */\n"))

It specifies insertion of text of this form:

    #ifndef lint
    static char vcid[] = "string";
    #endif /* lint */

Note that the text above starts with a blank line.

If you use more than one version header in a file, put them close together in the file. The mechanism in revert-buffer that preserves markers may not handle markers positioned between two version headers.

14.7.9 Customizing VC

There are many ways of customizing VC. The options you can set fall into four categories, described in the following sections.

14.7.9.1 Options for VC Backends

You can tell RCS and CVS whether to use locking for a file or not (see Section 14.7.1.2 [VC Concepts], page 119, for a description of locking). VC automatically recognizes what you have chosen, and behaves accordingly.

For RCS, the default is to use locking, but there is a mode called nonstrict locking in which you can check-in changes without locking the file first. Use ‘rcs -U’ to switch to non-strict locking for a particular file, see the ‘rcs’ manpage for details.

Under CVS, the default is not to use locking; anyone can change a work file at any time. However, there are ways to restrict this, resulting in behavior that resembles locking.

For one thing, you can set the CVSREAD environment variable to an arbitrary value. If this variable is defined, CVS makes your work files read-only by default. In Emacs, you must type C-x C-q to make the file writeable, so that editing works in fact similar as if locking was used. Note however, that no actual locking is performed, so several users can make their files writeable at the same time. When setting CVSREAD for the first time, make sure to check out all your modules anew, so that the file protections are set correctly.

Another way to achieve something similar to locking is to use the watch feature of CVS. If a file is being watched, CVS makes it read-only by default, and you must also use C-x C-q in Emacs to make it writable. VC calls cvs edit to make the file writeable, and CVS takes care to notify other developers of the fact that you intend to change the file. See the CVS documentation for details on using the watch feature.

You can turn off use of VC for CVS-managed files by setting the variable vc-handle-cvs to nil. If you do this, Emacs treats these files as if they were not registered, and the VC commands are not available for them. You must do all CVS operations manually.

14.7.9.2 VC Workfile Handling

Emacs normally does not save backup files for source files that are maintained with version control. If you want to make backup files even for files that use version control, set the variable vc-make-backup-files to a nonnil value.

Normally the work file exists all the time, whether it is locked or not. If you set vc-keep-workfiles to nil, then checking in a new version with C-x C-q deletes the work file; but any attempt to visit the file with Emacs creates it again. (With CVS, work files are always kept.)

Editing a version-controlled file through a symbolic link can be dangerous. It bypasses the version control system—you can edit the file without locking it, and fail to check your changes in. Also, your changes might overwrite those of another user. To protect against this, VC checks each symbolic link that you visit, to see if it points to a file under version control.

The variable vc-follow-symlinks controls what to do when a symbolic link points to a version-controlled file. If it is nil, VC only displays a warning message. If it is t, VC automatically follows the link, and visits the real file instead, telling you about this in the echo area. If the value is ask (the default), VC asks you each time whether to follow the link.

14.7.9.3 VC Status Retrieval

When deducing the locked/unlocked state of a file, VC first looks for an RCS version header string in the file (see Section 14.7.8.3 [Version Headers], page 132). If there is no header string, or if you are using SCCS, VC normally looks at the file permissions of the work file; this is fast. But there might be situations when the file permissions cannot be trusted. In this case the master file has to be consulted, which is rather expensive. Also the master file can only tell you if there’s any lock on the file, but not whether your work file really contains that locked version.

You can tell VC not to use version headers to determine lock status by setting vc-consult-headers to nil. VC then always uses the file permissions (if it can trust them), or else checks the master file.

You can specify the criterion for whether to trust the file permissions by setting the variable vc-mistrust-permissions. Its value can be t (always mistrust the file permissions and check the master file), nil (always trust the file permissions), or a function of one argument which makes the decision. The argument is the directory name of the ‘RCS’, ‘CVS’ or ‘SCCS’ subdirectory. A non-nil value from the function says to mistrust the file permissions. If you find that the file permissions of work files are changed erroneously, set vc-mistrust-permissions to t. Then VC always checks the master file to determine the file’s status.

14.7.9.4 VC Command Execution

   If vc-suppress-confirm is non-nil, then C-x C-q and C-x v i can save

the current buffer without asking, and C-x v u also operates without asking for confirmation. (This variable does not affect C-x v c; that operation is so drastic that it should always ask for confirmation.)

VC mode does much of its work by running the shell commands for RCS, CVS and SCCS. If vc-command-messages is non-nil, VC displays messages to indicate which shell commands it runs, and additional messages when the commands finish.

You can specify additional directories to search for version control programs by setting the variable vc-path. These directories are searched before the usual search path. But the proper files are usually found automatically.

14.8 File Directories

The file system groups files into directories. A directory listing is a list of all the files in a directory. Emacs provides commands to create and delete directories, and to make directory listings in brief format (file names only) and verbose format (sizes, dates, and authors included). There is also a directory browser called Dired; see Chapter 28 [Dired], page 321.

C-x C-d dir-or-pattern ^{⟨}\underline{\overline{RET}}^{⟩} Display a brief directory listing (list-directory).

C-u C-x C-d dir-or-pattern ^{⟨}\underline{\overline{RET}}^{⟩} Display a verbose directory listing.

M-x make-directory ^{⟨}\underline{\overline{RET}}^{⟩} dirname ^{⟨}\underline{\overline{RET}}^{⟩}

Create a new directory named dirname.

M-x delete-directory ^{⟨}\underline{\overline{RET}}^{⟩} dirname ^{⟨}\underline{\overline{RET}}^{⟩}

Delete the directory named dirname. It must be empty, or you get an error.

The command to display a directory listing is C-x C-d (list-directory). It reads using the minibuffer a file name which is either a directory to be listed or a wildcard-containing pattern for the files to be listed. For example,

    C-x C-d /u2/emacs/etc ^{⟨}\underline{\overline{RET}}^{⟩}

lists all the files in directory ‘/u2/emacs/etc’. Here is an example of specifying a file name pattern:

    C-x C-d /u2/emacs/src/*.c ^{⟨}\underline{\overline{RET}}^{⟩}

Normally, C-x C-d prints a brief directory listing containing just file names. A numeric argument (regardless of value) tells it to make a verbose listing including sizes, dates, and authors (like ‘ls -l’).

The text of a directory listing is obtained by running ls in an inferior process. Two Emacs variables control the switches passed to ls: listdirectory-brief-switches is a string giving the switches to use in brief

listings ("-CF" by default), and list-directory-verbose-switches is a

string giving the switches to use in a verbose listing ("-l" by default).

14.9 Comparing Files

The command M-x diff compares two files, displaying the differences in an Emacs buffer named ‘*Diff*’. It works by running the diff program, using options taken from the variable diff-switches, whose value should be a string.

The buffer ‘*Diff*’ has Compilation mode as its major mode, so you can use C-x ` to visit successive changed locations in the two source files. You can also move to a particular hunk of changes and type ^{⟨}\underline{\overline{RET}}^{⟩} or C-c C-c, or click Mouse-2 on it, to move to the corresponding source location. You can also use the other special commands of Compilation mode: ^{⟨}\underline{\overline{SPC}}^{⟩} and ^{⟨}\underline{\overline{DEL}}^{⟩} for scrolling, and M-p and M-n for cursor motion. See Section 23.1 [Compilation], page 269.

The command M-x diff-backup compares a specified file with its most recent backup. If you specify the name of a backup file, diff-backup compares it with the source file that it is a backup of.

The command M-x compare-windows compares the text in the current window with that in the next window. Comparison starts at point in each window, and each starting position is pushed on the mark ring in its respective buffer. Then point moves forward in each window, a character at a time, until a mismatch between the two windows is reached. Then the command is finished. For more information about windows in Emacs, Chapter 16 [Windows], page 149.

With a numeric argument, compare-windows ignores changes in whitespace. If the variable compare-ignore-case is non-nil, it ignores differences in case as well.

See also Section 22.14 [Emerge], page 251, for convenient facilities for merging two similar files.

14.10 Miscellaneous File Operations

Emacs has commands for performing many other operations on files. All operate on one file; they do not accept wildcard file names.

M-x view-file allows you to scan or read a file by sequential screenfuls. It reads a file name argument using the minibuffer. After reading the file into an Emacs buffer, view-file displays the beginning. You can then type ⟨\underline{\overline{SPC}}^{⟩} to scroll forward one windowful, or ^{⟨}\underline{\overline{DEL}}^{⟩} to scroll backward. Various other commands are provided for moving around in the file, but none for changing it; type ? while viewing for a list of them. They are mostly the same as normal Emacs cursor motion commands. To exit from viewing, type q. The commands for viewing are defined by a special major mode called View mode.

A related command, M-x view-buffer, views a buffer already present in Emacs. See Section 15.3 [Misc Buffer], page 143.

M-x insert-file inserts a copy of the contents of the specified file into the current buffer at point, leaving point unchanged before the contents and the mark after them.

M-x write-region is the inverse of M-x insert-file; it copies the contents of the region into the specified file. M-x append-to-file adds the text of the region to the end of the specified file. See Section 9.3 [Accumulating Text], page 68.

M-x delete-file deletes the specified file, like the rm command in the shell. If you are deleting many files in one directory, it may be more convenient to use Dired (see Chapter 28 [Dired], page 321).

M-x rename-file reads two file names old and new using the minibuffer, then renames file old as new. If a file named new already exists, you must confirm with yes or renaming is not done; this is because renaming causes the old meaning of the name new to be lost. If old and new are on different file systems, the file old is copied and deleted.

The similar command M-x add-name-to-file is used to add an additional name to an existing file without removing its old name. The new name must belong on the same file system that the file is on.

M-x copy-file reads the file old and writes a new file named new with the same contents. Confirmation is required if a file named new already exists, because copying has the consequence of overwriting the old contents of the file new.

M-x make-symbolic-link reads two file names target and linkname, then creates a symbolic link named linkname and pointing at target. The effect is that future attempts to open file linkname will refer to whatever file is named target at the time the opening is done, or will get an error if the name target is not in use at that time. This command does not expand the argument target, so that it allows you to specify a relative name as the target of the link.

Confirmation is required when creating the link if linkname is in use. Note that not all systems support symbolic links.

14.11 Accessing Compressed Files

Emacs comes with a library that can automatically uncompress compressed files when you visit them, and automatically recompress them if you alter them and save them. To enable this feature, type the command M-x

auto-compression-mode.

When automatic compression (which implies automatic uncompression as well) is enabled, Emacs recognizes compressed files by their file names. File names ending in ‘.gz’ indicate a file compressed with gzip. Other endings indicate other compression programs.

Automatic uncompression and compression apply to all the operations in which Emacs uses the contents of a file. This includes visiting it, saving it, inserting its contents into a buffer, loading it, and byte compiling it.

14.12 Remote Files

You can refer to files on other machines using a special file name syntax: /host:filename /user@host:filename

When you do this, Emacs uses the FTP program to read and write files on the specified host. It logs in through FTP using your user name or the name user. It may ask you for a password from time to time; this is used for logging in on host.

Normally, if you do not specify a user name in a remote file name, that means to use your own user name. But if you set the variable ange-ftp-default-user to a string, that string is used instead. (The Emacs package that implements FTP file access is called ange-ftp.)

You can entirely turn off the FTP file name feature by setting the variable

file-name-handler-alist to nil.

14.13 Quoted File Names

You can quote an absolute file name to prevent special characters and syntax in it from having their special effects. The way to do this is to add ‘/:’ at the beginning.

For example, you can quote a local file name which appears remote, to prevent it from being treated as a remote file name. Thus, if you have a directory named ‘/foo:’ and a file named ‘bar’ in it, you can refer to that file in Emacs as ‘/:/foo:/bar’.

‘/:’ can also prevent ‘~’ from being treated as a special character for a user’s home directory. For example, ‘/:/tmp/~hack’ refers to a file whose name is ‘~hack’ in directory ‘/tmp’.

Likewise, quoting with ‘/:’ is one way to enter in the minibuffer a file name that contains ‘$’. However, the ‘/:’ must be at the beginning of the buffer in order to quote ‘$’.

140 GNU Emacs Manual

15 Using Multiple Buffers

The text you are editing in Emacs resides in an object called a buffer. Each time you visit a file, a buffer is created to hold the file’s text. Each time you invoke Dired, a buffer is created to hold the directory listing. If you send a message with C-x m, a buffer named ‘*mail*’ is used to hold the text of the message. When you ask for a command’s documentation, that appears in a buffer called ‘*Help*’.

At any time, one and only one buffer is selected. It is also called the current buffer. Often we say that a command operates on “the buffer” as if there were only one; but really this means that the command operates on the selected buffer (most commands do).

When Emacs has multiple windows, each window has a chosen buffer which is displayed there, but at any time only one of the windows is selected and its chosen buffer is the selected buffer. Each window’s mode line displays the name of the buffer that the window is displaying (see Chapter 16 [Windows], page 149).

Each buffer has a name, which can be of any length, and you can select any buffer by giving its name. Most buffers are made by visiting files, and their names are derived from the files’ names. But you can also create an empty buffer with any name you want. A newly started Emacs has a buffer named ‘*scratch*’ which can be used for evaluating Lisp expressions in Emacs. The distinction between upper and lower case matters in buffer names.

Each buffer records individually what file it is visiting, whether it is modified, and what major mode and minor modes are in effect in it (see Chapter 19 [Major Modes], page 185). Any Emacs variable can be made local to a particular buffer, meaning its value in that buffer can be different from the value in other buffers. See Section 31.2.4 [Locals], page 393.

15.1 Creating and Selecting Buffers

C-x b buffer ^{⟨}\underline{\overline{RET}}^{⟩} Select or create a buffer named buffer (switch-to-buffer).

C-x 4 b buffer ^{⟨}\underline{\overline{RET}}^{⟩} Similar, but select buffer in another window (switch-to-

          buffer-other-window).

C-x 5 b buffer ^{⟨}\underline{\overline{RET}}^{⟩} Similar, but select buffer in a separate frame (switch-to-

          buffer-other-frame).

To select the buffer named bufname, type C-x b bufname ^{⟨}\underline{\overline{RET}}^{⟩}. This runs the command switch-to-buffer with argument bufname. You can use completion on an abbreviation for the buffer name you want (see Section 5.3 [Completion], page 41). An empty argument to C-x b specifies the most recently selected buffer that is not displayed in any window.

Most buffers are created by visiting files, or by Emacs commands that want to display some text, but you can also create a buffer explicitly by typing C-x b bufname ^{⟨}\underline{\overline{RET}}^{⟩}. This makes a new, empty buffer that is not visiting any file, and selects it for editing. Such buffers are used for making notes to yourself. If you try to save one, you are asked for the file name to use. The new buffer’s major mode is determined by the value of defaultmajor-mode (see Chapter 19 [Major Modes], page 185).

Note that C-x C-f, and any other command for visiting a file, can also be used to switch to an existing file-visiting buffer. See Section 14.2 [Visiting], page 106.

Emacs uses buffer names that start with a space for internal purposes. It treats these buffers specially in minor ways—for example, by default they do not record undo information. It is best to avoid using such buffer names yourself.

15.2 Listing Existing Buffers

C-x C-b List the existing buffers (list-buffers).

To display a list of all the buffers that exist, type C-x C-b. Each line in the list shows one buffer’s name, major mode and visited file. The buffers are listed in the order that they were current; the buffers that were current most recently come first.

‘*’ at the beginning of a line indicates the buffer is “modified.” If several buffers are modified, it may be time to save some with C-x s (see Section 14.3 [Saving], page 109). ‘%’ indicates a read-only buffer. ‘.’ marks the selected buffer. Here is an example of a buffer list:

      MR Buffer         Size  Mode           File
      -- ------         ----  ----           ----
     .*  emacs.tex      383402 Texinfo       /u2/emacs/man/emacs.tex
         *Help*         1287  Fundamental
         files.el       23076 Emacs-Lisp     /u2/emacs/lisp/files.el
       % RMAIL          64042 RMAIL          /u/rms/RMAIL
      *% man            747   Dired          /u2/emacs/man/
         net.emacs      343885 Fundamental   /u/rms/net.emacs
         fileio.c       27691 C              /u2/emacs/src/fileio.c
         NEWS           67340 Text           /u2/emacs/etc/NEWS
         *scratch*    0  Lisp Interaction

Note that the buffer ‘*Help*’ was made by a help request; it is not visiting any file. The buffer man was made by Dired on the directory

‘/u2/emacs/man/’.

15.3 Miscellaneous Buffer Operations

C-x C-q Toggle read-only status of buffer (vc-toggle-read-only).

M-x rename-buffer ^{⟨}\underline{\overline{RET}}^{⟩} name ^{⟨}\underline{\overline{RET}}^{⟩}

Change the name of the current buffer.

M-x rename-uniquely

Rename the current buffer by adding ‘<number>’ to the end.

M-x view-buffer ^{⟨}\underline{\overline{RET}}^{⟩} buffer ^{⟨}\underline{\overline{RET}}^{⟩} Scroll through buffer buffer.

A buffer can be read-only, which means that commands to change its contents are not allowed. The mode line indicates read-only buffers with ‘%%’ or ‘%*’ near the left margin. Read-only buffers are usually made by subsystems such as Dired and Rmail that have special commands to operate on the text; also by visiting a file whose access control says you cannot write it.

If you wish to make changes in a read-only buffer, use the command C-x C-q (vc-toggle-read-only). It makes a read-only buffer writable, and makes a writable buffer read-only. In most cases, this works by setting the variable buffer-read-only, which has a local value in each buffer and makes the buffer read-only if its value is non-nil. If the file is maintained with version control, C-x C-q works through the version control system to change the read-only status of the file as well as the buffer. See Section 14.7 [Version Control], page 118.

M-x rename-buffer changes the name of the current buffer. Specify the new name as a minibuffer argument. There is no default. If you specify a name that is in use for some other buffer, an error happens and no renaming is done.

M-x rename-uniquely renames the current buffer to a similar name with a numeric suffix added to make it both different and unique. This command does not need an argument. It is useful for creating multiple shell buffers: if you rename the ‘*Shell*’ buffer, then do M-x shell again, it makes a new shell buffer named ‘*Shell*’; meanwhile, the old shell buffer continues to exist under its new name. This method is also good for mail buffers, compilation buffers, and most Emacs features that create special buffers with particular names.

M-x view-buffer is much like M-x view-file (see Section 14.10 [Misc File Ops], page 137) except that it examines an already existing Emacs buffer. View mode provides commands for scrolling through the buffer conveniently but not for changing it. When you exit View mode with q, that switches back to the buffer (and the position) which was previously displayed in the window. Alternatively, if you exit View mode with e, the buffer and the value of point that resulted from your perusal remain in effect.

   The commands M-x append-to-buffer and M-x insert-buffer can be

used to copy text from one buffer to another. See Section 9.3 [Accumulating Text], page 68.

15.4 Killing Buffers

If you continue an Emacs session for a while, you may accumulate a large number of buffers. You may then find it convenient to kill the buffers you no longer need. On most operating systems, killing a buffer releases its space back to the operating system so that other programs can use it. Here are some commands for killing buffers:

C-x k bufname ^{⟨}\underline{\overline{RET}}^{⟩} Kill buffer bufname (kill-buffer).

M-x kill-some-buffers

Offer to kill each buffer, one by one.

C-x k (kill-buffer) kills one buffer, whose name you specify in the minibuffer. The default, used if you type just ^{⟨}\underline{\overline{RET}}^{⟩} in the minibuffer, is to kill the current buffer. If you kill the current buffer, another buffer is selected; one that has been selected recently but does not appear in any window now. If you ask to kill a file-visiting buffer that is modified (has unsaved editing), then you must confirm with yes before the buffer is killed.

The command M-x kill-some-buffers asks about each buffer, one by one. An answer of y means to kill the buffer. Killing the current buffer or a buffer containing unsaved changes selects a new buffer or asks for confirmation just like kill-buffer.

The buffer menu feature (see Section 15.5 [Several Buffers], page 145) is also convenient for killing various buffers.

If you want to do something special every time a buffer is killed, you can add hook functions to the hook kill-buffer-hook (see Section 31.2.3 [Hooks], page 392).

If you run one Emacs session for a period of days, as many people do, it can fill up with buffers that you used several days ago. The command M-x clean-buffer-list is a convenient way to purge them; it kills all the unmodified buffers that you have not used for a long time. An ordinary buffer is killed if it has not been displayed for three days; however, you can specify certain buffers that should never be killed automatically, and others that should be killed if they have been unused for a mere hour.

You can also have this buffer purging done for you, every day at midnight, by enabling Midnight mode. Midnight mode operates each day at midnight; at that time, it runs clean-buffer-list, or whichever functions you have placed in the normal hook midnight-hook (see Section 31.2.3 [Hooks], page 392).

To enable Midnight mode, use the Customization buffer to set the variable midnight-mode to t. See Section 31.2.2 [Easy Customization], page 386.

15.5 Operating on Several Buffers

The buffer-menu facility is like a “Dired for buffers”; it allows you to request operations on various Emacs buffers by editing an Emacs buffer containing a list of them. You can save buffers, kill them (here called deleting them, for consistency with Dired), or display them.

M-x buffer-menu

Begin editing a buffer listing all Emacs buffers.

The command buffer-menu writes a list of all Emacs buffers into the buffer ‘*Buffer List*’, and selects that buffer in Buffer Menu mode. The buffer is read-only, and can be changed only through the special commands described in this section. The usual Emacs cursor motion commands can be used in the ‘*Buffer List*’ buffer. The following commands apply to the buffer described on the current line.

d Request to delete (kill) the buffer, then move down. The request shows as a ‘D’ on the line, before the buffer name. Requested deletions take place when you type the x command.

C-d Like d but move up afterwards instead of down.

s Request to save the buffer. The request shows as an ‘S’ on the line. Requested saves take place when you type the x command. You may request both saving and deletion for the same buffer.

x Perform previously requested deletions and saves.

u Remove any request made for the current line, and move down.

⟨\underline{\overline{DEL}}^{⟩} Move to previous line and remove any request made for that line.

The d, C-d, s and u commands to add or remove flags also move down (or up) one line. They accept a numeric argument as a repeat count.

These commands operate immediately on the buffer listed on the current line:

~ Mark the buffer “unmodified.” The command ~ does this immediately when you type it.

% Toggle the buffer’s read-only flag. The command % does this immediately when you type it.

t Visit the buffer as a tags table. See Section 22.13.3 [Select Tags Table], page 247.

There are also commands to select another buffer or buffers:

q Quit the buffer menu—immediately display the most recent formerly visible buffer in its place.

⟨\underline{\overline{RET}}⟩

f Immediately select this line’s buffer in place of the ‘*Buffer List*’ buffer.

o Immediately select this line’s buffer in another window as if by C-x 4 b, leaving ‘*Buffer List*’ visible.

C-o Immediately display this line’s buffer in another window, but don’t select the window.

1 Immediately select this line’s buffer in a full-screen window.

2 Immediately set up two windows, with this line’s buffer in one, and the previously selected buffer (aside from the buffer ‘*Buffer List*’) in the other.

b Bury the buffer listed on this line.

m Mark this line’s buffer to be displayed in another window if you exit with the v command. The request shows as a ‘>’ at the beginning of the line. (A single buffer may not have both a delete request and a display request.)

v Immediately select this line’s buffer, and also display in other windows any buffers previously marked with the m command. If you have not marked any buffers, this command is equivalent to

          1.

All that buffer-menu does directly is create and switch to a suitable buffer, and turn on Buffer Menu mode. Everything else described above is implemented by the special commands provided in Buffer Menu mode. One consequence of this is that you can switch from the ‘*Buffer List*’ buffer to another Emacs buffer, and edit there. You can reselect the ‘*Buffer List*’ buffer later, to perform the operations already requested, or you can kill it, or pay no further attention to it.

The only difference between buffer-menu and list-buffers is that buffer-menu switches to the ‘*Buffer List*’ buffer in the selected window; list-buffers displays it in another window. If you run list-buffers (that is, type C-x C-b) and select the buffer list manually, you can use all of the commands described here.

The buffer ‘*Buffer List*’ is not updated automatically when buffers are created and killed; its contents are just text. If you have created, deleted or renamed buffers, the way to update ‘*Buffer List*’ to show what you have done is to type g (revert-buffer) or repeat the buffer-menu command.

15.6 Indirect Buffers

An indirect buffer shares the text of some other buffer, which is called the base buffer of the indirect buffer. In some ways it is the analogue, for buffers, of a symbolic link between files.

M-x make-indirect-buffer base-buffer ^{⟨}\underline{\overline{RET}}^{⟩} indirect-name ^{⟨}\underline{\overline{RET}}^{⟩} Create an indirect buffer named indirect-name whose base buffer is base-buffer.

The text of the indirect buffer is always identical to the text of its base buffer; changes made by editing either one are visible immediately in the other. But in all other respects, the indirect buffer and its base buffer are completely separate. They have different names, different values of point, different narrowing, different markers, different major modes, and different local variables.

An indirect buffer cannot visit a file, but its base buffer can. If you try to save the indirect buffer, that actually works by saving the base buffer. Killing the base buffer effectively kills the indirect buffer, but killing an indirect buffer has no effect on its base buffer.

One way to use indirect buffers is to display multiple views of an outline. See Section 21.8.4 [Outline Views], page 208.

148 GNU Emacs Manual

16 Multiple Windows

Emacs can split a frame into two or many windows. Multiple windows can display parts of different buffers, or different parts of one buffer. Multiple frames always imply multiple windows, because each frame has its own set of windows. Each window belongs to one and only one frame.

16.1 Concepts of Emacs Windows

Each Emacs window displays one Emacs buffer at any time. A single buffer may appear in more than one window; if it does, any changes in its text are displayed in all the windows where it appears. But the windows showing the same buffer can show different parts of it, because each window has its own value of point.

At any time, one of the windows is the selected window; the buffer this window is displaying is the current buffer. The terminal’s cursor shows the location of point in this window. Each other window has a location of point as well, but since the terminal has only one cursor there is no way to show where those locations are. When multiple frames are visible in X Windows, each frame has a cursor which appears in the frame’s selected window. The cursor in the selected frame is solid; the cursor in other frames is a hollow box.

Commands to move point affect the value of point for the selected Emacs window only. They do not change the value of point in any other Emacs window, even one showing the same buffer. The same is true for commands such as C-x b to change the selected buffer in the selected window; they do not affect other windows at all. However, there are other commands such as C-x 4 b that select a different window and switch buffers in it. Also, all commands that display information in a window, including (for example) C-h

f (describe-function) and C-x C-b (list-buffers), work by switching

buffers in a nonselected window without affecting the selected window.

When multiple windows show the same buffer, they can have different regions, because they can have different values of point. However, they all have the same value for the mark, because each buffer has only one mark position.

Each window has its own mode line, which displays the buffer name, modification status and major and minor modes of the buffer that is displayed in the window. See Section 1.3 [Mode Line], page 17, for full details on the mode line.

16.2 Splitting Windows

C-x 2 Split the selected window into two windows, one above the other

          (split-window-vertically).

C-x 3 Split the selected window into two windows positioned side by

          side (split-window-horizontally).
C-Mouse-2

In the mode line or scroll bar of a window, split that window.

The command C-x 2 (split-window-vertically) breaks the selected window into two windows, one above the other. Both windows start out displaying the same buffer, with the same value of point. By default the two windows each get half the height of the window that was split; a numeric argument specifies how many lines to give to the top window.

   C-x 3 (split-window-horizontally) breaks the selected window into

two side-by-side windows. A numeric argument specifies how many columns to give the one on the left. A line of vertical bars separates the two windows. Windows that are not the full width of the screen have mode lines, but they are truncated. On terminals where Emacs does not support highlighting, truncated mode lines sometimes do not appear in inverse video.

You can split a window horizontally or vertically by clicking C-Mouse-2 in the mode line or the scroll bar. The line of splitting goes through the place where you click: if you click on the mode line, the new scroll bar goes above the spot; if you click in the scroll bar, the mode line of the split window is side by side with your click.

When a window is less than the full width, text lines too long to fit are frequent. Continuing all those lines might be confusing. The variable

truncate-partial-width-windows can be set non-nil to force truncation

in all windows less than the full width of the screen, independent of the buffer being displayed and its value for truncate-lines. See Section 4.8 [Continuation Lines], page 35.

Horizontal scrolling is often used in side-by-side windows. See Chapter 11 [Display], page 77.

If split-window-keep-point is non-nil, the default, both of the windows resulting from C-x 2 inherit the value of point from the window that was split. This means that scrolling is inevitable. If this variable is nil, then C-x 2 tries to avoid shifting any text the screen, by putting point in each window at a position already visible in the window. It also selects whichever window contain the screen line that the cursor was previously on. Some users prefer the latter mode slow terminals.

16.3 Using Other Windows

C-x o Select another window (other-window). That is o, not zero.

C-M-v Scroll the next window (scroll-other-window).
M-x compare-windows

Find next place where the text in the selected window does not match the text in the next window.

Mouse-1 Mouse-1, in a window’s mode line, selects that window but does

          not move point in it (mouse-select-window).

To select a different window, click with Mouse-1 on its mode line. With the keyboard, you can switch windows by typing C-x o (other-window). That is an o, for ‘other’, not a zero. When there are more than two windows, this command moves through all the windows in a cyclic order, generally top to bottom and left to right. After the rightmost and bottommost window, it goes back to the one at the upper left corner. A numeric argument means to move several steps in the cyclic order of windows. A negative argument moves around the cycle in the opposite order. When the minibuffer is active, the minibuffer is the last window in the cycle; you can switch from the minibuffer window to one of the other windows, and later switch back and finish supplying the minibuffer argument that is requested. See Section 5.2 [Minibuffer Edit], page 40.

The usual scrolling commands (see Chapter 11 [Display], page 77) apply to the selected window only, but there is one command to scroll the next window. C-M-v (scroll-other-window) scrolls the window that C-x o would select. It takes arguments, positive and negative, like C-v. (In the minibuffer, C-M-v scrolls the window that contains the minibuffer help display, if any, rather than the next window in the standard cyclic order.)

The command M-x compare-windows lets you compare two files or buffers visible in two windows, by moving through them to the next mismatch. See Section 14.9 [Comparing Files], page 137, for details.

16.4 Displaying in Another Window

C-x 4 is a prefix key for commands that select another window (splitting the window if there is only one) and select a buffer in that window. Different C-x 4 commands have different ways of finding the buffer to select. C-x 4 b bufname ^{⟨}\underline{\overline{RET}}^{⟩}

Select buffer bufname in another window. This runs switch-

          to-buffer-other-window.

C-x 4 C-o bufname ^{⟨}\underline{\overline{RET}}^{⟩} Display buffer bufname in another window, but don’t select that buffer or that window. This runs display-buffer.

C-x 4 f filename ^{⟨}\underline{\overline{RET}}^{⟩} Visit file filename and select its buffer in another window. This runs find-file-other-window. See Section 14.2 [Visiting], page 106.

C-x 4 d directory ^{⟨}\underline{\overline{RET}}^{⟩} Select a Dired buffer for directory directory in another window. This runs dired-other-window. See Chapter 28 [Dired], page 321.

C-x 4 m Start composing a mail message in another window. This runs mail-other-window; its same-window analogue is C-x m (see Chapter 26 [Sending Mail], page 293).

C-x 4 . Find a tag in the current tags table, in another window. This runs find-tag-other-window, the multiple-window variant of M-. (see Section 22.13 [Tags], page 243).

C-x 4 r filename ^{⟨}\underline{\overline{RET}}^{⟩} Visit file filename read-only, and select its buffer in another win-

          dow. This runs find-file-read-only-other-window. See Sec-

tion 14.2 [Visiting], page 106.

16.5 Forcing Display in the Same Window

Certain Emacs commands switch to a specific buffer with special contents. For example, M-x shell switches to a buffer named ‘*Shell*’. By convention, all these commands are written to pop up the buffer in a separate window. But you can specify that certain of these buffers should appear in the selected window.

If you add a buffer name to the list same-window-buffer-names, the effect is that such commands display that particular buffer by switching to it in the selected window. For example, if you add the element "*grep*" to the list, the grep command will display its output buffer in the selected window.

The default value of same-window-buffer-names is not nil: it specifies buffer names ‘*info*’, ‘*mail*’ and ‘*shell*’ (as well as others used by more obscure Emacs packages). This is why M-x shell normally switches to the ‘*shell*’ buffer in the selected window. If you delete this element

from the value of same-window-buffer-names, the behavior of M-x shell

will change—it will pop up the buffer in another window instead.

You can specify these buffers more generally with the variable samewindow-regexps. Set it to a list of regular expressions; then any buffer whose name matches one of those regular expressions is displayed by switching to it in the selected window. (Once again, this applies only to buffers that normally get displayed for you in a separate window.) The default value of this variable specifies Telnet and rlogin buffers.

An analogous feature lets you specify buffers which should be displayed in their own individual frames. See Section 17.9 [Special Buffer Frames], page 161.

16.6 Deleting and Rearranging Windows

C-x 0 Delete the selected window (delete-window). The last character in this key sequence is a zero.

C-x 1 Delete all windows in the selected frame except the selected win-

          dow (delete-other-windows).

C-x 4 0 Delete the selected window and kill the buffer that was showing in it (kill-buffer-and-window). The last character in this key sequence is a zero.

C-x ^Make selected window taller (enlarge-window).

C-x }  Make selected window wider (enlarge-window-horizontally).
C-x {  Make selected window narrower (shrink-window-horizontally).

C-x - Shrink this window if its buffer doesn’t need so many lines

          (shrink-window-if-larger-than-buffer).

C-x + Make all windows the same height (balance-windows).

Drag-Mouse-1

Dragging a window’s mode line up or down with Mouse-1 changes window heights.

Mouse-2 Mouse-2 in a window’s mode line deletes all other windows in the

          frame (mouse-delete-other-windows).

Mouse-3 Mouse-3 in a window’s mode line deletes that window (mouse-

          delete-window).

To delete a window, type C-x 0 (delete-window). (That is a zero.) The space occupied by the deleted window is given to an adjacent window (but not the minibuffer window, even if that is active at the time). Once a window is deleted, its attributes are forgotten; only restoring a window configuration can bring it back. Deleting the window has no effect on the buffer it used to display; the buffer continues to exist, and you can select it in any window with C-x b.

   C-x 4 0 (kill-buffer-and-window) is a stronger command than C-x 0;

it kills the current buffer and then deletes the selected window.

C-x 1 (delete-other-windows) is more powerful in a different way; it deletes all the windows except the selected one (and the minibuffer); the selected window expands to use the whole frame except for the echo area.

You can also delete a window by clicking on its mode line with Mouse-2, and delete all the windows in a frame except one window by clicking on that window’s mode line with Mouse-3.

The easiest way to adjust window heights is with a mouse. If you press Mouse-1 on a mode line, you can drag that mode line up or down, changing the heights of the windows above and below it.

To readjust the division of space among vertically adjacent windows, use C-x ^(enlarge-window). It makes the currently selected window get one line bigger, or as many lines as is specified with a numeric argument. With a negative argument, it makes the selected window smaller. C-x } (enlarge-window-horizontally) makes the selected window wider by the specified number of columns. C-x { (shrink-window-horizontally) makes the selected window narrower by the specified number of columns.

When you make a window bigger, the space comes from one of its neighbors. If this makes any window too small, it is deleted and its space is given to an adjacent window. The minimum size is specified by the variables

window-min-height and window-min-width.
   The command C-x - (shrink-window-if-larger-than-buffer) re-

duces the height of the selected window, if it is taller than necessary to show the whole text of the buffer it is displaying. It gives the extra lines to other windows in the frame.

You can also use C-x + (balance-windows) to even out the heights of all the windows in the selected frame.

See Section 5.2 [Minibuffer Edit], page 40, for information about the Resize-Minibuffer mode, which automatically changes the size of the minibuffer window to fit the text in the minibuffer.

17 Frames and X Windows

When using the X Window System, you can create multiple windows at the X level in a single Emacs session. Each X window that belongs to Emacs displays a frame which can contain one or several Emacs windows. A frame initially contains a single general-purpose Emacs window which you can subdivide vertically or horizontally into smaller windows. A frame normally contains its own echo area and minibuffer, but you can make frames that don’t have these—they use the echo area and minibuffer of another frame.

Editing you do in one frame also affects the other frames. For instance, if you put text in the kill ring in one frame, you can yank it in another frame. If you exit Emacs through C-x C-c in one frame, it terminates all the frames. To delete just one frame, use C-x 5 0.

To avoid confusion, we reserve the word “window” for the subdivisions that Emacs implements, and never use it to refer to a frame.

Emacs compiled for MS-DOS emulates some aspects of the window system so that you can use many of the features described in this chapter. See Section C.1 [MS-DOS Input], page 455, for more information.

17.1 Mouse Commands for Editing

The mouse commands for selecting and copying a region are mostly compatible with the xterm program. You can use the same mouse commands for copying between Emacs and other X client programs.

If you select a region with any of these mouse commands, and then immediately afterward type the ^{⟨}\underline{\overline{DELETE}}^{⟩} function key, it deletes the region that you selected. The ^{⟨}\underline{\overline{BACKSPACE}}^{⟩} function key and the ASCII character ^{⟨}\underline{\overline{DEL}}^{⟩} do not do this; if you type any other key in between the mouse command and ^{⟨}\underline{\overline{DELETE}}^{⟩}, it does not do this. Mouse-1 Move point to where you click (mouse-set-point). This is nor-

mally the left button.

Drag-Mouse-1

Set the region to the text you select by dragging, and copy it to the kill ring (mouse-set-region). You can specify both ends of the region with this single command. If you move the mouse off the top or bottom of the window while dragging, the window scrolls at a steady rate until you move the mouse back into the window. This way, you can select regions that don’t fit entirely on the screen. The number of lines scrolled per step depends on how far away from the window edge the mouse has gone; the variable mouse-scroll-min-lines specifies a minimum step size.

Mouse-2 Yank the last killed text, where you click (mouse-yank-at-click). This is normally the middle button.

Mouse-3 This command, mouse-save-then-kill, has several functions depending on where you click and the status of the region. The most basic case is when you click Mouse-1 in one place and then Mouse-3 in another. This selects the text between those two positions as the region. It also copies the new region to the kill ring, so that you can copy it to someplace else. If you click Mouse-1 in the text, scroll with the scroll bar, and then click Mouse-3, it remembers where point was before scrolling (where you put it with Mouse-1), and uses that position as the other end of the region. This is so that you can select a region that doesn’t fit entirely on the screen. More generally, if you do not have a highlighted region, Mouse-3 selects the text between point and the click position as the region. It does this by setting the mark where point was, and moving point to where you click. If you have a highlighted region, or if the region was set just before by dragging button 1, Mouse-3 adjusts the nearer end of the region by moving it to where you click. The adjusted region’s text also replaces the old region’s text in the kill ring. If you originally specified the region using a double or triple Mouse-1, so that the region is defined to consist of entire words or lines, then adjusting the region with Mouse-3 also proceeds by entire words or lines. If you use Mouse-3 a second time consecutively, at the same place, that kills the region already selected.

Double-Mouse-1

This key sets the region around the word which you click on. If you click on a character with “symbol” syntax (such as underscore, in C mode), it sets the region around the symbol surrounding that character. If you click on a character with open-parenthesis or closeparenthesis syntax, it sets the region around the parenthetical grouping (sexp) which that character starts or ends. If you click on a character with string-delimiter syntax (such as a singlequote or doublequote in C), it sets the region around the string constant (using heuristics to figure out whether that character is the beginning or the end of it).

Double-Drag-Mouse-1

This key selects a region made up of the words you drag across.

Triple-Mouse-1

This key sets the region around the line you click on.

Triple-Drag-Mouse-1

This key selects a region made up of the lines you drag across.

The simplest way to kill text with the mouse is to press Mouse-1 at one end, then press Mouse-3 twice at the other end. See Section 9.1 [Killing], page 63. To copy the text into the kill ring without deleting it from the buffer, press Mouse-3 just once—or just drag across the text with Mouse-1. Then you can copy it elsewhere by yanking it.

To yank the killed or copied text somewhere else, move the mouse there and press Mouse-2. See Section 9.2 [Yanking], page 65. However, if mouseyank-at-point is non-nil, Mouse-2 yanks at point. Then it does not matter where you click, or even which of the frame’s windows you click on. The default value is nil. This variable also affects yanking the secondary selection.

To copy text to another X window, kill it or save it in the kill ring. Under X, this also sets the primary selection. Then use the “paste” or “yank” command of the program operating the other window to insert the text from the selection.

To copy text from another X window, use the “cut” or “copy” command of the program operating the other window, to select the text you want. Then yank it in Emacs with C-y or Mouse-2.

These cutting and pasting commands also work on MS-Windows.

When Emacs puts text into the kill ring, or rotates text to the front of the kill ring, it sets the primary selection in the X server. This is how other X clients can access the text. Emacs also stores the text in the cut buffer, but only if the text is short enough (x-cut-buffer-max specifies the maximum number of characters); putting long strings in the cut buffer can be slow.

The commands to yank the first entry in the kill ring actually check first for a primary selection in another program; after that, they check for text in the cut buffer. If neither of those sources provides text to yank, the kill ring contents are used.

17.2 Secondary Selection

The secondary selection is another way of selecting text using X. It does not use point or the mark, so you can use it to kill text without setting point or the mark.

M-Drag-Mouse-1

Set the secondary selection, with one end at the place where you press down the button, and the other end at the place where you release it (mouse-set-secondary). The highlighting appears and changes as you drag. If you move the mouse off the top or bottom of the window while dragging, the window scrolls at a steady rate until you move the mouse back into the window. This way, you can mark regions that don’t fit entirely on the screen.

M-Mouse-1

Set one endpoint for the secondary selection (mouse-start-

          secondary).
M-Mouse-3

Make a secondary selection, using the place specified with

          M-Mouse-1 as the other end (mouse-secondary-save-then-

kill). A second click at the same place kills the secondary selection just made.

M-Mouse-2

Insert the secondary selection where you click (mouse-yank-secondary). This places point at the end of the yanked text.

Double or triple clicking of M-Mouse-1 operates on words and lines, much

like Mouse-1.
   If mouse-yank-at-point is non-nil, M-Mouse-2 yanks at point. Then it

does not matter precisely where you click; all that matters is which window you click on. See Section 17.1 [Mouse Commands], page 155.

17.3 Following References with the Mouse

Some Emacs buffers display lists of various sorts. These include lists of files, of buffers, of possible completions, of matches for a pattern, and so on.

Since yanking text into these buffers is not very useful, most of them define Mouse-2 specially, as a command to use or view the item you click on.

For example, if you click Mouse-2 on a file name in a Dired buffer, you visit that file. If you click Mouse-2 on an error message in the ‘*Compilation*’ buffer, you go to the source code for that error message. If you click Mouse-2 on a completion in the ‘*Completions*’ buffer, you choose that completion.

You can usually tell when Mouse-2 has this special sort of meaning because the sensitive text highlights when you move the mouse over it.

17.4 Mouse Clicks for Menus

Mouse clicks modified with the ^{⟨}\underline{\overline{CTRL}}^{⟩} and ^{⟨}\underline{\overline{SHIFT}}^{⟩} keys bring up menus.

C-Mouse-1

This menu is for selecting a buffer.

C-Mouse-2

This menu is for specifying faces and other text properties for editing formatted text. See Section 21.11 [Formatted Text], page 213.

C-Mouse-3

This menu is mode-specific. For most modes, this menu has the same items as all the mode-specific menu-bar menus put together. Some modes may specify a different menu for this button.^{1}

S-mouse-1

This menu is for specifying the frame’s principal font.

17.5 Mode Line Mouse Commands

You can use mouse clicks on window mode lines to select and manipulate windows. Mouse-1 Mouse-1 on a mode line selects the window above. By dragging

Mouse-1 on the mode line, you can move it, thus changing the height of the windows above and below.

Mouse-2 Mouse-2 on a mode line expands that window to fill its frame. Mouse-3 Mouse-3 on a mode line deletes the window above.

C-Mouse-2

C-Mouse-2 on a mode line splits the window above horizontally, above the place in the mode line where you click.

C-Mouse-2 on a scroll bar splits the corresponding window vertically. See Section 16.2 [Split Window], page 150.

17.6 Creating Frames

The prefix key C-x 5 is analogous to C-x 4, with parallel subcommands. The difference is that C-x 5 commands create a new frame rather than just a new window in the selected frame (See Section 16.4 [Pop Up Window], page 151). If an existing visible or iconified frame already displays the requested material, these commands use the existing frame, after raising or deiconifying as necessary.

The various C-x 5 commands differ in how they find or create the buffer to select: C-x 5 2 Create a new frame (make-frame). C-x 5 b bufname ^{⟨}\underline{\overline{RET}}^{⟩}

Select buffer bufname in another frame. This runs switch-to-

          buffer-other-frame.

C-x 5 f filename ^{⟨}\underline{\overline{RET}}^{⟩} Visit file filename and select its buffer in another frame. This runs find-file-other-frame. See Section 14.2 [Visiting], page 106.

1 Some systems use Mouse-3 for a mode-specific menu. We took a survey of users, and found they preferred to keep Mouse-3 for selecting and killing regions. Hence the decision to use C-Mouse-3 for this menu.

C-x 5 d directory ^{⟨}\underline{\overline{RET}}^{⟩} Select a Dired buffer for directory directory in another frame. This runs dired-other-frame. See Chapter 28 [Dired], page 321.

C-x 5 m Start composing a mail message in another frame. This runs mail-other-frame. It is the other-frame variant of C-x m. See Chapter 26 [Sending Mail], page 293.

C-x 5 . Find a tag in the current tag table in another frame. This runs find-tag-other-frame, the multiple-frame variant of M-.. See Section 22.13 [Tags], page 243.

C-x 5 r filename ^{⟨}\underline{\overline{RET}}^{⟩} Visit file filename read-only, and select its buffer in another

          frame. This runs find-file-read-only-other-frame. See

Section 14.2 [Visiting], page 106.

You can control the appearance of new frames you create by setting the frame parameters in default-frame-alist. You can use the variable initial-frame-alist to specify parameters that affect only the initial frame. See section “Initial Parameters” in The Emacs Lisp Reference Manual, for more information.

17.7 Making and Using a Speedbar Frame

An Emacs frame can have a speedbar, which is a vertical window that serves as a scrollable menu of files you could visit and tags within those files. To create a speedbar, type M-x speedbar; this creates a speedbar window for the selected frame. From then on, you can click on a file name in the speedbar to visit that file in the corresponding Emacs frame, or click on a tag name to jump to that tag in the Emacs frame.

Initially the speedbar lists the immediate contents of the current directory, one file per line. Each line also has a box, ‘[+]’ or ‘<+>’, that you can click on with Mouse-2 to “open up” the contents of that item. If the line names a directory, opening it adds the contents of that directory to the speedbar display, underneath the directory’s own line. If the line lists an ordinary file, opening it up adds a list of the tags in that file to the speedbar display. When a file is opened up, the ‘[+]’ changes to ‘[-]’; you can click on that box to “close up” that file (hide its contents).

Some major modes, including Rmail mode, Info, and GUD, have specialized ways of putting useful items into the speedbar for you to select. For example, in Rmail mode, the speedbar shows a list of Rmail files, and lets you move the current message to another Rmail file by clicking on its ‘<M>’ box.

A speedbar belongs to one Emacs frame, and always operates on that frame. If you use multiple frames, you can make a speedbar for some or all of the frames; type M-x speedbar in any given frame to make a speedbar for it.

17.8 Multiple Displays

A single Emacs can talk to more than one X Windows display. Initially, Emacs uses just one display—the one specified with the DISPLAY environment variable or with the ‘--display’ option (see Section A.2 [Initial Options], page 434). To connect to another display, use the command make-

frame-on-display:
M-x make-frame-on-display ^{⟨}\underline{\overline{RET}}^{⟩} display ^{⟨}\underline{\overline{RET}}^{⟩}

Create a new frame on display display.

A single X server can handle more than one screen. When you open frames on two screens belonging to one server, Emacs knows they share a single keyboard, and it treats all the commands arriving from these screens as a single stream of input.

When you open frames on different X servers, Emacs makes a separate input stream for each server. This way, two users can type simultaneously on the two displays, and Emacs will not garble their input. Each server also has its own selected frame. The commands you enter with a particular X server apply to that server’s selected frame.

Despite these features, people using the same Emacs job from different displays can still interfere with each other if they are not careful. For example, if any one types C-x C-c, that exits the Emacs job for all of them!

17.9 Special Buffer Frames

You can make certain chosen buffers, for which Emacs normally creates a second window when you have just one window, appear in special frames of their own. To do this, set the variable special-display-buffer-names to a list of buffer names; any buffer whose name is in that list automatically gets a special frame, when an Emacs command wants to display it “in another window.”

For example, if you set the variable this way,

    (setq special-display-buffer-names
          '("*Completions*" "*grep*" "*tex-shell*"))

then completion lists, grep output and the TEX mode shell buffer get individual frames of their own. These frames, and the windows in them, are never automatically split or reused for any other buffers. They continue to show the buffers they were created for, unless you alter them by hand. Killing the special buffer deletes its frame automatically.

More generally, you can set special-display-regexps to a list of regular expressions; then a buffer gets its own frame if its name matches any of those regular expressions. (Once again, this applies only to buffers that normally get displayed for you in a separate window.)

The variable special-display-frame-alist specifies the frame parameters for these frames. It has a default value, so you don’t need to set it.

For those who know Lisp, an element of special-display-buffer-names or special-display-regexps can also be a list. Then the first element is the buffer name or regular expression; the rest of the list specifies how to create the frame. It can be an association list specifying frame parameter values; these values take precedence over parameter values specified in special-display-frame-alist. Alternatively, it can have this form:

(function args...) where function is a symbol. Then the frame is constructed by calling function; its first argument is the buffer, and its remaining arguments are args.

An analogous feature lets you specify buffers which should be displayed in the selected window. See Section 16.5 [Force Same Window], page 152. The same-window feature takes precedence over the special-frame feature; therefore, if you add a buffer name to special-display-buffer-names and it has no effect, check to see whether that feature is also in use for the same buffer name.

17.10 Setting Frame Parameters

This section describes commands for altering the display style and window management behavior of the selected frame.

M-x set-foreground-color ^{⟨}\underline{\overline{RET}}^{⟩} color ^{⟨}\underline{\overline{RET}}^{⟩}

Specify color color for the foreground of the selected frame.

M-x set-background-color ^{⟨}\underline{\overline{RET}}^{⟩} color ^{⟨}\underline{\overline{RET}}^{⟩}

Specify color color for the background of the selected frame. This changes the foreground color of the modeline face also, so that it remains in inverse video compared with the default.

M-x set-cursor-color ^{⟨}\underline{\overline{RET}}^{⟩} color ^{⟨}\underline{\overline{RET}}^{⟩}

Specify color color for the cursor of the selected frame.

M-x set-mouse-color ^{⟨}\underline{\overline{RET}}^{⟩} color ^{⟨}\underline{\overline{RET}}^{⟩}

Specify color color for the mouse cursor when it is over the selected frame.

M-x set-border-color ^{⟨}\underline{\overline{RET}}^{⟩} color ^{⟨}\underline{\overline{RET}}^{⟩}

Specify color color for the border of the selected frame.

M-x list-colors-display

Display the defined color names and show what the colors look like. This command is somewhat slow.

M-x auto-raise-mode

Toggle whether or not the selected frame should auto-raise. Auto-raise means that every time you move the mouse onto the frame, it raises the frame. Note that this auto-raise feature is implemented by Emacs itself. Some window managers also implement auto-raise. If you enable auto-raise for Emacs frames in your X window manager, it should work, but it is beyond Emacs’s control and therefore

auto-raise-mode has no effect on it.
M-x auto-lower-mode

Toggle whether or not the selected frame should auto-lower. Auto-lower means that every time you move the mouse off the frame, the frame moves to the bottom of the stack of X windows. The command auto-lower-mode has no effect on auto-lower implemented by the X window manager. To control that, you must use the appropriate window manager features.

M-x set-frame-font ^{⟨}\underline{\overline{RET}}^{⟩} font ^{⟨}\underline{\overline{RET}}^{⟩}

Specify font font as the principal font for the selected frame. The principal font is used for all text displayed in the frame, except when a face specifies a different font to use. See Section A.7 [Font X], page 441, for ways to list the available fonts on your system. You can also set a frame’s principal font through a pop-up menu. Press S-Mouse-1 to activate this menu.

In Emacs versions that use an X toolkit, the color-setting and font-setting functions don’t affect menus and the menu bar, since they are displayed by their own widget classes. To change the appearance of the menus and menu bar, you must use X resources (see Section A.13 [Resources X], page 446). See Section A.8 [Colors X], page 442, regarding colors. See Section A.7 [Font X], page 441, regarding choice of font.

For information on frame parameters and customization, see section “Frame Parameters” in The Emacs Lisp Reference Manual.

17.11 Scroll Bars

When using X, Emacs normally makes a scroll bar at the left of each Emacs window. The scroll bar runs the height of the window, and shows a moving rectangular inner box which represents the portion of the buffer currently displayed. The entire height of the scroll bar represents the entire length of the buffer.

You can use Mouse-2 (normally, the middle button) in the scroll bar to move or drag the inner box up and down. If you move it to the top of the scroll bar, you see the top of the buffer. If you move it to the bottom of the scroll bar, you see the bottom of the buffer.

The left and right buttons in the scroll bar scroll by controlled increments. Mouse-1 (normally, the left button) moves the line at the level where you click up to the top of the window. Mouse-3 (normally, the right button) moves the line at the top of the window down to the level where you click. By clicking repeatedly in the same place, you can scroll by the same distance over and over.

Aside from scrolling, you can also click C-Mouse-2 in the scroll bar to split a window vertically. The split occurs on the line where you click.

You can enable or disable Scroll Bar mode with the command M-x scroll-bar-mode. With no argument, it toggles the use of scroll bars. With an argument, it turns use of scroll bars on if and only if the argument is positive. This command applies to all frames, including frames yet to be created. You can use the X resource ‘verticalScrollBars’ to control the initial setting of Scroll Bar mode. See Section A.13 [Resources X], page 446.

To enable or disable scroll bars for just the selected frame, use the M-x

toggle-scroll-bar command.

17.12 Menu Bars

You can turn display of menu bars on or off with M-x menu-bar-mode. With no argument, this command toggles Menu Bar mode, a minor mode. With an argument, the command turns Menu Bar mode on if the argument is positive, off if the argument is not positive. You can use the X resource ‘menuBarLines’ to control the initial setting of Menu Bar mode. See Section A.13 [Resources X], page 446. Expert users often turn off the menu bar, especially on text-only terminals, where this makes one additional line available for text.

See Section 1.4 [Menu Bar], page 19 for information on how to invoke commands with the menu bar.

17.13 Using Multiple Typefaces

When using Emacs with X, you can set up multiple styles of displaying characters. The aspects of style that you can control are the type font, the foreground color, the background color, and whether to underline. Emacs on MS-DOS supports faces partially by letting you control the foreground and background colors of each face (see Appendix C [MS-DOS], page 455).

The way you control display style is by defining named faces. Each face can specify a type font, a foreground color, a background color, and an underline flag; but it does not have to specify all of them.

The style of display used for a given character in the text is determined by combining several faces. Any aspect of the display style that isn’t specified by overlays or text properties comes from the frame itself.

Enriched mode, the mode for editing formatted text, includes several commands and menus for specifying faces. See Section 21.11.4 [Format Faces], page 215, for how to specify the font for text in the buffer. See Section 21.11.5 [Format Colors], page 216, for how to specify the foreground and background color.

To alter the appearance of a face, use the customization buffer. See Section 31.2.2.3 [Face Customization], page 390. You can also use X resources to specify attributes of particular faces (see Section A.13 [Resources X], page 446).

To see what faces are currently defined, and what they look like, type M-x list-faces-display. It’s possible for a given face to look different in different frames; this command shows the appearance in the frame in which you type it. Here’s a list of the standardly defined faces:

default This face is used for ordinary text that doesn’t specify any other face.

modeline This face is used for mode lines. By default, it’s set up as the inverse of the default face. See Section 11.7 [Display Vars], page 81.

highlight

This face is used for highlighting portions of text, in various modes.

region This face is used for displaying a selected region (when Transient Mark mode is enabled—see below).

secondary-selection

This face is used for displaying a secondary selection (see Section 17.2 [Secondary Selection], page 157).

bold This face uses a bold variant of the default font, if it has one.

italic This face uses an italic variant of the default font, if it has one.

bold-italic

This face uses a bold italic variant of the default font, if it has one.

underline

This face underlines text.

When Transient Mark mode is enabled, the text of the region is highlighted when the mark is active. This uses the face named region; you can control the style of highlighting by changing the style of this face (see Section 31.2.2.3 [Face Customization], page 390). See Section 8.2 [Transient Mark], page 58, for more information about Transient Mark mode and activation and deactivation of the mark.

One easy way to use faces is to turn on Font Lock mode. This minor mode, which is always local to a particular buffer, arranges to choose faces according to the syntax of the text you are editing. It can recognize comments and strings in most languages; in several languages, it can also recognize and properly highlight various other important constructs. See Section 17.14 [Font Lock], page 166, for more information about Font Lock mode and syntactic highlighting.

You can print out the buffer with the highlighting that appears on your screen using the command ps-print-buffer-with-faces. See Section 30.5 [Postscript], page 370.

17.14 Font Lock mode

Font Lock mode is a minor mode, always local to a particular buffer, which highlights (or “fontifies”) using various faces according to the syntax of the text you are editing. It can recognize comments and strings in most languages; in several languages, it can also recognize and properly highlight various other important constructs—for example, names of functions being defined or reserved keywords.

The command M-x font-lock-mode turns Font Lock mode on or off according to the argument, and toggles the mode when it has no argument. The function turn-on-font-lock unconditionally enables Font Lock mode. This is useful in mode-hook functions. For example, to enable Font Lock mode whenever you edit a C file, you can do this:

    (add-hook 'c-mode-hook 'turn-on-font-lock)

To turn on Font Lock mode automatically in all modes which support it, use the function global-font-lock-mode, like this:

    (global-font-lock-mode t)

In Font Lock mode, when you edit the text, the highlighting updates automatically in the line that you changed. Most changes don’t affect the highlighting of subsequent lines, but occasionally they do. To rehighlight a

range of lines, use the command M-g M-g (font-lock-fontify-block).

In certain major modes, M-g M-g refontifies the entire current function. (The variable font-lock-mark-block-function controls how to find the current function.) In other major modes, M-g M-g refontifies 16 lines above and below point.

With a prefix argument n, M-g M-g refontifies n lines above and below point, regardless of the mode.

To get the full benefit of Font Lock mode, you need to choose a default font which has bold, italic, and bold-italic variants; or else you need to have a color or gray-scale screen.

The variable font-lock-maximum-decoration specifies the preferred level of fontification, for modes that provide multiple levels. Level 1 is the least amount of fontification; some modes support levels as high as 3. The normal default is “as high as possible.” You can specify an integer, which applies to all modes, or you can specify different numbers for particular major modes; for example, to use level 1 for C/C++ modes, and the default level otherwise, use this:

    (setq font-lock-maximum-decoration
          '((c-mode . 1) (c++-mode . 1)))

Fontification can be too slow for large buffers, so you can suppress it. The variable font-lock-maximum-size specifies a buffer size, beyond which buffer fontification is suppressed.

Comment and string fontification (or “syntactic” fontification) relies on analysis of the syntactic structure of the buffer text. For the purposes of speed, some modes including C mode and Lisp mode rely on a special convention: an open-parenthesis in the leftmost column always defines the beginning of a defun, and is thus always outside any string or comment. (See Section 22.4 [Defuns], page 224.) If you don’t follow this convention, then Font Lock mode can misfontify the text after an open-parenthesis in the leftmost column that is inside a string or comment.

   The variable font-lock-beginning-of-syntax-function (always

buffer-local) specifies how Font Lock mode can find a position guaranteed to be outside any comment or string. In modes which use the leftmost column parenthesis convention, the default value of the variable is beginning-of-defun—that tells Font Lock mode to use the convention. If you set this variable to nil, Font Lock no longer relies on the convention. This avoids incorrect results, but the price is that, in some cases, fontification for a changed text must rescan buffer text from the beginning of the buffer.

Font Lock highlighting patterns already exist for many modes, but you may want to fontify additional patterns. You can use the function fontlock-add-keywords, to add your own highlighting patterns for a particular mode. For example, to highlight ‘FIXME:’ words in C comments, use this:

    (font-lock-add-keywords
     'c-mode
     '(("\\<\\(FIXME\\):" 1 font-lock-warning-face t)))

17.15 Font Lock Support Modes

Font Lock support modes make Font Lock mode faster for large buffers. There are two support modes: Fast Lock mode and Lazy Lock mode. They use two different methods of speeding up Font Lock mode.

17.15.1 Fast Lock Mode

To make Font Lock mode faster for buffers visiting large files, you can use Fast Lock mode. Fast Lock mode saves the font information for each file in a separate cache file; each time you visit the file, it rereads the font information from the cache file instead of refontifying the text from scratch.

The command M-x fast-lock-mode turns Fast Lock mode on or off, according to the argument (with no argument, it toggles). You can also arrange to enable Fast Lock mode whenever you use Font Lock mode, like this:

    (setq font-lock-support-mode 'fast-lock-mode)

It is not worth writing a cache file for small buffers. Therefore, the variable fast-lock-minimum-size specifies a minimum file size for caching font information.

The variable fast-lock-cache-directories specifies where to put the cache files. Its value is a list of directories to try; "." means the same directory as the file being edited. The default value is ("." "~/.emacs-flc"), which means to use the same directory if possible, and otherwise the direc-

tory ‘~/.emacs-flc’.

The variable fast-lock-save-others specifies whether Fast Lock mode should save cache files for files that you do not own. A non-nil value means yes (and that is the default).

17.15.2 Lazy Lock Mode

To make Font Lock mode faster for large buffers, you can use Lazy Lock mode to reduce the amount of text that is fontified. In Lazy Lock mode, buffer fontification is demand-driven; it happens to portions of the buffer that are about to be displayed. And fontification of your changes is deferred; it happens only when Emacs has been idle for a certain short period of time.

The command M-x lazy-lock-mode turns Lazy Lock mode on or off, according to the argument (with no argument, it toggles). You can also arrange to enable Lazy Lock mode whenever you use Font Lock mode, like this:

    (setq font-lock-support-mode 'lazy-lock-mode)

It is not worth avoiding buffer fontification for small buffers. Therefore, the variable lazy-lock-minimum-size specifies a minimum buffer size for demand-driven buffer fontification. Buffers smaller than that are fontified all at once, as in plain Font Lock mode.

When you alter the buffer, Lazy Lock mode defers fontification of the text you changed. The variable lazy-lock-defer-time specifies how many seconds Emacs must be idle before it starts fontifying your changes. If the value is 0, then changes are fontified immediately, as in plain Font Lock mode.

Lazy Lock mode normally fontifies newly visible portions of the buffer before they are first displayed. However, if the value of lazy-lock-defer-on-scrolling is non-nil, newly visible text is fontified only when Emacs is

idle for lazy-lock-defer-time seconds.

In some modes, including C mode and Emacs Lisp mode, changes in one line’s contents can alter the context for subsequent lines, and thus change how they ought to be fontified. Ordinarily, you must type M-g M-g to refontify the subsequent lines. However, if you set the variable lazy-lock-defer-contextually to non-nil, Lazy Lock mode does this automatically,

after lazy-lock-defer-time seconds.

When Emacs is idle for a long time, Lazy Lock fontifies additional portions of the buffer, not yet displayed, in case you will display them later. This is called stealth fontification.

The variable lazy-lock-stealth-time specifies how many seconds Emacs has to be idle before stealth fontification starts. A value of nil means no stealth fontification. The variables lazy-lock-stealth-lines and lazy-lock-stealth-verbose specify the granularity and verbosity of stealth fontification.

17.15.3 Fast Lock or Lazy Lock?

Here is a simple guide to help you choose one of the Font Lock support modes.

• Fast Lock mode intervenes only during file visiting and buffer killing (and related events); therefore buffer editing and window scrolling are no faster or slower than in plain Font Lock mode.

• Fast Lock mode is slower at reading a cache file than Lazy Lock mode is at fontifying a window; therefore Fast Lock mode is slower at visiting a file than Lazy Lock mode.

• Lazy Lock mode intervenes during window scrolling to fontify text that scrolls onto the screen; therefore, scrolling is slower than in plain Font Lock mode.

• Lazy Lock mode doesn’t fontify during buffer editing (it defers fontification of changes); therefore, editing is faster than in plain Font Lock mode.

• Fast Lock mode can be fooled by a file that is kept under version control software; therefore buffer fontification may occur even when a cache file exists for the file.

• Fast Lock mode only works with a buffer visiting a file; Lazy Lock mode works with any buffer.

• Fast Lock mode generates cache files; Lazy Lock mode does not.

The variable font-lock-support-mode specifies which of these support modes to use; for example, to specify that Fast Lock mode is used for C/C++ modes, and Lazy Lock mode otherwise, set the variable like this:

    (setq font-lock-support-mode
          '((c-mode . fast-lock-mode) (c++-mode . fast-lock-mode)
            (t . lazy-lock-mode)))

17.16 Highlight Changes Mode

Use M-x highlight-changes-mode to enable a minor mode that uses faces (colors, typically) to indicate which parts of the buffer were changed most recently.

17.17 Miscellaneous X Window Features

The following commands let you create, delete and operate on frames:

C-z Iconify the selected Emacs frame (iconify-or-deiconify-frame). The normal meaning of C-z, to suspend Emacs, is not useful under a window system, so it has a different binding in that case. If you type this command on an Emacs frame’s icon, it deiconifies the frame.

C-x 5 0 Delete the selected frame (delete-frame). This is not allowed if there is only one frame.

C-x 5 o Select another frame, raise it, and warp the mouse to it so that it stays selected. If you repeat this command, it cycles through all the frames on your terminal.

17.18 Non-Window Terminals

If your terminal does not have a window system that Emacs supports, then it can display only one Emacs frame at a time. However, you can still create multiple Emacs frames, and switch between them. Switching frames on these terminals is much like switching between different window configurations.

Use C-x 5 2 to create a new frame and switch to it; use C-x 5 o to cycle through the existing frames; use C-x 5 0 to delete the current frame.

Each frame has a number to distinguish it. If your terminal can display only one frame at a time, the selected frame’s number n appears near the beginning of the mode line, in the form ‘Fn’.

‘Fn’ is actually the frame’s name. You can also specify a different name if you wish, and you can select a frame by its name. Use the command M-x set-frame-name ^{⟨}\underline{\overline{RET}}^{⟩} name ^{⟨}\underline{\overline{RET}}^{⟩} to specify a new name for the selected frame, and use M-x select-frame-by-name ^{⟨}\underline{\overline{RET}}^{⟩} name ^{⟨}\underline{\overline{RET}}^{⟩} to select a frame according to its name. The name you specify appears in the mode line when the frame is selected.

18 International Character Set Support

Emacs supports a wide variety of international character sets, including European variants of the Latin alphabet, as well as Chinese, Devanagari (Hindi and Marathi), Ethiopian, Greek, IPA, Japanese, Korean, Lao, Russian, Thai, Tibetan, and Vietnamese scripts. These features have been merged from the modified version of Emacs known as MULE (for “MULtilingual Enhancement to GNU Emacs”)

18.1 Introduction to International Character Sets

The users of these scripts have established many more-or-less standard coding systems for storing files. Emacs internally uses a single multibyte character encoding, so that it can intermix characters from all these scripts in a single buffer or string. This encoding represents each non-ASCII character as a sequence of bytes in the range 0200 through 0377. Emacs translates between the multibyte character encoding and various other coding systems when reading and writing files, when exchanging data with subprocesses, and (in some cases) in the C-q command (see below).

The command C-h h (view-hello-file) displays the file ‘etc/HELLO’, which shows how to say “hello” in many languages. This illustrates various scripts.

Keyboards, even in the countries where these character sets are used, generally don’t have keys for all the characters in them. So Emacs supports various input methods, typically one for each script or language, to make it convenient to type them.

The prefix key C-x ^{⟨}\underline{\overline{RET}}^{⟩} is used for commands that pertain to multibyte characters, coding systems, and input methods.

18.2 Enabling Multibyte Characters

You can enable or disable multibyte character support, either for Emacs as a whole, or for a single buffer. When multibyte characters are disabled in a buffer, then each byte in that buffer represents a character, even codes 0200 through 0377. The old features for supporting the European character sets, ISO Latin-1 and ISO Latin-2, work as they did in Emacs 19.

However, there is no need to turn off multibyte character support to use ISO Latin-1 or ISO Latin-2; the Emacs multibyte character set includes all the characters in these character sets, and Emacs can translate automatically to and from either of these ISO codes.

To edit a particular file in unibyte representation, visit it using findfile-literally. See Section 14.2 [Visiting], page 106. To convert a buffer in multibyte representation into a single-byte representation of the same characters, the easiest way is to save the contents in a file, kill the buffer, and find the file again with find-file-literally.

To turn off multibyte character support by default, start Emacs with the ‘--unibyte’ option (see Section A.2 [Initial Options], page 434), or set the environment variable ‘EMACS_UNIBYTE’.

The mode line indicates whether multibyte character support is enabled in the current buffer. If it is, there are two or more characters (most often two dashes) before the colon near the beginning of the mode line. When multibyte characters are not enabled, just one dash precedes the colon.

When multibyte characters are enabled, character codes 0240 (octal) through 0377 (octal) are not really legitimate in the buffer. The valid non-ASCII printing characters have codes that start from 0400.

If you type a self-inserting character in the invalid range 0240 through 0377, Emacs assumes you intended to use one of the ISO Latin-n character sets, and converts it to the Emacs code representing that Latin-n character. You select which ISO Latin character set to use though your choice of language environment (see below). If you do not specify a choice, the default is Latin-1.

The same thing happens when you use C-q to enter an octal code in this range.

18.3 Language Environments

All supported character sets are supported in Emacs buffers whenever multibyte characters are enabled; there is no need to select a particular language in order to display its characters in an Emacs buffer. However, it is important to select a language environment in order to set various defaults. The language environment really represents a choice of preferred script (more or less) rather that a choice of language.

The language environment controls which coding systems to recognize when reading text (see Section 18.7 [Recognize Coding], page 177). This applies to files, incoming mail, netnews, and any other text you read into Emacs. It may also specify the default coding system to use when you create a file. Each language environment also specifies a default input method.

The way to select a language environment is with the command M-x set-language-environment. It makes no difference which buffer is current when you use this command, because the effects apply globally to the Emacs session. The supported language environments include:

Chinese-BIG5, Chinese-CNS, Chinese-GB, Cyrillic-Alternativnyj, Cyrillic-ISO, Cyrillic-KOI8, Devanagari, English, Ethiopic, Greek, Hebrew, Japanese, Korean, Lao, Latin-1, Latin-2, Latin-3, Latin-4, Latin-5, Thai, Tibetan, and Vietnamese. Some operating systems let you specify the language you are using by set-

ting locale environment variables. Emacs handles one common special case of this: if your locale name for character types contains the string ‘8859-n’, Emacs automatically selects the corresponding language environment.

To display information about the effects of a certain language environment lang-env, use the command C-h L lang-env ^{⟨}\underline{\overline{RET}}^{⟩} (describelanguage-environment). This tells you which languages this language environment is useful for, and lists the character sets, coding systems, and input methods that go with it. It also shows some sample text to illustrate scripts used in this language environment. By default, this command describes the chosen language environment.

You can customize any language environment with the normal hook set-

language-environment-hook. The command set-language-environment

runs that hook after setting up the new language environment. The hook functions can test for a specific language environment by checking the vari-

able current-language-environment.

Before it starts to set up the new language environment, set-language-

environment first runs the hook exit-language-environment-hook. This

hook is useful for undoing customizations that were made with setlanguage-environment-hook. For instance, if you set up a special key binding in a specific language environment using set-language-environment-

hook, you should set up exit-language-environment-hook to restore the

normal binding for that key.

18.4 Input Methods

An input method is a kind of character conversion designed specifically for interactive input. In Emacs, typically each language has its own input method; sometimes several languages which use the same characters can share one input method. A few languages support several input methods.

The simplest kind of input method works by mapping ASCII letters into another alphabet. This is how the Greek and Russian input methods work.

A more powerful technique is composition: converting sequences of characters into one letter. Many European input methods use composition to produce a single non-ASCII letter from a sequence that consists of a letter followed by accent characters. For example, some methods convert the sequence 'a into a single accented letter. These input methods have no special commands of their own; all they do is compose sequences of printing characters.

The input methods for syllabic scripts typically use mapping followed by composition. The input methods for Thai and Korean work this way. First, letters are mapped into symbols for particular sounds or tone marks; then, sequences of these which make up a whole syllable are mapped into one syllable sign.

Chinese and Japanese require more complex methods. In Chinese input methods, first you enter the phonetic spelling of a Chinese word (in input method chinese-py, among others), or a sequence of portions of the character (input methods chinese-4corner and chinese-sw, and others).

Since one phonetic spelling typically corresponds to many different Chinese characters, you must select one of the alternatives using special Emacs commands. Keys such as C-f, C-b, C-n, C-p, and digits have special definitions in this situation, used for selecting among the alternatives. ^{⟨}\underline{\overline{TAB}}^{⟩} displays a buffer showing all the possibilities.

In Japanese input methods, first you input a whole word using phonetic spelling; then, after the word is in the buffer, Emacs converts it into one or more characters using a large dictionary. One phonetic spelling corresponds to many differently written Japanese words, so you must select one of them; use C-n and C-p to cycle through the alternatives.

Sometimes it is useful to cut off input method processing so that the characters you have just entered will not combine with subsequent characters. For example, in input method latin-1-postfix, the sequence e ' combines to form an ‘e’ with an accent. What if you want to enter them as separate characters?

One way is to type the accent twice; that is a special feature for entering the separate letter and accent. For example, e ' ' gives you the two characters ‘e'’. Another way is to type another letter after the e—something that won’t combine with that—and immediately delete it. For example, you could type e e ^{⟨}\underline{\overline{DEL}}^{⟩} ' to get separate ‘e’ and ‘'’.

Another method, more general but not quite as easy to type, is to use C-\ C-\ between two characters to stop them from combining. This is the

command C-\ (toggle-input-method) used twice.

C-\ C-\ is especially useful inside an incremental search, because it stops waiting for more characters to combine, and starts searching for what you have already entered.

   The variables input-method-highlight-flag and input-method-

verbose-flag control how input methods explain what is happening. If input-method-highlight-flag is non-nil, the partial sequence is highlighted in the buffer. If input-method-verbose-flag is non-nil, the list of possible characters to type next is displayed in the echo area (but not when you are in the minibuffer).

18.5 Selecting an Input Method

C-\ Enable or disable use of the selected input method. C-x ^{⟨}\underline{\overline{RET}}^{⟩} C-\ method ^{⟨}\underline{\overline{RET}}^{⟩}

Select a new input method for the current buffer.

C-h I method ^{⟨}\underline{\overline{RET}}^{⟩} C-h C-\ method ^{⟨}\underline{\overline{RET}}^{⟩}

Describe the input method method (describe-input-method). By default, it describes the current input method (if any). This description should give you the full details of how to use any particular input method.

M-x list-input-methods

Display a list of all the supported input methods.

To choose an input method for the current buffer, use C-x ^{⟨}\underline{\overline{RET}}^{⟩} C-\ (set-input-method). This command reads the input method name with the minibuffer; the name normally starts with the language environment that it is meant to be used with. The variable current-input-method records which input method is selected.

Input methods use various sequences of ASCII characters to stand for non-ASCII characters. Sometimes it is useful to turn off the input method temporarily. To do this, type C-\ (toggle-input-method). To reenable the input method, type C-\ again.

If you type C-\ and you have not yet selected an input method, it prompts for you to specify one. This has the same effect as using C-x ^{⟨}\underline{\overline{RET}}^{⟩} C-\ to specify an input method.

Selecting a language environment specifies a default input method for use in various buffers. When you have a default input method, you can select it in the current buffer by typing C-\. The variable default-input-method specifies the default input method (nil means there is none).

Some input methods for alphabetic scripts work by (in effect) remapping the keyboard to emulate various keyboard layouts commonly used for those scripts. How to do this remapping properly depends on your actual keyboard layout. To specify which layout your keyboard has, use the command M-x

quail-set-keyboard-layout.

To display a list of all the supported input methods, type M-x list-input-methods. The list gives information about each input method, including the string that stands for it in the mode line.

18.6 Coding Systems

Users of various languages have established many more-or-less standard coding systems for representing them. Emacs does not use these coding systems internally; instead, it converts from various coding systems to its own system when reading data, and converts the internal coding system to other coding systems when writing data. Conversion is possible in reading or writing files, in sending or receiving from the terminal, and in exchanging data with subprocesses.

Emacs assigns a name to each coding system. Most coding systems are used for one language, and the name of the coding system starts with the language name. Some coding systems are used for several languages; their names usually start with ‘iso’. There are also special coding systems noconversion, raw-text and emacs-mule which do not convert printing characters at all.

In addition to converting various representations of non-ASCII characters, a coding system can perform end-of-line conversion. Emacs handles three different conventions for how to separate lines in a file: newline, carriage-return linefeed, and just carriage-return.

C-h C coding ^{⟨}\underline{\overline{RET}}^{⟩} Describe coding system coding.

C-h C ^{⟨}\underline{\overline{RET}}^{⟩} Describe the coding systems currently in use.

M-x list-coding-systems

Display a list of all the supported coding systems.

The command C-h C (describe-coding-system) displays information about particular coding systems. You can specify a coding system name as argument; alternatively, with an empty argument, it describes the coding systems currently selected for various purposes, both in the current buffer and as the defaults, and the priority list for recognizing coding systems (see Section 18.7 [Recognize Coding], page 177).

To display a list of all the supported coding systems, type M-x list-coding-systems. The list gives information about each coding system, including the letter that stands for it in the mode line (see Section 1.3 [Mode Line], page 17).

Each of the coding systems that appear in this list—except for noconversion, which means no conversion of any kind—specifies how and whether to convert printing characters, but leaves the choice of end-of-line conversion to be decided based on the contents of each file. For example, if the file appears to use carriage-return linefeed between lines, that end-of-line conversion will be used.

Each of the listed coding systems has three variants which specify exactly what to do for end-of-line conversion:

...-unix Don’t do any end-of-line conversion; assume the file uses newline to separate lines. (This is the convention normally used on Unix and GNU systems.)

...-dos Assume the file uses carriage-return linefeed to separate lines, and do the appropriate conversion. (This is the convention normally used on Microsoft systems.)

...-mac Assume the file uses carriage-return to separate lines, and do the appropriate conversion. (This is the convention normally used on the Macintosh system.)

These variant coding systems are omitted from the list-coding-systems display for brevity, since they are entirely predictable. For example, the coding system iso-latin-1 has variants iso-latin-1-unix,

iso-latin-1-dos and iso-latin-1-mac.

The coding system raw-text is good for a file which is mainly ASCII text, but may contain byte values above 127 which are not meant to encode non-ASCII characters. With raw-text, Emacs copies those byte values un-

changed, and sets enable-multibyte-characters to nil in the current

buffer so that they will be interpreted properly. raw-text handles end-of-line conversion in the usual way, based on the data encountered, and has the usual three variants to specify the kind of end-of-line conversion to use.

In contrast, the coding system no-conversion specifies no character code conversion at all—none for non-ASCII byte values and none for end of line. This is useful for reading or writing binary files, tar files, and other files that must be examined verbatim. It, too, sets enable-multibyte-characters

to nil.

The easiest way to edit a file with no conversion of any kind is with the

M-x find-file-literally command. This uses no-conversion, and also

suppresses other Emacs features that might convert the file contents before you see them. See Section 14.2 [Visiting], page 106.

The coding system emacs-mule means that the file contains non-ASCII characters stored with the internal Emacs encoding. It handles end-of-line conversion based on the data encountered, and has the usual three variants to specify the kind of end-of-line conversion.

18.7 Recognizing Coding Systems

Most of the time, Emacs can recognize which coding system to use for any given file–once you have specified your preferences.

Some coding systems can be recognized or distinguished by which byte sequences appear in the data. However, there are coding systems that cannot be distinguished, not even potentially. For example, there is no way to distinguish between Latin-1 and Latin-2; they use the same byte values with different meanings.

Emacs handles this situation by means of a priority list of coding systems. Whenever Emacs reads a file, if you do not specify the coding system to use, Emacs checks the data against each coding system, starting with the first in priority and working down the list, until it finds a coding system that fits the data. Then it converts the file contents assuming that they are represented in this coding system.

The priority list of coding systems depends on the selected language environment (see Section 18.3 [Language Environments], page 172). For example, if you use French, you probably want Emacs to prefer Latin-1 to Latin-2; if you use Czech, you probably want Latin-2 to be preferred. This is one of the reasons to specify a language environment.

However, you can alter the priority list in detail with the command M-x prefer-coding-system. This command reads the name of a coding system from the minibuffer, and adds it to the front of the priority list, so that it is preferred to all others. If you use this command several times, each use adds one element to the front of the priority list.

Sometimes a file name indicates which coding system to use for the file. The variable file-coding-system-alist specifies this correspondence. There is a special function modify-coding-system-alist for adding elements to this list. For example, to read and write all ‘.txt’ files using the coding system china-iso-8bit, you can execute this Lisp expression:

     (modify-coding-system-alist 'file "\\.txt\\'" 'china-iso-8bit)

The first argument should be file, the second argument should be a regular expression that determines which files this applies to, and the third argument says which coding system to use for these files.

You can specify the coding system for a particular file using the ‘-*-...-*-’ construct at the beginning of a file, or a local variables list at the end (see Section 31.2.5 [File Variables], page 395). You do this by defining a value for the “variable” named coding. Emacs does not really have a variable coding; instead of setting a variable, it uses the specified coding system for the file. For example, ‘-*-mode: C; coding: latin-1;-*-’ specifies use of the Latin-1 coding system, as well as C mode. If you specify the coding explicitly in the file, that overrides file-coding-system-alist.

The variable auto-coding-alist is the strongest way to specify the coding system for certain patterns of file names; this variable even overrides ‘-*-coding:-*-’ tags in the file itself. Emacs uses this feature for tar and archive files, to prevent Emacs from being confused by a ‘-*-coding:-*-’ tag in a member of the archive and thinking it applies to the archive file as a whole.

Once Emacs has chosen a coding system for a buffer, it stores that coding system in buffer-file-coding-system and uses that coding system, by default, for operations that write from this buffer into a file. This includes the commands save-buffer and write-region. If you want to write files from this buffer using a different coding system, you can specify a different coding system for the buffer using set-buffer-file-coding-system (see Section 18.8 [Specify Coding], page 179).

When you send a message with Mail mode (see Chapter 26 [Sending Mail], page 293), Emacs has four different ways to determine the coding system to use for encoding the message text. It tries the buffer’s own value of buffer-file-coding-system, if that is non-nil. Otherwise, it uses the value of sendmail-coding-system, if that is non-nil. The third way is to use the default coding system for new files, which is controlled by your choice of language environment, if that is non-nil. If all of these three values are nil, Emacs encodes outgoing mail using the Latin-1 coding system.

When you get new mail in Rmail, each message is translated automatically from the coding system it is written in—as if it were a separate file. This uses the priority list of coding systems that you have specified.

For reading and saving Rmail files themselves, Emacs uses the coding system specified by the variable rmail-file-coding-system. The default value is nil, which means that Rmail files are not translated (they are read and written in the Emacs internal character code).

18.8 Specifying a Coding System

In cases where Emacs does not automatically choose the right coding system, you can use these commands to specify one:

C-x ^{⟨}\underline{\overline{RET}}^{⟩} f coding ^{⟨}\underline{\overline{RET}}^{⟩} Use coding system coding for the visited file in the current buffer.

C-x ^{⟨}\underline{\overline{RET}}^{⟩} c coding ^{⟨}\underline{\overline{RET}}^{⟩} Specify coding system coding for the immediately following command.

C-x ^{⟨}\underline{\overline{RET}}^{⟩} k coding ^{⟨}\underline{\overline{RET}}^{⟩} Use coding system coding for keyboard input.

C-x ^{⟨}\underline{\overline{RET}}^{⟩} t coding ^{⟨}\underline{\overline{RET}}^{⟩} Use coding system coding for terminal output.

C-x ^{⟨}\underline{\overline{RET}}^{⟩} p coding ^{⟨}\underline{\overline{RET}}^{⟩} Use coding system coding for subprocess input and output in the current buffer.

C-x ^{⟨}\underline{\overline{RET}}^{⟩} x coding ^{⟨}\underline{\overline{RET}}^{⟩} Use coding system coding for transferring selections to and from other programs through the window system.

C-x ^{⟨}\underline{\overline{RET}}^{⟩} X coding ^{⟨}\underline{\overline{RET}}^{⟩} Use coding system coding for transferring one selection—the next one—to or from the window system.

   The command C-x ^{⟨}\underline{\overline{RET}}^{⟩} f (set-buffer-file-coding-system) specifies

the file coding system for the current buffer—in other words, which coding system to use when saving or rereading the visited file. You specify which coding system using the minibuffer. Since this command applies to a file you have already visited, it affects only the way the file is saved.

Another way to specify the coding system for a file is when you visit the file. First use the command C-x ^{⟨}\underline{\overline{RET}}^{⟩} c (universal-coding-system-argument); this command uses the minibuffer to read a coding system name. After you exit the minibuffer, the specified coding system is used for the immediately following command.

So if the immediately following command is C-x C-f, for example, it reads the file using that coding system (and records the coding system for when the file is saved). Or if the immediately following command is C-x C-w, it writes the file using that coding system. Other file commands affected by a specified coding system include C-x C-i and C-x C-v, as well as the other-window variants of C-x C-f.

C-x ^{⟨}\underline{\overline{RET}}^{⟩} c also affects commands that start subprocesses, including M-x shell (see Section 30.2 [Shell], page 359).

However, if the immediately following command does not use the coding system, then C-x ^{⟨}\underline{\overline{RET}}^{⟩} c ultimately has no effect.

An easy way to visit a file with no conversion is with the M-x find-file-literally command. See Section 14.2 [Visiting], page 106.

   The variable default-buffer-file-coding-system specifies the choice

of coding system to use when you create a new file. It applies when you find a new file, and when you create a buffer and then save it in a file. Selecting a language environment typically sets this variable to a good choice of default coding system for that language environment.

   The command C-x ^{⟨}\underline{\overline{RET}}^{⟩} t (set-terminal-coding-system) specifies the

coding system for terminal output. If you specify a character code for terminal output, all characters output to the terminal are translated into that coding system.

This feature is useful for certain character-only terminals built to support specific languages or character sets—for example, European terminals that support one of the ISO Latin character sets. You need to specify the terminal coding system when using multibyte text, so that Emacs knows which characters the terminal can actually handle.

By default, output to the terminal is not translated at all, unless Emacs can deduce the proper coding system from your terminal type.

   The command C-x ^{⟨}\underline{\overline{RET}}^{⟩} k (set-keyboard-coding-system) specifies the

coding system for keyboard input. Character-code translation of keyboard input is useful for terminals with keys that send non-ASCII graphic characters—for example, some terminals designed for ISO Latin-1 or subsets of it.

By default, keyboard input is not translated at all.

There is a similarity between using a coding system translation for keyboard input, and using an input method: both define sequences of keyboard input that translate into single characters. However, input methods are designed to be convenient for interactive use by humans, and the sequences that are translated are typically sequences of ASCII printing characters. Coding systems typically translate sequences of non-graphic characters.

   The command C-x ^{⟨}\underline{\overline{RET}}^{⟩} x (set-selection-coding-system) specifies

the coding system for sending selected text to the window system, and for receiving the text of selections made in other applications. This command applies to all subsequent selections, until you override it by using the command

again. The command C-x ^{⟨}\underline{\overline{RET}}^{⟩} X (set-next-selection-coding-system)

specifies the coding system for the next selection made in Emacs or read by Emacs.

   The command C-x ^{⟨}\underline{\overline{RET}}^{⟩} p (set-buffer-process-coding-system) spec-

ifies the coding system for input and output to a subprocess. This command applies to the current buffer; normally, each subprocess has its own buffer, and thus you can use this command to specify translation to and from a particular subprocess by giving the command in the corresponding buffer.

By default, process input and output are not translated at all.

The variable file-name-coding-system specifies a coding system to use for encoding file names. If you set the variable to a coding system name (as a Lisp symbol or a string), Emacs encodes file names using that coding system for all file operations. This makes it possible to use non-ASCII characters in file names—or, at least, those non-ASCII characters which the specified coding system can encode.

If file-name-coding-system is nil, Emacs uses a default coding system determined by the selected language environment. In the default language environment, any non-ASCII characters in file names are not encoded specially; they appear in the file system using the internal Emacs representation.

Warning: if you change file-name-coding-system (or the language environment) in the middle of an Emacs session, problems can result if you have already visited files whose names were encoding using the earlier coding system and cannot be encoded (or are encoded differently) under the new coding system. If you try to save one of these buffers under the visited file name, saving may use the wrong file name, or it may get an error. If such a problem happens, use C-x C-w to specify a new file name for that buffer.

18.9 Fontsets

A font for X Windows typically defines shapes for one alphabet or script. Therefore, displaying the entire range of scripts that Emacs supports requires a collection of many fonts. In Emacs, such a collection is called a fontset. A fontset is defined by a list of fonts, each assigned to handle a range of character codes.

Each fontset has a name, like a font. The available X fonts are defined by the X server; fontsets, however, are defined within Emacs itself. Once you have defined a fontset, you can use it within Emacs by specifying its name, anywhere that you could use a single font. Of course, Emacs fontsets can use only the fonts that the X server supports; if certain characters appear on the screen as hollow boxes, this means that the fontset in use for them has no font for those characters.

Emacs creates two fontsets automatically: the standard fontset and the startup fontset. The standard fontset is most likely to have fonts for a wide variety of non-ASCII characters; however, this is not the default for Emacs to use. (By default, Emacs tries to find a font which has bold and italic variants.) You can specify use of the standard fontset by starting Emacs in this way:

    emacs -fn fontset-standard

A fontset does not necessarily specify a font for all character codes. If a fontset specifies no font for a certain character, or if it specifies a font that does not exist on your system, then it cannot display that character. It will display an empty box instead.

The fontset height and width are determined by the ASCII characters (that is, by the font used for ASCII characters in that fontset). If another font in the fontset has a different height, or a different width, then characters assigned to that font are clipped to the fontset’s size. If highlight-wrong-size-font is non-nil, a box is displayed around these wrong-size characters as well.

18.10 Defining fontsets

Emacs creates a standard fontset automatically according to the value of

standard-fontset-spec. This fontset’s name is
    -*-fixed-medium-r-normal-*-16-*-*-*-*-*-fontset-standard

or just ‘fontset-standard’ for short.

Bold, italic, and bold-italic variants of the standard fontset are created automatically. Their names have ‘bold’ instead of ‘medium’, or ‘i’ instead of ‘r’, or both.

If you specify a default ASCII font with the ‘Font’ resource or the ‘-fn’ argument, Emacs generates a fontset from it automatically. This is the startup fontset and its name is fontset-startup. It does this by replacing the foundry, family, add style, and average width fields of the font name with ‘*’, replacing charset registry field with ‘fontset’, and replacing charset encoding field with ‘startup’, then using the resulting string to specify a fontset.

For instance, if you start Emacs this way,

    emacs -fn "*courier-medium-r-normal--14-140-*-iso8859-1"

Emacs generates the following fontset and uses it for the initial X window frame:

    -*-*-medium-r-normal-*-14-140-*-*-*-*-fontset-startup

With the X resource ‘Emacs.Font’, you can specify a fontset name just like an actual font name. But be careful not to specify a fontset name in a wildcard resource like ‘Emacs*Font’—that wildcard specification applies to various other purposes, such as menus, and menus cannot handle fontsets.

You can specify additional fontsets using X resources named ‘Fontset-n’, where n is an integer starting from 0. The resource value should have this form:

fontpattern, [charsetname:fontname]…

fontpattern should have the form of a standard X font name, except for the last two fields. They should have the form ‘fontset-alias’.

The fontset has two names, one long and one short. The long name is fontpattern. The short name is ‘fontset-alias’. You can refer to the fontset by either name.

The construct ‘charset:font’ specifies which font to use (in this fontset) for one particular character set. Here, charset is the name of a character set, and font is the font to use for that character set. You can use this construct any number of times in defining one fontset.

For the other character sets, Emacs chooses a font based on fontpattern. It replaces ‘fontset-alias’ with values that describe the character set. For the ASCII character font, ‘fontset-alias’ is replaced with ‘ISO8859-1’.

In addition, when several consecutive fields are wildcards, Emacs collapses them into a single wildcard. This is to prevent use of auto-scaled fonts. Fonts made by scaling larger fonts are not usable for editing, and scaling a smaller font is not useful because it is better to use the smaller font in its own size, which Emacs does.

Thus if fontpattern is this,

    -*-fixed-medium-r-normal-*-24-*-*-*-*-*-fontset-24

the font specification for ASCII characters would be this:

    -*-fixed-medium-r-normal-*-24-*-ISO8859-1

and the font specification for Chinese GB2312 characters would be this:

    -*-fixed-medium-r-normal-*-24-*-gb2312*-*

You may not have any Chinese font matching the above font specification. Most X distributions include only Chinese fonts that have ‘song ti’ or ‘fangsong ti’ in family field. In such a case, ‘Fontset-n’ can be specified as below:

     Emacs.Fontset-0: -*-fixed-medium-r-normal-*-24-*-*-*-*-*-fontset-24,\
             chinese-gb2312:-*-*-medium-r-normal-*-24-*-gb2312*-*

Then, the font specifications for all but Chinese GB2312 characters have ‘fixed’ in the family field, and the font specification for Chinese GB2312 characters has a wild card ‘*’ in the family field.

The function that processes the fontset resource value to create the fontset

is called create-fontset-from-fontset-spec. You can also call this func-

tion explicitly to create a fontset.

See Section A.7 [Font X], page 441, for more information about font naming in X.

18.11 Single-byte European Character Support

The ISO 8859 Latin-n character sets define character codes in the range 160 to 255 to handle the accented letters and punctuation needed by various European languages. If you disable multibyte characters with ‘--unibyte’, Emacs can still handle one of these character codes at a time. To specify which of these codes to use, invoke M-x set-language-environment and specify a suitable language environment such as ‘Latin-n’.

Emacs can also display those characters, provided the terminal or font in use supports them. This works automatically. Alternatively, if you are using a window system, Emacs can also display single-byte characters through fontsets, in effect by displaying the equivalent multibyte characters according to the current language environment. To request this, set the variable

unibyte-display-via-language-environment to a non-nil value.

If your terminal does not support display of the Latin-1 character set, Emacs can display these characters as ASCII sequences which at least give you a clear idea of what the characters are. To do this, load the library iso-ascii. Similar libraries for other Latin-n character sets could be implemented, but we don’t have them yet.

There are three different ways you can input single-byte non-ASCI characters characters:

• If your keyboard can generate character codes 128 and up, representing non-ASCII characters, execute the following expression to enable Emacs to understand them:

        (set-input-mode (car (current-input-mode))
                        (nth 1 (current-input-mode))
                        0)

• You can use an input method for the selected language environment. When you use an input method in a unibyte buffer, the non-ASCII character you specify with it is converted to unibyte.

• For Latin-1 only, you can load the library iso-transl to turn the key C-x 8 into a “compose character” prefix for entry of non-ASCII Latin-1 printing characters. C-x 8 is good for insertion (in the minibuffer as well as other buffers), for searching, and in any other context where a key sequence is allowed. The ^{⟨}\underline{\overline{ALT}}^{⟩} modifier key, if you have one, serves the same purpose as C-x 8; use ^{⟨}\underline{\overline{ALT}}^{⟩} together with an accent character to modify the following letter.

19 Major Modes

Emacs provides many alternative major modes, each of which customizes Emacs for editing text of a particular sort. The major modes are mutually exclusive, and each buffer has one major mode at any time. The mode line normally shows the name of the current major mode, in parentheses (see Section 1.3 [Mode Line], page 17).

The least specialized major mode is called Fundamental mode. This mode has no mode-specific redefinitions or variable settings, so that each Emacs command behaves in its most general manner, and each option is in its default state. For editing text of a specific type that Emacs knows about, such as Lisp code or English text, you should switch to the appropriate major mode, such as Lisp mode or Text mode.

Selecting a major mode changes the meanings of a few keys to become more specifically adapted to the language being edited. The ones that are changed frequently are ^{⟨}\underline{\overline{TAB}}^{⟩}, ^{⟨}\underline{\overline{DEL}}^{⟩}, and C-j. The prefix key C-c normally contains mode-specific commands. In addition, the commands which handle comments use the mode to determine how comments are to be delimited. Many major modes redefine the syntactical properties of characters appearing in the buffer. See Section 31.6 [Syntax], page 411.

The major modes fall into three major groups. Lisp mode (which has several variants), C mode, Fortran mode and others are for specific programming languages. Text mode, Nroff mode, TEX mode and Outline mode are for editing English text. The remaining major modes are not intended for use on users’ files; they are used in buffers created for specific purposes by Emacs, such as Dired mode for buffers made by Dired (see Chapter 28 [Dired], page 321), Mail mode for buffers made by C-x m (see Chapter 26 [Sending Mail], page 293), and Shell mode for buffers used for communicating with an inferior shell process (see Section 30.2.2 [Interactive Shell], page 360).

Most programming-language major modes specify that only blank lines separate paragraphs. This is to make the paragraph commands useful. (See Section 21.3 [Paragraphs], page 195.) They also cause Auto Fill mode to use the definition of ^{⟨}\underline{\overline{TAB}}^{⟩} to indent the new lines it creates. This is because most lines in a program are usually indented. (See Chapter 20 [Indentation], page 189.)

19.1 How Major Modes are Chosen

You can select a major mode explicitly for the current buffer, but most of the time Emacs determines which mode to use based on the file name or on special text in the file.

Explicit selection of a new major mode is done with a M-x command. From the name of a major mode, add -mode to get the name of a command to select that mode. Thus, you can enter Lisp mode by executing M-x

lisp-mode.

When you visit a file, Emacs usually chooses the right major mode based on the file’s name. For example, files whose names end in ‘.c’ are edited in C mode. The correspondence between file names and major modes is controlled by the variable auto-mode-alist. Its value is a list in which each element has this form,

(regexp . mode-function) or this form,

(regexp mode-function flag) For example, one element normally found in the list has the form ("\\.c\\'" . c-mode), and it is responsible for selecting C mode for files whose names end in ‘.c’. (Note that ‘\\’ is needed in Lisp syntax to include a ‘\’ in the string, which is needed to suppress the special meaning of ‘.’ in regexps.) If the element has the form (regexp mode-function flag) and flag is non-nil, then after calling function, the suffix that matched regexp is discarded and the list is searched again for another match.

You can specify which major mode should be used for editing a certain file by a special sort of text in the first nonblank line of the file. The mode name should appear in this line both preceded and followed by ‘-*-’. Other text may appear on the line as well. For example,

    ;-*-Lisp-*-

tells Emacs to use Lisp mode. Such an explicit specification overrides any defaulting based on the file name. Note how the semicolon is used to make Lisp treat this line as a comment.

Another format of mode specification is

    -*- mode: modename;-*-

which allows you to specify local variables as well, like this:

    -*- mode: modename; var: value; ... -*-

See Section 31.2.5 [File Variables], page 395, for more information about this.

When a file’s contents begin with ‘#!’, it can serve as an executable shell command, which works by running an interpreter named on the file’s first line. The rest of the file is used as input to the interpreter.

When you visit such a file in Emacs, if the file’s name does not specify a major mode, Emacs uses the interpreter name on the first line to choose a mode. If the first line is the name of a recognized interpreter program, such as ‘perl’ or ‘tcl’, Emacs uses a mode appropriate for programs for that interpreter. The variable interpreter-mode-alist specifies the correspondence between interpreter program names and major modes.

When you visit a file that does not specify a major mode to use, or when you create a new buffer with C-x b, the variable default-major-mode specifies which major mode to use. Normally its value is the symbol fundamental-mode, which specifies Fundamental mode. If default-major-mode is nil, the major mode is taken from the previously selected buffer.

If you change the major mode of a buffer, you can go back to the major mode Emacs would choose automatically: use the command M-x normal-mode to do this. This is the same function that find-file calls to choose the major mode. It also processes the file’s local variables list if any.

   The commands C-x C-w and set-visited-file-name change to a new

major mode if the new file name implies a mode (see Section 14.3 [Saving], page 109). However, this does not happen if the buffer contents specify a major mode, and certain “special” major modes do not allow the mode to change. You can turn off this mode-changing feature by setting change-

major-mode-with-file-name to nil.

188 GNU Emacs Manual

20 Indentation

This chapter describes the Emacs commands that add, remove, or adjust indentation.

⟨\underline{\overline{TAB}}^{⟩} Indent current line “appropriately” in a mode-dependent fashion.

C-j Perform ^{⟨}\underline{\overline{RET}}^{⟩} followed by ^{⟨}\underline{\overline{TAB}}^{⟩} (newline-and-indent).

M-^Merge two lines (delete-indentation). This would cancel out the effect of C-j.

C-M-o Split line at point; text on the line after point becomes a new line indented to the same column that it now starts in (split-line).

M-m Move (forward or back) to the first nonblank character on the

          current line (back-to-indentation).

C-M-\ Indent several lines to same column (indent-region).

C-x ^{⟨}\underline{\overline{TAB}}^{⟩} Shift block of lines rigidly right or left (indent-rigidly).

M-i Indent from point to the next prespecified tab stop column (tab-

          to-tab-stop).
M-x indent-relative

Indent from point to under an indentation point in the previous line.

Most programming languages have some indentation convention. For Lisp code, lines are indented according to their nesting in parentheses. The same general idea is used for C code, though many details are different.

Whatever the language, to indent a line, use the ^{⟨}\underline{\overline{TAB}}^{⟩} command. Each major mode defines this command to perform the sort of indentation appropriate for the particular language. In Lisp mode, ^{⟨}\underline{\overline{TAB}}^{⟩} aligns the line according to its depth in parentheses. No matter where in the line you are when you type ^{⟨}\underline{\overline{TAB}}^{⟩}, it aligns the line as a whole. In C mode, ^{⟨}\underline{\overline{TAB}}^{⟩} implements a subtle and sophisticated indentation style that knows about many aspects of C syntax.

In Text mode, ^{⟨}\underline{\overline{TAB}}^{⟩} runs the command tab-to-tab-stop, which indents to the next tab stop column. You can set the tab stops with M-x

edit-tab-stops.

20.1 Indentation Commands and Techniques

To move over the indentation on a line, do M-m (back-to-indentation). This command, given anywhere on a line, positions point at the first nonblank character on the line.

To insert an indented line before the current line, do C-a C-o ^{⟨}\underline{\overline{TAB}}^{⟩}. To make an indented line after the current line, use C-e C-j.

If you just want to insert a tab character in the buffer, you can type C-q ⟨\underline{\overline{TAB}}^{⟩}.

C-M-o (split-line) moves the text from point to the end of the line vertically down, so that the current line becomes two lines. C-M-o first moves point forward over any spaces and tabs. Then it inserts after point a newline and enough indentation to reach the same column point is on. Point remains before the inserted newline; in this regard, C-M-o resembles C-o.

To join two lines cleanly, use the M-^(delete-indentation) command. It deletes the indentation at the front of the current line, and the line boundary as well, replacing them with a single space. As a special case (useful for Lisp code) the single space is omitted if the characters to be joined are consecutive open parentheses or closing parentheses, or if the junction follows another newline. To delete just the indentation of a line, go to the beginning of the line and use M-\ (delete-horizontal-space), which deletes all spaces and tabs around the cursor.

If you have a fill prefix, M-^deletes the fill prefix if it appears after the newline that is deleted. See Section 21.5.3 [Fill Prefix], page 200.

There are also commands for changing the indentation of several lines at once. C-M-\ (indent-region) applies to all the lines that begin in the region; it indents each line in the “usual” way, as if you had typed ^{⟨}\underline{\overline{TAB}}^{⟩} at the beginning of the line. A numeric argument specifies the column to indent to, and each line is shifted left or right so that its first nonblank character appears in that column. C-x ^{⟨}\underline{\overline{TAB}}^{⟩} (indent-rigidly) moves all of the lines in the region right by its argument (left, for negative arguments). The whole group of lines moves rigidly sideways, which is how the command gets its name.

M-x indent-relative indents at point based on the previous line (actually, the last nonempty line). It inserts whitespace at point, moving point, until it is underneath an indentation point in the previous line. An indentation point is the end of a sequence of whitespace or the end of the line. If point is farther right than any indentation point in the previous line, the whitespace before point is deleted and the first indentation point then applicable is used. If no indentation point is applicable even then, indent-

relative runs tab-to-tab-stop (see next section).

indent-relative is the definition of ^{⟨}\underline{\overline{TAB}}^{⟩} in Indented Text mode. See Chapter 21 [Text], page 193.

See Section 21.11.6 [Format Indentation], page 217, for another way of specifying the indentation for part of your text.

20.2 Tab Stops

For typing in tables, you can use Text mode’s definition of ^{⟨}\underline{\overline{TAB}}^{⟩}, tabto-tab-stop. This command inserts indentation before point, enough to reach the next tab stop column. If you are not in Text mode, this command can be found on the key M-i.

You can specify the tab stops used by M-i. They are stored in a variable called tab-stop-list, as a list of column-numbers in increasing order.

The convenient way to set the tab stops is with M-x edit-tab-stops, which creates and selects a buffer containing a description of the tab stop settings. You can edit this buffer to specify different tab stops, and then type C-c C-c to make those new tab stops take effect. edit-tab-stops records which buffer was current when you invoked it, and stores the tab stops back in that buffer; normally all buffers share the same tab stops and changing them in one buffer affects all, but if you happen to make tab-stop-list local in one buffer then edit-tab-stops in that buffer will edit the local settings.

Here is what the text representing the tab stops looks like for ordinary tab stops every eight columns.

            :       :       :       :       :       :
    0         1         2         3         4
    0123456789012345678901234567890123456789012345678
    To install changes, type C-c C-c

The first line contains a colon at each tab stop. The remaining lines are present just to help you see where the colons are and know what to do.

Note that the tab stops that control tab-to-tab-stop have nothing to do with displaying tab characters in the buffer. See Section 11.7 [Display Vars], page 81, for more information on that.

20.3 Tabs vs. Spaces

Emacs normally uses both tabs and spaces to indent lines. If you prefer, all indentation can be made from spaces only. To request this, set indenttabs-mode to nil. This is a per-buffer variable; altering the variable affects only the current buffer, but there is a default value which you can change as well. See Section 31.2.4 [Locals], page 393.

There are also commands to convert tabs to spaces or vice versa, always preserving the columns of all nonblank text. M-x tabify scans the region for sequences of spaces, and converts sequences of at least three spaces to tabs if that can be done without changing indentation. M-x untabify changes all tabs in the region to appropriate numbers of spaces.

192 GNU Emacs Manual

21 Commands for Human Languages

The term text has two widespread meanings in our area of the computer field. One is data that is a sequence of characters. Any file that you edit with Emacs is text, in this sense of the word. The other meaning is more restrictive: a sequence of characters in a human language for humans to read (possibly after processing by a text formatter), as opposed to a program or commands for a program.

Human languages have syntactic/stylistic conventions that can be supported or used to advantage by editor commands: conventions involving words, sentences, paragraphs, and capital letters. This chapter describes Emacs commands for all of these things. There are also commands for filling, which means rearranging the lines of a paragraph to be approximately equal in length. The commands for moving over and killing words, sentences and paragraphs, while intended primarily for editing text, are also often useful for editing programs.

Emacs has several major modes for editing human-language text. If the file contains text pure and simple, use Text mode, which customizes Emacs in small ways for the syntactic conventions of text. Outline mode provides special commands for operating on text with an outline structure. See Section 21.8 [Outline Mode], page 204.

For text which contains embedded commands for text formatters, Emacs has other major modes, each for a particular text formatter. Thus, for input to TEX, you would use TEX mode (see Section 21.9 [TeX Mode], page 208). For input to nroff, use Nroff mode.

Instead of using a text formatter, you can edit formatted text in WYSI-WYG style (“what you see is what you get”), with Enriched mode. Then the formatting appears on the screen in Emacs while you edit. See Section 21.11 [Formatted Text], page 213.

21.1 Words

Emacs has commands for moving over or operating on words. By convention, the keys for them are all Meta characters.

M-f Move forward over a word (forward-word).

M-b Move backward over a word (backward-word).

M-d Kill up to the end of a word (kill-word).

M-^{⟨}\underline{\overline{DEL}}^{⟩} Kill back to the beginning of a word (backward-kill-word).

M-@ Mark the end of the next word (mark-word).

M-t Transpose two words or drag a word across other words

          (transpose-words).

Notice how these keys form a series that parallels the character-based C-f, C-b, C-d, ^{⟨}\underline{\overline{DEL}}^{⟩} and C-t. M-@ is cognate to C-@, which is an alias for C-^{⟨}\underline{\overline{SPC}}^{⟩}.

   The commands M-f (forward-word) and M-b (backward-word) move for-

ward and backward over words. These Meta characters are thus analogous to the corresponding control characters, C-f and C-b, which move over single characters in the text. The analogy extends to numeric arguments, which serve as repeat counts. M-f with a negative argument moves backward, and M-b with a negative argument moves forward. Forward motion stops right after the last letter of the word, while backward motion stops right before the first letter.

M-d (kill-word) kills the word after point. To be precise, it kills everything from point to the place M-f would move to. Thus, if point is in the middle of a word, M-d kills just the part after point. If some punctuation comes between point and the next word, it is killed along with the word. (If you wish to kill only the next word but not the punctuation before it, simply do M-f to get the end, and kill the word backwards with M-^{⟨}\underline{\overline{DEL}}^{⟩}.) M-d takes arguments just like M-f.

M-^{⟨}\underline{\overline{DEL}}^{⟩} (backward-kill-word) kills the word before point. It kills everything from point back to where M-b would move to. If point is after the space in ‘FOO, BAR’, then ‘FOO, ’ is killed. (If you wish to kill just ‘FOO’, and not the comma and the space, use M-b M-d instead of M-^{⟨}\underline{\overline{DEL}}^{⟩}.)

M-t (transpose-words) exchanges the word before or containing point with the following word. The delimiter characters between the words do not move. For example, ‘FOO, BAR’ transposes into ‘BAR, FOO’ rather than ‘BAR FOO,’. See Section 13.2 [Transpose], page 99, for more on transposition and on arguments to transposition commands.

To operate on the next n words with an operation which applies between point and mark, you can either set the mark at point and then move over the words, or you can use the command M-@ (mark-word) which does not move point, but sets the mark where M-f would move to. M-@ accepts a numeric argument that says how many words to scan for the place to put the mark. In Transient Mark mode, this command activates the mark.

The word commands’ understanding of syntax is completely controlled by the syntax table. Any character can, for example, be declared to be a word delimiter. See Section 31.6 [Syntax], page 411.

21.2 Sentences

The Emacs commands for manipulating sentences and paragraphs are mostly on Meta keys, so as to be like the word-handling commands.

M-a Move back to the beginning of the sentence (backward-

          sentence).

M-e Move forward to the end of the sentence (forward-sentence).

M-k Kill forward to the end of the sentence (kill-sentence).

C-x ^{⟨}\underline{\overline{DEL}}^{⟩} Kill back to the beginning of the sentence (backward-kill-

          sentence).
   The commands M-a and M-e (backward-sentence and forward-

sentence) move to the beginning and end of the current sentence, respectively. They were chosen to resemble C-a and C-e, which move to the beginning and end of a line. Unlike them, M-a and M-e if repeated or given numeric arguments move over successive sentences.

Moving backward over a sentence places point just before the first character of the sentence; moving forward places point right after the punctuation that ends the sentence. Neither one moves over the whitespace at the sentence boundary.

Just as C-a and C-e have a kill command, C-k, to go with them, so M-a and M-e have a corresponding kill command M-k (kill-sentence) which kills from point to the end of the sentence. With minus one as an argument it kills back to the beginning of the sentence. Larger arguments serve as a repeat count. There is also a command, C-x ^{⟨}\underline{\overline{DEL}}^{⟩} (backward-kill-sentence), for killing back to the beginning of a sentence. This command is useful when you change your mind in the middle of composing text.

The sentence commands assume that you follow the American typist’s convention of putting two spaces at the end of a sentence; they consider a sentence to end wherever there is a ‘.’, ‘?’ or ‘!’ followed by the end of a line or two spaces, with any number of ‘)’, ‘]’, ‘'’, or ‘"’ characters allowed in between. A sentence also begins or ends wherever a paragraph begins or ends.

The variable sentence-end controls recognition of the end of a sentence. It is a regexp that matches the last few characters of a sentence, together with the whitespace following the sentence. Its normal value is

    "[.?!][]\"')]*\\($\\|\t\\|  \\)[ \t\n]*"

This example is explained in the section on regexps. See Section 12.5 [Regexps], page 88.

If you want to use just one space between sentences, you should set

sentence-end to this value:
    "[.?!][]\"')]*\\($\\|\t\\| \\)[ \t\n]*"

You should also set the variable sentence-end-double-space to nil so that the fill commands expect and leave just one space at the end of a sentence. Note that this makes it impossible to distinguish between periods that end sentences and those that indicate abbreviations.

21.3 Paragraphs

The Emacs commands for manipulating paragraphs are also Meta keys.

M-{ Move back to previous paragraph beginning (backward-paragraph).

M-} Move forward to next paragraph end (forward-paragraph).

M-h Put point and mark around this or next paragraph (mark-

          paragraph).

M-{ moves to the beginning of the current or previous paragraph, while M-} moves to the end of the current or next paragraph. Blank lines and text-formatter command lines separate paragraphs and are not considered part of any paragraph. In Fundamental mode, but not in Text mode, an indented line also starts a new paragraph. (If a paragraph is preceded by a blank line, these commands treat that blank line as the beginning of the paragraph.)

In major modes for programs, paragraphs begin and end only at blank lines. This makes the paragraph commands continue to be useful even though there are no paragraphs per se.

When there is a fill prefix, then paragraphs are delimited by all lines which don’t start with the fill prefix. See Section 21.5 [Filling], page 197.

When you wish to operate on a paragraph, you can use the command M-h (mark-paragraph) to set the region around it. Thus, for example, M-h C-w kills the paragraph around or after point. The M-h command puts point at the beginning and mark at the end of the paragraph point was in. In Transient Mark mode, it activates the mark. If point is between paragraphs (in a run of blank lines, or at a boundary), the paragraph following point is surrounded by point and mark. If there are blank lines preceding the first line of the paragraph, one of these blank lines is included in the region.

The precise definition of a paragraph boundary is controlled by the

variables paragraph-separate and paragraph-start. The value of

paragraph-start is a regexp that should match any line that either starts or separates paragraphs. The value of paragraph-separate is another regexp that should match only lines that separate paragraphs without being part of any paragraph (for example, blank lines). Lines that start a new paragraph and are contained in it must match only paragraph-start, not paragraphseparate. For example, in Fundamental mode, paragraph-start is "[

\t\n\f]" and paragraph-separate is "[ \t\f]*$".

Normally it is desirable for page boundaries to separate paragraphs. The default values of these variables recognize the usual separator for pages.

21.4 Pages

Files are often thought of as divided into pages by the formfeed character (ASCII control-L, octal code 014). When you print hardcopy for a file, this character forces a page break; thus, each page of the file goes on a separate page on paper. Most Emacs commands treat the page-separator character just like any other character: you can insert it with C-q C-l, and delete it with ^{⟨}\underline{\overline{DEL}}^{⟩}. Thus, you are free to paginate your file or not. However, since pages are often meaningful divisions of the file, Emacs provides commands to move over them and operate on them.

C-x [ Move point to previous page boundary (backward-page).

C-x ] Move point to next page boundary (forward-page).

C-x C-p Put point and mark around this page (or another page) (mark-

          page).

C-x l Count the lines in this page (count-lines-page).

The C-x [ (backward-page) command moves point to immediately after the previous page delimiter. If point is already right after a page delimiter, it skips that one and stops at the previous one. A numeric argument serves as a repeat count. The C-x ] (forward-page) command moves forward past the next page delimiter.

The C-x C-p command (mark-page) puts point at the beginning of the current page and the mark at the end. The page delimiter at the end is included (the mark follows it). The page delimiter at the front is excluded (point follows it). C-x C-p C-w is a handy way to kill a page to move it elsewhere. If you move to another page delimiter with C-x [ and C-x ], then yank the killed page, all the pages will be properly delimited once again. The reason C-x C-p includes only the following page delimiter in the region is to ensure that.

A numeric argument to C-x C-p is used to specify which page to go to, relative to the current one. Zero means the current page. One means the next page, and −1 means the previous one.

The C-x l command (count-lines-page) is good for deciding where to break a page in two. It prints in the echo area the total number of lines in the current page, and then divides it up into those preceding the current line and those following, as in

    Page has 96 (72+25) lines

Notice that the sum is off by one; this is correct if point is not at the beginning of a line.

The variable page-delimiter controls where pages begin. Its value is a regexp that matches the beginning of a line that separates pages. The normal value of this variable is "^\f", which matches a formfeed character at the beginning of a line.

21.5 Filling Text

Filling text means breaking it up into lines that fit a specified width. Emacs does filling in two ways. In Auto Fill mode, inserting text with selfinserting characters also automatically fills it. There are also explicit fill commands that you can use when editing text leaves it unfilled. When you edit formatted text, you can specify a style of filling for each portion of the text (see Section 21.11 [Formatted Text], page 213).

21.5.1 Auto Fill Mode

Auto Fill mode is a minor mode in which lines are broken automatically when they become too wide. Breaking happens only when you type a ^{⟨}\underline{\overline{SPC}}^{⟩} or ^{⟨}\underline{\overline{RET}}^{⟩}.

M-x auto-fill-mode

Enable or disable Auto Fill mode.

⟨\underline{\overline{SPC}}⟩

⟨\underline{\overline{RET}}^{⟩} In Auto Fill mode, break lines when appropriate.

M-x auto-fill-mode turns Auto Fill mode on if it was off, or off if it was on. With a positive numeric argument it always turns Auto Fill mode on, and with a negative argument always turns it off. You can see when Auto Fill mode is in effect by the presence of the word ‘Fill’ in the mode line, inside the parentheses. Auto Fill mode is a minor mode which is enabled or disabled for each buffer individually. See Section 31.1 [Minor Modes], page 383.

In Auto Fill mode, lines are broken automatically at spaces when they get longer than the desired width. Line breaking and rearrangement takes place only when you type ^{⟨}\underline{\overline{SPC}}^{⟩} or ^{⟨}\underline{\overline{RET}}^{⟩}. If you wish to insert a space or newline without permitting line-breaking, type C-q ^{⟨}\underline{\overline{SPC}}^{⟩} or C-q C-j (recall that a newline is really a control-J). Also, C-o inserts a newline without line breaking.

Auto Fill mode works well with programming-language modes, because it indents new lines with ^{⟨}\underline{\overline{TAB}}^{⟩}. If a line ending in a comment gets too long, the text of the comment is split into two comment lines. Optionally, new comment delimiters are inserted at the end of the first line and the beginning of the second so that each line is a separate comment; the variable commentmulti-line controls the choice (see Section 22.7 [Comments], page 237).

Adaptive filling (see the following section) works for Auto Filling as well as for explicit fill commands. It takes a fill prefix automatically from the second or first line of a paragraph.

Auto Fill mode does not refill entire paragraphs; it can break lines but cannot merge lines. So editing in the middle of a paragraph can result in a paragraph that is not correctly filled. The easiest way to make the paragraph properly filled again is usually with the explicit fill commands.

Many users like Auto Fill mode and want to use it in all text files. The section on init files says how to arrange this permanently for yourself. See Section 31.7 [Init File], page 411.

21.5.2 Explicit Fill Commands

M-q Fill current paragraph (fill-paragraph).

C-x f  Set the fill column (set-fill-column).
M-x fill-region

Fill each paragraph in the region (fill-region).

M-x fill-region-as-paragraph

Fill the region, considering it as one paragraph.

M-s Center a line.

To refill a paragraph, use the command M-q (fill-paragraph). This operates on the paragraph that point is inside, or the one after point if point is between paragraphs. Refilling works by removing all the line-breaks, then inserting new ones where necessary.

To refill many paragraphs, use M-x fill-region, which divides the region into paragraphs and fills each of them.

M-q and fill-region use the same criteria as M-h for finding paragraph boundaries (see Section 21.3 [Paragraphs], page 195). For more control, you can use M-x fill-region-as-paragraph, which refills everything between point and mark. This command deletes any blank lines within the region, so separate blocks of text end up combined into one block.

A numeric argument to M-q causes it to justify the text as well as filling it. This means that extra spaces are inserted to make the right margin line up exactly at the fill column. To remove the extra spaces, use M-q with no argument. (Likewise for fill-region.) Another way to control justification, and choose other styles of filling, is with the justification text property; see Section 21.11.7 [Format Justification], page 218.

The command M-s (center-line) centers the current line within the current fill column. With an argument n, it centers n lines individually and moves past them.

The maximum line width for filling is in the variable fill-column. Altering the value of fill-column makes it local to the current buffer; until that time, the default value is in effect. The default is initially 70. See Section 31.2.4 [Locals], page 393. The easiest way to set fill-column is to use the command C-x f (set-fill-column). With a numeric argument, it uses that as the new fill column. With just C-u as argument, it sets fill-column to the current horizontal position of point.

Emacs commands normally consider a period followed by two spaces or by a newline as the end of a sentence; a period followed by just one space indicates an abbreviation and not the end of a sentence. To preserve the distinction between these two ways of using a period, the fill commands do not break a line after a period followed by just one space.

If the variable sentence-end-double-space is nil, the fill commands expect and leave just one space at the end of a sentence. Ordinarily this variable is t, so the fill commands insist on two spaces for the end of a sentence, as explained above. See Section 21.2 [Sentences], page 194.

If the variable colon-double-space is non-nil, the fill commands put two spaces after a colon.

21.5.3 The Fill Prefix

To fill a paragraph in which each line starts with a special marker (which might be a few spaces, giving an indented paragraph), you can use the fill prefix feature. The fill prefix is a string that Emacs expects every line to start with, and which is not included in filling. You can specify a fill prefix explicitly; Emacs can also deduce the fill prefix automatically (see Section 21.5.4 [Adaptive Fill], page 201).

C-x .  Set the fill prefix (set-fill-prefix).

M-q Fill a paragraph using current fill prefix (fill-paragraph).

M-x fill-individual-paragraphs

Fill the region, considering each change of indentation as starting a new paragraph.

M-x fill-nonuniform-paragraphs

Fill the region, considering only paragraph-separator lines as starting a new paragraph.

To specify a fill prefix, move to a line that starts with the desired prefix, put point at the end of the prefix, and give the command C-x . (set-fill-prefix). That’s a period after the C-x. To turn off the fill prefix, specify an empty prefix: type C-x . with point at the beginning of a line.

When a fill prefix is in effect, the fill commands remove the fill prefix from each line before filling and insert it on each line after filling. Auto Fill mode also inserts the fill prefix automatically when it makes a new line. The C-o command inserts the fill prefix on new lines it creates, when you use it at the beginning of a line (see Section 4.7 [Blank Lines], page 34). Conversely, the command M-^deletes the prefix (if it occurs) after the newline that it deletes (see Chapter 20 [Indentation], page 189).

For example, if fill-column is 40 and you set the fill prefix to ‘;; ’, then M-q in the following text

    ;; This is an
    ;; example of a paragraph
    ;; inside a Lisp-style comment.

produces this:

    ;; This is an example of a paragraph
    ;; inside a Lisp-style comment.

Lines that do not start with the fill prefix are considered to start paragraphs, both in M-q and the paragraph commands; this gives good results for paragraphs with hanging indentation (every line indented except the first one). Lines which are blank or indented once the prefix is removed also separate or start paragraphs; this is what you want if you are writing multi-paragraph comments with a comment delimiter on each line.

   You can use M-x fill-individual-paragraphs to set the fill prefix for

each paragraph automatically. This command divides the region into paragraphs, treating every change in the amount of indentation as the start of a new paragraph, and fills each of these paragraphs. Thus, all the lines in one “paragraph” have the same amount of indentation. That indentation serves as the fill prefix for that paragraph.

   M-x fill-nonuniform-paragraphs is a similar command that divides

the region into paragraphs in a different way. It considers only paragraphseparating lines (as defined by paragraph-separate) as starting a new paragraph. Since this means that the lines of one paragraph may have different amounts of indentation, the fill prefix used is the smallest amount of indentation of any of the lines of the paragraph. This gives good results with styles that indent a paragraph’s first line more or less that the rest of the paragraph.

The fill prefix is stored in the variable fill-prefix. Its value is a string, or nil when there is no fill prefix. This is a per-buffer variable; altering the variable affects only the current buffer, but there is a default value which you can change as well. See Section 31.2.4 [Locals], page 393.

The indentation text property provides another way to control the amount of indentation paragraphs receive. See Section 21.11.6 [Format Indentation], page 217.

21.5.4 Adaptive Filling

The fill commands can deduce the proper fill prefix for a paragraph automatically in certain cases: either whitespace or certain punctuation characters at the beginning of a line are propagated to all lines of the paragraph.

If the paragraph has two or more lines, the fill prefix is taken from the paragraph’s second line, but only if it appears on the first line as well.

If a paragraph has just one line, fill commands may take a prefix from that line. The decision is complicated because there are three reasonable things to do in such a case:

• Use the first line’s prefix on all the lines of the paragraph.

• Indent subsequent lines with whitespace, so that they line up under the text that follows the prefix on the first line, but don’t actually copy the prefix from the first line.

• Don’t do anything special with the second and following lines.

All three of these styles of formatting are commonly used. So the fill commands try to determine what you would like, based on the prefix that appears and on the major mode. Here is how.

If the prefix found on the first line matches adaptive-fill-first-line-regexp, or if it appears to be a comment-starting sequence (this depends on the major mode), then the prefix found is used for filling the paragraph, provided it would not act as a paragraph starter on subsequent lines.

Otherwise, the prefix found is converted to an equivalent number of spaces, and those spaces are used as the fill prefix for the rest of the lines, provided they would not act as a paragraph starter on subsequent lines.

In Text mode, and other modes where only blank lines and page delimiters separate paragraphs, the prefix chosen by adaptive filling never acts as a paragraph starter, so it can always be used for filling.

The variable adaptive-fill-regexp determines what kinds of line beginnings can serve as a fill prefix: any characters at the start of the line that match this regular expression are used. If you set the variable adaptivefill-mode to nil, the fill prefix is never chosen automatically.

You can specify more complex ways of choosing a fill prefix automatically by setting the variable adaptive-fill-function to a function. This function is called with point after the left margin of a line, and it should return the appropriate fill prefix based on that line. If it returns nil, that means it sees no fill prefix in that line.

21.6 Case Conversion Commands

Emacs has commands for converting either a single word or any arbitrary range of text to upper case or to lower case.

M-l Convert following word to lower case (downcase-word).

M-u Convert following word to upper case (upcase-word).

M-c Capitalize the following word (capitalize-word).

C-x C-l Convert region to lower case (downcase-region).

C-x C-u Convert region to upper case (upcase-region).

The word conversion commands are the most useful. M-l (downcaseword) converts the word after point to lower case, moving past it. Thus, repeating M-l converts successive words. M-u (upcase-word) converts to all capitals instead, while M-c (capitalize-word) puts the first letter of the word into upper case and the rest into lower case. All these commands convert several words at once if given an argument. They are especially convenient for converting a large amount of text from all upper case to mixed case, because you can move through the text using M-l, M-u or M-c on each word as appropriate, occasionally using M-f instead to skip a word.

When given a negative argument, the word case conversion commands apply to the appropriate number of words before point, but do not move point. This is convenient when you have just typed a word in the wrong case: you can give the case conversion command and continue typing.

If a word case conversion command is given in the middle of a word, it applies only to the part of the word which follows point. This is just like what M-d (kill-word) does. With a negative argument, case conversion applies only to the part of the word before point.

The other case conversion commands are C-x C-u (upcase-region) and C-x C-l (downcase-region), which convert everything between point and mark to the specified case. Point and mark do not move.

The region case conversion commands upcase-region and downcaseregion are normally disabled. This means that they ask for confirmation if you try to use them. When you confirm, you may enable the command, which means it will not ask for confirmation again. See Section 31.4.10 [Disabling], page 409.

21.7 Text Mode

When you edit files of text in a human language, it’s more convenient to use Text mode rather than Fundamental mode. To enter Text mode, type

M-x text-mode.

In Text mode, only blank lines and page delimiters separate paragraphs. As a result, paragraphs can be indented, and adaptive filling determines what indentation to use when filling a paragraph. See Section 21.5.4 [Adaptive Fill], page 201.

Text mode defines ^{⟨}\underline{\overline{TAB}}^{⟩} to run indent-relative (see Chapter 20 [Indentation], page 189), so that you can conveniently indent a line like the previous line. When the previous line is not indented, indent-relative runs tab-to-tab-stop, which uses Emacs tab stops that you can set (see Section 20.2 [Tab Stops], page 190).

Text mode turns off the features concerned with comments except when you explicitly invoke them. It changes the syntax table so that periods are not considered part of a word, while apostrophes, backspaces and underlines are considered part of words.

If you indent the first lines of paragraphs, then you should use Paragraph-Indent Text mode rather than Text mode. In this mode, you do not need to have blank lines between paragraphs, because the first-line indentation is sufficient to start a paragraph; however paragraphs in which every line is indented are not supported. Use M-x paragraph-indent-text-mode to enter this mode.

Text mode, and all the modes based on it, define M-^{⟨}\underline{\overline{TAB}}^{⟩} as the command ispell-complete-word, which performs completion of the partial word in the buffer before point, using the spelling dictionary as the space of possible words. See Section 13.4 [Spelling], page 101.

Entering Text mode runs the hook text-mode-hook. Other major modes related to Text mode also run this hook, followed by hooks of their own; this includes Paragraph-Indent Text mode, Nroff mode, TEX mode, Outline mode, and Mail mode. Hook functions on text-mode-hook can look at the value of major-mode to see which of these modes is actually being entered. See Section 31.2.3 [Hooks], page 392.

21.8 Outline Mode

Outline mode is a major mode much like Text mode but intended for editing outlines. It allows you to make parts of the text temporarily invisible so that you can see the outline structure. Type M-x outline-mode to switch to Outline mode as the major mode of the current buffer.

When Outline mode makes a line invisible, the line does not appear on the screen. The screen appears exactly as if the invisible line were deleted, except that an ellipsis (three periods in a row) appears at the end of the previous visible line (only one ellipsis no matter how many invisible lines follow).

Editing commands that operate on lines, such as C-n and C-p, treat the text of the invisible line as part of the previous visible line. Killing an entire visible line, including its terminating newline, really kills all the following invisible lines along with it.

Outline minor mode provides the same commands as the major mode, Outline mode, but you can use it in conjunction with other major modes. Type M-x outline-minor-mode to enable the Outline minor mode in the current buffer. You can also specify this in the text of a file, with a file local variable of the form ‘mode: outline-minor’ (see Section 31.2.5 [File Variables], page 395).

The major mode, Outline mode, provides special key bindings on the C-c prefix. Outline minor mode provides similar bindings with C-c @ as the prefix; this is to reduce the conflicts with the major mode’s special commands. (The variable outline-minor-mode-prefix controls the prefix used.)

Entering Outline mode runs the hook text-mode-hook followed by the hook outline-mode-hook (see Section 31.2.3 [Hooks], page 392).

21.8.1 Format of Outlines

Outline mode assumes that the lines in the buffer are of two types: heading lines and body lines. A heading line represents a topic in the outline. Heading lines start with one or more stars; the number of stars determines the depth of the heading in the outline structure. Thus, a heading line with one star is a major topic; all the heading lines with two stars between it and the next one-star heading are its subtopics; and so on. Any line that is not a heading line is a body line. Body lines belong with the preceding heading line. Here is an example:

    * Food
    This is the body,

    which says something about the topic of food.
    ** Delicious Food
    This is the body of the second-level header.
    ** Distasteful Food
    This could have
    a body too, with
    several lines.
    *** Dormitory Food
    * Shelter
    Another first-level topic with its header line.

A heading line together with all following body lines is called collectively an entry. A heading line together with all following deeper heading lines and their body lines is called a subtree.

You can customize the criterion for distinguishing heading lines by setting the variable outline-regexp. Any line whose beginning has a match for this regexp is considered a heading line. Matches that start within a line (not at the left margin) do not count. The length of the matching text determines the level of the heading; longer matches make a more deeply nested level. Thus, for example, if a text formatter has commands ‘@chapter’, ‘@section’ and ‘@subsection’ to divide the document into chapters and sections, you could make those lines count as heading lines by setting outline-regexp to

‘"@chap\\|@\\(sub\\)*section"’. Note the trick: the two words ‘chapter’

and ‘section’ are equally long, but by defining the regexp to match only ‘chap’ we ensure that the length of the text matched on a chapter heading is shorter, so that Outline mode will know that sections are contained in chapters. This works as long as no other command starts with ‘@chap’.

It is possible to change the rule for calculating the level of a heading line by setting the variable outline-level. The value of outline-level should be a function that takes no arguments and returns the level of the current heading. Some major modes such as C, Nroff, and Emacs Lisp mode set this variable in order to work with Outline minor mode.

21.8.2 Outline Motion Commands

Outline mode provides special motion commands that move backward and forward to heading lines.

C-c C-n Move point to the next visible heading line (outline-next-

          visible-heading).

C-c C-p Move point to the previous visible heading line (outline-

          previous-visible-heading).

C-c C-f Move point to the next visible heading line at the same level as

          the one point is on (outline-forward-same-level).

C-c C-b Move point to the previous visible heading line at the same level

          (outline-backward-same-level).

C-c C-u Move point up to a lower-level (more inclusive) visible heading

          line (outline-up-heading).
   C-c C-n (outline-next-visible-heading) moves down to the next
heading line. C-c C-p (outline-previous-visible-heading) moves sim-

ilarly backward. Both accept numeric arguments as repeat counts. The names emphasize that invisible headings are skipped, but this is not really a special feature. All editing commands that look for lines ignore the invisible lines automatically.

More powerful motion commands understand the level structure of head-

ings. C-c C-f (outline-forward-same-level) and C-c C-b (outline-

backward-same-level) move from one heading line to another visible heading at the same depth in the outline. C-c C-u (outline-up-heading) moves backward to another heading that is less deeply nested.

21.8.3 Outline Visibility Commands

The other special commands of outline mode are used to make lines visible or invisible. Their names all start with hide or show. Most of them fall into pairs of opposites. They are not undoable; instead, you can undo right past them. Making lines visible or invisible is simply not recorded by the undo mechanism.

C-c C-t Make all body lines in the buffer invisible (hide-body).

C-c C-a Make all lines in the buffer visible (show-all).

C-c C-d Make everything under this heading invisible, not including this heading itself (hide-subtree).

C-c C-s Make everything under this heading visible, including body, subheadings, and their bodies (show-subtree).

C-c C-l Make the body of this heading line, and of all its subheadings, invisible (hide-leaves).

C-c C-k Make all subheadings of this heading line, at all levels, visible

          (show-branches).

C-c C-i Make immediate subheadings (one level down) of this heading line

          visible (show-children).

C-c C-c Make this heading line’s body invisible (hide-entry).

C-c C-e Make this heading line’s body visible (show-entry).

C-c C-q Hide everything except the top n levels of heading lines (hide-

          sublevels).

C-c C-o Hide everything except for the heading or body that point is in, plus the headings leading up from there to the top level of the

          outline (hide-other).

Two commands that are exact opposites are C-c C-c (hide-entry) and C-c C-e (show-entry). They are used with point on a heading line, and apply only to the body lines of that heading. Subheadings and their bodies are not affected.

Two more powerful opposites are C-c C-d (hide-subtree) and C-c C-s (show-subtree). Both expect to be used when point is on a heading line, and both apply to all the lines of that heading’s subtree: its body, all its subheadings, both direct and indirect, and all of their bodies. In other words, the subtree contains everything following this heading line, up to and not including the next heading of the same or higher rank.

Intermediate between a visible subtree and an invisible one is having all the subheadings visible but none of the body. There are two commands for doing this, depending on whether you want to hide the bodies or make the subheadings visible. They are C-c C-l (hide-leaves) and C-c C-k (show-

branches).
   A little weaker than show-branches is C-c C-i (show-children). It

makes just the direct subheadings visible—those one level down. Deeper subheadings remain invisible, if they were invisible.

Two commands have a blanket effect on the whole file. C-c C-t (hidebody) makes all body lines invisible, so that you see just the outline structure. C-c C-a (show-all) makes all lines visible. These commands can be thought of as a pair of opposites even though C-c C-a applies to more than just body lines.

The command C-c C-q (hide-sublevels) hides all but the top level headings. With a numeric argument n, it hides everything except the top n levels of heading lines.

The command C-c C-o (hide-other) hides everything except the heading or body text that point is in, plus its parents (the headers leading up from there to top level in the outline).

You can turn off the use of ellipses at the ends of visible lines by setting selective-display-ellipses to nil. Then there is no visible indication of the presence of invisible lines.

When incremental search finds text that is hidden by Outline mode, it makes that part of the buffer visible. If you exit the search at that position, the text remains visible.

21.8.4 Viewing One Outline in Multiple Views

You can display two views of a single outline at the same time, in different windows. To do this, you must create an indirect buffer using M-x make-indirect-buffer. The first argument of this command is the existing outline buffer name, and its second argument is the name to use for the new indirect buffer. See Section 15.6 [Indirect Buffers], page 147.

Once the indirect buffer exists, you can display it in a window in the normal fashion, with C-x 4 b or other Emacs commands. The Outline mode commands to show and hide parts of the text operate on each buffer independently; as a result, each buffer can have its own view. If you want more than two views on the same outline, create additional indirect buffers.

21.9 TEX Mode

TEX is a powerful text formatter written by Donald Knuth; it is also free, like GNU Emacs. LaTEX is a simplified input format for TEX, implemented by TEX macros; it comes with TEX. SliTEX is a special form of LaTEX.

Emacs has a special TEX mode for editing TEX input files. It provides facilities for checking the balance of delimiters and for invoking TEX on all or part of the file.

TEX mode has three variants, Plain TEX mode, LaTEX mode, and SliTEX mode (these three distinct major modes differ only slightly). They are designed for editing the three different formats. The command M-x tex-mode looks at the contents of the buffer to determine whether the contents appear to be either LaTEX input or SliTEX input; if so, it selects the appropriate mode. If the file contents do not appear to be LaTEX or SliTEX, it selects Plain TEX mode. If the contents are insufficient to determine this, the variable tex-default-mode controls which mode is used.

When M-x tex-mode does not guess right, you can use the commands

M-x plain-tex-mode, M-x latex-mode, and M-x slitex-mode to select ex-

plicitly the particular variants of TEX mode.

21.9.1 TEX Editing Commands

Here are the special commands provided in TEX mode for editing the text of the file.

" Insert, according to context, either ‘``’ or ‘"’ or ‘''’ (tex-

          insert-quote).

C-j Insert a paragraph break (two newlines) and check the previous paragraph for unbalanced braces or dollar signs (tex-

          terminate-paragraph).

M-x validate-tex-region

Check each paragraph in the region for unbalanced braces or dollar signs.

C-c { Insert ‘{}’ and position point between them (tex-insert-

          braces).

C-c } Move forward past the next unmatched close brace (up-list).

In TEX, the character ‘"’ is not normally used; we use ‘``’ to start a quotation and ‘''’ to end one. To make editing easier under this formatting convention, TEX mode overrides the normal meaning of the key " with a command that inserts a pair of single-quotes or backquotes (tex-insert-quote). To be precise, this command inserts ‘``’ after whitespace or an open brace, ‘"’ after a backslash, and ‘''’ after any other character.

If you need the character ‘"’ itself in unusual contexts, use C-q to insert it. Also, " with a numeric argument always inserts that number of ‘"’ characters. You can turn off the feature of " expansion by eliminating that binding in the local map (See Section 31.4 [Key Bindings], page 400).

In TEX mode, ‘$’ has a special syntax code which attempts to understand the way TEX math mode delimiters match. When you insert a ‘$’ that is meant to exit math mode, the position of the matching ‘$’ that entered math mode is displayed for a second. This is the same feature that displays the open brace that matches a close brace that is inserted. However, there is no way to tell whether a ‘$’ enters math mode or leaves it; so when you insert a ‘$’ that enters math mode, the previous ‘$’ position is shown as if it were a match, even though they are actually unrelated.

TEX uses braces as delimiters that must match. Some users prefer to keep braces balanced at all times, rather than inserting them singly. Use C-c { (tex-insert-braces) to insert a pair of braces. It leaves point between the two braces so you can insert the text that belongs inside. Afterward, use the command C-c } (up-list) to move forward past the close brace.

There are two commands for checking the matching of braces. C-j (texterminate-paragraph) checks the paragraph before point, and inserts two newlines to start a new paragraph. It prints a message in the echo area if any mismatch is found. M-x validate-tex-region checks a region, paragraph by paragraph. When it finds a paragraph that contains a mismatch, it displays point at the beginning of the paragraph for a few seconds and sets the mark at that spot. Scanning continues until the whole buffer has been checked or until you type another key. Afterward, you can use the mark ring to find the last several paragraphs that had mismatches (see Section 8.5 [Mark Ring], page 61).

Note that Emacs commands count square brackets and parentheses in TEX mode, not just braces. This is not strictly correct for the purpose of checking TEX syntax. However, parentheses and square brackets are likely to be used in text as matching delimiters and it is useful for the various motion commands and automatic match display to work with them.

21.9.2 LaTEX Editing Commands

LaTEX mode, and its variant, SliTEX mode, provide a few extra features not applicable to plain TEX.

C-c C-o Insert ‘\begin’ and ‘\end’ for LaTEX block and position point on a line between them (tex-latex-block).

C-c C-e Close the innermost LaTEX block not yet closed (tex-close-

          latex-block).

In LaTEX input, ‘\begin’ and ‘\end’ commands are used to group blocks of text. To insert a ‘\begin’ and a matching ‘\end’ (on a new line following the ‘\begin’), use C-c C-o (tex-latex-block). A blank line is inserted between the two, and point is left there. You can use completion when you enter the block type; to specify additional block type names beyond the standard list, set the variable latex-block-names. For example, here’s how

to add ‘theorem’, ‘corollary’, and ‘proof’:
    (setq latex-block-names '("theorem" "corollary" "proof"))

In LaTEX input, ‘\begin’ and ‘\end’ commands must balance. You can use C-c C-e (tex-close-latex-block) to insert automatically a matching ‘\end’ to match the last unmatched ‘\begin’. It indents the ‘\end’ to match the corresponding ‘\begin’. It inserts a newline after ‘\end’ if point is at the beginning of a line.

21.9.3 TEX Printing Commands

You can invoke TEX as an inferior of Emacs on either the entire contents of the buffer or just a region at a time. Running TEX in this way on just one chapter is a good way to see what your changes look like without taking the time to format the entire file.

C-c C-r Invoke TEX on the current region, together with the buffer’s

          header (tex-region).

C-c C-b Invoke TEX on the entire current buffer (tex-buffer).

C-c ^{⟨}\underline{\overline{TAB}}^{⟩} Invoke BibTEX on the current file (tex-bibtex-file).

C-c C-f Invoke TEX on the current file (tex-file).

C-c C-l Recenter the window showing output from the inferior TEX so that the last line can be seen (tex-recenter-output-buffer).

C-c C-k Kill the TEX subprocess (tex-kill-job).

C-c C-p Print the output from the last C-c C-r, C-c C-b, or C-c C-f com-

          mand (tex-print).

C-c C-v Preview the output from the last C-c C-r, C-c C-b, or C-c C-f command (tex-view).

C-c C-q Show the printer queue (tex-show-print-queue).

You can pass the current buffer through an inferior TEX by means of C-c C-b (tex-buffer). The formatted output appears in a temporary file; to print it, type C-c C-p (tex-print). Afterward, you can use C-c C-q (tex-show-print-queue) to view the progress of your output towards being printed. If your terminal has the ability to display TEX output files, you can preview the output on the terminal with C-c C-v (tex-view).

You can specify the directory to use for running TEX by setting the variable tex-directory. "." is the default value. If your environment variable TEXINPUTS contains relative directory names, or if your files contains ‘\input’ commands with relative file names, then tex-directory must be "." or you will get the wrong results. Otherwise, it is safe to specify some other directory, such as "/tmp".

If you want to specify which shell commands are used in the inferior TEX, you can do so by setting the values of the variables tex-run-

command, latex-run-command, slitex-run-command, tex-dvi-print-
command, tex-dvi-view-command, and tex-show-queue-command. You

must set the value of tex-dvi-view-command for your particular terminal; this variable has no default value. The other variables have default values that may (or may not) be appropriate for your system.

Normally, the file name given to these commands comes at the end of the command string; for example, ‘latex filename’. In some cases, however, the file name needs to be embedded in the command; an example is when you need to provide the file name as an argument to one command whose output is piped to another. You can specify where to put the file name with ‘*’ in the command string. For example,

    (setq tex-dvi-print-command "dvips -f * | lpr")

The terminal output from TEX, including any error messages, appears in a buffer called ‘*tex-shell*’. If TEX gets an error, you can switch to this buffer and feed it input (this works as in Shell mode; see Section 30.2.2 [Interactive Shell], page 360). Without switching to this buffer you can scroll it so that its last line is visible by typing C-c C-l.

Type C-c C-k (tex-kill-job) to kill the TEX process if you see that its output is no longer useful. Using C-c C-b or C-c C-r also kills any TEX process still running.

You can also pass an arbitrary region through an inferior TEX by typing C-c C-r (tex-region). This is tricky, however, because most files of TEX input contain commands at the beginning to set parameters and define macros, without which no later part of the file will format correctly. To solve this problem, C-c C-r allows you to designate a part of the file as containing essential commands; it is included before the specified region as part of the input to TEX. The designated part of the file is called the header.

To indicate the bounds of the header in Plain TEX mode, you insert two special strings in the file. Insert ‘%**start of header’ before the header, and ‘%**end of header’ after it. Each string must appear entirely on one line, but there may be other text on the line before or after. The lines containing the two strings are included in the header. If ‘%**start of header’ does not appear within the first 100 lines of the buffer, C-c C-r assumes that there is no header.

In LaTEX mode, the header begins with ‘\documentstyle’ and ends with ‘\begin{document}’. These are commands that LaTEX requires you to use in any case, so nothing special needs to be done to identify the header.

The commands (tex-buffer) and (tex-region) do all of their work in a temporary directory, and do not have available any of the auxiliary files needed by TEX for cross-references; these commands are generally not suitable for running the final copy in which all of the cross-references need to be correct.

When you want the auxiliary files for cross references, use C-c C-f (texfile) which runs TEX on the current buffer’s file, in that file’s directory. Before running TEX, it offers to save any modified buffers. Generally, you need to use (tex-file) twice to get the cross-references right.

Large TEX documents are often split into several files—one main file, plus subfiles. Running TEX on a subfile typically does not work; you have to run it on the main file. In order to make tex-file useful when you are editing a subfile, you can set the variable tex-main-file to the name of the main file. Then tex-file runs TEX on that file.

The most convenient way to use tex-main-file is to specify it in a local variable list in each of the subfiles. See Section 31.2.5 [File Variables], page 395.

For LaTEX files, you can use BibTEX to process the auxiliary file for the current buffer’s file. BibTEX looks up bibliographic citations in a data base and prepares the cited references for the bibliography section. The com-

mand C-c TAB (tex-bibtex-file) runs the shell command (tex-bibtex-

command) to produce a ‘.bbl’ file for the current buffer’s file. Generally, you need to do C-c C-f (tex-file) once to generate the ‘.aux’ file, then do C-c

TAB (tex-bibtex-file), and then repeat C-c C-f (tex-file) twice more

to get the cross-references correct.

Entering any kind of TEX mode runs the hooks text-mode-hook and

tex-mode-hook. Then it runs either plain-tex-mode-hook or latex-mode-

hook, whichever is appropriate. For SliTEX files, it calls slitex-mode-hook. Starting the TEX shell runs the hook tex-shell-hook. See Section 31.2.3 [Hooks], page 392.

21.10 Nroff Mode

Nroff mode is a mode like Text mode but modified to handle nroff commands present in the text. Invoke M-x nroff-mode to enter this mode. It differs from Text mode in only a few ways. All nroff command lines are considered paragraph separators, so that filling will never garble the nroff commands. Pages are separated by ‘.bp’ commands. Comments start with backslash-doublequote. Also, three special commands are provided that are not in Text mode:

M-n Move to the beginning of the next line that isn’t an nroff command (forward-text-line). An argument is a repeat count.

M-p  Like M-n but move up (backward-text-line).

M-? Prints in the echo area the number of text lines (lines that are not nroff commands) in the region (count-text-lines).

The other feature of Nroff mode is that you can turn on Electric Nroff mode. This is a minor mode that you can turn on or off with M-x electric-nroff-mode (see Section 31.1 [Minor Modes], page 383). When the mode is on, each time you use ^{⟨}\underline{\overline{RET}}^{⟩} to end a line that contains an nroff command that opens a kind of grouping, the matching nroff command to close that grouping is automatically inserted on the following line. For example, if you are at the beginning of a line and type . ( b ^{⟨}\underline{\overline{RET}}^{⟩}, this inserts the matching command ‘.)b’ on a new line following point.

If you use Outline minor mode with Nroff mode (see Section 21.8 [Outline Mode], page 204), heading lines are lines of the form ‘.H’ followed by a number (the header level).

Entering Nroff mode runs the hook text-mode-hook, followed by the hook nroff-mode-hook (see Section 31.2.3 [Hooks], page 392).

21.11 Editing Formatted Text

Enriched mode is a minor mode for editing files that contain formatted text in WYSIWYG fashion, as in a word processor. Currently, formatted text in Enriched mode can specify fonts, colors, underlining, margins, and types of filling and justification. In the future, we plan to implement other formatting features as well.

Enriched mode is a minor mode (see Section 31.1 [Minor Modes], page 383). Typically it is used in conjunction with Text mode (see Section 21.7 [Text Mode], page 203). However, you can also use it with other major modes such as Outline mode and Paragraph-Indent Text mode.

Potentially, Emacs can store formatted text files in various file formats. Currently, only one format is implemented: text/enriched format, which is defined by the MIME protocol. See section “Format Conversion” in the Emacs Lisp Reference Manual, for details of how Emacs recognizes and converts file formats.

The Emacs distribution contains a formatted text file that can serve as an example. Its name is ‘etc/enriched.doc’. It contains samples illustrating all the features described in this section. It also contains a list of ideas for future enhancements.

21.11.1 Requesting to Edit Formatted Text

Whenever you visit a file that Emacs saved in the text/enriched format, Emacs automatically converts the formatting information in the file into Emacs’s own internal format (text properties), and turns on Enriched mode.

To create a new file of formatted text, first visit the nonexistent file, then type M-x enriched-mode before you start inserting text. This command turns on Enriched mode. Do this before you begin inserting text, to ensure that the text you insert is handled properly.

More generally, the command enriched-mode turns Enriched mode on if it was off, and off if it was on. With a prefix argument, this command turns Enriched mode on if the argument is positive, and turns the mode off otherwise.

When you save a buffer while Enriched mode is enabled in it, Emacs automatically converts the text to text/enriched format while writing it into the file. When you visit the file again, Emacs will automatically recognize the format, reconvert the text, and turn on Enriched mode again.

Normally, after visiting a file in text/enriched format, Emacs refills each paragraph to fit the specified right margin. You can turn off this refilling, to

save time, by setting the variable enriched-fill-after-visiting to nil

or to ask.

However, when visiting a file that was saved from Enriched mode, there is no need for refilling, because Emacs saves the right margin settings along with the text.

You can add annotations for saving additional text properties, which Emacs normally does not save, by adding to enriched-translations. Note that the text/enriched standard requires any non-standard annotations to have names starting with ‘x-’, as in ‘x-read-only’. This ensures that they will not conflict with standard annotations that may be added later.

21.11.2 Hard and Soft Newlines

In formatted text, Emacs distinguishes between two different kinds of newlines, hard newlines and soft newlines.

Hard newlines are used to separate paragraphs, or items in a list, or anywhere that there should always be a line break regardless of the margins. The ^{⟨}\underline{\overline{RET}}^{⟩} command (newline) and C-o (open-line) insert hard newlines.

Soft newlines are used to make text fit between the margins. All the fill commands, including Auto Fill, insert soft newlines—and they delete only soft newlines.

Although hard and soft newlines look the same, it is important to bear the difference in mind. Do not use ^{⟨}\underline{\overline{RET}}^{⟩} to break lines in the middle of filled paragraphs, or else you will get hard newlines that are barriers to further filling. Instead, let Auto Fill mode break lines, so that if the text or the margins change, Emacs can refill the lines properly. See Section 21.5.1 [Auto Fill], page 198.

On the other hand, in tables and lists, where the lines should always remain as you type them, you can use ^{⟨}\underline{\overline{RET}}^{⟩} to end lines. For these lines, you may also want to set the justification style to unfilled. See Section 21.11.7 [Format Justification], page 218.

21.11.3 Editing Format Information

There are two ways to alter the formatting information for a formatted text file: with keyboard commands, and with the mouse.

The easiest way to add properties to your document is by using the Text Properties menu. You can get to this menu in two ways: from the Edit menu in the menu bar, or with C-mouse-2 (hold the ^{⟨}\underline{\overline{CTRL}}^{⟩} key and press the middle mouse button).

Most of the items in the Text Properties menu lead to other submenus. These are described in the sections that follow. Some items run commands directly:

Remove Properties

Delete from the region all the text properties that the Text Properties menu works with (facemenu-remove-props).

Remove All

Delete all text properties from the region (facemenu-remove-

          all).
List Properties

List all the text properties of the character following point

          (list-text-properties-at).
Display Faces

Display a list of all the defined faces.

Display Colors

Display a list of all the defined colors.

21.11.4 Faces in Formatted Text

The Faces submenu lists various Emacs faces including bold, italic, and underline. Selecting one of these adds the chosen face to the region. See Section 17.13 [Faces], page 164. You can also specify a face with these keyboard commands:

M-g d Set the region, or the next inserted character, to the default face

          (facemenu-set-default).

M-g b Set the region, or the next inserted character, to the bold face

          (facemenu-set-bold).

M-g i Set the region, or the next inserted character, to the italic face

          (facemenu-set-italic).

M-g l Set the region, or the next inserted character, to the bold-italic

          face (facemenu-set-bold-italic).

M-g u Set the region, or the next inserted character, to the underline

          face (facemenu-set-underline).

M-g o face ^{⟨}\underline{\overline{RET}}^{⟩} Set the region, or the next inserted character, to the face face

          (facemenu-set-face).

If you use these commands with a prefix argument—or, in Transient Mark mode, if the region is not active—then these commands specify a face to use for your next self-inserting input. See Section 8.2 [Transient Mark], page 58. This applies to both the keyboard commands and the menu commands.

Enriched mode defines two additional faces: excerpt and fixed. These correspond to codes used in the text/enriched file format.

The excerpt face is intended for quotations. This face is the same as italic unless you customize it (see Section 31.2.2.3 [Face Customization], page 390).

The fixed face is meant to say, “Use a fixed-width font for this part of the text.” Emacs currently supports only fixed-width fonts; therefore, the fixed annotation is not necessary now. However, we plan to support variable width fonts in future Emacs versions, and other systems that display text/enriched format may not use a fixed-width font as the default. So if you specifically want a certain part of the text to use a fixed-width font, you should specify the fixed face for that part.

The fixed face is normally defined to use a different font from the default. However, different systems have different fonts installed, so you may need to customize this.

If your terminal cannot display different faces, you will not be able to see them, but you can still edit documents containing faces. You can even add faces and colors to documents. They will be visible when the file is viewed on a terminal that can display them.

21.11.5 Colors in Formatted Text

You can specify foreground and background colors for portions of the text. There is a menu for specifying the foreground color and a menu for specifying the background color. Each color menu lists all the colors that you have used in Enriched mode in the current Emacs session.

If you specify a color with a prefix argument—or, in Transient Mark mode, if the region is not active—then it applies to your next self-inserting input. See Section 8.2 [Transient Mark], page 58. Otherwise, the command applies to the region.

Each color menu contains one additional item: ‘Other’. You can use this item to specify a color that is not listed in the menu; it reads the color name with the minibuffer. To display list of available colors and their names, use the ‘Display Colors’ menu item in the Text Properties menu (see Section 21.11.3 [Editing Format Info], page 215).

Any color that you specify in this way, or that is mentioned in a formatted text file that you read in, is added to both color menus for the duration of the Emacs session.

There are no key bindings for specifying colors, but you can do so

with the extended commands M-x facemenu-set-foreground and M-x

facemenu-set-background. Both of these commands read the name of the color with the minibuffer.

21.11.6 Indentation in Formatted Text

When editing formatted text, you can specify different amounts of indentation for the right or left margin of an entire paragraph or a part of a paragraph. The margins you specify automatically affect the Emacs fill commands (see Section 21.5 [Filling], page 197) and line-breaking commands.

The Indentation submenu provides a convenient interface for specifying these properties. The submenu contains four items:

Indent More

Indent the region by 4 columns (increase-left-margin). In Enriched mode, this command is also available on C-x ^{⟨}\underline{\overline{TAB}}^{⟩}; if you supply a numeric argument, that says how many columns to add to the margin (a negative argument reduces the number of columns).

Indent Less

Remove 4 columns of indentation from the region.

Indent Right More

Make the text narrower by indenting 4 columns at the right margin.

Indent Right Less

Remove 4 columns of indentation from the right margin.

You can use these commands repeatedly to increase or decrease the indentation.

The most common way to use these commands is to change the indentation of an entire paragraph. However, that is not the only use. You can change the margins at any point; the new values take effect at the end of the line (for right margins) or the beginning of the next line (for left margins).

This makes it possible to format paragraphs with hanging indents, which means that the first line is indented less than subsequent lines. To set up a hanging indent, increase the indentation of the region starting after the first word of the paragraph and running until the end of the paragraph.

Indenting the first line of a paragraph is easier. Set the margin for the whole paragraph where you want it to be for the body of the paragraph, then indent the first line by inserting extra spaces or tabs.

Sometimes, as a result of editing, the filling of a paragraph becomes messed up—parts of the paragraph may extend past the left or right margins. When this happens, use M-q (fill-paragraph) to refill the paragraph.

The variable standard-indent specifies how many columns these commands should add to or subtract from the indentation. The default value is 4. The overall default right margin for Enriched mode is controlled by the variable fill-column, as usual.

The fill prefix, if any, works in addition to the specified paragraph indentation: C-x . does not include the specified indentation’s whitespace in the new value for the fill prefix, and the fill commands look for the fill prefix after the indentation on each line. See Section 21.5.3 [Fill Prefix], page 200.

21.11.7 Justification in Formatted Text

When editing formatted text, you can specify various styles of justification for a paragraph. The style you specify automatically affects the Emacs fill commands.

The Justification submenu provides a convenient interface for specifying the style. The submenu contains five items:

Flush Left

This is the most common style of justification (at least for English). Lines are aligned at the left margin but left uneven at the right.

Flush Right

This aligns each line with the right margin. Spaces and tabs are added on the left, if necessary, to make lines line up on the right.

Full This justifies the text, aligning both edges of each line. Justified text looks very nice in a printed book, where the spaces can all be adjusted equally, but it does not look as nice with a fixedwidth font on the screen. Perhaps a future version of Emacs will be able to adjust the width of spaces in a line to achieve elegant justification.

Center This centers every line between the current margins.

None This turns off filling entirely. Each line will remain as you wrote it; the fill and auto-fill functions will have no effect on text which has this setting. You can, however, still indent the left margin. In unfilled regions, all newlines are treated as hard newlines (see Section 21.11.2 [Hard and Soft Newlines], page 214) .

In Enriched mode, you can also specify justification from the keyboard using the M-j prefix character:

M-j l  Make the region left-filled (set-justification-left).
M-j r  Make the region right-filled (set-justification-right).

M-j f Make the region fully-justified (set-justification-full).

M-j c
M-S  Make the region centered (set-justification-center).
M-j u  Make the region unfilled (set-justification-none).

Justification styles apply to entire paragraphs. All the justificationchanging commands operate on the paragraph containing point, or, if the region is active, on all paragraphs which overlap the region.

The default justification style is specified by the variable defaultjustification. Its value should be one of the symbols left, right, full,

center, or none.

21.11.8 Setting Other Text Properties

The Other Properties menu lets you add or remove three other useful text

properties: read-only, invisible and intangible. The intangible prop-

erty disallows moving point within the text, the invisible text property hides text from display, and the read-only property disallows alteration of the text.

Each of these special properties has a menu item to add it to the region. The last menu item, ‘Remove Special’, removes all of these special properties from the text in the region.

Currently, the invisible and intangible properties are not saved in the text/enriched format. The read-only property is saved, but it is not a standard part of the text/enriched format, so other editors may not respect it.

21.11.9 Forcing Enriched Mode

Normally, Emacs knows when you are editing formatted text because it recognizes the special annotations used in the file that you visited. However, there are situations in which you must take special actions to convert file contents or turn on Enriched mode:

• When you visit a file that was created with some other editor, Emacs may not recognize the file as being in the text/enriched format. In this case, when you visit the file you will see the formatting commands rather than the formatted text. Type M-x format-decode-buffer to translate it.

• When you insert a file into a buffer, rather than visiting it. Emacs does the necessary conversions on the text which you insert, but it does not enable Enriched mode. If you wish to do that, type M-x enriched-mode. The command format-decode-buffer translates text in various formats

into Emacs’s internal format. It asks you to specify the format to translate from; however, normally you can type just ^{⟨}\underline{\overline{RET}}^{⟩}, which tells Emacs to guess the format.

If you wish to look at text/enriched file in its raw form, as a sequence of characters rather than as formatted text, use the M-x find-file-literally command. This visits a file, like find-file, but does not do format conversion. It also inhibits character code conversion (see Section 18.6 [Coding Systems], page 175) and automatic uncompression (see Section 14.11 [Compressed Files], page 138). To disable format conversion but allow character code conversion and/or automatic uncompression if appropriate, use format-find-file with suitable arguments.

22 Editing Programs

Emacs has many commands designed to understand the syntax of programming languages such as Lisp and C. These commands can

• Move over or kill balanced expressions or sexps (see Section 22.2 [Lists], page 222).

• Move over or mark top-level expressions—defuns, in Lisp; functions, in C (see Section 22.4 [Defuns], page 224).

• Show how parentheses balance (see Section 22.6 [Matching], page 236).

• Insert, kill or align comments (see Section 22.7 [Comments], page 237).

• Follow the usual indentation conventions of the language (see Section 22.5 [Program Indent], page 225).

The commands for words, sentences and paragraphs are very useful in editing code even though their canonical application is for editing human language text. Most symbols contain words (see Section 21.1 [Words], page 193); sentences can be found in strings and comments (see Section 21.2 [Sentences], page 194). Paragraphs per se don’t exist in code, but the paragraph commands are useful anyway, because programming language major modes define paragraphs to begin and end at blank lines (see Section 21.3 [Paragraphs], page 195). Judicious use of blank lines to make the program clearer will also provide useful chunks of text for the paragraph commands to work on.

The selective display feature is useful for looking at the overall structure of a function (see Section 11.4 [Selective Display], page 79). This feature causes only the lines that are indented less than a specified amount to appear on the screen.

22.1 Major Modes for Programming Languages

Emacs also has major modes for the programming languages Lisp, Scheme (a variant of Lisp), Awk, C, C++, Fortran, Icon, Java, Objective-C, Pascal, Perl and Tcl. There is also a major mode for makefiles, called Makefile mode. An second alternative mode for Perl is called CPerl mode.

Ideally, a major mode should be implemented for each programming language that you might want to edit with Emacs; but often the mode for one language can serve for other syntactically similar languages. The language modes that exist are those that someone decided to take the trouble to write.

There are several forms of Lisp mode, which differ in the way they interface to Lisp execution. See Section 23.6 [Executing Lisp], page 276.

Each of the programming language major modes defines the ^{⟨}\underline{\overline{TAB}}^{⟩} key to run an indentation function that knows the indentation conventions of that language and updates the current line’s indentation accordingly. For example, in C mode ^{⟨}\underline{\overline{TAB}}^{⟩} is bound to c-indent-line. C-j is normally defined to do ^{⟨}\underline{\overline{RET}}^{⟩} followed by ^{⟨}\underline{\overline{TAB}}^{⟩}; thus, it too indents in a mode-specific fashion.

In most programming languages, indentation is likely to vary from line to line. So the major modes for those languages rebind ^{⟨}\underline{\overline{DEL}}^{⟩} to treat a tab as if it were the equivalent number of spaces (using the command backwarddelete-char-untabify). This makes it possible to rub out indentation one column at a time without worrying whether it is made up of spaces or tabs. Use C-b C-d to delete a tab character before point, in these modes.

Programming language modes define paragraphs to be separated only by blank lines, so that the paragraph commands remain useful. Auto Fill mode, if enabled in a programming language major mode, indents the new lines which it creates.

Turning on a major mode runs a normal hook called the mode hook, which is the value of a Lisp variable. Each major mode has a mode hook, and the hook’s name is always made from the mode command’s name by adding ‘-hook’. For example, turning on C mode runs the hook c-mode-hook, while turning on Lisp mode runs the hook lisp-mode-hook. See Section 31.2.3 [Hooks], page 392.

22.2 Lists and Sexps

By convention, Emacs keys for dealing with balanced expressions are usually Control-Meta characters. They tend to be analogous in function to their Control and Meta equivalents. These commands are usually thought of as pertaining to expressions in programming languages, but can be useful with any language in which some sort of parentheses exist (including human languages).

These commands fall into two classes. Some deal only with lists (parenthetical groupings). They see nothing except parentheses, brackets, braces (whichever ones must balance in the language you are working with), and escape characters that might be used to quote those.

The other commands deal with expressions or sexps. The word ‘sexp’ is derived from s-expression, the ancient term for an expression in Lisp. But in Emacs, the notion of ‘sexp’ is not limited to Lisp. It refers to an expression in whatever language your program is written in. Each programming language has its own major mode, which customizes the syntax tables so that expressions in that language count as sexps.

Sexps typically include symbols, numbers, and string constants, as well as anything contained in parentheses, brackets or braces.

In languages that use prefix and infix operators, such as C, it is not possible for all expressions to be sexps. For example, C mode does not recognize ‘foo + bar’ as a sexp, even though it is a C expression; it recognizes ‘foo’ as one sexp and ‘bar’ as another, with the ‘+’ as punctuation between them. This is a fundamental ambiguity: both ‘foo + bar’ and ‘foo’ are legitimate choices for the sexp to move over if point is at the ‘f’. Note that ‘(foo + bar)’ is a single sexp in C mode.

Some languages have obscure forms of expression syntax that nobody has bothered to make Emacs understand properly.

22.3 List And Sexp Commands

C-M-f Move forward over a sexp (forward-sexp).

C-M-b Move backward over a sexp (backward-sexp).

C-M-k Kill sexp forward (kill-sexp).

C-M-^{⟨}\underline{\overline{DEL}}^{⟩} Kill sexp backward (backward-kill-sexp).

C-M-u Move up and backward in list structure (backward-up-list).

C-M-d Move down and forward in list structure (down-list).

C-M-n Move forward over a list (forward-list).

C-M-p Move backward over a list (backward-list).

C-M-t Transpose expressions (transpose-sexps).

C-M-@ Put mark after following expression (mark-sexp).

To move forward over a sexp, use C-M-f (forward-sexp). If the first significant character after point is an opening delimiter (‘(’ in Lisp; ‘(’, ‘[’ or ‘{’ in C), C-M-f moves past the matching closing delimiter. If the character begins a symbol, string, or number, C-M-f moves over that.

The command C-M-b (backward-sexp) moves backward over a sexp. The detailed rules are like those above for C-M-f, but with directions reversed. If there are any prefix characters (single-quote, backquote and comma, in Lisp) preceding the sexp, C-M-b moves back over them as well. The sexp commands move across comments as if they were whitespace in most modes.

C-M-f or C-M-b with an argument repeats that operation the specified number of times; with a negative argument, it moves in the opposite direction.

Killing a whole sexp can be done with C-M-k (kill-sexp) or C-M-^{⟨}\underline{\overline{DEL}}^{⟩} (backward-kill-sexp). C-M-k kills the characters that C-M-f would move over, and C-M-^{⟨}\underline{\overline{DEL}}^{⟩} kills the characters that C-M-b would move over.

The list commands move over lists, as the sexp commands do, but skip blithely over any number of other kinds of sexps (symbols, strings, etc.).

They are C-M-n (forward-list) and C-M-p (backward-list). The main

reason they are useful is that they usually ignore comments (since the comments usually do not contain any lists).

C-M-n and C-M-p stay at the same level in parentheses, when that’s possible. To move up one (or n) levels, use C-M-u (backward-up-list). C-M-u moves backward up past one unmatched opening delimiter. A positive argument serves as a repeat count; a negative argument reverses direction of motion and also requests repetition, so it moves forward and up one or more levels.

To move down in list structure, use C-M-d (down-list). In Lisp mode, where ‘(’ is the only opening delimiter, this is nearly the same as searching for a ‘(’. An argument specifies the number of levels of parentheses to go down.

A somewhat random-sounding command which is nevertheless handy is C-M-t (transpose-sexps), which drags the previous sexp across the next one. An argument serves as a repeat count, and a negative argument drags backwards (thus canceling out the effect of C-M-t with a positive argument). An argument of zero, rather than doing nothing, transposes the sexps ending after point and the mark.

To set the region around the next sexp in the buffer, use C-M-@ (marksexp), which sets mark at the same place that C-M-f would move to. C-M-@ takes arguments like C-M-f. In particular, a negative argument is useful for putting the mark at the beginning of the previous sexp.

The list and sexp commands’ understanding of syntax is completely controlled by the syntax table. Any character can, for example, be declared to be an opening delimiter and act like an open parenthesis. See Section 31.6 [Syntax], page 411.

22.4 Defuns

In Emacs, a parenthetical grouping at the top level in the buffer is called a defun. The name derives from the fact that most top-level lists in a Lisp file are instances of the special form defun, but any top-level parenthetical grouping counts as a defun in Emacs parlance regardless of what its contents are, and regardless of the programming language in use. For example, in C, the body of a function definition is a defun.

C-M-a Move to beginning of current or preceding defun (beginning-of-

          defun).

C-M-e Move to end of current or following defun (end-of-defun).

C-M-h Put region around whole current or following defun (mark-defun).

The commands to move to the beginning and end of the current defun

are C-M-a (beginning-of-defun) and C-M-e (end-of-defun).

If you wish to operate on the current defun, use C-M-h (mark-defun) which puts point at the beginning and mark at the end of the current or next defun. For example, this is the easiest way to get ready to move the defun to a different place in the text. In C mode, C-M-h runs the function c-mark-function, which is almost the same as mark-defun; the difference is that it backs up over the argument declarations, function name and returned data type so that the entire C function is inside the region. See Section 8.4 [Marking Objects], page 60.

Emacs assumes that any open-parenthesis found in the leftmost column is the start of a defun. Therefore, never put an open-parenthesis at the left margin in a Lisp file unless it is the start of a top-level list. Never put an open-brace or other opening delimiter at the beginning of a line of C code unless it starts the body of a function. The most likely problem case is when you want an opening delimiter at the start of a line inside a string. To avoid trouble, put an escape character (‘\’, in C and Emacs Lisp, ‘/’ in some other Lisp dialects) before the opening delimiter. It will not affect the contents of the string.

In the remotest past, the original Emacs found defuns by moving upward a level of parentheses until there were no more levels to go up. This always required scanning all the way back to the beginning of the buffer, even for a small function. To speed up the operation, Emacs was changed to assume that any ‘(’ (or other character assigned the syntactic class of openingdelimiter) at the left margin is the start of a defun. This heuristic is nearly always right and avoids the costly scan; however, it mandates the convention described above.

22.5 Indentation for Programs

The best way to keep a program properly indented is to use Emacs to reindent it as you change it. Emacs has commands to indent properly either a single line, a specified number of lines, or all of the lines inside a single parenthetical grouping.

Emacs also provides a Lisp pretty-printer in the library pp. This program reformats a Lisp object with indentation chosen to look nice.

22.5.1 Basic Program Indentation Commands

⟨\underline{\overline{TAB}}^{⟩} Adjust indentation of current line.

C-j Equivalent to ^{⟨}\underline{\overline{RET}}^{⟩} followed by ^{⟨}\underline{\overline{TAB}}^{⟩} (newline-and-indent).

The basic indentation command is ^{⟨}\underline{\overline{TAB}}^{⟩}, which gives the current line the correct indentation as determined from the previous lines. The function that ^{⟨}\underline{\overline{TAB}}^{⟩} runs depends on the major mode; it is lisp-indent-line in Lisp mode, c-indent-line in C mode, etc. These functions understand different syntaxes for different languages, but they all do about the same thing. ^{⟨}\underline{\overline{TAB}}^{⟩} in any programming-language major mode inserts or deletes whitespace at the beginning of the current line, independent of where point is in the line. If point is inside the whitespace at the beginning of the line, ^{⟨}\underline{\overline{TAB}}^{⟩} leaves it at the end of that whitespace; otherwise, ^{⟨}\underline{\overline{TAB}}^{⟩} leaves point fixed with respect to the characters around it.

Use C-q ^{⟨}\underline{\overline{TAB}}^{⟩} to insert a tab at point.

When entering lines of new code, use C-j (newline-and-indent), which is equivalent to a ^{⟨}\underline{\overline{RET}}^{⟩} followed by a ^{⟨}\underline{\overline{TAB}}^{⟩}. C-j creates a blank line and then gives it the appropriate indentation.

⟨\underline{\overline{TAB}}^{⟩} indents the second and following lines of the body of a parenthetical grouping each under the preceding one; therefore, if you alter one line’s indentation to be nonstandard, the lines below will tend to follow it. This behavior is convenient in cases where you have overridden the standard result of ^{⟨}\underline{\overline{TAB}}^{⟩} because you find it unaesthetic for a particular line.

Remember that an open-parenthesis, open-brace or other opening delimiter at the left margin is assumed by Emacs (including the indentation routines) to be the start of a function. Therefore, you must never have an opening delimiter in column zero that is not the beginning of a function, not even inside a string. This restriction is vital for making the indentation commands fast; you must simply accept it. See Section 22.4 [Defuns], page 224, for more information on this.

22.5.2 Indenting Several Lines

When you wish to reindent several lines of code which have been altered or moved to a different level in the list structure, you have several commands available.

C-M-q Reindent all the lines within one list (indent-sexp).

C-u ^{⟨}\underline{\overline{TAB}}^{⟩} Shift an entire list rigidly sideways so that its first line is properly indented.

C-M-\ Reindent all lines in the region (indent-region).

You can reindent the contents of a single list by positioning point before the beginning of it and typing C-M-q (indent-sexp in Lisp mode, c-indent-exp in C mode; also bound to other suitable commands in other modes). The indentation of the line the sexp starts on is not changed; therefore, only the relative indentation within the list, and not its position, is changed. To correct the position as well, type a ^{⟨}\underline{\overline{TAB}}^{⟩} before the C-M-q.

If the relative indentation within a list is correct but the indentation of its first line is not, go to that line and type C-u ^{⟨}\underline{\overline{TAB}}^{⟩}. ^{⟨}\underline{\overline{TAB}}^{⟩} with a numeric argument reindents the current line as usual, then reindents by the same amount all the lines in the grouping starting on the current line. In other words, it reindents the whole grouping rigidly as a unit. It is clever, though, and does not alter lines that start inside strings, or C preprocessor lines when in C mode.

Another way to specify the range to be reindented is with the region. The command C-M-\ (indent-region) applies ^{⟨}\underline{\overline{TAB}}^{⟩} to every line whose first character is between point and mark.

22.5.3 Customizing Lisp Indentation

The indentation pattern for a Lisp expression can depend on the function called by the expression. For each Lisp function, you can choose among several predefined patterns of indentation, or define an arbitrary one with a Lisp program.

The standard pattern of indentation is as follows: the second line of the expression is indented under the first argument, if that is on the same line as the beginning of the expression; otherwise, the second line is indented underneath the function name. Each following line is indented under the previous line whose nesting depth is the same.

If the variable lisp-indent-offset is non-nil, it overrides the usual indentation pattern for the second line of an expression, so that such lines are always indented lisp-indent-offset more columns than the containing list.

The standard pattern is overridden for certain functions. Functions whose names start with def always indent the second line by lisp-body-indent extra columns beyond the open-parenthesis starting the expression.

The standard pattern can be overridden in various ways for individual functions, according to the lisp-indent-function property of the function name. There are four possibilities for this property:

nil This is the same as no property; the standard indentation pattern is used.

defun The pattern used for function names that start with def is used for this function also.

a number, number The first number arguments of the function are distinguished arguments; the rest are considered the body of the expression. A line in the expression is indented according to whether the first argument on it is distinguished or not. If the argument is part of the body, the line is indented lisp-body-indent more columns than the open-parenthesis starting the containing expression. If the argument is distinguished and is either the first or second argument, it is indented twice that many extra columns. If the argument is distinguished and not the first or second argument, the standard pattern is followed for that line.

a symbol, symbol symbol should be a function name; that function is called to calculate the indentation of a line within this expression. The function receives two arguments:

state The value returned by parse-partial-sexp (a Lisp primitive for indentation and nesting computation) when it parses up to the beginning of this line. pos The position at which the line being indented begins.

It should return either a number, which is the number of columns of indentation for that line, or a list whose car is such a number. The difference between returning a number and returning a list is that a number says that all following lines at the same nesting level should be indented just like this one; a list says that following lines might call for different indentations. This makes a difference when the indentation is being computed by C-M-q; if the value is a number, C-M-q need not recalculate indentation for the following lines until the end of the list.

22.5.4 Commands for C Indentation

Here are the commands for indentation in C mode and related modes:

C-c C-q Reindent the current top-level function definition or aggregate type declaration (c-indent-defun).

C-M-q Reindent each line in the balanced expression that follows point (c-indent-exp). A prefix argument inhibits error checking and warning messages about invalid syntax.

⟨\underline{\overline{TAB}}^{⟩} Reindent the current line, and/or in some cases insert a tab char-

acter (c-indent-command).

If c-tab-always-indent is t, this command always reindents the current line and does nothing else. This is the default. If that variable is nil, this command reindents the current line only if point is at the left margin or in the line’s indentation; otherwise, it inserts a tab (or the equivalent number of spaces,

if indent-tabs-mode is nil).

Any other value (not nil or t) means always reindent the line, and also insert a tab if within a comment, a string, or a preprocessor directive.

C-u ^{⟨}\underline{\overline{TAB}}^{⟩} Reindent the current line according to its syntax; also rigidly reindent any other lines of the expression that starts on the current line. See Section 22.5.2 [Multi-line Indent], page 226.

To reindent the whole current buffer, type C-x h C-M-\. This first selects the whole buffer as the region, then reindents that region.

To reindent the current block, use C-M-u C-M-q. This moves to the front of the block and then reindents it all.

22.5.5 Customizing C Indentation

C mode and related modes use a simple yet flexible mechanism for customizing indentation. The mechanism works in two steps: first it classifies the line syntactically according to its contents and context; second, it associates each kind of syntactic construct with an indentation offset which you can customize.

22.5.5.1 Step 1—Syntactic Analysis

In the first step, the C indentation mechanism looks at the line you are currently indenting and determines the syntactic components of the construct on that line. It builds a list of these syntactic components, where each component on the list contains a syntactic symbol and a relative buffer position. Syntactic symbols describe grammatical elements, for example statement and substatement; others describe locations amidst grammatical elements, for example class-open and knr-argdecl.

Conceptually, a line of C code is always indented relative to the indentation of some line higher up in the buffer. This is represented by the relative buffer positions in the syntactic component list.

Here is an example. Suppose we have the following code in a C++ mode buffer (the line numbers don’t actually appear in the buffer):

    1: void swap (int& a, int& b)
    2: {
    3:   int tmp = a;
    4:   a = b;
    5:   b = tmp;
    6: }

If you type C-c C-s (which runs the command c-show-syntactic-information) on line 4, it shows the result of the indentation mechanism for that line:

    ((statement . 32))

This indicates that the line is a statement and it is indented relative to buffer position 32, which happens to be the ‘i’ in int on line 3. If you move the cursor to line 3 and type C-c C-s, it displays this:

    ((defun-block-intro . 28))

This indicates that the int line is the first statement in a block, and is indented relative to buffer position 28, which is the brace just after the function header. Here is another example:

    1: int add (int val, int incr, int doit)
    2: {
    3:   if (doit)
    4:     {
    5:       return (val + incr);
    6:     }
    7:   return (val);

    8: }

Typing C-c C-s on line 4 displays this:

    ((substatement-open . 43))

This says that the brace opens a substatement block. By the way, a substatement indicates the line after an if, else, while, do, switch, and for statements.

Within the C indentation commands, after a line has been analyzed syntactically for indentation, the variable c-syntactic-context contains a list that describes the results. Each element in this list is a a syntactic component: a cons cell containing a syntactic symbol and (optionally) its corresponding buffer position. There may be several elements in a component list; typically only one element has a buffer position.

22.5.5.2 Step 2—Indentation Calculation

The C indentation mechanism calculates the indentation for the current line using the list of syntactic components, c-syntactic-context, derived from syntactic analysis. Each component is a cons cell that contains a syntactic symbol and may also contain a buffer position.

Each component contributes to the final total indentation of the line in two ways. First, the syntactic symbol identifies an element of c-offsets-alist, which is an association list mapping syntactic symbols into indentation offsets. Each syntactic symbol’s offset adds to the total indentation. Second, if the component includes a buffer position, the column number of that position adds to the indentation. All these offsets and column numbers, added together, give the total indentation.

The following examples demonstrate the workings of the C indentation mechanism:

    1: void swap (int& a, int& b)
    2: {
    3:   int tmp = a;
    4:   a = b;
    5:   b = tmp;
    6: }

Suppose that point is on line 3 and you type ^{⟨}\underline{\overline{TAB}}^{⟩} to reindent the line. As explained above (see Section 22.5.5.1 [Syntactic Analysis], page 229), the syntactic component list for that line is:

    ((defun-block-intro . 28))

In this case, the indentation calculation first looks up defun-block-intro in the c-offsets-alist alist. Suppose that it finds the integer 2; it adds this to the running total (initialized to zero), yielding a updated total indentation of 2 spaces.

The next step is to find the column number of buffer position 28. Since the brace at buffer position 28 is in column zero, this adds 0 to the running total. Since this line has only one syntactic component, the total indentation for the line is 2 spaces.

    1: int add (int val, int incr, int doit)
    2: {
    3:   if (doit)
    4:     {
    5:       return(val + incr);
    6:     }
    7:   return(val);
    8: }

If you type ^{⟨}\underline{\overline{TAB}}^{⟩} on line 4, the same process is performed, but with different data. The syntactic component list for this line is:

    ((substatement-open . 43))

Here, the indentation calculation’s first job is to look up the symbol

substatement-open in c-offsets-alist. Let’s assume that the offset for

this symbol is 2. At this point the running total is 2 (0 + 2 = 2). Then it adds the column number of buffer position 43, which is the ‘i’ in if on line 3. This character is in column 2 on that line. Adding this yields a total indentation of 4 spaces.

If a syntactic symbol in the analysis of a line does not appear in coffsets-alist, it is ignored; if in addition the variable c-strict-syntax-p is non-nil, it is an error.

22.5.5.3 Changing Indentation Style

There are two ways to customize the indentation style for the C modes. First, you can select one of several predefined styles, each of which specifies offsets for all the syntactic symbols. For more flexibility, you can customize the handling of individual syntactic symbols. See Section 22.5.5.4 [Syntactic Symbols], page 232, for a list of all defined syntactic symbols.

M-x c-set-style ^{⟨}\underline{\overline{RET}}^{⟩} style ^{⟨}\underline{\overline{RET}}^{⟩} Select predefined indentation style style. Type ? when entering style to see a list of supported styles; to find out what a style looks like, select it and reindent some C code.

C-c C-o symbol ^{⟨}\underline{\overline{RET}}^{⟩} offset ^{⟨}\underline{\overline{RET}}^{⟩} Set the indentation offset for syntactic symbol symbol (c-set-offset). The second argument offset specifies the new indentation offset.

The c-offsets-alist variable controls the amount of indentation to give to each syntactic symbol. Its value is an association list, and each element of the list has the form (syntactic-symbol . offset). By changing the offsets for various syntactic symbols, you can customize indentation in fine detail. To change this alist, use c-set-offset (see below).

Each offset value in c-offsets-alist can be an integer, a function or variable name, or one of the following symbols: +, -, ++, or --, indicating positive or negative multiples of the variable c-basic-offset. Thus, if you want to change the levels of indentation to be 3 spaces instead of 2 spaces,

set c-basic-offset to 3.

Using a function as the offset value provides the ultimate flexibility in customizing indentation. The function is called with a single argument containing the cons of the syntactic symbol and the relative indent point. The function should return an integer offset.

The command C-c C-o (c-set-offset) is the easiest way to set offsets, both interactively or in your ‘~/.emacs’ file. First specify the syntactic symbol, then the offset you want. See Section 22.5.5.4 [Syntactic Symbols], page 232, for a list of valid syntactic symbols and their meanings.

22.5.5.4 Syntactic Symbols

Here is a table of valid syntactic symbols for C mode indentation, with their syntactic meanings. Normally, most of these symbols are assigned

offsets in c-offsets-alist.

string Inside a multi-line string.

c Inside a multi-line C style block comment.

defun-open

On a brace that opens a function definition.

defun-close

On a brace that closes a function definition.

defun-block-intro

In the first line in a top-level defun.

class-open

On a brace that opens a class definition.

class-close

On a brace that closes a class definition.

inline-open

On a brace that opens an in-class inline method.

inline-close

On a brace that closes an in-class inline method.

extern-lang-open

On a brace that opens an external language block.

extern-lang-close

On a brace that closes an external language block.

func-decl-cont

The region between a function definition’s argument list and the defun opening brace (excluding K&R function definitions). In C, you cannot put anything but whitespace and comments between them; in C++ and Java, throws declarations and other things can appear in this context.

knr-argdecl-intro

On the first line of a K&R C argument declaration.

knr-argdecl

In one of the subsequent lines in a K&R C argument declaration.

topmost-intro

On the first line in a topmost construct definition.

topmost-intro-cont

On the topmost definition continuation lines.

member-init-intro

On the first line in a member initialization list.

member-init-cont

On one of the subsequent member initialization list lines.

inher-intro

On the first line of a multiple inheritance list.

inher-cont

On one of the subsequent multiple inheritance lines.

block-open

On a statement block open brace.

block-close

On a statement block close brace.

brace-list-open

On the opening brace of an enum or static array list.

brace-list-close

On the closing brace of an enum or static array list.

brace-list-intro

On the first line in an enum or static array list.

brace-list-entry

On one of the subsequent lines in an enum or static array list.

statement

On an ordinary statement.

statement-cont

On a continuation line of a statement.

statement-block-intro

On the first line in a new statement block.

statement-case-intro

On the first line in a case “block.”

statement-case-open

On the first line in a case block starting with brace.

substatement

On the first line after an if, while, for, do, or else.

substatement-open

On the brace that opens a substatement block.

case-label
          On a case or default label.
access-label
          On a C++ private, protected, or public access label.

label On any ordinary label.

do-while-closure

On the while that ends a do-while construct.

else-clause

On the else of an if-else construct.

comment-intro

On a line containing only a comment introduction.

arglist-intro

On the first line in an argument list.

arglist-cont

On one of the subsequent argument list lines when no arguments follow on the same line as the arglist opening parenthesis.

arglist-cont-nonempty

On one of the subsequent argument list lines when at least one argument follows on the same line as the arglist opening parenthesis.

arglist-close

On the closing parenthesis of an argument list.

stream-op

On one of the lines continuing a stream operator construct.

inclass On a construct that is nested inside a class definition.

inextern-lang

On a construct that is nested inside an external language block.

cpp-macro

On the start of a cpp macro.

friend On a C++ friend declaration.

objc-method-intro

On the first line of an Objective-C method definition.

objc-method-args-cont

On one of the lines continuing an Objective-C method definition.

objc-method-call-cont

On one of the lines continuing an Objective-C method call.

22.5.5.5 Variables for C Indentation

This section describes additional variables which control the indentation behavior of C mode and related mode.

c-offsets-alist

Association list of syntactic symbols and their indentation offsets. You should not set this directly, only with c-set-offset. See Section 22.5.5.3 [Changing Indent Style], page 231, for details.

c-style-alist

Variable for defining indentation styles; see below.

c-basic-offset

Amount of basic offset used by + and - symbols in c-offsets-

          alist.
c-special-indent-hook

Hook for user-defined special indentation adjustments. This hook is called after a line is indented by C mode and related modes.

The variable c-style-alist specifies the predefined indentation styles. Each element has form (name variable-setting...), where name is the name of the style. Each variable-setting has the form (variable . value); variable is one of the customization variables used by C mode, and value is the value for that variable when using the selected style.

When variable is c-offsets-alist, that is a special case: value is appended to the front of the value of c-offsets-alist instead of replacing that value outright. Therefore, it is not necessary for value to specify each and every syntactic symbol—only those for which the style differs from the default.

The indentation of lines containing only comments is also affected by the variable c-comment-only-line-offset (see Section 22.15.5 [Comments in C], page 260).

22.5.5.6 C Indentation Styles

A C style is a collection of indentation style customizations. Emacs comes with several predefined indentation styles for C code including gnu,

k&r, bsd, stroustrup, linux, python, java, whitesmith, ellemtel, and

cc-mode. The default style is gnu.

To choose the style you want, use the command M-x c-set-style. Specify a style name as an argument (case is not significant in C style names). The chosen style only affects newly visited buffers, not those you are already editing.

To define a new C indentation style, call the function c-add-style: (c-add-style name values use-now)

Here name is the name of the new style (a string), and values is an alist whose elements have the form (variable . value). The variables you specify should be among those documented in Section 22.5.5.5 [Variables for C Indent], page 235.

If use-now is non-nil, c-add-style switches to the new style after defining it.

22.6 Automatic Display Of Matching Parentheses

The Emacs parenthesis-matching feature is designed to show automatically how parentheses match in the text. Whenever you type a self-inserting character that is a closing delimiter, the cursor moves momentarily to the location of the matching opening delimiter, provided that is on the screen. If it is not on the screen, some text near it is displayed in the echo area. Either way, you can tell what grouping is being closed off.

In Lisp, automatic matching applies only to parentheses. In C, it applies to braces and brackets too. Emacs knows which characters to regard as matching delimiters based on the syntax table, which is set by the major mode. See Section 31.6 [Syntax], page 411.

If the opening delimiter and closing delimiter are mismatched—such as in ‘[x)’—a warning message is displayed in the echo area. The correct matches are specified in the syntax table.

Three variables control parenthesis match display. blink-matching-paren turns the feature on or off; nil turns it off, but the default is t to turn match display on. blink-matching-delay says how many seconds to wait; the default is 1, but on some systems it is useful to specify a fraction of a second. blink-matching-paren-distance specifies how many characters back to search to find the matching opening delimiter. If the match is not found in that far, scanning stops, and nothing is displayed. This is to prevent scanning for the matching delimiter from wasting lots of time when there is no match. The default is 12,000.

When using X Windows, you can request a more powerful alternative kind of automatic parenthesis matching by enabling Show Paren mode. This mode turns off the usual kind of matching parenthesis display and instead uses highlighting to show what matches. Whenever point is after a close parenthesis, the close parenthesis and its matching open parenthesis are both highlighted; otherwise, if point is before an open parenthesis, the matching close parenthesis is highlighted. (There is no need to highlight the open parenthesis after point because the cursor appears on top of that character.) Use the command M-x show-paren-mode to enable or disable this mode.

22.7 Manipulating Comments

Because comments are such an important part of programming, Emacs provides special commands for editing and inserting comments.

22.7.1 Comment Commands

The comment commands insert, kill and align comments.

M-; Insert or align comment (indent-for-comment).

C-x ;  Set comment column (set-comment-column).
C-u - C-x ;

Kill comment on current line (kill-comment).

C-M-j Like ^{⟨}\underline{\overline{RET}}^{⟩} followed by inserting and aligning a comment (indent-

          new-comment-line).
M-x comment-region

Add or remove comment delimiters on all the lines in the region.

The command that creates a comment is M-; (indent-for-comment). If there is no comment already on the line, a new comment is created, aligned at a specific column called the comment column. The comment is created by inserting the string Emacs thinks comments should start with (the value of comment-start; see below). Point is left after that string. If the text of the line extends past the comment column, then the indentation is done to a suitable boundary (usually, at least one space is inserted). If the major mode has specified a string to terminate comments, that is inserted after point, to keep the syntax valid.

M-; can also be used to align an existing comment. If a line already contains the string that starts comments, then M-; just moves point after it and reindents it to the conventional place. Exception: comments starting in column 0 are not moved.

Some major modes have special rules for indenting certain kinds of comments in certain contexts. For example, in Lisp code, comments which start with two semicolons are indented as if they were lines of code, instead of at the comment column. Comments which start with three semicolons are supposed to start at the left margin. Emacs understands these conventions by indenting a double-semicolon comment using ^{⟨}\underline{\overline{TAB}}^{⟩}, and by not changing the indentation of a triple-semicolon comment at all.

    ;; This function is just an example
    ;;; Here either two or three semicolons are appropriate.
    (defun foo (x)
    ;;; And now, the first part of the function:
      ;; The following line adds one.
      (1+ x))           ; This line adds one.

In C code, a comment preceded on its line by nothing but whitespace is indented like a line of code.

Even when an existing comment is properly aligned, M-; is still useful for moving directly to the start of the comment.

C-u - C-x ; (kill-comment) kills the comment on the current line, if there is one. The indentation before the start of the comment is killed as well. If there does not appear to be a comment in the line, nothing is done. To reinsert the comment on another line, move to the end of that line, do C-y, and then do M-; to realign it. Note that C-u - C-x ; is not a distinct key; it is C-x ; (set-comment-column) with a negative argument. That command is programmed so that when it receives a negative argument it calls kill-comment. However, kill-comment is a valid command which you could bind directly to a key if you wanted to.

22.7.2 Multiple Lines of Comments

If you are typing a comment and wish to continue it on another line, you can use the command C-M-j (indent-new-comment-line). This terminates the comment you are typing, creates a new blank line afterward, and begins a new comment indented under the old one. When Auto Fill mode is on, going past the fill column while typing a comment causes the comment to be continued in just this fashion. If point is not at the end of the line when C-M-j is typed, the text on the rest of the line becomes part of the new comment line.

To turn existing lines into comment lines, use the M-x comment-region command. It adds comment delimiters to the lines that start in the region, thus commenting them out. With a negative argument, it does the opposite—it deletes comment delimiters from the lines in the region.

With a positive argument, comment-region duplicates the last character of the comment start sequence it adds; the argument specifies how many copies of the character to insert. Thus, in Lisp mode, C-u 2 M-x comment-region adds ‘;;’ to each line. Duplicating the comment delimiter is a way of calling attention to the comment. It can also affect how the comment is indented. In Lisp, for proper indentation, you should use an argument of two, if between defuns, and three, if within a defun.

The variable comment-padding specifies how many spaces commentregion should insert on each line between the comment delimiter and the line’s original text. The default is 1.

22.7.3 Options Controlling Comments

The comment column is stored in the variable comment-column. You can set it to a number explicitly. Alternatively, the command C-x ; (setcomment-column) sets the comment column to the column point is at. C-u C-x ; sets the comment column to match the last comment before point in the buffer, and then does a M-; to align the current line’s comment under the previous one. Note that C-u - C-x ; runs the function kill-comment as described above.

The variable comment-column is per-buffer: setting the variable in the normal fashion affects only the current buffer, but there is a default value which you can change with setq-default. See Section 31.2.4 [Locals], page 393. Many major modes initialize this variable for the current buffer.

The comment commands recognize comments based on the regular expression that is the value of the variable comment-start-skip. Make sure this regexp does not match the null string. It may match more than the comment starting delimiter in the strictest sense of the word; for example, in C mode the value of the variable is "/\\*+ *", which matches extra stars and spaces after the ‘/*’ itself. (Note that ‘\\’ is needed in Lisp syntax to include a ‘\’ in the string, which is needed to deny the first star its special meaning in regexp syntax. See Section 12.5 [Regexps], page 88.)

When a comment command makes a new comment, it inserts the value of comment-start to begin it. The value of comment-end is inserted after point, so that it will follow the text that you will insert into the comment.

In C mode, comment-start has the value "/* " and comment-end has the

value " */".

   The variable comment-multi-line controls how C-M-j (indent-new-

comment-line) behaves when used inside a comment. If comment-multi-line is nil, as it normally is, then the comment on the starting line is terminated and a new comment is started on the new following line. If comment-multi-line is not nil, then the new following line is set up as part of the same comment that was found on the starting line. This is done by not inserting a terminator on the old line, and not inserting a starter on the new line. In languages where multi-line comments work, the choice of value for this variable is a matter of taste.

The variable comment-indent-function should contain a function that will be called to compute the indentation for a newly inserted comment or for aligning an existing comment. It is set differently by various major modes. The function is called with no arguments, but with point at the beginning of the comment, or at the end of a line if a new comment is to be inserted. It should return the column in which the comment ought to start.

For example, in Lisp mode, the indent hook function bases its decision on how many semicolons begin an existing comment, and on the code in the preceding lines.

22.8 Editing Without Unbalanced Parentheses

M-( Put parentheses around next sexp(s) (insert-parentheses).

M-) Move past next close parenthesis and reindent (move-past-

          close-and-reindent).
   The commands M-( (insert-parentheses) and M-) (move-past-close-

and-reindent) are designed to facilitate a style of editing which keeps parentheses balanced at all times. M-( inserts a pair of parentheses, either together as in ‘()’, or, if given an argument, around the next several sexps. It leaves point after the open parenthesis. The command M-) moves past the close parenthesis, deleting any indentation preceding it (in this example there is none), and indenting with C-j after it.

For example, instead of typing ( F O O ), you can type M-( F O O, which has the same effect except for leaving the cursor before the close parenthesis.

M-( may insert a space before the open parenthesis, depending on the syntax class of the preceding character. Set parens-require-spaces to nil value if you wish to inhibit this.

22.9 Completion for Symbol Names

Usually completion happens in the minibuffer. But one kind of completion is available in all buffers: completion for symbol names.

The character M-^{⟨}\underline{\overline{TAB}}^{⟩} runs a command to complete the partial symbol before point against the set of meaningful symbol names. Any additional characters determined by the partial name are inserted at point.

If the partial name in the buffer has more than one possible completion and they have no additional characters in common, a list of all possible completions is displayed in another window.

In most programming language major modes, M-^{⟨}\underline{\overline{TAB}}^{⟩} runs the command complete-symbol, which provides two kinds of completion. Normally it does completion based on a tags table (see Section 22.13 [Tags], page 243); with a numeric argument (regardless of the value), it does completion based on the names listed in the Info file indexes for your language. Thus, to complete the name of a symbol defined in your own program, use M-^{⟨}\underline{\overline{TAB}}^{⟩} with no argument; to complete the name of a standard library function, use C-u M-^{⟨}\underline{\overline{TAB}}^{⟩}. Of course, Info-based completion works only if there is an Info file for the standard library functions of your language, and only if it is installed at your site.

In Emacs-Lisp mode, the name space for completion normally consists of nontrivial symbols present in Emacs—those that have function definitions, values or properties. However, if there is an open-parenthesis immediately before the beginning of the partial symbol, only symbols with function definitions are considered as completions. The command which implements this

is lisp-complete-symbol.

In Text mode and related modes, M-^{⟨}\underline{\overline{TAB}}^{⟩} completes words based on the spell-checker’s dictionary. See Section 13.4 [Spelling], page 101.

22.10 Which Function Mode

Which Function mode is a minor mode that displays the current function name in the mode line, as you move around in a buffer.

To enable (or disable) Which Function mode, use the command M-x which-function-mode. This command is global; it applies to all buffers, both existing ones and those yet to be created. However, this only affects certain major modes, those listed in the value of which-func-modes. (If the value is t, then Which Function mode applies to all major modes that know how to support it—which are the major modes that support Imenu.)

22.11 Documentation Commands

As you edit Lisp code to be run in Emacs, the commands C-h f

(describe-function) and C-h v (describe-variable) can be used to print

documentation of functions and variables that you want to call. These commands use the minibuffer to read the name of a function or variable to document, and display the documentation in a window.

For extra convenience, these commands provide default arguments based on the code in the neighborhood of point. C-h f sets the default to the function called in the innermost list containing point. C-h v uses the symbol name around or adjacent to point as its default.

For Emacs Lisp code, you can also use Eldoc mode. This minor mode constantly displays in the echo area the argument list for the function being called at point. (In other words, it finds the function call that point is contained in, and displays the argument list of that function.) Eldoc mode applies in Emacs Lisp and Lisp Interaction modes only. Use the command M-x eldoc-mode to enable or disable this feature.

For C, Lisp, and other languages, you can use C-h C-i (info-lookup-symbol) to view the Info documentation for a symbol. You specify the symbol with the minibuffer; by default, it uses the symbol that appears in the buffer at point. The major mode determines where to look for documentation for the symbol—which Info files and which indices. You can also use M-x info-lookup-file to look for documentation for a file name.

You can read the “man page” for an operating system command, library function, or system call, with the M-x manual-entry command. It runs the man program to format the man page, and runs it asynchronously if your system permits, so that you can keep on editing while the page is being formatted. (MS-DOS and MS-Windows 3 do not permit asynchronous subprocesses, so on these systems you cannot edit while Emacs waits for man to exit.) The result goes in a buffer named ‘*Man topic*’. These buffers use a special major mode, Man mode, that facilitates scrolling and examining other manual pages. For details, type C-h m while in a man page buffer.

For a long man page, setting the faces properly can take substantial time. By default, Emacs uses faces in man pages if Emacs can display different fonts or colors. You can turn off use of faces in man pages by setting the

variable Man-fontify-manpage-flag to nil.

If you insert the text of a man page into an Emacs buffer in some other fashion, you can use the command M-x Man-fontify-manpage to perform the same conversions that M-x manual-entry does.

Eventually the GNU project hopes to replace most man pages with betterorganized manuals that you can browse with Info. See Section 7.7 [Misc Help], page 55. Since this process is only partially completed, it is still useful to read manual pages.

22.12 Change Logs

The Emacs command C-x 4 a adds a new entry to the change log file for

the file you are editing (add-change-log-entry-other-window).

A change log file contains a chronological record of when and why you have changed a program, consisting of a sequence of entries describing individual changes. Normally it is kept in a file called ‘ChangeLog’ in the same directory as the file you are editing, or one of its parent directories. A single ‘ChangeLog’ file can record changes for all the files in its directory and all its subdirectories.

A change log entry starts with a header line that contains your name, your email address (taken from the variable user-mail-address), and the current date and time. Aside from these header lines, every line in the change log starts with a space or a tab. The bulk of the entry consists of items, each of which starts with a line starting with whitespace and a star. Here are two entries, both dated in May 1993, each with two items:

     1993-05-25  Richard Stallman  <rms@gnu.org>
             * man.el: Rename symbols `man-*' to `Man-*'.
             (manual-entry): Make prompt string clearer.
             * simple.el (blink-matching-paren-distance):
             Change default to 12,000.

     1993-05-24  Richard Stallman  <rms@gnu.org>
             * vc.el (minor-mode-map-alist): Don't use it if it's void.
             (vc-cancel-version): Doc fix.

(Previous Emacs versions used a different format for the date.)

One entry can describe several changes; each change should have its own item. Normally there should be a blank line between items. When items are related (parts of the same change, in different places), group them by leaving no blank line between them. The second entry above contains two items grouped in this way.

C-x 4 a visits the change log file and creates a new entry unless the most recent entry is for today’s date and your name. It also creates a new item for the current file. For many languages, it can even guess the name of the function or other object that was changed.

The change log file is visited in Change Log mode. In this major mode, each bunch of grouped items counts as one paragraph, and each entry is considered a page. This facilitates editing the entries. C-j and auto-fill indent each new line like the previous line; this is convenient for entering the contents of an entry.

Version control systems are another way to keep track of changes in your program and keep a change log. See Section 14.7.3.3 [Log Buffer], page 122.

22.13 Tags Tables

A tags table is a description of how a multi-file program is broken up into files. It lists the names of the component files and the names and positions of the functions (or other named subunits) in each file. Grouping the related files makes it possible to search or replace through all the files with one command. Recording the function names and positions makes possible the M-. command which finds the definition of a function by looking up which of the files it is in.

Tags tables are stored in files called tags table files. The conventional name for a tags table file is ‘TAGS’.

Each entry in the tags table records the name of one tag, the name of the file that the tag is defined in (implicitly), and the position in that file of the tag’s definition.

Just what names from the described files are recorded in the tags table depends on the programming language of the described file. They normally include all functions and subroutines, and may also include global variables, data types, and anything else convenient. Each name recorded is called a tag.

22.13.1 Source File Tag Syntax

Here is how tag syntax is defined for the most popular languages:

• In C code, any C function or typedef is a tag, and so are definitions of struct, union and enum. #define macro definitions and enum constants are also tags, unless you specify ‘--no-defines’ when making the tags table. Similarly, global variables are tags, unless you specify

‘--no-globals’. Use of ‘--no-globals’ and ‘--no-defines’ can make

the tags table file much smaller.

• In C++ code, in addition to all the tag constructs of C code, member functions are also recognized, and optionally member variables if you use the ‘--members’ option. Tags for variables and functions in classes are named ‘class::variable’ and ‘class::function’.

• In Java code, tags include all the constructs recognized in C++, plus the extends and implements constructs. Tags for variables and functions in classes are named ‘class.variable’ and ‘class.function’.

• In LaTEX text, the argument of any of the commands \chapter,

\section, \subsection, \subsubsection, \eqno, \label, \ref,
\cite, \bibitem, \part, \appendix, \entry, or \index, is a tag.

Other commands can make tags as well, if you specify them in the environment variable TEXTAGS before invoking etags. The value of this environment variable should be a colon-separated list of command names. For example,

    TEXTAGS="def:newcommand:newenvironment"
    export TEXTAGS

specifies (using Bourne shell syntax) that the commands ‘\def’,

‘\newcommand’ and ‘\newenvironment’ also define tags.

• In Lisp code, any function defined with defun, any variable defined with defvar or defconst, and in general the first argument of any expression that starts with ‘(def’ in column zero, is a tag.

• In Scheme code, tags include anything defined with def or with a construct whose name starts with ‘def’. They also include variables set with set! at top level in the file. Several other languages are also supported:

• In assembler code, labels appearing at the beginning of a line, followed by a colon, are tags.

• In Bison or Yacc input files, each rule defines as a tag the nonterminal it constructs. The portions of the file that contain C code are parsed as C code.

• In Cobol code, tags are paragraph names; that is, any word starting in column 8 and followed by a period.

• In Erlang code, the tags are the functions, records, and macros defined in the file.

• In Fortran code, functions, subroutines and blockdata are tags.

• In Objective C code, tags include Objective C definitions for classes, class categories, methods, and protocols.

• In Pascal code, the tags are the functions and procedures defined in the file.

• In Perl code, the tags are the procedures defined by the sub keyword.

• In Postscript code, the tags are the functions.

• In Prolog code, a tag name appears at the left margin.

You can also generate tags based on regexp matching (see Section 22.13.2 [Create Tags Table], page 245) to handle other formats and languages.

22.13.2 Creating Tags Tables

The etags program is used to create a tags table file. It knows the syntax of several languages, as described in the previous section. Here is how to

run etags:

etags inputfiles... The etags program reads the specified files, and writes a tags table named ‘TAGS’ in the current working directory. etags recognizes the language used in an input file based on its file name and contents. You can specify the language with the ‘--language=name’ option, described below.

If the tags table data become outdated due to changes in the files described in the table, the way to update the tags table is the same way it was made in the first place. It is not necessary to do this often.

If the tags table fails to record a tag, or records it for the wrong file, then Emacs cannot possibly find its definition. However, if the position recorded in the tags table becomes a little bit wrong (due to some editing in the file that the tag definition is in), the only consequence is a slight delay in finding the tag. Even if the stored position is very wrong, Emacs will still find the tag, but it must search the entire file for it.

So you should update a tags table when you define new tags that you want to have listed, or when you move tag definitions from one file to another, or when changes become substantial. Normally there is no need to update the tags table after each edit, or even every day.

One tags table can effectively include another. Specify the included tags file name with the ‘--include=file’ option when creating the file that is to include it. The latter file then acts as if it contained all the files specified in the included file, as well as the files it directly contains.

If you specify the source files with relative file names when you run etags, the tags file will contain file names relative to the directory where the tags file was initially written. This way, you can move an entire directory tree containing both the tags file and the source files, and the tags file will still refer correctly to the source files.

If you specify absolute file names as arguments to etags, then the tags file will contain absolute file names. This way, the tags file will still refer to the same files even if you move it, as long as the source files remain in the same place. Absolute file names start with ‘/’, or with ‘device:/’ on MS-DOS and MS-Windows.

When you want to make a tags table from a great number of files, you may have problems listing them on the command line, because some systems have a limit on its length. The simplest way to circumvent this limit is to tell etags to read the file names from its standard input, by typing a dash in place of the file names, like this:

    find . -name "*.[chCH]" -print | etags -

Use the option ‘--language=name’ to specify the language explicitly. You can intermix these options with file names; each one applies to the file names that follow it. Specify ‘--language=auto’ to tell etags to resume guessing the language from the file names and file contents. Specify ‘--language=none’ to turn off language-specific processing entirely; then etags recognizes tags by regexp matching alone. ‘etags --help’ prints the list of the languages etags knows, and the file name rules for guessing the language.

The ‘--regex’ option provides a general way of recognizing tags based on regexp matching. You can freely intermix it with file names. Each ‘--regex’ option adds to the preceding ones, and applies only to the following files. The syntax is:

--regex=/tagregexp[/nameregexp]/ where tagregexp is used to match the lines to tag. It is always anchored, that is, it behaves as if preceded by ‘^’. If you want to account for indentation, just match any initial number of blanks by beginning your regular expression with ‘[ \t]*’. In the regular expressions, ‘\’ quotes the next character, and ‘\t’ stands for the tab character. Note that etags does not handle the other C escape sequences for special characters.

The syntax of regular expressions in etags is the same as in Emacs, augmented with the interval operator, which works as in grep and ed. The syntax of an interval operator is ‘\{m,n\}’, and its meaning is to match the preceding expression at least m times and up to n times.

You should not match more characters with tagregexp than that needed to recognize what you want to tag. If the match is such that more characters than needed are unavoidably matched by tagregexp, you may find useful to add a nameregexp, in order to narrow the tag scope. You can find some examples below.

The ‘-R’ option deletes all the regexps defined with ‘--regex’ options. It applies to the file names following it, as you can see from the following example:

    etags --regex=/reg1/ voo.doo --regex=/reg2/ \
        bar.ber -R --lang=lisp los.er

Here etags chooses the parsing language for ‘voo.doo’ and ‘bar.ber’ according to their contents. etags also uses reg1 to recognize additional tags in ‘voo.doo’, and both reg1 and reg2 to recognize additional tags in ‘bar.ber’. etags uses the Lisp tags rules, and no regexp matching, to recognize tags in

‘los.er’.

Here are some more examples. The regexps are quoted to protect them from shell interpretation.

• Tag the DEFVAR macros in the emacs source files:

         --regex='/[ \t]*DEFVAR_[A-Z_\t(]+"\([^"]+\)"/'

• Tag VHDL files (this example is a single long line, broken here for formatting reasons):

         --language=none
         --regex='/[ \t]*\(ARCHITECTURE\|CONFIGURATION\) +[^]* +OF/'
         --regex='/[ \t]*\(ATTRIBUTE\|ENTITY\|FUNCTION\|PACKAGE\
         \( BODY\)?\|PROCEDURE\|PROCESS\|TYPE\)[ \t]+\([^\t(]+\)/\3/'

• Tag Tcl files (this last example shows the usage of a nameregexp):

         --lang=none --regex='/proc[ \t]+\([^\t]+\)/\1/'

For a list of the other available etags options, execute etags --help.

22.13.3 Selecting a Tags Table

Emacs has at any time one selected tags table, and all the commands for working with tags tables use the selected one. To select a tags table, type M-x visit-tags-table, which reads the tags table file name as an argument. The name ‘TAGS’ in the default directory is used as the default file name.

All this command does is store the file name in the variable tags-file-name. Emacs does not actually read in the tags table contents until you try to use them. Setting this variable yourself is just as good as using visit-tags-table. The variable’s initial value is nil; that value tells all the commands for working with tags tables that they must ask for a tags table file name to use.

Using visit-tags-table when a tags table is already loaded gives you a choice: you can add the new tags table to the current list of tags tables, or start a new list. The tags commands use all the tags tables in the current list. If you start a new list, the new tags table is used instead of others. If you add the new table to the current list, it is used as well as the others. When the tags commands scan the list of tags tables, they don’t always start at the beginning of the list; they start with the first tags table (if any) that describes the current file, proceed from there to the end of the list, and then scan from the beginning of the list until they have covered all the tables in the list.

You can specify a precise list of tags tables by setting the variable tagstable-list to a list of strings, like this:

    (setq tags-table-list
          '("~/emacs" "/usr/local/lib/emacs/src"))

This tells the tags commands to look at the ‘TAGS’ files in your ‘~/emacs’ directory and in the ‘/usr/local/lib/emacs/src’ directory. The order depends on which file you are in and which tags table mentions that file, as explained above.

   Do not set both tags-file-name and tags-table-list.

22.13.4 Finding a Tag

The most important thing that a tags table enables you to do is to find the definition of a specific tag.

M-. tag ^{⟨}\underline{\overline{RET}}^{⟩} Find first definition of tag (find-tag).

C-u M-. Find next alternate definition of last tag specified.

C-u - M-. Go back to previous tag found.

C-M-. pattern ^{⟨}\underline{\overline{RET}}^{⟩} Find a tag whose name matches pattern (find-tag-regexp).

C-u C-M-.Find the next tag whose name matches the last pattern used.

C-x 4 . tag ^{⟨}\underline{\overline{RET}}^{⟩} Find first definition of tag, but display it in another window

          (find-tag-other-window).

C-x 5 . tag ^{⟨}\underline{\overline{RET}}^{⟩} Find first definition of tag, and create a new frame to select the

          buffer (find-tag-other-frame).

M-* Pop back to where you previously invoked M-. and friends.

M-. (find-tag) is the command to find the definition of a specified tag. It searches through the tags table for that tag, as a string, and then uses the tags table info to determine the file that the definition is in and the approximate character position in the file of the definition. Then findtag visits that file, moves point to the approximate character position, and searches ever-increasing distances away to find the tag definition.

If an empty argument is given (just type ^{⟨}\underline{\overline{RET}}^{⟩}), the sexp in the buffer before or around point is used as the tag argument. See Section 22.2 [Lists], page 222, for info on sexps.

You don’t need to give M-. the full name of the tag; a part will do. This is because M-. finds tags in the table which contain tag as a substring. However, it prefers an exact match to a substring match. To find other tags that match the same substring, give find-tag a numeric argument, as in C-u M-.; this does not read a tag name, but continues searching the tags table’s text for another tag containing the same substring last used. If you have a real ^{⟨}\underline{\overline{META}}^{⟩} key, M-0 M-. is an easier alternative to C-u M-..

Like most commands that can switch buffers, find-tag has a variant that displays the new buffer in another window, and one that makes a new frame for it. The former is C-x 4 ., which invokes the command find-tag-other-

window. The latter is C-x 5 ., which invokes find-tag-other-frame.

To move back to places you’ve found tags recently, use C-u - M-.; more generally, M-. with a negative numeric argument. This command can take you to another buffer. C-x 4 . with a negative argument finds the previous tag location in another window.

As well as going back to places you’ve found tags recently, you can go back to places from where you found them. Use M-*, which invokes the command pop-tag-mark, for this. Typically you would find and study the definition of something with M-. and then return to where you were with

M-*.

Both C-u - M-. and M-* allow you to retrace your steps to a depth de-

termined by the variable find-tag-marker-ring-length.

The command C-M-. (find-tag-regexp) visits the tags that match a specified regular expression. It is just like M-. except that it does regexp matching instead of substring matching.

22.13.5 Searching and Replacing with Tags Tables

The commands in this section visit and search all the files listed in the selected tags table, one by one. For these commands, the tags table serves only to specify a sequence of files to search.

M-x tags-search ^{⟨}\underline{\overline{RET}}^{⟩} regexp ^{⟨}\underline{\overline{RET}}^{⟩} Search for regexp through the files in the selected tags table.

M-x tags-query-replace ^{⟨}\underline{\overline{RET}}^{⟩} regexp ^{⟨}\underline{\overline{RET}}^{⟩} replacement ^{⟨}\underline{\overline{RET}}^{⟩} Perform a query-replace-regexp on each file in the selected tags table.

M-, Restart one of the commands above, from the current location of

          point (tags-loop-continue).

M-x tags-search reads a regexp using the minibuffer, then searches for matches in all the files in the selected tags table, one file at a time. It displays the name of the file being searched so you can follow its progress. As soon as it finds an occurrence, tags-search returns.

Having found one match, you probably want to find all the rest. To find one more match, type M-, (tags-loop-continue) to resume the tagssearch. This searches the rest of the current buffer, followed by the remaining files of the tags table.

   M-x tags-query-replace performs a single query-replace-regexp

through all the files in the tags table. It reads a regexp to search for and a string to replace with, just like ordinary M-x query-replace-regexp. It searches much like M-x tags-search, but repeatedly, processing matches according to your input. See Section 12.7 [Replace], page 93, for more information on query replace.

It is possible to get through all the files in the tags table with a single invocation of M-x tags-query-replace. But often it is useful to exit temporarily, which you can do with any input event that has no special query replace meaning. You can resume the query replace subsequently by typing M-,; this command resumes the last tags search or replace command that you did.

The commands in this section carry out much broader searches than the find-tag family. The find-tag commands search only for definitions of tags that match your substring or regexp. The commands tags-search and tags-query-replace find every occurrence of the regexp, as ordinary search commands and replace commands do in the current buffer.

These commands create buffers only temporarily for the files that they have to search (those which are not already visited in Emacs buffers). Buffers in which no match is found are quickly killed; the others continue to exist.

It may have struck you that tags-search is a lot like grep. You can also run grep itself as an inferior of Emacs and have Emacs show you the matching lines one by one. This works much like running a compilation; finding the source locations of the grep matches works like finding the compilation errors. See Section 23.1 [Compilation], page 269.

22.13.6 Tags Table Inquiries

M-x list-tags ^{⟨}\underline{\overline{RET}}^{⟩} file ^{⟨}\underline{\overline{RET}}^{⟩} Display a list of the tags defined in the program file ‘file’.

M-x tags-apropos ^{⟨}\underline{\overline{RET}}^{⟩} regexp ^{⟨}\underline{\overline{RET}}^{⟩} Display a list of all tags matching regexp.

M-x list-tags reads the name of one of the files described by the selected tags table, and displays a list of all the tags defined in that file. The “file name” argument is really just a string to compare against the file names recorded in the tags table; it is read as a string rather than as a file name. Therefore, completion and defaulting are not available, and you must enter the file name the same way it appears in the tags table. Do not include a directory as part of the file name unless the file name recorded in the tags table includes a directory.

M-x tags-apropos is like apropos for tags (see Section 7.3 [Apropos], page 52). It reads a regexp, then finds all the tags in the selected tags table whose entries match that regexp, and displays the tag names found.

You can also perform completion in the buffer on the name space of tag names in the current tags tables. See Section 22.9 [Symbol Completion], page 240.

22.14 Merging Files with Emerge

It’s not unusual for programmers to get their signals crossed and modify the same program in two different directions. To recover from this confusion, you need to merge the two versions. Emerge makes this easier. See also Section 14.9 [Comparing Files], page 137, for commands to compare in a more manual fashion, and section “Emerge” in The Ediff Manual.

22.14.1 Overview of Emerge

To start Emerge, run one of these four commands:

M-x emerge-files

Merge two specified files.

M-x emerge-files-with-ancestor

Merge two specified files, with reference to a common ancestor.

M-x emerge-buffers

Merge two buffers.

M-x emerge-buffers-with-ancestor

Merge two buffers with reference to a common ancestor in a third buffer.

The Emerge commands compare two files or buffers, and display the comparison in three buffers: one for each input text (the A buffer and the B buffer), and one (the merge buffer) where merging takes place. The merge buffer shows the full merged text, not just the differences. Wherever the two input texts differ, you can choose which one of them to include in the merge buffer.

The Emerge commands that take input from existing buffers use only the accessible portions of those buffers, if they are narrowed (see Section 30.8 [Narrowing], page 374).

If a common ancestor version is available, from which the two texts to be merged were both derived, Emerge can use it to guess which alternative is right. Wherever one current version agrees with the ancestor, Emerge presumes that the other current version is a deliberate change which should be kept in the merged version. Use the ‘with-ancestor’ commands if you want to specify a common ancestor text. These commands read three file or buffer names—variant A, variant B, and the common ancestor.

After the comparison is done and the buffers are prepared, the interactive merging starts. You control the merging by typing special merge commands in the merge buffer. The merge buffer shows you a full merged text, not just differences. For each run of differences between the input texts, you can choose which one of them to keep, or edit them both together.

The merge buffer uses a special major mode, Emerge mode, with commands for making these choices. But you can also edit the buffer with ordinary Emacs commands.

At any given time, the attention of Emerge is focused on one particular difference, called the selected difference. This difference is marked off in the three buffers like this:

    vvvvvvvvvvvvvvvvvvvv

text that differs

    ^^^^^^^^^^^^^^^^^^^^

Emerge numbers all the differences sequentially and the mode line always shows the number of the selected difference.

Normally, the merge buffer starts out with the A version of the text. But when the A version of a difference agrees with the common ancestor, then the B version is initially preferred for that difference.

Emerge leaves the merged text in the merge buffer when you exit. At that point, you can save it in a file with C-x C-w. If you give a numeric

argument to emerge-files or emerge-files-with-ancestor, it reads the

name of the output file using the minibuffer. (This is the last file name those commands read.) Then exiting from Emerge saves the merged text in the output file.

Normally, Emerge commands save the output buffer in its file when you exit. If you abort Emerge with C-], the Emerge command does not save the output buffer, but you can save it yourself if you wish.

22.14.2 Submodes of Emerge

You can choose between two modes for giving merge commands: Fast mode and Edit mode. In Fast mode, basic merge commands are single characters, but ordinary Emacs commands are disabled. This is convenient if you use only merge commands. In Edit mode, all merge commands start with the prefix key C-c C-c, and the normal Emacs commands are also available. This allows editing the merge buffer, but slows down Emerge operations.

Use e to switch to Edit mode, and C-c C-c f to switch to Fast mode. The mode line indicates Edit and Fast modes with ‘E’ and ‘F’.

Emerge has two additional submodes that affect how particular merge commands work: Auto Advance mode and Skip Prefers mode.

If Auto Advance mode is in effect, the a and b commands advance to the next difference. This lets you go through the merge faster as long as you simply choose one of the alternatives from the input. The mode line indicates Auto Advance mode with ‘A’.

If Skip Prefers mode is in effect, the n and p commands skip over differences in states prefer-A and prefer-B (see Section 22.14.3 [State of Difference], page 253). Thus you see only differences for which neither version is presumed “correct.” The mode line indicates Skip Prefers mode with ‘S’.

Use the command s a (emerge-auto-advance-mode) to set or clear Auto Advance mode. Use s s (emerge-skip-prefers-mode) to set or clear Skip Prefers mode. These commands turn on the mode with a positive argument, turns it off with a negative or zero argument, and toggle the mode with no argument.

22.14.3 State of a Difference

In the merge buffer, a difference is marked with lines of ‘v’ and ‘^’ characters. Each difference has one of these seven states: A The difference is showing the A version. The a command always

produces this state; the mode line indicates it with ‘A’.

B The difference is showing the B version. The b command always produces this state; the mode line indicates it with ‘B’.

default-A default-B The difference is showing the A or the B state by default, because

you haven’t made a choice. All differences start in the default-A state (and thus the merge buffer is a copy of the A buffer), except those for which one alternative is “preferred” (see below). When you select a difference, its state changes from default-A or default-B to plain A or B. Thus, the selected difference never has state default-A or default-B, and these states are never displayed in the mode line. The command d a chooses default-A as the default state, and d b chooses default-B. This chosen default applies to all differences which you haven’t ever selected and for which no alternative is preferred. If you are moving through the merge sequentially, the differences you haven’t selected are those following the selected one. Thus, while moving sequentially, you can effectively make the A version the default for some sections of the merge buffer and the B version the default for others by using d a and d b between sections.

prefer-A prefer-B The difference is showing the A or B state because it is preferred.

This means that you haven’t made an explicit choice, but one alternative seems likely to be right because the other alternative agrees with the common ancestor. Thus, where the A buffer agrees with the common ancestor, the B version is preferred, because chances are it is the one that was actually changed. These two states are displayed in the mode line as ‘A*’ and ‘B*’.

combined The difference is showing a combination of the A and B states, as a result of the x c or x C commands. Once a difference is in this state, the a and b commands don’t do anything to it unless you give them a numeric argument. The mode line displays this state as ‘comb’.

22.14.4 Merge Commands

Here are the Merge commands for Fast mode; in Edit mode, precede them

with C-c C-c:

p Select the previous difference. n Select the next difference. a Choose the A version of this difference. b Choose the B version of this difference. C-u n j Select difference number n. . Select the difference containing point. You can use this command

in the merge buffer or in the A or B buffer.

q Quit—finish the merge. C-] Abort—exit merging and do not save the output. f Go into Fast mode. (In Edit mode, this is actually C-c C-c f.) e Go into Edit mode. l Recenter (like C-l) all three windows.

- Specify part of a prefix numeric argument. digit Also specify part of a prefix numeric argument. d a Choose the A version as the default from here down in the merge

buffer.

d b Choose the B version as the default from here down in the merge

buffer.

c a Copy the A version of this difference into the kill ring. c b Copy the B version of this difference into the kill ring. i a Insert the A version of this difference at point. i b Insert the B version of this difference at point. m Put point and mark around the difference. ^Scroll all three windows down (like M-v). v Scroll all three windows up (like C-v). < Scroll all three windows left (like C-x <). > Scroll all three windows right (like C-x >). | Reset horizontal scroll on all three windows. x 1 Shrink the merge window to one line. (Use C-u l to restore it to

full size.) x c Combine the two versions of this difference (see Section 22.14.6

[Combining in Emerge], page 255).

x f Show the names of the files/buffers Emerge is operating on, in a Help window. (Use C-u l to restore windows.)

x j Join this difference with the following one. (C-u x j joins this difference with the previous one.)

x s Split this difference into two differences. Before you use this command, position point in each of the three buffers at the place where you want to split the difference.

x t Trim identical lines off the top and bottom of the difference. Such lines occur when the A and B versions are identical but differ from the ancestor version.

22.14.5 Exiting Emerge

The q command (emerge-quit) finishes the merge, storing the results into the output file if you specified one. It restores the A and B buffers to their proper contents, or kills them if they were created by Emerge and you haven’t changed them. It also disables the Emerge commands in the merge buffer, since executing them later could damage the contents of the various buffers.

C-] aborts the merge. This means exiting without writing the output file. If you didn’t specify an output file, then there is no real difference between aborting and finishing the merge.

If the Emerge command was called from another Lisp program, then its return value is t for successful completion, or nil if you abort.

22.14.6 Combining the Two Versions

Sometimes you want to keep both alternatives for a particular difference. To do this, use x c, which edits the merge buffer like this:

    #ifdef NEW

version from A buffer

    #else /* not NEW */

version from B buffer

    #endif /* not NEW */

While this example shows C preprocessor conditionals delimiting the two alternative versions, you can specify the strings to use by setting the variable

emerge-combine-versions-template to a string of your choice. In the

string, ‘%a’ says where to put version A, and ‘%b’ says where to put version B. The default setting, which produces the results shown above, looks like this:

    "#ifdef NEW\n%a#else /* not NEW */\n%b#endif /* not NEW */\n"

22.14.7 Fine Points of Emerge

During the merge, you mustn’t try to edit the A and B buffers yourself. Emerge modifies them temporarily, but ultimately puts them back the way they were.

You can have any number of merges going at once—just don’t use any one buffer as input to more than one merge at once, since the temporary changes made in these buffers would get in each other’s way.

Starting Emerge can take a long time because it needs to compare the files fully. Emacs can’t do anything else until diff finishes. Perhaps in the future someone will change Emerge to do the comparison in the background when the input files are large—then you could keep on doing other things with Emacs until Emerge is ready to accept commands.

After setting up the merge, Emerge runs the hook emerge-startup-hook (see Section 31.2.3 [Hooks], page 392).

22.15 C Modes and Java Mode

This section describes special features available in C, C++, Objective-C and Java modes.

22.15.1 C Mode Motion Commands

This section describes commands for moving point, in C mode and related modes. C-c C-u Move point back to the containing preprocessor conditional, leav-

ing the mark behind. A prefix argument acts as a repeat count. With a negative argument, move point forward to the end of the containing preprocessor conditional. When going backwards, #elif is treated like #else followed by #if. When going forwards, #elif is ignored.

C-c C-p Move point back over a preprocessor conditional, leaving the mark behind. A prefix argument acts as a repeat count. With a negative argument, move forward.

C-c C-n Move point forward across a preprocessor conditional, leaving the mark behind. A prefix argument acts as a repeat count. With a negative argument, move backward.

M-a Move point to the beginning of the innermost C statement (cbeginning-of-statement). If point is already at the beginning of a statement, move to the beginning of the preceding statement. With prefix argument n, move back n − 1 statements. If point is within a string or comment, or next to a comment (only whitespace between them), this command moves by sentences instead of statements.

When called from a program, this function takes three optional arguments: the numeric prefix argument, a buffer position limit (don’t move back before that place), and a flag that controls whether to do sentence motion when inside of a comment.

M-e Move point to the end of the innermost C statement; like M-a except that it moves in the other direction (c-end-of-statement).

M-x c-backward-into-nomenclature

Move point backward to beginning of a C++ nomenclature section or word. With prefix argument n, move n times. If n is negative, move forward. C++ nomenclature means a symbol name in the style of NamingSymbolsWithMixedCaseAndNoUnderlines; each capital letter begins a section or word. In the GNU project, we recommend using underscores to separate words within an identifier in C or C++, rather than using case distinctions.

M-x c-forward-into-nomenclature

Move point forward to end of a C++ nomenclature section or word. With prefix argument n, move n times.

22.15.2 Electric C Characters

In C mode and related modes, certain printing characters are “electric”— in addition to inserting themselves, they also reindent the current line and may insert newlines. This feature is controlled by the variable c-auto-newline. The “electric” characters are {, }, :, #, ;, ,, <, >, / and *.

Electric characters insert newlines only when the auto-newline feature is enabled (indicated by ‘/a’ in the mode line after the mode name). This feature is controlled by the variable c-auto-newline. You can turn this feature on or off with the command C-c C-a:

C-c C-a Toggle the auto-newline feature (c-toggle-auto-state). With a prefix argument, this command turns the auto-newline feature on if the argument is positive, and off if it is negative.

The colon character is electric because that is appropriate for a single colon. But when you want to insert a double colon in C++, the electric behavior of colon is inconvenient. You can insert a double colon with no reindentation or newlines by typing C-c ::

C-c : Insert a double colon scope operator at point, without reindenting the line or adding any newlines (c-scope-operator).

The electric # key reindents the line if it appears to be the beginning of a preprocessor directive. This happens when the value of c-electric-pound-behavior is (alignleft). You can turn this feature off by setting

c-electric-pound-behavior to nil.

The variable c-hanging-braces-alist controls the insertion of newlines before and after inserted braces. It is an association list with elements of the following form: (syntactic-symbol . nl-list). Most of the syntactic symbols that appear in c-offsets-alist are meaningful here as well.

The list nl-list may contain either of the symbols before or after, or both; or it may be nil. When a brace is inserted, the syntactic context it defines is looked up in c-hanging-braces-alist; if it is found, the nl-list is used to determine where newlines are inserted: either before the brace, after, or both. If not found, the default is to insert a newline both before and after braces.

The variable c-hanging-colons-alist controls the insertion of newlines before and after inserted colons. It is an association list with elements of the following form: (syntactic-symbol . nl-list). The list nl-list may contain either of the symbols before or after, or both; or it may be nil.

When a colon is inserted, the syntactic symbol it defines is looked up in this list, and if found, the nl-list is used to determine where newlines are inserted: either before the brace, after, or both. If the syntactic symbol is not found in this list, no newlines are inserted.

Electric characters can also delete newlines automatically when the autonewline feature is enabled. This feature makes auto-newline more acceptable, by deleting the newlines in the most common cases where you do not want them. Emacs can recognize several cases in which deleting a newline might be desirable; by setting the variable c-cleanup-list, you can specify which of these cases that should happen. The variable’s value is a list of symbols, each describing one case for possible deletion of a newline. Here are the meaningful symbols, and their meanings:

brace-else-brace

Clean up ‘} else {’ constructs by placing the entire construct on a single line. The clean-up occurs when you type the ‘{’ after the else, but only if there is nothing but white space between the braces and the else.

brace-elseif-brace

Clean up ‘} else if (...) {’ constructs by placing the entire construct on a single line. The clean-up occurs when you type the ‘{’, if there is nothing but white space between the ‘}’ and ‘{’ aside from the keywords and the if-condition.

empty-defun-braces

Clean up empty defun braces by placing the braces on the same line. Clean-up occurs when you type the closing brace.

defun-close-semi

Clean up the semicolon after a struct or similar type declaration, by placing the semicolon on the same line as the closing brace. Clean-up occurs when you type the semicolon.

list-close-comma

Clean up commas following braces in array and aggregate initializers. Clean-up occurs when you type the comma.

scope-operator

Clean up double colons which may designate a C++ scope operator, by placing the colons together. Clean-up occurs when you type the second colon, but only when the two colons are separated by nothing but whitespace.

22.15.3 Hungry Delete Feature in C

When the hungry-delete feature is enabled (indicated by ‘/h’ or ‘/ah’ in the mode line after the mode name), a single ^{⟨}\underline{\overline{DEL}}^{⟩} command deletes all preceding whitespace, not just one space. To turn this feature on or off, use

C-c C-d:

C-c C-d Toggle the hungry-delete feature (c-toggle-hungry-state). With a prefix argument, this command turns the hungry-delete feature on if the argument is positive, and off if it is negative.

C-c C-t Toggle the auto-newline and hungry-delete features, both at once

(c-toggle-auto-hungry-state).

The variable c-hungry-delete-key controls whether the hungry-delete feature is enabled.

22.15.4 Other Commands for C Mode

C-M-h Put mark at the end of a function definition, and put point at the

beginning (c-mark-function).

M-q Fill a paragraph, handling C and C++ comments (c-fill-paragraph). If any part of the current line is a comment or within a comment, this command fills the comment or the paragraph of it that point is in, preserving the comment indentation and comment delimiters.

C-c C-e Run the C preprocessor on the text in the region, and show the result, which includes the expansion of all the macro calls (cmacro-expand). The buffer text before the region is also included in preprocessing, for the sake of macros defined there, but the output from this part isn’t shown. When you are debugging C code that uses macros, sometimes it is hard to figure out precisely how the macros expand. With this command, you don’t have to figure it out; you can see the expansions.

C-c C-\ Insert or align ‘\’ characters at the ends of the lines of the region (c-backslash-region). This is useful after writing or editing a C macro definition. If a line already ends in ‘\’, this command adjusts the amount of whitespace before it. Otherwise, it inserts a new ‘\’. However, the last line in the region is treated specially; no ‘\’ is inserted on that line, and any ‘\’ there is deleted.

M-x cpp-highlight-buffer

Highlight parts of the text according to its preprocessor conditionals. This command displays another buffer named ‘*CPP Edit*’, which serves as a graphic menu for selecting how to display particular kinds of conditionals and their contents. After changing various settings, click on ‘[A]pply these settings’ (or go to that buffer and type a) to rehighlight the C mode buffer accordingly.

C-c C-s Display the syntactic information about the current source line

(c-show-syntactic-information). This is the information

that directs how the line is indented.

22.15.5 Comments in C Modes

C mode and related modes use a number of variables for controlling comment format.

c-comment-only-line-offset

Extra offset for line which contains only the start of a comment. It can be either an integer or a cons cell of the form (non-anchored-offset . anchored-offset), where non-anchored-offset is the amount of offset given to non-column-zero anchored comment-only lines, and anchored-offset is the amount of offset to give column-zero anchored comment-only lines. Just an integer as value is equivalent to (val . 0).

c-comment-start-regexp

This buffer-local variable specifies how to recognize the start of a comment.

c-hanging-comment-ender-p

If this variable is nil, c-fill-paragraph leaves the comment terminator of a block comment on a line by itself. The default value is t, which puts the comment-end delimiter ‘*/’ at the end of the last line of the comment text.

c-hanging-comment-starter-p

If this variable is nil, c-fill-paragraph leaves the starting delimiter of a block comment on a line by itself. The default value is t, which puts the comment-start delimiter ‘/*’ at the beginning of the first line of the comment text.

22.16 Fortran Mode

Fortran mode provides special motion commands for Fortran statements and subprograms, and indentation commands that understand Fortran conventions of nesting, line numbers and continuation statements. Fortran mode has its own Auto Fill mode that breaks long lines into proper Fortran continuation lines.

Special commands for comments are provided because Fortran comments are unlike those of other languages. Built-in abbrevs optionally save typing when you insert Fortran keywords.

Use M-x fortran-mode to switch to this major mode. This command runs the hook fortran-mode-hook (see Section 31.2.3 [Hooks], page 392).

22.16.1 Motion Commands

Fortran mode provides special commands to move by subprograms (functions and subroutines) and by statements. There is also a command to put the region around one subprogram, convenient for killing it or moving it.

C-M-a Move to beginning of subprogram (beginning-of-fortran-

          subprogram).
C-M-e Move to end of subprogram (end-of-fortran-subprogram).

C-M-h Put point at beginning of subprogram and mark at end (mark-

          fortran-subprogram).

C-c C-n Move to beginning of current or next statement (fortran-next-

          statement).

C-c C-p Move to beginning of current or previous statement (fortran-

          previous-statement).

22.16.2 Fortran Indentation

Special commands and features are needed for indenting Fortran code in order to make sure various syntactic entities (line numbers, comment line indicators and continuation line flags) appear in the columns that are required for standard Fortran.

22.16.2.1 Fortran Indentation Commands

⟨\underline{\overline{TAB}}^{⟩} Indent the current line (fortran-indent-line).

C-j Indent the current and start a new indented line (fortran-

          indent-new-line).

C-M-j Break the current line and set up a continuation line.

M-^Join this line to the previous line.

C-M-q Indent all the lines of the subprogram point is in (fortran-

          indent-subprogram).

Fortran mode redefines ^{⟨}\underline{\overline{TAB}}^{⟩} to reindent the current line for Fortran (fortran-indent-line). This command indents line numbers and continuation markers to their required columns, and independently indents the body of the statement based on its nesting in the program.

The key C-j runs the command fortran-indent-new-line, which reindents the current line then makes and indents a new line. This command is useful to reindent the closing statement of ‘do’ loops and other blocks before starting a new line.

   The key C-M-q runs fortran-indent-subprogram, a command to rein-

dent all the lines of the Fortran subprogram (function or subroutine) containing point.

The key C-M-j runs fortran-split-line, which splits a line in the appropriate fashion for Fortran. In a non-comment line, the second half becomes a continuation line and is indented accordingly. In a comment line, both halves become separate comment lines.

M-^runs the command fortran-join-line, which is more or less the inverse of fortran-split-line. It joins the current line to the previous line in a suitable way for Fortran code.

22.16.2.2 Continuation Lines

Most modern Fortran compilers allow two ways of writing continuation lines. If the first non-space character on a line is in column 5, then that line is a continuation of the previous line. We call this fixed format. (In GNU Emacs we always count columns from 0.) The variable fortrancontinuation-string specifies what character to put on column 5. A line that starts with a tab character followed by any digit except ‘0’ is also a continuation line. We call this style of continuation tab format.

Fortran mode can make either style of continuation line, but you must specify which one you prefer. The value of the variable indent-tabs-mode controls the choice: nil for fixed format, and non-nil for tab format. You can tell which style is presently in effect by the presence or absence of the string ‘Tab’ in the mode line.

If the text on a line starts with the conventional Fortran continuation marker ‘$’, or if it begins with any non-whitespace character in column 5, Fortran mode treats it as a continuation line. When you indent a continuation line with ^{⟨}\underline{\overline{TAB}}^{⟩}, it converts the line to the current continuation style. When you split a Fortran statement with C-M-j, the continuation marker on